A Complete Bibliography of Publications in ACM Transactions on Programming Languages and Systems

(TOPLAS)

Preston Briggs Tera Computer Company

2815 Eastlake East Seattle, WA 98102

USA

Tel: +1 206 325-0800

E-mail: preston@tera.com

and

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

11 October 2024 Version 2.155

Title word cross-reference

(k) [ADGM91, BL94b, KM81]. 2 [Dam03]. 3 [SRW02]. ^{+} [Han81a]. ^{T M} [Bla03]._{/} [AW82].

∥ [DDDCG02]. A [DES12]. R [JMSY92]. RLin [VR95]. ℓ [ADG^{+}94]. O(n log n) [Pet82]. ϕ [CF95, DR05]. π [ABL03].

-calculus [ABL03]. -Composition [Ras24].

-Exclusion [ADG^{+}94]. -function [DR05]. -Nodes [CF95]. -Tree [Han81a]. -valued [SRW02].

11 [ND16]. 16 [TGT20].

2 [SZLY21]. 2018 [Ahm20]. 2020 [Mül21]. 2021 [Yos22]. 256 [App15].

40 [TGT20].

568 [Han81b].

8 [Ano18].

90 [DP99]. 95 [WJS^{+}00].

Abstract [ABG24, BGL93, BK11, CMB^{+}95, CFG^{+}97, DGG97, DC22, DHH^{+}24, DLR16, ELS^{+}14, EO80, GS98, GMP24a, HL82, JPP91, KRR18, Lan80, LO94, LV94, LM18, LR13, Loe87, MSJ94, MXZ^{+}22, MP88, SS98, She91, SCS24, TY21, Wei89, van88, ABS09, BDL^{+}08, BdlBH99, Leu04, RM07, SYYH07, SJ03]. Abstraction [BNNN22, CGL94, CL94, Der85, GMH81, GKM20, SM81, BMR05, BBTS07, GMZ00, LN02, LH08, MOSS96, PR07, Ell82]. abstractions [BCF04]. Access [ABLP93, BCC04, KS83, Mis86, NBG13, SR21, HR02, HO07, KSK07, PHP02, PSS05]. Access-Right [KS83]. accessed [RR05]. Accessing [CB80]. Accommodating [GMP24a]. Account [ABC^{+}21]. Accumulation [Bir84, Bir85]. Accumulators [Cam89]. Accuracy [LVRG21, CEG07, HDH02]. accurate [CG04, VBLG04, VALG05]. ACE [Le 88]. ACM [Ano18, TGT20, Kro82, Kro83, Kro84, Kro85, Kro86, Kro87, Kro88, Kro89, Kro90, Kro91, Kro92]. Across [NSTD^{+}15]. Action [BKS88]. Actions [Lam94, LS83]. Active [SR21]. Activity [Bar81, MTG80].

Actor [TCP^{+}17]. Acyclic [BE94, JF81]. Ad [MDCB91, PS08]. Ada [Bak82, Dil90, Hil88, LP80, WJS^{+}00]. Adaptation [Dha91]. Adaptive [ABH06, HOYY18, PXL95, TCVB14, UJ92, RD03]. adaptors [YS97]. Addendum [Bir85]. Adding [ACW90, BN94]. Addition [CBMO19]. Addressing [Hol87, ZP10]. Adequacy [KKSD94, Wey83]. adjusting [ABB^{+}09]. Adversities [GMP24a]. advice [WKD04]. Æminium [SNS^{+}14]. Affine [BAC16, BCEM15, CFNH18, DG19, ELS^{+}14, VJB12]. Affix [GF85]. agents [BCC04]. aggregate [LSLR05]. Ahead [BLH12, DP82]. al [Ano18, TGT20]. Alarms [LLK^{+}17]. Algebra [Koz97, Wil82a, KBC^{+}99]. Algebraic [BP82, BWP87, CIJGP18, CGG^{+}19, Jen97, Lin93, SV20, JB06, SP07]. Algorithm [AB81, Bak82, BB79, BAC16, BP82, Dan23, DSW82, Dha91, DP93, GHS83, HL22, Hua90, Hud91, JJCO19, LV94, LY98, Lei90, LT79, LH91, MM82, MC82a, Pet82, SH89, TB98, Wis79, BKRW98, BH99, DR05, DVD07, JNZ06, Van96a, Van96b, Han81b, BKRW05]. Algorithmic [BP82, CFNH18, GM12, JZX^{+}24, Loe87]. Algorithms [Apt86, BA84, CIJGP18, CGG^{+}19, CS95, CN83, GLO88, KRS84, KKM90, Kro82, Kro83, Kro84, Kro85, Kro86, Kro87, Kro88, Kro89, Kro90, Kro91, Kro92, MM89, RD87, RH87, RP88, TM93, WW95, Apt00, DAS98, GC01, ZGZ05]. Alias [Hor97, HBCC99, RRSY08]. Aliasing [Boe85, Ram94, RLS^{+}01]. All-Purpose [Spo86]. Allocating [ZP07]. Allocation [BB79, Bre89, BCT94, CH90, CS95, FLBB89, GSO94, LCBS19, Rob79, SH89, CGS^{+}03, HCS10, LGAT00, PS99, PF96, RDG08, SRM10, TP04]. Alma [ABPS98]. Alma-O [ABPS98]. almost [Duc08, Ram99]. Alternative [Gho93, GH80, Zav85]. Alway [Gri79]. ambients [BCC04, LS03, MH06]. Ambiguity [Tho94]. amortized [HAH12].

Amulet [VHM^{+}01]. Analyses [AC94, CC95, CFM94, TN19, KSV96, SJ03]. Analysis [AKNP17, ABE^{+}05, AD98, Bac84, BNN18, BC85b, Blo94, BE13, Bur90a, CFNH18, CFG19, CDK^{+}18, CMN91, DKKL18, DL93, Deb95, DP97, DC22, DAW88, FPS19, FJK^{+}17, GNS^{+}15, GKM20, GDF23, GJ93, HP96, HL22, HOYY18, Hil88, Hor97, HP24, ISY88, Jen97, JJCO19, KD94, LLK^{+}17, LTMS20, LR13, LHR19, LWR21, McG82, MRGP20, MWB94, MOS07b, OHL^{+}14, OLH^{+}16, Pal95, PO95, PCC85, PP91, PW94, PW98, Pur91, RTD83, RTP17, RRB19, RP88, SJW23, SR95, SSS83, SGL98, SS13, ABB^{+}09, BDFZ09, BAL07, Bla03, Blu99, BCG^{+}07, CSW06, Cha02, CGS^{+}03, CKT86, DDV99, DGS97, FF99, GHB^{+}96, GJ05, GZ04, GCRN11, HAM^{+}05, HPMS00, HBCC99, HVDH07, HAH12, IK05, JLRS10, KBC^{+}99, KK07, KSK07, LP00, LH08, MPM03, NS13, PHP02, Pal98, PKH07, Ram00, Rep00]. analysis [RSL10, RD97, RRSY08, RR03, RR05, RLS^{+}01, SRW98, SRW02, STS03, SdSCP13, SS96, ST00a, WP10, WJ98, ZSD09, dHB^{+}96]. analyze [DMM01, VBLG04]. analyzer [SMP10]. Analyzing [AW85, BEF^{+}16, CFP^{+}04, GMM99]. And/Or [Har80]. Annotations [Bur84]. AOP [DES12]. APL [Bud84, GD82, Hob84]. Applicability [DAW88, How80, LS98]. applicable [Gom92]. Application [CD79, DF80, DF81, LBN17, LR13]. Applications [BLRS12, Bou88, MRGP20, RKW^{+}23, SR21, BALP06, CMLC06, NR06]. Applicative [AC94, KS86]. Applying [JZX^{+}24]. apprentice [MP02]. Approach [AKNP17, ABR81, AR80, BAC16, BP82, Bur90a, CH90, CD79, DS90, Ell82, ES97, FT94, GGL15, Har80, Hes88, HP24, KKW14, Lam79, Lam80, Lee86, LTMS20, MW80, MDCB91, ND16, OA88, Sam80, Spo86, SM81, SNS^{+}14, Bou05, CRN^{+}08, DHM^{+}12, FGM^{+}07a, JLRS10, KV00, LP80, MBT09, PSS05, PCJD08, RC03, SP07, WS97]. Approximation [ABG24]. approximations [BGP99]. Apt [Moi83]. architected [ZP07]. Architecture [Wal92]. Architectures [Han94, KPF95, NSTD^{+}15, PAS^{+}15]. Arising [Bac84]. Arithmetic [Fis80, GNS^{+}15, Hen83, LdR81, MOS07b]. ARM [FKW98, ADG^{+}21]. Armada [LCK^{+}22]. Armed [ADG^{+}21]. Array [CGST95, CG95, LS79, Per79, PW98, JB06, LSLR05, NI05, PHP02, RMH06, RR05, ZCG^{+}07]. array-valued [RMH06]. Arrays [BBC16]. Article [Ano18, TGT20]. ASF [VHKO02]. aspect [DWWW08, WKD04]. aspect-oriented [DWWW08, WKD04]. AspectML [DWWW08]. Aspects [Bor81, Set83]. assembly [AAR^{+}10, MWCG99]. Assertions [BKB80]. Assessing [BDH^{+}16, Wey83]. Assets [COE^{+}20]. Assignment [BM94, CFR^{+}91, GL80, GPF08, LDK^{+}96]. Assisted [HCHP92]. Assisting [Fea82]. Associated [PPS79]. associativity [Cha02]. Associons [Rem81]. assume [HQRT02]. assume-guarantee [HQRT02]. Assumptions [ES97]. AST [GVC15]. Asynchronous [Bag89, GLO88, Mis86, GM12, HR02]. ATL [WSH15]. Atomic [WL85, Wei90, AE01]. Atomicity [JLP^{+}14, Wei89, FFLQ08]. Attacks [SBE^{+}19]. Attribute [CP95, Hud91, JP81, Jon90, Kat84, KR79, MK94, RD87, WW95, Boy96, CP96, Wu04]. Attributes [HT86]. Author [Ano86a, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano94, Ano95, Ano98, Bur91]. authorization [FGM07b]. Authors [Ano82, Ano83, Ano84, Ano85, Ano86b, Ano87, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b]. auto [ZP10]. auto-addressing [ZP10]. Automata [BMW91, CBMO19, ES97, Pro95, KV00].

Automata-Theoretic [ES97, KV00]. Automated [GRSK^{+}11, KZC15, KF00, LCK^{+}22, SSFZ^{+}23, Sok87, JNGG10]. Automatic [AKNP17, AK87, Ano02a, BBC16, Cat80, CES86, DS90, KK98, Le 88, LK02, LS04, MS83, PZJ05, RH87, SSS81, SLC03, She91, VS22, Wat91, Wha94, ABHI11, ATD08, BdlBH99, CRN^{+}08, ZCG^{+}07]. Automatically [Slo95]. Automating [GKL94, MTSS09]. Avoidance [FGL94]. Aware [BPRB23, MQ05]. Axiomatic [AR80, App94a, Boe85, Sou84, YB87, YB88, LP80]. Axioms [Mis86].

B [Han81a]. backpropagator [PS08]. Backtracking [Lin79, VR95, FM87a]. Backward [DL18, Mye18]. Balanced [AS80, PB80, vHK00]. Barrier [CHMY19]. Base [NcS20, LS98]. Based [BPP16, BGL93, Bur90a, CGJ^{+}97a, CI84, CP95, CH90, CPS93, DVLM15, DLR16, EGP14, GG85, HT86, JTM98, Kai89, KH92, KR79, LFF14, PW98, RTD83, SR95, SGL98, Ste18, SNS^{+}14, TY18, Wat94, WGS92, vPS81, BFG08, BMR01, BHM^{+}07, BCG^{+}07, CTT07, DDV99, Eug07, FF99, HBJ98, JZX^{+}24, JKT23, KBC^{+}99, KK07, KC01, LP00, LH08, LGAT00, MTSS09, ML21, MTK21, MH06, Pal98, PPT08, PCJD08, SP97, SP07, SMP10, SYN06, BDP14, WGS93, WM12]. Basic [CGG^{+}19, GLR83]. Bayesian [HOYY18]. BDD [LH08]. BDD-based [LH08]. Be [Bee94, Coh91, Wir91, CG04, LP99]. Behavior [KLS92, GMM99, VBLG04]. Behavioral [LN15, LW94]. Behavioural [SV20]. Being [Cop94]. benefits [GMP^{+}00]. Better [Gri79, Lam88]. between [BS88]. Beyond [GSW95]. BI [BBTS07]. BI-hyperdoctrines [BBTS07]. Bidirectional [DP93, MMR95, FGM^{+}07a, GPWZ08]. Big [KEA^{+}24]. binaries [STSP05]. Binary [Sip82, SCR24, DDD05, MMM^{+}07, RC03, YF09]. binding [ACE96]. Birrell [MDJ05]. Bisimulation [FDY12, MH06, San09]. bisimulation-based [MH06]. Bisimulations [SV19, SKS11]. Bit [CDK^{+}18, KD94, KK07]. Bit-Precise [CDK^{+}18]. bitvector [KSV96]. Bliss [GNS^{+}15]. Block [LS81, Mur91]. Block-Structured [LS81]. Blockchain [COE^{+}20]. Blocked [FTJ95]. Blocks [Jag94]. Boolean [XA07]. Bootstrapping [App94a]. Borrowing [Pea21]. Both [KZC15]. bottlenecks [RD03]. Bottom [BGL93, GCRN11]. Bottom-Up [BGL93, GCRN11]. bound [KK07, NI05]. Bounded [ADG^{+}94, ITF^{+}22, MXZ^{+}22, LLOY23]. Bounds [CP17, FNBG20, ISIRS22, PW94, BP12, CEI^{+}07, RR05, SS05a]. Box [WLBF16]. boxed [BCC04]. Branch [CGJ^{+}97a, CEG07, YUW02, YS99]. Branches [WZ91, RC03]. Branching [CBMO19]. Broad [DAW88]. Buddy [Kau84]. Buffer [Žic94]. bugs [HCS10]. Building [Jag94]. BURS [Pro95]. Bus [Pur94]. Bytecode [SA99, BDL^{+}08, CSCM00, FM99, GPF08, KR01, Qia00, SMP10, WR08]. Byzantine [LSP82].

’C [PHEK99, BR97, HSS^{+}14, MRGP20, ND16, PKH07, PFH11, Ven95]. C# [BCF04]. C/C [ND16]. C0 [DMZ^{+}24]. C11 [JP17]. Cache [GMM99, KLS92, MMM^{+}07, SS96, VBLG04]. Caching [ABM93, FK85, KS86, LST98]. Calculational [Bou06]. calculi [ABS09]. Calculus [ABLP93, BKL^{+}97, BN94, Gom92, Kob98, LBMTT22, Miq19, MRG88, Nel89, Oho95, WM95, ABL03, AH10, BG22, Bou05, Bou06, BCC04, DES12, HR02, IPW01, Jay04, TA08a, KPT99]. Call [DP97, GL80, dBH21, GC01, HL05, KK07, SW97a]. Call-by-Value [dBH21, HL05, SW97a].

Calls [BNN18, Coh83, Coh85, FF08]. Can [Boe85, Coh91, Wir91, CG04]. Capabilities [SDB20, WCM00]. capability [TA08a]. Capturing [BGOL^{+}23]. Carlo [FL15]. carrying [AM01]. Case [CFG19, FTJ95, WW95, BdlBH99, KF03]. Cats [AMT14, ADG^{+}21]. Causally [LB22]. Cause [Cas95]. CCP [EGM01]. CCured [NCH^{+}05]. Cedar [SZBH86]. Cells [ISY88]. Centered [CHY12]. Centers [KRS84]. Centralized [HM84]. centric [DHM^{+}12]. Certificate [BGKR09, BK11]. certified [STSP05]. CFLOBDDs [SCR24]. CHAD [VS22]. Chaining [LS80]. Chains [HS94]. challenge [MP02]. change [BA08, CP96, Lee09]. Changes [Ber94, MTSS09]. changing [MP07]. channel [SCP23]. Chariots [PB97]. CHC [MTK21]. CHC-based [MTK21]. Check [AP94]. checked [KN06]. Checker [CMPP23, NP08]. Checking [Car95, CGL94, ES97, FF08, GL94, ND16, AY01, ACM11, BGP99, FFLQ08, HQRT02, JJD98, KF10, KV00, NI05, SG04, VJB12, YMW97]. Checks [CG95, CEI^{+}07]. Chocola [SDD21]. Choice [BN94, JCMM19]. Choral [GMP24b]. Choreographic [GMP24b]. CIRCAL [Mil85]. Circuits [HRL^{+}23, LLOY23]. Circular [Jon90, Pet82]. Circularity [WW95, Wu04]. Clarification [PA86a]. Class [CBMO19, HKMN94, Han92, SJ03, SDTF13, HS11, MH04, ST00a]. Classes [SDTF13, WT11, HHPW96, HMS06]. Classical [JSB^{+}12, Miq19]. Classification [KZC15]. Classifying [GSW95]. Claus [WP10]. Cliché [Wat94]. Cliché-Based [Wat94]. client [RKW^{+}23]. Clique [GSO94]. Closure [Pal95, SW97b, SA00]. CLP [DHM00, GLMM05, JMSY92, KMM^{+}98, VR95]. Clustering [LLK^{+}17]. Clusters [BGH^{+}13, HBG^{+}09]. coalescing [GA96, Hai05, PM04]. Coalgebraic [KBP22]. Coalgebras [KBP22]. Code [AGT89, BHM^{+}19, Cat80, Cop94, DF84, FGL94, GF85, Hen82, HG83, JSB^{+}12, KRS94, LR13, LCK^{+}22, ND16, Rob79, TvS82, VP23, Wan82, AM01, DEMD00, Hai98, HBG^{+}09, HK07, JNZ06, LDK^{+}96, MSRR00, ME97, Oho07, PHEK99, WS97, vHK00, CM93, Pem83, WST85]. Coder [SBS22]. Cohen [Coh85]. coherence [SS96]. coinduction [San09]. Collect [JCMM19]. Collecting [HY91]. Collection [BA84, CN83, DSW82, Lan80, TM93, URJ18, WLBF16, BALP06, HDH02, PBK^{+}07, Piq96]. Collector [YW22, BBYG^{+}05, LP06, TSBR08]. Coloring [BCT94, CH90, GSO94]. combination [BCG^{+}07]. Combinator [FRW90, KLS92]. Combinatorial [LCBS19]. Combinators [FGM^{+}07a, KS88, KS89]. Combinatory [RS97, VS22]. Combining [Ber94, BP82, CC95, CMB^{+}95]. Come [LH91]. Comments [AB94, KS79, LaL84, NN86, Sor89]. Communicating [AFdR80, GC86, HM84, MW84, MC82b, Moi83, Oss83, PP91, Pur91, Sou84, Ber80, KS79]. Communication [Ang89, CHY12, FJK^{+}17, FY85, Gel85, Hua90, LH91, MB83, vPS81, KBC^{+}99, Mil85, SWU10, WM12]. Communication-Centered [CHY12]. Communications [RS84b]. Commutativity [RD97, Apt00, Cha02]. Compact [BC79, Sip82, Wad90]. Compactification [RH87]. Compacting [CN83]. Compaction [CP17, Wis79, BP12, DDD05, DEMD00]. Comparative [GDF23, WCW90, WCW91]. Comparing [Hai05]. Comparison [CN83]. Compartmentalized [WLBF16]. CompCert [BDP14]. Compensation [FGL94]. Compilation [ABC^{+}21, DLR16, FKW98, FL91, JLP^{+}14, JF81, Oho95, PAS^{+}15, PG21, PKWC24, Sit79, KMM^{+}98, LST02, LDM07, SYN06].

Compile [ABR81, GW99, Hol87, Tra08]. Compile-Time [ABR81, Hol87, GW99, Tra08]. Compiler [ABC^{+}21, App94a, BPRB23, Bud84, CM86b, DK17, DEMD00, FT94, FGL94, JSB^{+}12, NN24, Rei83, Slo95, Son87, Wha94, YBL16, Ano02a, CMLC06, DHS09, GMM99, KN06, PE08, PHEK99, SYK^{+}05, VHKO02]. Compiler-Driven [YBL16]. Compiler-Level [BPRB23]. Compilers [BDFH97, DDH84, HP96, Han94, BGKR09, RD97, SYN06]. Compiling [Cha93, CH87, Fis80, Set83, VHKO02]. Complementation [CFG^{+}97]. Complete [BDFH97, WM95]. Completeness [LBN17, TB95, WGS92, dBH21, Wu04, WGS93]. completion [KR01]. Complexity [BG22, BEF^{+}16, Deb95, ISIRS22, Le 88, dBH21, RRSY08, SSD09]. component [LS98, YS97]. component-base [LS98]. Componential [FF99]. Components [CIJGP18]. Composability [PKWC24]. Composable [SDD21]. Composing [AL93, HKMN94]. Composite [Fea87]. Composition [Ras24, AH10, Pau01]. Compositional [DSFG21, Fos96, Jon94, JTM98, LFF14, SZLY21, ZSO21]. Compositionally [Dan23]. compressed [DAS98]. Compression [BMW91, CSCM00, DKV07]. Computability [HMS06]. Computable [PK82]. Computation [AC94, BG22, BOV85, DP82, HS94, LST98, PB97, ABB^{+}09, AE01, DR05, LK02, SWU10, SGL97, Hal85]. computational [ATD08, SSD09]. Computations [DW89, MC82b, Ras24, VSS94, YS91, LSLR05, Mon08, YF98]. Computer [HCHP92, Wol92]. Computer-Assisted [HCHP92]. Computers [Fis80, LK02]. Computing [ANP89, CFR^{+}91, CF95, KM81, HVB^{+}99, MMG00]. Concept [Tur86, ST00a]. Concepts [Eug07]. Concerning [Sha82]. Concrete [Bar81, DHH^{+}24]. Concurrency [ADG^{+}21, BG89a, Lam90, LMM21, SDD21, Wei89, BCF04, Mil85, TA08a, CPS93]. Concurrent [BC91, Car95, CIJGP18, Cla80, CES86, CPS93, CFM94, DSFG21, DGMP97, DHH^{+}24, FT94, Hal85, HSP83, HW90, Her93, JTM98, Kar84, Lam83, LFF14, LCK^{+}22, MSM^{+}16, OL82, Pet83a, Pet83b, RY88, SZLY21, Sku95, SNS^{+}14, AE98, AE01, AAE04, BBYG^{+}05, BGP99, CSW06, JPS^{+}08, RS97, SRM10, YS10]. Concurrent-by-Default [SNS^{+}14]. Concurrent-Program [FT94]. Condensation [JTM98]. Condition [HW90]. Conditional [Boy96, GGSV22, WZ91, Dam03]. Conditioning [OGJ^{+}18]. conditions [KWL09]. Conference [Wol92]. confined [GPV07]. Conflict [Cas95]. Conjecture [KPS92, Sag86]. Conjoining [AL95]. Conquer [JZX^{+}24]. conservative [Hai05]. considered [Gor04]. Consistency [FSH23]. Consistent [LB22, XBOS20]. Constant [CIJGP18, CGG^{+}19, Coh91, WZ91, Wir91]. Constrained [BG89a, DAW88, PS96, Žic94, LPP01]. Constraint [Bor81, DGMP97, DDV99, HLH19, NSTD^{+}15, Pal95, PW98, Ste18, Apt00, BMR01, DPPR00, FH04, GHB^{+}96, HPMS00, SS08, SS09, SP07, SSD09, dHB^{+}96]. Constraint-Based [PW98, Ste18, DDV99, SP07]. Constraint-Oriented [Bor81]. Constraint-Solving [NSTD^{+}15]. Constraints [AKP94, DFR15, HG83, Mye90, BA08, RM10, TFK^{+}11, Van96b, VHM^{+}01, Van96a]. Construct [Ans87, BS83, Kat93]. Construction [ADGM91, HIT97, LaL81, MB83, RH87, SL92, CMS03, GC01]. Constructions [LBMTT22]. Constructive [Loe87]. Constructs [AR84, DJP^{+}16, Par90]. Containerless [Ste22]. Context [GHR80, KBP22, LTMS20, LWR21, Ode93, Pad19, PK80, Ram00, RTD83, SCR24, VP23, Rep00]. Context-Dependent [Ode93, RTD83]. Context-Free [GHR80, Pad19, KBP22]. Context-Free-Language [SCR24]. Context-sensitive [Ram00, VP23, Rep00]. Context-Unbounded [LWR21]. Contexts [Ode93]. Contextual [TDA^{+}23]. Continuation [BDM15, Wan82]. Continuation-Passing [BDM15]. Continuations [BDM15, HF87]. Continuous [KF03]. Contract [RKW^{+}23, DF11]. Contract-client [RKW^{+}23]. Contracts [SIG17, SDTF13, CGP09]. Contravariance [Cas95]. Control [ABLP93, Bur84, CL94, CFR^{+}91, DP97, FM87b, HBS22, Kat93, Lam88, Lin79, NBG13, PB97, PBR^{+}15, RSPS23, Set83, SCP23, SS13, Tur84, Wat83, Wei89, BCM99, BCC04, HO07, PSS05, RAB^{+}07, Zho96]. Control-flow [SCP23]. Controlled [Min84, Tho94, JC97]. Controlling [BALP06, LaL81, LMD98]. Convention [AF84]. Convergence [Bar85]. Conversion [CS87, SW97b, SA00, YK97]. Cooperating [GLR83, NO79]. Cooperation [BKS88]. Coordinating [JS94]. Coordination [GLR83]. copying [TSBR08]. Coq [HRV^{+}23]. core [IPW01]. Cores [YGRBA23]. Coroutine [Sam80]. Coroutines [LS81, DI09]. Correct [DGMP97, Hen86, JP17, SS88, AAD^{+}07]. Correction [FA93]. Correctness [ABC^{+}21, Apt86, CM86b, FRW90, Gom92, HW90, Lam79, Lam80, ML21, Oss83, dBJPT24, San96]. correlated [YS99]. Correspondence [BS88, Bur90b, Bur91, Coh91, CM93, DS88, Ell82, FA93, Fra81, Hen83, LaL83, LaL84, Moh81, Moi83, MS88, NN86, Par90, Pem83, Sor89, SM82, Tan83, Tic88, Vol91, WST85, Wir91, YB88]. Corrigenda [WCW91]. Corrigendum [Ano18, BKRW05, DF81, Fra80a, KS89, Lam80, Pur91, QG95, TGT20, Van96a, Wal81, WGS93]. Cost [AB81, Bac84, DL93, Hai98, Han81a, ZGZ05, VALG05]. Cost-optimal [Hai98]. costs [GMP^{+}00]. Counting [Bal94, LP06]. Counts [Bob80, Wis79]. Coupled [ACW90]. Covariance [Cas95]. covariant [PZJ05]. Creating [Mye90]. criteria [Hai05]. Critical [PS93]. Critique [GM81]. Cross [Ano18, FTJ95, GSS^{+}18]. Cross-Interferences [FTJ95]. Cross-Language [Ano18, GSS^{+}18]. Cryptographic [App15, HRV^{+}23]. CS [CD79]. CSim [SZLY21]. CSP [AF84, Bag89, BS83, Fid93, Hua90, LS84, Žic94]. CSP-Like [Hua90]. CSS [HLH19]. currency [DS98]. Curry [LR19]. Curry-Style [LR19]. Custom [DJP^{+}16]. CV3 [CZ84]. Cycle [BG89b, PBK^{+}07]. Cycles [FRW90]. Cyclic [RY88].

D. [Bur91]. Data [AMT14, ANP89, AM85, Bac84, BNNN22, BC85b, BL87, Bur90a, Cha93, CS87, Deb89, DP93, DD85, Ell82, EO80, FL81, GMH81, GEGP17, HL82, Her93, Hes88, Hol87, Jen97, JJCO19, KH92, Kam83, KZC15, KK98, KD94, KEA^{+}24, LaL89, LO94, LN02, LVRG21, Loe87, Mal82, MMR95, MCT96, PP91, QG95, Ras24, RCRH95, RP88, SSS81, Sku95, SGL98, SM81, TWW82, WL85, Wei89, Wei90, Wet82, Wey83, CFP^{+}04, DHM^{+}12, DGS97, HBJ98, KBC^{+}99, KF00, LK02, Rep00, SP07, VALG05, YUW02, ZGZ05, Pur91]. data-centric [DHM^{+}12]. Data-Driven [BL87, CS87, JJCO19]. Data-Flow [BC85b, Bur90a, Wet82, RP88, KBC^{+}99]. data-independence [Rep00]. data-member [KF00]. Data-Parallel [Cha93, Ras24, HBJ98]. Database [Bar85, CB80]. Databases [SR21]. Dataflow [Deb95, DFR15, MWB94, SS13, SS96, Van96a, Van96b, VHM^{+}01]. datalog [LS09, ZSS20]. datatypes [MBC04]. De/ [Ras24]. Deadlock [CHMY19, Hua90, Kob98]. Deadlock-Free [Kob98]. Deadlocks [FJK^{+}17]. Dealing [GLMM05, GG85]. Debugging [CMN91, CM93, Cop94, Hen82, WST85, ZSS20]. Decentralized [RKW^{+}23]. Decidable [LB22]. Deciding [GGL15]. Decision [MTG80, NO79, SCR24]. decisions [MTSS09]. Declarative [ABPS98, TCVB14, Bou05, MME^{+}10]. Decompilation [BB94]. Decomposing [BDL^{+}08]. Decomposition [JZX^{+}24, LK02]. Decomposition-based [JZX^{+}24]. decrease [LDK^{+}96]. Deducing [TB95]. deduction [LMD98]. Deductive [MW80]. Deep [YW22]. Default [SNS^{+}14, LMM21]. Deferring [MTSS09]. Defined [Wal92, Wal80, Wal81]. Defining [Ode93]. definite [RKRR04]. Definition [Bou92, BWP87, CI84, CD79, Fid93, HS94, WCW90, WCW91, Wol94]. Definition-Use [HS94]. Definitions [BS86, Wil82b, Dam03, VHKO02, Sij89]. Delay [BG89b]. Delayed [KPF95, RC03]. Delayed-Load [KPF95]. Delaying [Kau84]. Deleting [GP81]. Delimited [BDM15]. Demand [FPS19, GSW95, PA85, PA86a, PA86b, PF96, SR95, DGS97]. Demand-Driven [GSW95, PA85, PA86a, PA86b, FPS19, PF96, DGS97]. Demanded [SCS24]. Denali [JNZ06]. Denotational [AB94, FA93, Gud92, MSJ94, NF89, Nie85, Sch85, dBB85]. Dependence [BGH^{+}13, CFR^{+}91, FOW87, HBG^{+}09, HRB90, ML21, PB97, PW98, Wol94, RAB^{+}07]. Dependence-based [ML21]. Dependences [PW94]. Dependencies [Deb89, HBS22, CSS99]. Dependency [Blu99]. Dependent [LS80, Miq19, NBG13, Ode93, RTD83, Rob79]. dequeues [Chi05]. DeRemer [Sag86]. Derivation [BKB80, Cat80, DSW82, Gie83, HIT97, Kna90, TM93, Ano02a]. Derivative [JKT23]. Derivative-based [JKT23]. Deriving [Wan82, Bou06]. Describing [AW85]. Description [McG82]. Descriptions [Boe85, BKL^{+}97, Cat80, Ano02a]. Descriptors [Hol87]. Design [ABG24, BPP16, BCD^{+}15, BO94, DF80, DF81, DC22, FT94, HM84, KKM90, LDM07, ML80, NN24, RCS93, RL98, SYK^{+}05, Bou05, MTSS09, CMLC06]. design-pattern-based [MTSS09]. Designing [LaL89, ALZ03]. Designs [AW85]. destructive [SRW98]. Detect [ISY88]. Detecting [GSW95, HCS10, Sch85]. Detection [CM86a, Hua90, MC82a, MC82b, TM93, AFF06, HDH02, PFH11, PCJD08, XA07]. Determinacy [TK94]. determination [DS98]. Determining [MF88]. determinism [TA08a]. Deterministic [KR79, Mye18, YGRBA23, DL18, Tar07]. Development [BKB80, Col84, Fea87, Jon83, ML80, PPS79, Wil82a]. Diagnosis [BF87]. Diagrams [SCR24]. Dialect [Mul92]. Dialects [CP95]. dialogue [BCM99]. DIB [FM87a]. difference [BA08]. Differencing [PK82, RSL10]. Differential [BKOZB13, TDA^{+}23, ZP07]. Differentiation [Sha82, VS22]. Difficulties [ABG24]. Diffusing [MC82b]. Dijkstra [BN94, Nel89]. Dimensional [Hil88, Ras24, GPWZ08]. direct [YK97]. Directed [BDJ13, DMM88, Gud92, Han94, Set83, SYYH07, OKN06]. Direction [Dar90]. Directly [Hob84]. Director [KS88, KS89]. Directory [Han81b]. Discipline [VMLY22, FGM07b]. Disciplines [SS84]. Discovering [FJK^{+}17]. discovery [PZJ05]. Discrete [Bar81]. Discrete-Event [Bar81]. Disintegration [NcS20]. Disjunctive [Jen97, JJCO19]. dispatch [DAS98, MFRW09]. dispatching [GZ07]. Distance [Wol94, ZSD09]. distribute [CRN^{+}08]. Distributed [ABLP93, AF84, Apt86, AW85, BKS88, BCEM15, Bur84, CJK95, CM86a, CBDGF95, CS95, DAW88, Dug99, FLBB89, Fra80b, GHS83, HSG17, Hua90, HM84, Jon94, Kat93, KK98, KRS84, KKM90, Lam84, LS83, MC82a, RCRH95, SS84, Sch82, TM93, TCP^{+}17, Zav85, ABL03, FM87a, HVB^{+}99, KGMO04, LK02, MDJ05, Piq96, Fra80a, Moh81, VHB^{+}97]. Distributed-Memory [KK98, RCRH95]. div [Bou92]. Dive [YW22]. Divergence [SdSCP13]. Divide [JZX^{+}24]. Divide-and-Conquer-like [JZX^{+}24]. DJ [DR05, SGL96, SGL98, UM02]. DJ-graphs [UM02]. DLLs [Dug02]. do [SS05a]. Documentation [MH86]. does [DMP96]. dolce [MP10a]. Dollars [HL22]. Domain [LM18, Tra08, RM07, SS05a]. Domains [ABG24, CMB^{+}95, ELS^{+}14, GS98, FH04, GLMM05]. dominance [Ano02b, DVD07]. dominator [SGL97]. Dominators [LT79, Ano02b, BKRW98, BKRW05]. Don’t [AKNP17]. Down [HL22, SZLY21]. DRAM [KEA^{+}24]. drf [MSM^{+}16]. Drinking [CM84, MS88]. Drive [PK80]. Driven [BL87, CS87, GF85, GSW95, JJCO19, PA85, PA86a, PA86b, TGT18, YBL16, DGS97, FPS19, PF96, TGT20, YMW97]. Dually [MT08]. Dummy [Lam88]. During [BKB80]. DyC [GMP^{+}00]. Dyck [LZR22]. Dynamic [ACPP91, AGT89, ASF17, BNNN22, BB79, BDM15, Bre89, CGG^{+}19, CHMY19, CTT07, DS98, Dug99, HSS^{+}14, HN05, Kai89, KR79, RCRH95, Ven95, WR08, dBB85, ACE96, BP12, CEI^{+}07, DDDCG02, GZ07, MMM^{+}07, PHEK99, SJP12, SHB^{+}07, SYK^{+}05, SYN06, WKD04, ZGZ05].

eager [FKW00]. Earley [Lei90]. Early [AB81]. ECCS [CDFP89]. Edge [DP93]. Editing [FL81, HT86, Nix85]. Editor [FM87b, DeM83, Mye18, Per90, Rep86, Wol92]. Editorial [AP07, App93, AG93, AF94, MP07, Mye19, Pal11a, Pal11b, Pal12, Pal13, Pal15, FP02, OP04]. Editors [DMM88, MM89, RTD83, Wat94]. EDO [OKN06]. Effect [Gor21, RLS^{+}01]. Effective [BS83, Col84, JB20, KKN06, NI05, PE08, WJ98, YUW02]. Effectiveness [BdlBH99, SH89]. Effects [Boe85, MXZ^{+}22, SV20, TA08b]. Efficient [AKBLN89, ADGM91, BB79, BGH^{+}13, Bre89, Cam89, CS95, DP82, DMM88, GZ05, GZ07, GLR83, GLO88, GSO94, HVB^{+}99, HS94, HSS^{+}14, HIT97, JP81, Jon90, KKM90, KRS88, KPF95, MVV^{+}01, MM82, NB99, NI05, PHP02, PXL95, PG21, PKH07, PA85, PA86b, RH87, SS08, SA00, SS88, TN19, WG98, YUW02, BCP08, GB99, KSV96, LPSO04, LS09, PBK^{+}07, TP04, VWJB10, YF98, PA86a, SS09]. Efficiently [Bal94, CFR^{+}91, CF95]. Eiffel [ACE96]. elaboration [KR01]. Election [Hua93]. Eliminating [BT93, Coh83, Coh85, RD03]. Elimination [DP93, SCP23, SGL98, KKN06, KCL^{+}99]. Elimination-Based [SGL98]. embedded [BCP08, CSCM00, HK07, Rhi03, SRM10, TP04, ZP10]. Embedding [HF87]. Empirical [BHK07, BDH^{+}16]. Empowering [JSB^{+}12]. Emulator [ML80]. Enabled [ADG^{+}94]. Encapsulating [GPV07]. Encapsulation [AR84, RSPS23, DDM11]. Enclaved [BNV^{+}21]. Encoding [Hob84, GZ05, ZP07]. Encodings [BC79, TY21]. End [BDP14, CSCM00]. enforcement [HMS06]. Enforcing [CEI^{+}07, RKW^{+}23]. engines [SS08, SS09]. enhanced [GH97]. Entries [LaL84, SS82]. Enumeration [BB94, JJD98]. Environment [CO90, SZBH86, CKT86]. Environmental [SKS11, SV19]. Environments [BS86, GKL94, HK85, HT86, Kai89, dJKVS12]. Epochs [Sol92]. equalities [FMoPS11]. Equality [Pal98]. Equality-based [Pal98]. Equations [HO82, Bou06, GS11, GMM99].

Equiprobable [PB80]. Equivalence [SV20, VJB12, VSS94]. Equivalent [PO95, NP08]. Erlang [TCP^{+}17]. Erratum [SS09]. Error [AB81, Bac84, BN99, BF87, FL15, KC01, LaL84, MF88, MYD95, PK80, Ric85, SSS83, SS82, Wet82, ZMVPJ17, dJKVS12, Jef03, XA07]. Errors [AWW95, SBB^{+}19, Wha94, CPRT02, JNGG10]. Escape [Bla03, CGS^{+}03]. ESOP [Ahm20, Mül21, Yos22]. ESOP’05 [Sag07]. Essential [DES12]. Esterel [Tar07]. Estimation [SBB^{+}19]. Estimations [LVRG21]. Eta [DMP96]. Eta-expansion [DMP96]. Euclid [HW82]. Euclidean [Bou92]. Evaluating [BLH12]. Evaluation [AFV98, Bur84, CGST95, CK93, Gri82, Hud91, Jon90, LV94, LLOY23, PA85, PA86a, PA86b, RD87, RL98, Slo95, SG90, WCW90, WCW91, ZSS20, ADR06, CP96, CG04, GJ05, LDM07, Leu04, ST00b, SYK^{+}05]. Evaluations [BDH^{+}16]. Evaluator [Gom92, JP81, KR79, Le 88]. Evaluators [CP95]. Event [Bar81, BPRB23, YMW97]. event-driven [YMW97]. Events [Bal94]. Ever [Gri79]. Evidence [CGJ^{+}97a]. Evidence-Based [CGJ^{+}97a]. Example [CM86a, Mye90, Nix85]. Examples [Oss83, Jef03]. Exception [YB85, YB87, YB88, LS98, LP80, OKN06]. Exception-Directed [OKN06]. Exception-Handling [YB85, LS98]. Exceptional [WN08]. Exceptions [ASF17, Hau96, LP00]. Exclusion [ADG^{+}94, LH91, ABHI11]. Executable [Hob84]. executables [YF09]. eXecute [BR10]. Execution [BNV^{+}21, CS87, DMZ^{+}24, Dil90, GJ93, JW17, JNGG10, JF81, SS98, SS88, BALP06, GPA^{+}01, TSY00, YF98]. Exemplars [LaL89]. Exemplified [DGL^{+}79]. Exercise [Kna90, Mis81]. Exhaustive [Bur90a, LLOY23]. Existential [MP88]. existing [LS98]. expansion [DMP96]. Expansions [SBB^{+}19]. Experience [FSS83, Wal92]. experiences [Eug07]. Experimental [LV94, SSS83, Ven95, ABB^{+}09, BGP99]. Experiments [Tur84]. Explanation [Mis81]. Exploiting [KOE^{+}06, KEA^{+}24, VP23]. exploring [WS97]. exponential [Wu04]. exponential-time [Wu04]. Expression [GP81, YB87, YB88, HVP05]. Expression-Oriented [GP81, YB87, YB88]. Expressions [BG89b, CGST95, CC97, DAW88, Fis80, Geo84, Gri82, Hen83, HY91, KS83, LdR81, PK82, Sha82, Sit79, Wat91, Dam03, NN86]. Expressive [MFRW09]. Expressiveness [WGS92, WGS93, PS96]. Extended [ABC^{+}21, CBMO19, KGMO04]. Extending [CEW14, CMS03, MSRR00, MK94]. Extensibility [LCK^{+}22]. Extensible [HSG17, Sto04, ATD08, MBC04]. Extension [Bur90b, Coh91, WSH15, Wir91, ALZ03, KKN06, LS08]. Extensions [Wir88]. Extent [MF88]. External [Wal80, Wal81]. Extracting [GP95]. extraction [TSL^{+}02]. extrapolation [WM12]. Extrema [Pet82].

F [MWCG99]. Facets [ASF17]. factoring [DRSS96]. Failure [BN99, Dar90, GHH^{+}19, Kar84]. Failure-Free [Kar84]. Fair [BN94, PR07]. Fairness [ES97, OA88, TB95, AH98]. Families [LaL89]. Fashioned [AL94]. Fast [ADR06, DAS98, FMoPS11, HVDH07, LT79, LZR22, SR95, DR05, PE08, TP04, VBLG04, DVLM15]. Faster [CGG^{+}19]. Fault [CS95, Lam84, LJ99, AAE04]. Fault-Tolerance [LJ99]. Fault-Tolerant [CS95, Lam84, AAE04]. FD} [GLMM05]. FeatherTrait [LS08]. Featherweight [IPW01, LST02, LS08]. Feature [ASAVF19, AH10]. Feature-Specific [ASAVF19]. Feeding [PA86a]. Fence [AKNP17]. Few [HL22]. Fickle} [DDDCG02, AAD^{+}07]. field [PKH07]. field-sensitive [PKH07]. fields [PZJ05]. FIFO [FLBB89]. Final [Kam83]. Finding [KRS84, KKM90, LT79]. Fine [DSFG21, HL22, PBR^{+}15, DNS^{+}06]. Fine-Grained [HL22, PBR^{+}15, DSFG21, DNS^{+}06]. fingerprinting [CTT07]. Finitary [AH98]. Finite [ACW90, BLH12, CES86, Dan23, GC86, PK82, PP91, Pur91, RSL10, Zav85]. Finite-State [ACW90, BLH12, CES86]. Finite-State-Machine [Zav85]. First [ADG^{+}94, Bre89, DP97, HKMN94, Han92, JPP91, JS94, LH91, MH04, MPLM23, SDTF13, HMY^{+}24]. First-Class [HKMN94, Han92, SDTF13, MH04]. First-Come-First-Served [LH91]. First-Enabled [ADG^{+}94]. First-Fit [Bre89]. First-In [ADG^{+}94]. First-Order [DP97, JPP91, JS94, MPLM23]. Fit [Bre89]. Fixed [SS98]. Fixed-Order [SS98]. Fixpoint [AC94, Qia00]. Flash [KEA^{+}24]. Flexible [AD98, Hud91, MSM^{+}16, WG98, Wil82b, dJKVS12, IV06, KGMO04]. Floating [CK94, Fat82, SBB^{+}19, Hau96, Mon08]. Floating-Point [CK94, Fat82, SBB^{+}19, Hau96, Mon08]. flop [MMG00]. Flow [AR80, AD98, ASF17, Bac84, BC85b, Bur90a, DP97, DP93, FJKA06, Hor97, KD94, MMR95, NBG13, PO95, PP91, PBR^{+}15, Pur91, RSPS23, Set83, SGL98, SS13, Wet82, DGS97, HR02, HY07, KBC^{+}99, Pal98, PS03, RRSY08, RP88, SCP23, TZ07, WJ98]. Flow-Insensitive [Hor97, FJKA06]. Flowback [CMN91]. Flowgraph [LT79]. Flows [Kna90]. Floyd [Yin11]. Fly [CF95, BA84, LP06, PBK^{+}07, URJ18]. Focusing [EKD23]. fold [RKRR04]. Folklore [LY98]. Font [FK85]. Foo [FA93]. ForeC [YGRBA23]. foreign [FF08]. Foreword [Mye17, Mye18]. Form [AK87, BOV85, BM94, CFR^{+}91, GSW95, Pal95, PG21, GPF08, KCL^{+}99]. Forma} [ZCG^{+}07]. Formal [ADG^{+}21, BS86, BDP14, CB80, CD79, Fid93, Gie83, HIT97, Kna90, Lee86, Mal82, MH86, SSFZ^{+}23, Sha82, WP10]. Formalism [Pea21]. Formalisms [PCC85]. Formalization [BPP16]. Formally [SP97]. Format [Wat83]. Forms [DS83]. formulas [RSL10]. formulations [RS97]. Fortran [AK87, DP99]. Foundation [KRR18, Ban11, RAB^{+}07, Rhi03]. Foundational [HRV^{+}23, AM01]. Foundations [GTWA14, LW93, AAR^{+}10]. Fractal [MPM03]. fractional [Boy10]. Frames [MPLM23, SJP12]. Framework [BGL93, Gie83, HRV^{+}23, JW17, KRR18, NSZS13, NSTD^{+}15, OHL^{+}14, SGL98, TN19, ATD08, DGS97, GMM99, GZ04, GC01, Leu04, PS08, RKRR04, TP04, VBLG04, XA07, ZCG^{+}07, ZP10, vHK00]. Frameworks [KEA^{+}24, MMR95, KK07]. Framing [BNN18, BNNN22]. Francez [Fra81, Moh81, Moi83]. Free [AP94, GEGP17, GHR80, Her91, Kar84, Kob98, Pad19, SCR24, JJD98, KBP22, KSV96]. freedom [KS10]. Fresh [GMN^{+}21]. frontiers [Ano02b]. full [GB99]. Fully [JPP91, TY21]. function [DR05, FF08]. Functional [AFV98, Ban87, Blo94, Bou05, Bur84, DW89, FL91, ISY88, JPP91, WM95, Web95, Wil82a, ABH06, Bou06, DWWW08, DF98, PS08, San96, SP97]. Functions [AKP94, AK82, Bou92, PB80, SM89, TY21, Lee09, MBC04, MB99, MT08, PPT08]. Further [CM93]. Fusion [JB20, LGAT00]. Fusion-based [LGAT00].

G. [Tic88]. Garbage [BA84, CN83, DSW82, ISY88, JCMM19, TM93, URJ18, WLBF16, Wis79, YW22, BBYG^{+}05, BALP06, HDH02, LP06, Piq96, TSBR08]. Garnet [VHM^{+}01]. General [BGL93, CHMY19, HSS^{+}14]. General-Purpose [HSS^{+}14]. Generalization [Nel89, LMD98]. Generalized [Ans87, BS83, GKM20, KD94, Lin79]. Generalizing [DB85]. Generals [LSP82]. Generate [Son87]. Generated [Slo95, dJKVS12]. Generating [HBM^{+}06, HT86, Jef03, LR13, JNZ06]. Generation [AGT89, AS80, BOV85, BM94, DS83, DS90, GF85, GVC15, HKR92, HKR94, Pro95, Rei83, Rob79, She91, ST00b, UJ92, DAS98, MSRR00, PHEK99]. Generative [Gel85]. Generator [JKT23, PPS79]. Generators [Cat80, GHK81]. Generic [LV94, DDM11]. generics [IV06]. Geometry [CR87]. Georgeff [NN86]. GJ [IPW01]. Glanville [MSRR00]. Global [Bac84, Dha91, GHB^{+}96, OHL^{+}14, PK80, PS92, Sch85, Sor89, dHB^{+}96, CS04, KBC^{+}99, DS88]. Globally [DHH^{+}24]. GLR [SJ06]. Goal [Dar90, Gud92, SYYH07]. Goal-Directed [Gud92, SYYH07]. Goal-Oriented [Dar90]. Goto [CF94]. GPU [BCD^{+}15]. Gradual [DMZ^{+}24, TGT18, TGT20]. Gradualizing [LBMTT22]. Graham [MSRR00]. Graham-Glanville [MSRR00]. Grained [HL22, PBR^{+}15, DSFG21, DNS^{+}06]. Grammar [CI84, CP95, GF85, JP81, KR79, Web95]. Grammar-Based [CI84]. Grammars [BS88, JKT23, Jon90, Kat84, LaL81, RD87, RH87, Tai79, WW95, Boy96, CP96, Wu04]. Grammatic [Tho94]. Grammers [BB94, MK94]. Granularity [RRB19]. Graph [Ass00, Bee94, BCT94, CFR^{+}91, FOW87, KKSD94, KLS92, LZR22, MC82a, Son87, CTT07, GC01]. graph-based [CTT07]. Graphic [Mal82]. graphical [VHM^{+}01]. Graphs [GKM20, HRB90, KPS92, Kna90, SGL98, DR05, JC97, KSK07, SGL96, UM02]. grid [VWJB10]. Grimmer [Ano18]. groundness [CSS99]. Grover [BH99]. growth [BALP06]. Guarantee [GEGP17, LFF14, SZLY21, HQRT02]. guarantees [LS09]. guard [MP07]. guarded [SP07]. Guardians [LS83]. Guards [Ber80]. Guest [FP02, OP04, DeM83, Per90, Rep86, Wol92]. Guide [App94a, BDH^{+}16]. Guided [OLH^{+}16]. guiding [VALG05].

Hackers [App94a]. Hancock [CFP^{+}04]. handle [VJB12]. Handling [CCEG23, Hau96, LdR81, Piq96, SSS83, UM02, YB85, YB87, YB88, CRN^{+}08, LS98, LP80, SSD09, Hen83]. Hard [Hor97]. Hardware [BKL^{+}97, Mis86]. Hare [Dan23]. harmful [Gor04]. Hashing [PB80, Duc08]. Haskell [GRSK^{+}11, HHPW96]. Heap [KSK07, BALP06, KF00, YS10]. heap-manipulating [YS10]. Heavily [BG89a]. Hennessy [CM93, WST85]. Herding [AMT14]. Heuristic [SL92]. hiding [LN02, OYR09]. hierarchic [AG04]. Hierarchical [BA99, CP95, CD79, AY01, CP96]. hierarchically [MBC04]. hierarchies [ST00a, Van96a, Van96b]. hierarchy [KF00]. High [Cam89, Fat82, MSM^{+}16, URJ18, CMS03, VWJB10]. High-Level [Cam89, Fat82, CMS03, VWJB10]. High-Performance [URJ18]. Higher [AC94, AD98, CJK95, DJP^{+}16, FPS19, SV19, BBTS07, DF11, SKS11, SP97]. Higher-Order [AC94, AD98, CJK95, DJP^{+}16, FPS19, SV19, BBTS07, DF11, SKS11, SP97]. Highly [Her93, Sku95]. Hoare [Apt81, GM81, LS84, Sok87, Yin11, dBH21]. Hoc [MDCB91]. Holistic [ZMVPJ17]. Homeostasis [NN24]. Homomorphic [LLOY23, VS22]. Homomorphisms [HIT97, Ras24]. HOP [BLRS12]. Hot [VP23]. Hybrid [KF10, KS10]. Hyperball [LM18]. hyperdoctrines [BBTS07].

I-Structures [ANP89]. I/O [Car95]. Icon [GHK81, Gri82]. id [Bee94]. idempotency [KOE^{+}06]. Identical [FLBB89].

Identification [BGH^{+}13, SBE^{+}19]. Identifiers [SSFZ^{+}23]. identify [MMM^{+}07]. Identifying [Ram99, SGL96]. Idioms [PP94]. IDL [Lam87]. IEEE [Fat82]. Ignorance [GNS^{+}15]. Illustrative [Oss83]. Immutability [RSPS23]. Impact [BHM^{+}19, OLH^{+}16, CKT86]. Imperative [AB20, ABPS98, DFR15, Gro06]. Implementation [AKBLN89, AOC^{+}88, BCD^{+}15, Bou88, Bre89, BS83, CM86b, GMH81, Gaz83, Lin93, MDCB91, NN24, PXL95, RL98, WL85, CMLC06, FM87a, GB99, LDM07, LPSO04, Tra08, Zho96]. Implementations [dBJPT24, BBF^{+}11, BFGT08, DF98]. Implemented [DB85]. Implementing [BR97, Her93, HW82, Sku95]. Implications [Fat82]. Implicit [BH05b, SJP12]. Implicit-signal [BH05b]. Improve [VP23, KF00]. Improved [GHR80, Mur91, KK07]. Improvement [MS83, San96]. Improvements [BCT94]. Improving [CK94, CMB^{+}95, MCT96, SSFZ^{+}23, TCP^{+}17, WS97]. impure [Pip97]. Incomplete [MRGP20, GLMM05]. Incremental [Bur90a, CP95, DMM88, GM79, HKR92, HKR94, HPMS00, Hud91, Kai89, Lar95, LST98, LHR19, PS92, RTD83, RP88, SGL97, WG98, YS91, BBYG^{+}05, CP96, Van96a, Van96b]. Incrementally [QL91]. Independence [DHM00, GGSV22, Rep00]. Independent [ML80, Mul92]. Index [Ano86a, Ano88a, Ano89a, Ano90a, Ano91a, Ano92a, Ano94, Ano95, Ano98]. indexed [AM01]. indices [RR05]. Indirect [Piq96, CEG07, YK97]. Induction [GSW95, Sit79]. Inductive [LBMTT22]. inefficiencies [MMM^{+}07]. Inessential [SS82, LaL84]. Inference [CEW14, Deb89, Hen93, LO94, LY98, MRGP20, Pad19, SR21, TB98, Wey83, FFLQ08, JB06, PM06, PT00, PS03, Van06]. Inferring [FNBG20]. Influence [FTJ95].

Information [AR80, Ano82, Ano83, Ano84, Ano85, Ano86b, Ano87, Ano88b, Ano89b, Ano90b, Ano91b, Ano92b, ASF17, BC85b, HR02, NBG13, PBR^{+}15, PS03, RSPS23, GS99, HY07, LN02, OYR09, TZ07]. Information-Flow [BC85b, TZ07]. infrastructure [SWU10]. Inheritance [LN15, WT11]. initialization [FM99]. Injection [SBE^{+}19]. Input [BS83, SJW23, vPS81]. Input-Output [BS83]. Inputs [PA86a]. Insensitive [Hor97, FJKA06]. Insertion [AKNP17, GJ05]. inspection [CF04, FG03]. Instantiation [Der85]. Instead [Lam84, Rem81]. Instruction [KPF95, LCBS19]. Instructions [LS80, PS93, RF97, Rob79, LPP01]. Integer [BAGM12, BEF^{+}16, FNBG20, HSD22, BGP99]. Integrated [SS13]. Integrating [HPR89, WJS^{+}00]. Integration [CO90, Leu04]. Intensional [HP24, STS03]. Interaction [WSH15, WT11, van88, BCM99]. Interactions [GDF23, JS94]. Interactive [ACS84, BS86, SCS24]. Interconnectability [TY18]. Interface [Win87, van88]. Interfaces [DS90, Mye90, TLHL11, WT11]. Interferences [FTJ95]. Interfering [Jon83]. Interleaved [LZR22]. Interleaved-Dyck [LZR22]. Intermediate [Lam87, Pem83, TvS82]. Internal [Han81a]. International [Wol92]. Interoperability [Ano18, GSS^{+}18]. interoperable [BFGT08]. Interpretation [BGL93, CFG^{+}97, DC22, DLR16, GMP24a, KRR18, LV94, MSJ94, SCS24, BDL^{+}08, BdlBH99, DGG97, Leu04, SYYH07]. Interpretation-Based [DLR16]. Interpretations [CMB^{+}95, HY91, SJ03]. Interpreters [LR13, CEG07]. Interprocedural [Bur90a, BT93, DP97, HAM^{+}05, HS94, HBCC99, HRB90, LWR21, ML21, NR06, SH89, CKT86, DVD07, DGS97, FMoPS11, JLRS10, KK07, RLS^{+}01]. Interprocess [RS84b]. Interprocessor [Ang89]. Interruptible [BNV^{+}21]. intersection [Dam03]. Intertask [FY85]. Interval [Bur90a, GNS^{+}15, FH04]. Interval-Based [Bur90a]. Introduction [Ahm20, DeM83, HCW82, Mül21, Per90, Rep86, Sag07, Wol92, Yos22]. Invariant [BKB80]. Invariants [Cla80, GEGP17]. Irreducible [Hav97, UM02]. irregular [YF98]. Irrelevant [GP81]. Iso [LBN17, ZZO22]. Iso-Recursive [LBN17, ZZO22]. Isolation [Wha94]. Isomorph [JJD98]. Isomorph-free [JJD98]. Issue [Ahm20, Ano18, TGT20, Yos22, Sag07]. Issues [BO94]. Iterable [Gor21]. Iterated [GA96]. Iteration [Cam89, MOSS96, GS11, JLF02, Qia00]. Iterative [Ans87, Par90, DR05, JNGG10, LS04].

Jade [RL98]. Jam [ALZ03]. Java [AFF06, ALZ03, AAD^{+}07, BH05a, Bla03, BALP06, CGS^{+}03, CMS03, CSCM00, FFLQ08, FM99, GPF08, IPW01, KKN06, KGMO04, KN06, KR01, LST02, LP06, LS08, Loc13, MVV^{+}01, MME^{+}10, MFRW09, MMG00, NR06, OKN06, Qia00, RRB19, SLC03, SMP10, SBE^{+}19, SA99, SYK^{+}05, TN19, TSL^{+}02, WR08]. Java-like [KN06]. JavaCOP [MME^{+}10]. JavaGI [WT11]. join [WKD04]. JR [KGMO04]. Jump [LS80, RS84a]. Just [DLR16, TN19, SYK^{+}05]. Just-In-Time [TN19, DLR16, SYK^{+}05]. JVM [HO07].

KaffeOS [BH05a]. Kaiser [Tic88]. Kernels [BCD^{+}15, ATD08]. Keys [PB80]. Kilbury [Lei90]. Kitsune [HSS^{+}14]. Kleene [Koz97]. Knot [MC82a]. knowledge [GLMM05].

labels [Sto04]. Laboratory [Bor81]. LaLonde [Hen83, LaL83]. LALR [DP82, KM81, PCC85]. Lambda [Geo84, Gom92, NN86, PS08]. Laminar [PBR^{+}15]. Lamport [Ang89, Pet83b]. Language [ACPP91, AOC^{+}88, Ano18, ABPS98, BS86, BPP16, BO94, Bor81, BC91, DVLM15, Fat82, Fea87, FFF^{+}18, GSS^{+}18, Gud92, Hal85, HSG17, JMSY92, JPP91, Kai89, LVRG21, McG82, Per79, PPS79, RTD83, RCS93, RKW^{+}23, SCR24, Spo86, SNS^{+}14, SDD21, Tur84, Wet82, Win87, YS91, YB87, dJKVS12, van88, Bou05, BSvGF03, CFP^{+}04, DWWW08, DF98, FM99, Gro06, HBJ98, KN06, LP99, MF09, MWCG99, PPT08, PHEK99, Tra08, VHKO02, HCW82, YB88]. Language-Based [Kai89, RTD83]. Languages [Ano18, AR84, AD98, Bar81, BL94b, BHM^{+}19, Blo94, BM94, BWP87, CDFP89, CMPP23, DHH^{+}24, Dug99, Fos96, FL91, HU96, Lee86, LR19, LMM21, MSM^{+}16, Mur91, Ryu16, SV19, TK94, TGT20, AAR^{+}10, ACM11, DHM00, GW99, RS97, Rhi03, SRW98, SKS11, SP97, SWU10, Wol92]. Larch [Win87]. Large [GLR83, HSD22, MK94, MH86, WCW90, WCW91, ZSS20]. Large-scale [ZSS20]. Lattice [AKBLN89, MMR95, FH04]. Lauer [GM81]. Layered [HP24]. Layout [KK98, LVV^{+}83, GPWZ08, KF00]. Lazy [ABM93, FKW00, HKR94, Hud91, ITF^{+}22, TCVB14, Chi05]. LCF [Sok87]. lead [SS05a]. Leader [Hua93, KKM90]. leak [HDH02]. learned [VHM^{+}01]. Learning [CGJ^{+}97a, HOYY18, JJCO19, SR21]. Least [AB81, Bac84]. Least-Cost [AB81, Bac84]. Left [FKW98]. Left-Linear [FKW98]. legacy [NCH^{+}05]. length [SMP10]. Lessons [URJ18, VHM^{+}01]. Let [LY98]. Let-Polymorphic [LY98]. Level [BPRB23, Cam89, Fat82, GP95, YBL16, CMS03, VWJB10]. Lexical [HKR92, SJW23]. libraries [Dug02]. LIFE [AKP94]. lifetime [HBM^{+}06]. Lifetimes [Pea21]. Lightweight [Pea21, SW97b]. Like [Hua90, JZX^{+}24, KN06]. Limitations [CP17]. Limits [ABG24]. Linchpin [BGH^{+}13]. Linda [Gel85]. Line [Bal94]. Linear [BL94b, FKW98, PS99, RS84a, TDA^{+}23, YR94, BKRW98, BKRW05, FMoPS11, KBC^{+}99, Ram99, Rep98, RM10]. Linear-Time [YR94, BKRW98, BKRW05]. Linearity [KPT99]. Linearizability [HW90, DSW11]. Linearization [SCP23]. Linguistic [LS83, Wei90, FGM^{+}07a]. Link [DDD05]. Link-time [DDD05]. Linking [QL91, Dug02]. LIPS [CDFP89]. LISF [GCRN11]. LISP [Mul92, Pip97, SH89, Wat83]. List [BC79, HIT97, Kau84, Sij89]. listing [MDJ05]. Lists [Dan23]. Literature [Oss83]. Live [DSFG21, MWB94]. Live-Structure [MWB94]. Liveness [ACW90, GC86, OL82, RY88, HDH02]. LL [BF87]. LLVM [HL22]. Load [KPF95]. Loaded [BG89a]. Local [BDFZ09, CBDGF95, HMY^{+}24, PT00, SDB20, TSBR08, Wei89, Dam03, San96]. Local-first [HMY^{+}24]. Locality [BAC16, MCT96, VALG05, ZSD09]. Locally [AB81, Bac84, DHH^{+}24, Min84]. locating [JNGG10]. Locator [ZMVPJ17]. Lock [GEGP17, KS10]. Lock-Free [GEGP17]. lock-freedom [KS10]. locking [AFF06]. LOCKSMITH [PFH11]. Logic [AS89, AFV98, Apt81, BNNN22, BGL93, BL87, BCD90, BDJ13, BMPT94, CS04, CES86, CFM94, DW89, Deb89, DL93, Deb95, DJP^{+}16, JPP91, Kar84, LS84, Lam94, MW84, MSJ94, MMG92, MPLM23, PZ22, SS98, Sok87, TK94, TB95, dBH21, BBTS07, BMR01, BCG^{+}07, BdlBH99, CU08, CG86, CSS99, DDV99, DPPR00, GHB^{+}96, GW99, HVB^{+}99, HPMS00, KWL09, LMD98, Leu04, PM06, RKRR04, SRW02, Yin11, dHB^{+}96]. Logical [BNN18, GGL15, GS98, TY18, RSL10, Tar07]. Look [DP82, GMN^{+}21]. Look-Ahead [DP82]. Lookahead [KM81, MF88]. Loop [BAC16, CS87, MCT96, Sit79, RKSR12]. Loops [BAGM12, Boo82, CK94, DB85, FTJ95, Hav97, Wat91, Ano02b, LS04, LSLR05, Ram99, RDG08, SGL96, UM02]. LoRe [HMY^{+}24]. low [CSCM00]. low-end [CSCM00]. Lower [FNBG20, PW94]. LR [LaL84, ADGM91, BL94b, BF87, CPRT02, DMM88, Jef03, JP17, KC01, LaL81, SS82, ST00b]. LR-based [KC01].

M [Bur91, Mul92]. M-LISP [Mul92].

Machine [CGJ^{+}97a, Cat80, GNS^{+}15, Gie83, Han94, JJCO19, LR13, ML80, RF97, SS98, SDB20, Wal92, Zav85, Ano02a, CEG07, CF04, HK07, KN06, Oho07, RRB19]. machine-checked [KN06]. Machine-Code [LR13]. Machine-Independent [ML80]. Machine-Learning [JJCO19]. Machine-Specific [Gie83]. machinery [FKW00]. Machines [ACW90, Bee94, CGST95, GC86, KK98, PS93, PP91, Rob79, RCRH95, AY01, AG04, ABE^{+}05, ABS09, TSY00, Pur91]. Made [LM22]. Madsen [Ell82, SM82]. Magma2 [Tur84]. Maintenance [GKL94]. Making [JC97, Loc13]. malware [PCJD08]. Managed [KEA^{+}24]. Management [JP81, Mur91, SDB20, van88, BP12, WCM00, Zho96]. Managing [Bob80]. Manifest [SIG17]. manipulating [YS10]. Manipulation [DVLM15]. Manipulations [SBS22]. many [AE98]. massive [BHK07]. Massively [CGST95]. Matching [AC96, AGT89, CP95, KPS92, ADR06, Van06]. Matching-Based [CP95]. materializations [RMH06]. Mathematical [Ban11, Hes88, LW93]. MATLAB [DP99]. MATLAB(R) [JB06]. Matrix [FTJ95]. Matrix-Vector [FTJ95]. Maximal [BG89b, Rep98]. Maximal-munch [Rep98]. Maximization [GLO88]. Maximum [Kna90]. May [Hor97]. May-Alias [Hor97]. MCALIB [FL15]. Measures [NcS20]. Measuring [FL15]. Mechanically [DSW11]. Mechanism [CO90, YB85, DNS^{+}06]. Mechanisms [Rei83, HMS06]. Mechanizing [Pau01]. Median [Com80]. Medians [KRS84]. megaflops [MMG00]. member [KF00]. Memory [AMT14, CK94, Cha93, CBMO19, FSH23, KZC15, KK98, KRS88, LB22, MSM^{+}16, Mis86, RCRH95, SS88, ABHI11, BP12, GMM99, GW99, JNGG10, KF00, LK02, Loc13, QR00, RR05, TSY00, TP04, VBLG04, WCM00, MMM^{+}07]. memory-efficient [TP04]. memory-hierarchy [KF00]. Merge [Ber94]. Merlin [HBM^{+}06]. Message [CSW06, SS84, VMLY22, Gor04]. Messages [BB79, Jef03]. meta [Tra08]. meta-programming [Tra08]. Metalevel [Jag94]. Metaprogramming [CI84]. Method [BNN18, BCD90, BF87, HL82, Jon83, Loe87, JJD98]. Methodology [Ban87, Her93, Sku95]. Methods [DAW88, KM81]. METRIC [MMM^{+}07]. Mezzo [BPP16]. Microanalysis [HCHP92]. Microcode [MV87]. Microprocessors [BNV^{+}21]. Middle [BDP14]. Middle-End [BDP14]. Might [Bee94]. migration [Piq96]. Minification [HLH19]. Minimal [FKW98, IPW01]. Minimization [RS84a]. minimizing [RMH06]. Minimum [GHS83]. Minimum-Weight [GHS83]. Mining [AMT14]. Misled [Cop94]. miss [GMM99]. Mitigating [KEA^{+}24]. Mixin [HL05, RD13]. mixins [ALZ03]. ML [Blu99, CBMO19, HM93, HT04, PS03, RD13, Spo86]. Mobile [LS03, VHB^{+}97, BCC04, KS10, SWU10]. mod [Bou92]. Modalities [SV20]. mode [PS08, ZP10]. Model [AY01, Ang89, BK11, BL87, BGP99, CMPP23, CGL94, DLR16, ES97, GS98, GG85, GL94, HMY^{+}24, Han81a, HW82, Hol87, JB20, JJCO19, KH92, MSM^{+}16, MMG92, ND16, VSS94, ACM11, AM01, AE01, JJD98, JPS^{+}08, KN06, KV00, Loc13, NP08, QR00, SG04, VWJB10, VALG05, YMW97]. Model-Checking [ES97, BGP99]. Modeling [AF84]. Modelling [AMT14, ADG^{+}21]. Models [FSH23, GJ93, KZC15, dBJPT24]. Modern [BCF04, LMM21, YW22, RAB^{+}07]. Modes [Deb89]. modest [LS08]. Modification [Lei90, RLS^{+}01]. Modula [EO80]. Modular [AG04, BMPT94, CDK^{+}18, EMH20, GL94, HRV^{+}23, JBK18, Jag94, KKM90, LN15, MBC04, Wei89, YB85, dJKVS12, KV00, MFRW09, MOS07b]. modularity [BA99]. Module [PAS^{+}15, RD13]. Modules [CL95, HW82, Lam83, HL05]. Modulo [FSH23]. Monadic [DG19, MH04]. Monitors [BLH12, BH05b]. Monolingual [HK85]. Monte [FL15]. Morel [Sor89, Dha91, DS88]. Morphing [HS11]. Morris [Wis79]. Mostly [YF09, BBYG^{+}05]. Motion [KRS94, Hai98]. MPI [FJK^{+}17, TSY00]. Multi [Ano18, GSS^{+}18, ITF^{+}22, Ras24, YGRBA23, MF09]. Multi-Cores [YGRBA23]. Multi-Dimensional [Ras24]. Multi-Language [Ano18, GSS^{+}18, MF09]. Multi-threaded [ITF^{+}22]. Multialgebraic [WM95]. multidimensional [RDG08]. MultiJava [CMLC06]. Multilisp [Hal85]. multimethod [DAS98]. Multimethods [CL95]. Multiparty [JS94]. Multiple [ASF17, NSTD^{+}15, SJW23]. Multiply [FTJ95]. Multiprocess [Lam79, Lam80]. Multiprocessing [ABR81]. Multiprocessor [GP81]. Multiprocessors [Cha93, KRS88]. Multisource [MMR95]. Multithreaded [EGP14, JBK18, JSB^{+}12, KKW14, NR06]. Multivariate [HAH12]. Multiway [Cha87, Van96a, Van96b]. munch [Rep98]. Mutandis} [SHB^{+}07]. Mutatis [SHB^{+}07]. Mutual [LH91, ABHI11]. Mutual-Exclusion [LH91]. Myths [Gor04].

n [CKT86]. Naming [BDP93]. Natural [GZ04, dJKVS12, ACE96]. Neighborhood [BG89a]. Neighborhood-Constrained [BG89a]. Nested [Cha93, DDMP22, NB99, ACM11]. Nesting [Hav97, Boy10]. Nests [BAC16]. Net [JTM98]. Network [WGS92, WGS93]. Networks [CGJ97b, GC86, KRS84, dBB85]. Newtonian [RTP17]. Nicholson [FA93]. No [Ano18, TGT20]. node [JC97, UM02]. Nodes [CF95, Han81a]. Nomadic [SWU10]. Nominal [CU08]. Non [CFG19, DL18, LLK^{+}17, Mye18, BS88]. non- [BS88]. Non-Deterministic [Mye18, DL18]. Non-polynomial [CFG19]. Non-Statistical [LLK^{+}17]. Noncanonical [Tai79]. Noncorrecting [Ric85]. Nondeterminate [TK94]. Nondeterminism [Ber80, CCEG23, Hes88, WM95]. Nondeterministic [QG95, MT08]. Noninterfering [HPR89]. nonnumerical [ME97]. Nonprocedural [PPS79]. nonrectangular [JLF02]. nonscalars [CRN^{+}08]. Nonsequentiality [Bar81]. Nonstrict [Blo94]. Nonterminating [ML21]. Nontermination [PM06]. normal [LMD98]. Normalize [CRN^{+}08]. norms [BCG^{+}07]. Notation [Rem81, Wil82b]. Note [Com80, CM93, MS88, WST85, Coh85, Pal11b, YK97]. Notes [Sku95]. Nothing [BDH^{+}16]. Notion [LW94]. NP [Hor97]. NP-Hard [Hor97]. NQLALR [BS88]. nulled [SJ06]. Number [Ste22]. Numbers [GLR83]. numeric [Hau96].

O [ABPS98, Car95]. Object [DF84, GMP24b, HU96, KH92, Ryu16, Ste22, WCW90, WCW91, BSvGF03, DMM01, DDDCG02, FM99, GPWZ08, HBM^{+}06, JPS^{+}08, LPSO04, Piq96, WJS^{+}00]. Object-Based [KH92]. Object-Oriented [HU96, Ryu16, Ste22, GMP24b, BSvGF03, DMM01, JPS^{+}08, WJS^{+}00]. Objects [AM85, CJK95, HF87, HW90, Her93, SM89, VHB^{+}97, Wal80, Wal81, Win84, GPV07, HBJ98, KF00, Sto04, WJS^{+}00, Sku95]. obligations [DSW11]. Observability [Gaz83]. Observation [LWR21]. Observations [Sha82]. Obsidian [COE^{+}20]. Occur [AP94]. Occur-Check-Free [AP94]. Octagons [GMN^{+}21]. Off [SBB^{+}19]. Offline [CG04, GJ05]. Old [AL94]. Old-Fashioned [AL94]. Omnisemantics [CCEG23]. On-Line [Bal94]. On-The-Fly [CF95, BA84, URJ18, LP06, PBK^{+}07]. One [Bak82, BG89b, VHM^{+}01]. One-Pass [Bak82]. one-way [VHM^{+}01]. online [CG04, HVDH07]. only [PZJ05]. OO [RSPS23]. Opacity [QG95]. OpenJDK [YW22]. Operating [HLH^{+}23, HM84, BCP08]. Operational [BLRS12, Han94, MF09]. Operations [AKBLN89, CK94, Lee86, LS79]. Operator [CMPP23, CSV01, Hen83, LdR81]. Operators [Ive79, She91]. Optimal [BOV85, CGST95, FK85, JCMM19, KRS94, Lar95, PB97, Hai98, JNZ06, KSV96, MSRR00]. optimality [CP96]. Optimally [BL94a]. Optimistic [PM04]. Optimization [BPRB23, Bee94, BBC16, Blo94, BAC16, BT93, DF84, DP97, DDH84, Dha91, DS88, FOW87, HG83, HOYY18, Pem83, PP94, RRB19, SS82, Sor89, TvS82, Web95, Ass00, BHK07, KBC^{+}99, KF03, PE08, TVA07, ZP10, CG95, LaL84, OKN06]. Optimization-Aware [BPRB23]. Optimizations [CC95, JSB^{+}12, CGS^{+}03, CKT86, GMP^{+}00, SYK^{+}05]. optimize [DMM01, VBLG04]. Optimized [CM93, Cop94, Hen82, WST85, DS98, UM02]. Optimizer [DF80, FSS83, HRL^{+}23, DF81]. Optimizers [Gie83]. Optimizing [CEG07, KMM^{+}98, LLOY23, LSLR05, ML80, NSZS13, QR00, BGKR09]. Or-Parallel [GJ93]. orchestration [PE08]. Order [AC94, AD98, Bur84, CJK95, DP97, DJP^{+}16, JPP91, JS94, SS98, BBTS07, DF11, FPS19, MPLM23, SKS11, SV19, SP97]. Ordered [SCR24]. Ordering [FSH23, GS99]. Organization [Han81a]. Oriented [Bor81, Dar90, Ell82, FFF^{+}18, GTWA14, GKL94, GP81, HU96, Ryu16, SM81, Ste22, Tur84, YB87, YB88, BSvGF03, DWWW08, DMM01, GMP24b, JPS^{+}08, WKD04, WP10, WJS^{+}00]. origins [San09]. OSI [CDFP89]. Output [Ber80, BS83]. overflow [KOE^{+}06]. overhead [BP12, SS96]. overlays [SWU10]. Overload [Bak82]. overloading [SS05b]. Overview [AOC^{+}88]. ownership [DDM11, SS96]. Oz [VHB^{+}97].

Package [Hil88]. Paper [GM81]. Paradigms [JZX^{+}24]. Parallel [ANP89, BG22, BOV85, BO94, BE13, Cha93, CGST95, CMN91, CL94, DS83, Fos96, GLO88, GJ93, GPA^{+}01, HCHP92, HIT97, JF81, Kna90, LHR19, Mis94, NSZS13, OA88, Rao94, Ras24, SS88, VMLY22, YGRBA23, dBJPT24, BBYG^{+}05, CG86, GB99, HBJ98, KSV96, LK02, MVV^{+}01, RR03, YF98]. Parallelism [Bur84, GP95, KSV96, NB99, PW94, TCVB14, YBL16]. Parallelization [BAC16, BDJ13, PP94, BdlBH99, HAM^{+}05]. Parallelizing [HP96, ME97, RD97]. Parameter [Gaz83, Zho96]. Parameterization [TWW82]. Parameterized [CGJ97b, CK93, Gaz83, RKSR12]. Parametric [HFC09, MMG92, SRW02, IV06]. Parametricities [DPP22]. Parenthesis [AS80]. Parlog [CG86]. Parsed [Wad90]. Parser [DDH84, JKT23, JP17, LaL84, SS82]. Parsers [BN99, LaL81, MYD95, PK80, CPRT02, SJ06, ST00b]. Parsing [CH87, DMM88, Fis80, GM79, Lar95, RH87, Sam80, SJW23, WG98, KC01]. Part [LaL81, PA85, PA86a, PA86b, Apt81]. Partial [AFV98, CP17, CK93, DS88, Gom92, KCL^{+}99, SCP23, Sor89, ADR06, BP12, CG04, GJ05, LMD98, Leu04, ST00b]. Partially [BLH12, Kob98, VP23, RRSY08]. partially-flow-sensitive [RRSY08]. partitioning [RM07, YF09]. Parts [Son87]. Pascal [LS79]. Pass [Bak82, BM94]. Passing [BDM15, Gaz83, SS84, VMLY22, CSW06, Gor04, Zho96]. Passive [AKP94]. Passport [SSFZ^{+}23]. past [PM09]. Path [Blo94, CIJGP18, SMP10]. path-length [SMP10]. Patient [FFF^{+}18]. Patient-Oriented [FFF^{+}18]. Pattern [EGP14, ADR06, Jay04, MTSS09, Van06]. Pattern-Based [EGP14]. Patterns [GH80]. PDS [Han81b]. PEAK [PE08]. Peephole [DF80, DF81, Pem83, TvS82]. PegaSys [MH86]. Pennello [Sag86]. Perfect [Duc08]. Performance [HU96, MSM^{+}16, PB80, URJ18, VP23, KF00, PE08]. Performed [Coh91, Wir91]. Permission [BPP16, SNS^{+}14]. Permission-Based [BPP16, SNS^{+}14]. permissions [Boy10]. Persistent [AM85]. Petri [JTM98]. Petri-Net-Based [JTM98]. Phases [Bar81]. Philosopher [CM84]. Philosophers [MS88]. Pi [BG22, HR02, KPT99]. Pi-calculus [BG22, HR02, KPT99]. pict [SWU10]. Pictures [MH86]. Pipeline [HG83]. Pipelined [BG89b, LPP01, RDG08]. pipelining [ME97]. pitfalls [Mon08]. PL [CD79, CZ84, FFF^{+}18]. PL/CS [CD79]. PL/CV3 [CZ84]. place [GW99]. Placement [DP93, GS99, vHK00]. Platform [TCP^{+}17]. pluggable [MME^{+}10]. Plurals [Ste22]. Pluto [BAC16]. Point [CK94, Fat82, SBB^{+}19, GJ05, Hau96, Mon08]. Pointer [LTMS20, LHR19, LS79, RR03, SDB20, HBCC99, HVDH07, PKH07, RLS^{+}01]. Pointers [SS13, RR05]. Points [GKM20, WKD04]. Points-to [GKM20]. Pointwise [VSS94]. Policies [NBG13, BDFZ09, FGM07b]. Policy [Kro82, Kro83, Kro84, Kro85, Kro86, Kro87, Kro88, Kro89, Kro90, Kro91, Kro92, UJ92, BFG08]. policy-based [BFG08]. Polyhedra [GVC15]. Polyhedral [GVC15, QR00]. POLYLITH [Pur94]. PolyMage [JB20]. Polymorphic [BMR05, Dug99, Gor21, HT04, Hen93, KTU93, LO94, LY98, Oho95, SIG17, SV96, TY21, WJ98, BSvGF03, DWWW08]. Polymorphism [Bur90b, MDCB91, HFC09]. polynomial [BAL07, CFG19]. PolyTOIL [BSvGF03]. polyvariance [LMD98]. Polyvariant [AC94, WJ98]. POP [FFF^{+}18]. POP-PL [FFF^{+}18]. Portable [DDH84, Han81b, HK07]. Porting [HLH^{+}23]. Possibly [JP17, ML21]. Postfix [DS83]. Postpass [HG83]. Power [TWW82, SSD09]. Powerlist [Mis94]. PPMexe [DKV07]. PQ [GZ05]. PQ-encoding [GZ05]. Practical [AD98, BAC16, BF87, CP17, Dha91, LR19, ND16, PBR^{+}15, SS13, TSL^{+}02, WC97, Bou05, DR05, DVD07, DGS97, JNZ06, PFH11]. Practice [KRS94, Ryu16, Bla03, DRSS96]. Pragmatic [BDH^{+}16]. Pragmatics [Gom92]. Pre [OLH^{+}16]. Pre-Analysis [OLH^{+}16]. Precedence [CMPP23, Hen83, LdR81]. Precise [CDK^{+}18, FJK^{+}17, GKM20, Hor97, TN19, PHP02]. Precise-Yet-Efficient [TN19]. precision [ZGZ05]. Precondition [Boo82]. Predicate [Lam90, BMR05, Bou05, Bou06, MFRW09, MMS96, PR07]. Predicates [CBDGF95, Lam88]. predictable [SHB^{+}07, HK07]. Prediction [CGJ^{+}97a, CEG07, YS99]. Predictive [FJK^{+}17]. Prepaging [FK85]. Prescription [FFF^{+}18]. Presence [AWW95, CF94, KTU93]. preserving [DHS09, LST02]. Pressure [KEA^{+}24]. pretenuring [BHM^{+}07]. Pretty [Chi05]. Prettyprinter [Wat83]. Prettyprinting [Opp80]. Primitive [App15]. principals [TZ07]. Principled [LTMS20]. Principles [Bou88, DRSS96]. printing [Chi05]. Priority [CH90, Fid93]. Priority-Based [CH90]. Prisma [RKW^{+}23]. Privacy [BKOZB13, LVRG21, TDA^{+}23]. Privileges [Min84]. Probabilistic [AB20, BKOZB13, CFNH18, DG19, HSP83, MMS96, OGJ^{+}18, Rao94, SV19, BH99, PPT08]. Problem [ADG^{+}94, CM84, DS88, Gho93, LSP82, MS88, Pet82, Pet83b, PB97, Sor89, FGM^{+}07a, Wu04]. Problems [Bac84, CFNH18, DP93, MMR95, SRW98]. Procedural [HF87, Lin93, VSS94]. Procedure [CDK^{+}18, GS99, GL80]. Procedure-Modular [CDK^{+}18]. Procedures [AM85, Kat84, NO79]. Process [Kob98, vPS81, WP10]. process-oriented [WP10]. Processes [AFdR80, Bag89, FDY12, HM84, KS79, MW84, MC82b, Oss83, RY88, Sou84, TY18, dBB85, AE98, KS10, Ber80, Moi83]. Processing [GH80, HSG17, Rei83]. Processor [BG89b, Bud84]. Processors [GLR83, Per79, LPP01, ZP10]. Product [EMH20, RTP17]. Production [Wad90]. Productivity [Sij89]. Profile [BHM^{+}07, YUW02]. Profile-based [BHM^{+}07]. Profiles [VP23]. Profiling [ASAVF19, BL94a, BPRB23, SP97]. Program [Bal94, BNNN22, Bar85, BAL07, BKB80, Col84, DKKL18, Der85, FPS19, Fea82, FOW87, FT94, FL91, HSP83, HKR94, HSD22, HLH^{+}23, Jen97, JJCO19, KKW14, KWL09, Lam83, Lam88, LLOY23, LFF14, LWR21, MS83, MW80, Mis81, Nie85, PP94, PPS79, Rem81, RTP17, SBS22, TSY00, Wat94, Wey83, ZSD09, Ass00, DDD05, GZ04, KF03, LH08, NS13, Pau01, RAB^{+}07, SLC03, WZ07, WN08, YF09, DKV07]. Programming [AGT89, Ano18, AR84, ABPS98, BS86, BPP16, BHM^{+}19, BL87, Bir84, Bor81, BMPT94, BWP87, BCEM15, CHY12, COE^{+}20, CL94, Dar90, DFR15, DGL^{+}79, DHH^{+}24, Dug99, FFF^{+}18, Fos96, FL15, GTWA14, GMP24b, HMY^{+}24, Har80, HK85, HO82, Kai89, KH92, Lee86, LVV^{+}83, LMM21, LVRG21, MK94, Mye90, OGJ^{+}18, Pet83b, RCS93, SS84, Ste22, SNS^{+}14, SZBH86, TK94, TGT20, YGRBA23, ZSO21, ABH06, BMR01, Bou06, BdlBH99, CU08, CG86, CKT86, DWWW08, DPPR00, GW99, HBJ98, JPS^{+}08, KGMO04, MVV^{+}01, MTSS09, MQ05, Tra08, VWJB10, WKD04, WJS^{+}00, Bir85, SWU10]. Programming-in-the-Large [MK94]. Programs [AWW95, AK87, AFV98, AB20, AR80, AP94, AC94, BL94a, Ban87, BGL93, BC85a, BC85b, Ber94, BCD90, BE94, BE13, BEF^{+}16, CR87, CB80, CM86a, Cha93, CFNH18, CFG19, CEW14, CMN91, Cla80, CFM94, CS87, DSFG21, DL18, DGMP97, DW89, Deb89, DL93, Deb95, DP97, Dil90, EMH20, EGP14, FJK^{+}17, FNBG20, GG85, GM81, Har80, HCHP92, HPR89, How80, HIT97, ISY88, ITF^{+}22, JBK18, JW17, Jon83, JF81, Kna90, Lam79, LS83, MSJ94, ML21, MTK21, MRGP20, MH86, Mye18, NSZS13, OA88, OL82, PS92, QL91, Rao94, SS98, Sch82, SSS81, SS88, TOUH21, TN19, VMLY22, Ven95, Wad90, Web95, Wil82a, AE01, AAE04, BCG^{+}07, CSW06, CSS99, DP99, DDV99, DS98, DMM01, EGM01, GM12, GHB^{+}96, GH97, GPA^{+}01, Hau96, HPMS00, JPS^{+}08, KSV96]. programs [LMD98, Leu04, LS09, MF09, NR06, PM06, RKRR04, RR03, San96, VJB12, WM12, YS10, Yin11, dHB^{+}96, Bur84, Lam80]. PROLOG [LV94, AP94, AB94, BC91, CH87, FA93, GPA^{+}01, MWB94, NF89, Zho96]. Promotion [Bir84, Bir85]. Proof [AFdR80, BDJ13, FRW90, GL80, Moi83, Sag86, SS84, Sok87, WGS92, WGS93, AM01, DSW11, Oho07]. proof-carrying [AM01]. Proof-Directed [BDJ13]. Proofs [Apt86, BC85a, CM86b, HRV^{+}23, JW17, LY98, Oss83, GRSK^{+}11]. Propagation [SR95, WZ91, Apt00, CP96, SS05a, SS08, SS09]. Properties [ACW90, AS89, CIJGP18, Kar84, LM18, OL82, RY88, TB95, Wei89, YS10]. Prophecy [LM22]. Proposed [Fat82]. prossima [MP10b]. Protected [PAS^{+}15, WJS^{+}00]. Protocol [SL92, YS97]. Protocols [MB83, RKW^{+}23, BFGT08, SS96]. Prototype [WCW90, WCW91]. Prototypes [HW82]. Provably [SDB20, GB99]. Provenance [ZSS20]. provenly [AAD^{+}07]. Proving [DGMP97, GC86, Hen86, Kar84, Lam79, Lam80, OL82, dBJPT24]. Pruning [BN99]. PSG [BS86]. PSO [FSH23]. publish [Eug07]. publish/subscribe [Eug07]. Pure [BNN18, HU96, Pip97, Tar07]. Purpose [App94b, HSS^{+}14, Spo86]. Pushdown [CBMO19, JKT23]. PYE [TN19].

qualifiers [FJKA06]. Qualitative [CFNH18]. Quality [BHM^{+}19]. Quantification [Vol91, Bur91]. Quantified [Gro06, STS03]. Quantitative [CFNH18]. Quantum [FDY12, HRL^{+}23, BH99, Yin11]. Queries [Bal94, CGG^{+}19]. Queuing [BB79]. Quiescence [CM86a].

R [CKT86, KMM^{+}98, AW82]. R. [Tic88]. race [AFF06, PFH11]. Races [KZC15]. Random [AS80]. Randomized [TOUH21]. Range [CG95]. Rank [Dam03]. Ranking [Lee09, TOUH21]. Ratio [CK94]. rational [GS11]. rationale [CMLC06]. Re [Ras24]. Reach [FKW98]. Reachability [LZR22, NS13, TOUH21]. Reactive [DFR15, AG04, DGG97]. read [AE01, PZJ05]. read-only [PZJ05]. read/write [AE01]. Readable [Spo86]. Reading [Pet83a]. Real [AL94, MMG92, RS84b, GH97, HK07, LS98, YMW97]. Real-Time [MMG92, RS84b, GH97, HK07, LS98, YMW97]. realities [Gor04]. Reals [DK17]. Reasoning [BKOZB13, BLRS12, BDP93, BP82, BH99, CB80, DSFG21, Lam88, LN15, Rao94, SDB20, dBH21, TSBR08]. receive [Gor04]. receptive [ABL03]. Recipe [AL94]. reclassification [DDDCG02]. recognition [ATD08]. Recognizer [GHR80]. Recognizing [BL94b]. Recombination [Kau84]. Recombination-Delaying [Kau84]. Recompilation [BT93, SK88, Tic86, Tic88]. Reconciling [HU96]. Reconstruction [YR94]. Record [LS79, Oho95]. Recovery [AB81, ACS84, Bac84, BF87, GHH^{+}19, PK80, Ric85, dJKVS12]. recurrences [VJB12]. Recursion [AK82, Col84, Hen93, KTU93, Mis94, YK97]. Recursive [AC93, AK82, Ban87, CFG19, Coh83, Coh85, LBN17, Sij89, ZZO22, ABE^{+}05, AM01, CF04, Dug02, Pal98]. Recursively [BE13]. Reduce [BN99, MYD95, BALP06, KOE^{+}06, SS96]. reduced [SG04]. Reducible [Hav97, JC97]. Reduction [Bee94, Bur84, FRW90, Geo84, KLS92, Mul92, NN86, CSV01]. Redundancies [Sor89, DS88]. redundancy [KCL^{+}99]. Redundant [Coh83, Coh85]. Reentrant [Bob80]. Reexamination [CG95]. Refactoring [Ste18, TFK^{+}11]. Reference [Bob80, Pea21, Wis79, KSK07, KOE^{+}06, LP06, MDJ05]. reference-counting [LP06]. References [Han92, TGT18, TGT20, SV96]. Referencing [LS81]. Referential [QG95]. Refinement [BBF^{+}11, BKL^{+}97, BCEM15, CM86a, DGL^{+}79, EKD23, GEGP17, JLP^{+}14, MRG88, SL92, AG04, QG95]. reflecting [HS11]. reflection [SW97a]. Regeneration [SR21]. Region [TB98, SYN06]. region-based [SYN06]. regions [RR05]. Register [BCT94, CH90, GSO94, JLF02, LCBS19, RDG08, SH89, GA96, HCS10, LGAT00, PM04, PS99, PF96, TP04]. registers [ZP07]. Regular [CC97, HVP05, KBP22, LaL81]. relating [ABC^{+}21]. Relation [LBN17, MTG80]. Relational [BNNN22, BKOZB13, CB80, GS98, TLHL11, JJD98, JLRS10]. Relations [ELS^{+}14, HT86, LH08]. Relationship [BS88]. Reliability [LM18, WN08]. Reliably [TCP^{+}17]. Rely [GEGP17, LFF14, SZLY21]. Rely-Guarantee [GEGP17, SZLY21]. Rely-Guarantee-Based [LFF14]. Remembrances [PM09]. Remote [BCP08, SG90]. Removal [AK82]. Rendezvous [Cha87]. Renvoise [Sor89, Dha91, DS88]. Reoptimization [PS92]. reordering [YUW02]. Repair [BN99, MF88, MYD95, KC01]. Repairing [CPRT02]. Replacement [MM89]. Replicate [RB94]. replication [RD03]. Reply [Bur91, Fra81, LaL83, Tan83, Wir91, SM82]. Representation [DGL^{+}79, Mul92, SM89, Wad90, Wan82, Mil85]. Representation-Independent [Mul92]. Representations [Lam87, RF97, Wal80, Wal81, BGP99]. Reproduction [BHM^{+}19]. Repulsing [TOUH21]. reshaping [ZCG^{+}07]. Resilient [GHH^{+}19, WL85]. Resolution [ABR81, Bak82]. Resolved [SIG17]. Resource [CS95, Cla80, IK05, MQ05, BDFZ09, CEI^{+}07, HR02, HAH12]. Resources [And81, FLBB89]. Respect [Gaz83]. Response [Tic88]. Responsibility [DC22]. Responsiveness [HU96]. Restores [Wis79]. Result [TB95]. Results [Ven95, BGP99, SYYH07]. Retargetable [DF80, DF81, MV87]. Retention [LS81]. Rethinking [LHR19]. retrofitting [NCH^{+}05]. Return [SDB20]. reuse [DNS^{+}06, GW99, ZSD09]. Reversal [ACS84]. Reverse [PS08]. Reverse-mode [PS08]. Revised [SIG17]. Revision [FM87b]. revisited [MDJ05, Zho96]. Revisiting [DI09, ZZO22]. Rewrite [FKW98, Ass00]. Rewriting [KKSD94, BCM99, DDD05, FKW00, GRSK^{+}11, MMM^{+}07]. Right [KS83, LaL81, SJ06]. Rigorous [SBB^{+}19]. Rings [BP89, Hua93]. RISC [PS93]. Rivieres [Hen83]. RMI [MVV^{+}01]. Robust [LS83, PKWC24]. Robustly [PG21]. Robustness [GMP24a]. Roever [Moi83]. role [Apt00]. Roman [PB97]. Round [SBB^{+}19]. Round-Off [SBB^{+}19]. Rounding [FL15]. Row [MM89]. RSMs [CGG^{+}19]. rule [HQRT02]. Rules [GL80, JTM98, SS84, LS09, SSD09]. Run [ISY88, TZ07, GMP^{+}00]. Run-Time [ISY88, TZ07, GMP^{+}00]. Runtime [Ano18, BLH12, BEF^{+}16, FNBG20, GSS^{+}18, ISIRS22, TCVB14, BH05a, TSY00]. Rust [MTK21, Pea21]. RustHorn [MTK21].

S [HCW82]. S/SL [HCW82]. Safe [AWW95, Dug02, HMY^{+}24, JW17, LMM21, PG21, SDB20, AFF06, BSvGF03, LS03, Loc13, NCH^{+}05, SA00, ZCG^{+}07, MH06, SHB^{+}07]. Safe-by-default [LMM21]. safe-for-space [SA00]. Safer [COE^{+}20]. safety [FF08, YS10]. same [SS05a]. sampling [PPT08]. Santa [WP10]. Sapphire [URJ18]. Sather [MOSS96]. Satin [VWJB10]. satisfaction [DF11]. Satisfiability [FSH23, XA07]. satisfying [Van96a, Van96b]. Saturn [XA07]. Scalability [TCP^{+}17]. Scalable [FT94, GKM20, ZSS20, XA07]. ScalaExtrap [WM12]. scale [ZSS20]. Scaling [TCP^{+}17]. scan [PS99]. Scanners [HKR92]. Scanning [GVC15]. Scavengers [UJ92]. Schanuel [KPS92]. schedulability [GH97]. schedule [TVA07]. Scheduler [TCVB14]. schedules [MH04]. Scheduling [BG89b, FGL94, KR79, KPF95, LPP01, LJ99, LCBS19, NB99, NSTD^{+}15, PS93, TCVB14, Ban11, ME97, YF98]. schema [RLS^{+}01]. Scheme [Mur91, YR94, IV06, WC97]. Schemes [Son87, TM93]. Schorr [BP82]. Schwanke [Tic88]. Scientific [How80]. Scope [App94b]. Scratchpad [SRM10]. Screen [MM89]. SDF [VHKO02]. Search [Dar90, LLOY23, BH99, SS05a]. Searching [CC97]. Section [Mül21, Wol92]. Secure [ABC^{+}21, BCEM15, PAS^{+}15, PG21, BBF^{+}11, HY07]. Securely [RB94]. Securing [BNV^{+}21]. Security [TGT18, TGT20, BFGT08, BFG08]. see [BR10]. Selection [DF84, SSS81]. Selective [LTMS20, Min84, OLH^{+}16, ME97]. Self [BP89, DHS09, Gho93, Gom92, NN24, ABB^{+}09]. self-adjusting [ABB^{+}09]. Self-applicable [Gom92]. Self-Stabilization [Gho93, DHS09]. Self-Stabilizing [BP89, NN24]. Semantic [AAR^{+}10, AWW95, GGL15, LCK^{+}22, ML21, MH06, HCW82]. Semantics [ABHI11, Ans87, AB94, AW82, BGL93, Ber94, BLRS12, Bou88, Boy10, CPS93, CD79, DHH^{+}24, FA93, GM81, Gud92, Han94, JPP91, Kai89, Mul92, NF89, Set83, Sou84, WM95, Wan82, dBB85, ACE96, BMR01, Bou06, GZ04, MF09, PCJD08, SWU10, SJ03, Tar07, WKD04]. Semantics-Based [BGL93, CPS93, PCJD08]. Semantics-Directed [Han94, Set83]. Semaphore [CR87]. Semiring [BMR01]. Semiring-based [BMR01]. Send [Gor04]. Send-receive [Gor04]. Sensitive [OLH^{+}16, HBS22, PKH07, Ram00, Rep00, RRSY08, VP23]. Sensitivity [FL15, KRR18, LTMS20]. Separating [DDM11, Ste22]. Separation [BDJ13, DJP^{+}16, OYR09, BBTS07, PZ22]. Separators [GSO94]. Sequences [GSW95, LWR21]. Sequent [ABS09, Miq19]. Sequential [AFdR80, Ber80, Gor21, GLR83, HM84, KS79, MC82b, Moi83, Sou84]. Sequentialization [ITF^{+}22]. Series [Wat91]. Served [LH91]. server [LDM07]. servers [BBYG^{+}05]. service [CMS03]. Services [CHY12, RB94, BFG08, CGP09].

Session [DDMP22, KBP22, Pad19, TY18]. Session-Based [TY18]. Sessions [TY21]. Set [Sha82, FF99]. set-based [FF99]. SETL [DGL^{+}79, FSS83, SSS81]. Sets [DP82, DPPR00]. Setting [Lin79, Nie85, HL05]. SHA [App15]. SHA-256 [App15]. shape [GCRN11, JLRS10, JB06, SRW98, SRW02]. shape-analysis [SRW98]. shaping [HS11]. Share [SS88]. Shared [Cha93, FLBB89, KH92, KRS88, LB22, Pet83b, Dug02, HBJ98, TSY00, BC91]. Shared-Memory [Cha93, TSY00]. Sharing [CSS99, Lam87]. SHErrLoc [ZMVPJ17]. Shift [BN99, MYD95]. Shift-Reduce [BN99, MYD95]. Short [Sag86]. Should [LP99]. Side [Boe85, SCP23, KWL09, RLS^{+}01, TA08b]. Side-channel [SCP23]. side-effect [RLS^{+}01]. sign [KKN06]. signal [BH05b]. Signatures [BR97]. Signedness [GNS^{+}15]. similar [AE98]. Simple [Boe85, GLO88, JP17, LM22, SH89]. simpler [BKRW98, BKRW05]. Simplification [LZR22, NO79]. Simula [Lan80]. Simulating [KKSD94]. Simulation [AMT14, Bar81, Bor81, LFF14, HQRT02]. sin} [Lam90]. Single [BM94, CFR^{+}91, JBK18, GPF08]. Single-Assignment [BM94]. Single-Pass [BM94]. Single-Threaded [JBK18]. Sit [AKNP17]. situations [WN08]. Size [BA08, BEF^{+}16, JB20, Lee09, LDK^{+}96]. Size-change [BA08, Lee09]. Sized [DG19]. Sketches [HSD22]. Slicing [AB20, AHJR14, CF94, DL18, GH97, HRB90, ML21, Mye18, Ven95, WZ07, BHK07, GZ07, NR06, RAB^{+}07, WR08, ZGZ05]. SLR [BS88, Tai79]. Small [BNV^{+}21, FLBB89, LH91, Pet83b]. Smart [Tic86]. Smarter [SK88, Tic88]. Smooth [CCEG23, JF81]. Soft [WC97]. Software [ACM11, AW85, Ber94, DAW88, HMY^{+}24, HSS^{+}14, How80, JW17, PXL95, PPS79, Pur94, Wal92, YBL16, CTT07, HN05, LS98, ME97, NCH^{+}05, RDG08, SHB^{+}07, SRM10]. Software-Defined [Wal92]. Soininen [LaL84]. Soisalon [LaL84]. Soisalon-Soininen [LaL84]. Solution [ADG^{+}94, DS88, Gho93, Pet83b, Sor89, WP10]. Solving [GS11, HLH19, HSD22, NSTD^{+}15, SRW98]. Some [AB94, AK82, Sha82, Sor89]. Sometimes [Gri79]. Sound [LLK^{+}17, LCK^{+}22, RSPS23]. Soundness [Sok87]. source [HBG^{+}09]. Space [BP12, BB79, FLBB89, JP81, NB99, RD87, YF98, LS09, SS05a, SA00]. Space-Efficient [JP81, NB99]. Space/time [YF98]. Space/time-efficient [YF98]. spaces [JLF02]. Span [LS80, Rob79]. Span-Dependent [LS80, Rob79]. Spanning [GHS83]. Sparse [OHL^{+}14]. Spatial [NSTD^{+}15]. Special [Ahm20, Mül21, Wol92, Yos22, Sag07]. Specialization [AHJR14, BCP08, GJ05, HT04, SLC03]. specialization-point [GJ05]. Specializing [BCD90]. Specific [ASAVF19, Gie83, Tra08]. Specification [BCM99, CDFP89, EO80, Fea87, GMH81, Jon94, Kam83, LN15, Lin93, LJ99, Loe87, Mal82, Mor88, PPS79, RY88, TWW82, LP99, LPSO04]. Specificational [MB99]. Specifications [AL93, AL95, BNN18, CES86, DB85, Gaz83, Loe87, MW84, MB83, Rei83, Sch85, Win87, Zav85, Žic94, vPS81, JJD98, YS97]. Specifying [GM81, Lam83, RF97]. Speculation [YBL16, GB99]. speculative [KOE^{+}06]. SPL [HSG17]. Split [Com80]. splitting [JC97, UM02, WJ98]. SPMD [WM12]. Squeezers [ISIRS22]. SR [AOC^{+}88]. SSA [BDP14, GSW95, KCL^{+}99]. SSA-Based [BDP14]. SSProve [HRV^{+}23]. Stabilization [Gho93, DHS09]. Stabilizing [BP89, NN24]. Stack [CGS^{+}03, FG03, LaL81, SDB20, CF04, Zho96].

Stack-Controlling [LaL81]. Standard [Fat82, HM93, Qia00, Blu99]. State [ACW90, BLH12, CES86, GC86, PP91, Pur91, Zav85, AY01, ABE^{+}05, MBT09]. Statement [Ell82, Mor88, SM81]. Statement-Oriented [Ell82, SM81]. Statements [CF94]. States [ADGM91, CBMO19]. Static [AKNP17, AC94, BM94, CGJ^{+}97a, CF94, CFR^{+}91, DL18, Deb89, HOYY18, LLK^{+}17, LST98, MRGP20, MOS07a, Mye18, PW94, SBE^{+}19, YS99, ZMVPJ17, CEI^{+}07, GPF08, GZ04, HO07, PSS05, PFH11, RSL10, VJB12, WCM00, YF09, AFF06, FFLQ08]. Statically [ACPP91]. Statistical [LLK^{+}17]. Statistics [Lan80]. Staveren [Pem83]. Steensgaard [Ell82, SM82]. Steensgaard-Madsen [Ell82, SM82]. stencil [LS04]. Step [Col84, TVA07]. Steps [Jon83]. Stepwise [CM86a, SL92]. Stevenson [Pem83]. Storage [BBC16, Bre89, JP81, KEA^{+}24, LDK^{+}96, Mur91, Rob79, Sip82, KOE^{+}06, TVA07]. Strategies [Bir84, Bir85, Geo84, NN86]. Strategy [Bre89, PK80, WSH15, ZSS20, GS11]. Stratified [SS98]. Stream [HSG17]. streams [CFP^{+}04]. strength [CSV01]. Strict [Bee94]. Strictness [Bee94, SR95]. String [GH80]. Strings [AS80, KS88, KS89, ADR06, KK07]. Strong [KZC15, PZ22]. Strong-separation [PZ22]. Structural [SZBH86, MTSS09]. Structurally [HS11]. Structure [BC79, GKL94, HM93, Mis94, MWB94, She91, HY07]. Structure-Oriented [GKL94]. Structured [BM94, CHY12, GD82, Har80, LS81, Mur91, RR03]. Structures [ANP89, Bob80, FL81, GEGP17, RCRH95, SSS81, LPSO04, RAB^{+}07]. Study [BHM^{+}19, FTJ95, BHK07, BdlBH99, DF98, KF03, LS98]. Style [BDM15, LR19]. Sublanguage [DGL^{+}79]. Sublinear [RD87]. Sublinear-Space [RD87].

Submodule [MB83]. Subroutines [SA99]. subscribe [Eug07]. Subscript [CG95]. Subsequence [Han92]. Subset [BL87]. Substrings [BL94b, Han92]. subtype [Duc08, KR01]. Subtypes [Vol91, Bur91]. Subtyping [AC96, AC93, GGL15, LN15, LR19, LBN17, LW94, XBOS20, ZZO22, GZ05, IV06]. Subtyping-Relation [LBN17]. SUIF [HAM^{+}05]. Summarization [SCS24]. Supercompiler [Tur86]. Superimposition [Kat93]. Supermartingales [TOUH21]. Support [Bal94, DS90, Fea87, LS83, MK94, Wei90, TSY00]. Supporting [RCRH95]. Supports [ABPS98]. Suppression [DS88, FGL94, Sor89, JNGG10]. Survey [Apt81, GPA^{+}01]. Suspension [CFM94]. Symbol [ABR81, Rei83]. Symbolic [DMZ^{+}24, Dil90, HP96, Hal85, Hen82, NcS20, RR05, SBB^{+}19, YMW97, BGP99, MPM03, CM93, WST85]. Symmetric [FY85]. Symmetry [ES97, SG04]. Synchronisation [CHMY19]. Synchronization [Bag89, DJP^{+}16, Her91, KRS88, RS84b, Sch82, CGS^{+}03, DHM^{+}12, Ram00, RD03]. synchronization-sensitive [Ram00]. Synchronizing [And81]. Synchronous [CS87, TLHL11, YGRBA23]. synchrony [CS04]. Syntactic [BF87, GMZ00, MF88, PK80, Wil82b]. Syntax [DMM88, Ode93, Ric85, SSS83, BMR01, CPRT02, Jef03, HCW82]. Syntax-Directed [DMM88]. Syntax-Error-Handling [SSS83]. Syntax/Semantic [HCW82]. Synthesis [AE98, AE01, AAE04, Ban87, BDJ13, BKL^{+}97, Cla80, DKKL18, HLH^{+}23, JZX^{+}24, LLOY23, MW80, MW84, MV87, SBS22]. System [AFdR80, AW85, BS86, Bou88, CB80, Fea82, GD82, GP81, Han81b, HM84, JMSY92, LR13, ML80, Moi83, MH86, PO95, RD13, SA99, WC97, dBJPT24, BH05a, FH04, FM99, HO07, JB06, KS10, MTSS09, NP08, PE08, STSP05, MWCG99]. Systematic [DC22, DF98, PSS05]. Systems [ABLP93, Ano18, AR84, ACS84, BKS88, BG89a, BDP93, CI84, CDFP89, CBDGF95, CIJGP18, CES86, CPS93, CBMO19, DL18, DAW88, Fea87, FKW98, Gor21, Hen86, HLH^{+}23, Jag94, Jon94, JTM98, Kar84, Kat93, Kau84, Lam84, LW93, Mis86, Mye18, SZLY21, TGT20, WGS92, WGS93, WCW90, van88, Ass00, AE98, BCP08, BCM99, BGP99, CSCM00, DGG97, GS11, TP04, TZ07, YMW97, WCW91]. Systolic [Hen86].

T [Žic94]. Table [BMW91, PK80, DAS98]. Table-Drive [PK80]. Tabled [SS98]. Tables [ADGM91, DDH84]. TaDA [DSFG21]. Tail [DP97, CF04]. Tail-Call [DP97]. tail-recursive [CF04]. Tailored [Kau84]. Tailored-List [Kau84]. Tale [TY21]. Tanenbaum [Pem83, Tan83]. Target [Wan82]. Task [GP95, NSZS13, RRB19, HBJ98]. task- [HBJ98]. Task-Level [GP95]. Task-Parallel [NSZS13]. Tasking [Dil90]. Tasks [GP81]. Taylor [SBB^{+}19]. tcc [PHEK99]. Technical [BS88, Bur90b, Bur91, Coh91, CM93, DS88, Ell82, FA93, Fra81, Hen83, LaL83, LaL84, Moh81, Moi83, MS88, NN86, Par90, Pem83, Sor89, SM82, Tan83, Tic88, Vol91, WST85, Wir91, YB88, MMG00]. Technique [AWW95, BN99, BCD^{+}15, JSB^{+}12, KKM90, SSS81, SSS83, JNGG10, KBC^{+}99, RD97, SYN06]. Techniques [AK82, CMN91, DP99, GLR83, How80, TWW82, WCW90, WCW91, BHK07, DDD05, DEMD00, LS98, MSRR00, SS96, TSL^{+}02]. technology [LS98]. Temporal [AS89, CBDGF95, CES86, Kar84, Lam94, MW84, GS99, KWL09]. temporal-ordering [GS99]. temporaries [RMH06]. Ten [Apt81]. Tensor [RTP17, SBS22]. Tentative [Jon83]. Tenuring [UJ92]. TeraHeap [KEA^{+}24]. Term [KKSD94, MBT09, GRSK^{+}11].

Termination [AF84, Apt86, BAGM12, BCG^{+}07, CFNH18, CDK^{+}18, DSFG21, DG19, Fra80b, GJ05, HSP83, JBK18, MC82b, TM93, BAL07, BA08, DDV99, GRSK^{+}11, Lee09, PR07, SMP10, Fra80a, Moh81]. Test [Wey83, WW95, Duc08]. Testing [AMT14, GMH81, TK94]. Tests [Coh91, Koz97, Wir91, GZ05]. Text [CC97]. TF [SBS22]. TF-Coder [SBS22]. Their [Kam83, LaL84, SS82, PS96]. Theoretic [ES97, Sha82, KV00]. Theoretical [KRR18]. Theories [NSTD^{+}15, Bou06]. Theory [AB20, CZ84, FSH23, HP24, KD94, KRS94, NBG13, Ryu16, TLHL11, CGP09, MH06, Oho07, Pau01, SS05b, Bla03, FG03]. ThingLab [Bor81]. things [PM09]. Thinking [WLBF16]. Thinning [Web95]. Third [Wol92]. ThisType [Ryu16]. Thread [YBL16]. Thread-Level [YBL16]. Threaded [JBK18, ITF^{+}22, TSY00]. Three [DPP22, Oss83]. Tichy [Tic88]. Tierless [RKW^{+}23]. Tile [JB20]. tiling [JLF02, LS04, RKSR12]. Time [AL94, ABR81, BL94b, BLH12, Coh91, DLR16, HBS22, Hol87, ISY88, Jef85, Lam84, LLOY23, MMG92, PS93, RS84a, RS84b, TN19, Wir91, YR94, Žic94, BAL07, BALP06, BKRW98, BKRW05, DDD05, GH97, GMP^{+}00, GB99, GW99, HK07, LS98, LPP01, LS09, Mil85, Ram99, Rep98, SYK^{+}05, Tra08, TZ07, Wu04, YMW97, LW93]. Time-bounded [LLOY23]. Time-Constrained [Žic94, LPP01]. Time-Critical [PS93]. time-efficient [GB99, YF98]. Time-sensitive [HBS22]. Timed [Žic94]. Timeout [Lam84]. Timing [LJ99]. tokenization [Rep98]. Tolerance [LJ99]. Tolerant [CS95, Lam84, AAE04]. Tool [CPS93]. Toolkit [BDFH97]. toolkits [VHM^{+}01]. Tools [van88]. Top [SZLY21]. Top-down [SZLY21]. TOPLAS [Ano18, TGT20, MP10a, MP10b]. topology [DDM11]. Tortoise [Dan23]. Total [San96].

Trace [ABC^{+}21, FGL94, WGS92, Ban11, RM07, SJ03, WGS93, WM12]. Trace-Based [WGS92, WGS93, WM12]. Trace-relating [ABC^{+}21]. traces [HBM^{+}06, WR08]. Tracing [BL94a, DLR16, MMM^{+}07]. tradeoffs [ZGZ05]. Trailing [VR95]. Traits [DNS^{+}06]. Transactional [URJ18, ABHI11, CFP^{+}04]. Transactions [Ano18, HKMN94, TGT20]. Transducer [DVLM15]. Transducer-Based [DVLM15]. Transformation [BKB80, Fea82, FL91, NSZS13, Wat91, RKRR04, San96, TSY00, WZ07]. Transformational [BDFH97, Bir84, Bir85, DSW82, OA88, RC03]. Transformations [Bar85, EGM01, Geo84, LdR81, LFF14, MS83, MCT96, Nie85, FGM^{+}07a, KWL09, MOS07a, VALG05, WS97, Hen83, NN86]. Transformers [Lam90, MMS96, MBT09]. TransformGen [GKL94]. Transforming [AWW95, BE94]. Transition [PR07, dBJPT24]. Translation [AK87, BK11, Kat84, Son87, AAD^{+}07, BGKR09, DP99, RC03]. Transmission [HL82]. Transparently [JSB^{+}12]. Transport [Min84]. transpose [CRN^{+}08]. Traversals [LPSO04]. Treatment [YB87, YB88]. Tree [AGT89, BOV85, BMW91, DVLM15, DS83, Han81a, Hen83, LdR81, FGM^{+}07a]. Trees [Com80, GHS83, MTG80, Sip82, Wad90, ACM11, SGL97]. Treewidth [CIJGP18, CGG^{+}19]. trick [DMP96]. Truth [BDH^{+}16]. TSL [LR13]. TSO [FSH23]. tuning [GMM99, PE08]. Tuples [Rem81]. Tutorial [GM81]. Two [BO94, CDFP89, DPP22, GPWZ08, TY21, FMoPS11]. Two-dimensional [GPWZ08]. two-variable [FMoPS11]. Type [Bur90b, Car95, CEW14, Coh91, CZ84, Dug02, Eug07, HHPW96, HM93, Hen93, HP24, KPS92, KTU93, KR01, Lan80, LO94, LST02, LY98, LP00, MRGP20, MP88, NBG13, Pad19, PO95, SA99, SM89, Ste22, TWW82, TGT18, TGT20, Van06, VMLY22, Wal80, WT11, Wir88, WC97, BSvGF03, BCG^{+}07, FJKA06, FGM07b, FM99, FF08, GZ07, GMZ00, HO07, HDH02, HY07, KF10, KS10, NP08, NCH^{+}05, PT00, STSP05, TFK^{+}11, TZ07, Wal81, Wir91]. Type-based [Eug07, LP00, BCG^{+}07]. Type-Driven [TGT18, TGT20]. Type-Extension [Coh91, Wir91]. Type-Graphs [KPS92]. Type-preserving [LST02]. Type-Safe [Dug02, BSvGF03, NCH^{+}05]. Typechecking [CL95, MBC04]. Typed [ACPP91, Geo84, GDF23, Kob98, NN86, WCM00, AAR^{+}10, LP99, MWCG99]. Typed-Untyped [GDF23]. Types [AFF06, AC93, BG22, BGOL^{+}23, BB94, BCEM15, DDMP22, DPP22, DD85, EO80, FFLQ08, GEGP17, HL82, Hes88, Jen97, Kam83, KBP22, LaL89, LO94, LBN17, Loe87, Mal82, Miq19, MP88, TDA^{+}23, WL85, Wei89, Wei90, AM01, BBF^{+}11, Dam03, DDM11, DMM01, Gro06, GPV07, HVP05, IV06, MME^{+}10, PS96, Pal98, STS03, SP07]. Typestate [COE^{+}20, GTWA14]. Typestate-Oriented [GTWA14]. Typing [ACPP91, DG19, Dug99, EKD23, GGSV22, RM10, SV96].

ultimate [PS08]. Ultracomputers [Sch80]. Unassigned [Win84]. Unbounded [LWR21, BGP99]. uncaught [LP00]. Undecidability [Ram94, Rep00, Cha02]. undecidable [Ram00]. Under-Approximation [ABG24]. Understandable [MSM^{+}16]. Understanding [ST00a]. Undo [Lee86]. unfold [RKRR04]. unfold/fold [RKRR04]. Unidirectional [Pet82]. Unification [MM82, DRSS96]. Unified [VSS94]. Uniform [AS80, BP89, Hua93, AH10, HY07]. Uniformly [DB85]. unifying [TVA07]. unique [Van06]. UNITY [Pau01, TB95].

Universal [DPP22, PKWC24]. universe [DDM11]. Unnecessary [BT93]. Untrusted [JW17]. Untyped [GDF23]. Update [Hud91, FGM^{+}07a, GW99]. Updating [HSS^{+}14, HN05, SRW98, SHB^{+}07]. Upper [PW94]. Usage [MS83, BDFZ09, IK05, QR00]. Use [FOW87, GH80, HS94, LaL84, PPS79, She91, SS82, CC97]. usefulness [HDH02]. User [ACS84, DS90, Mye90, Wal80, Wal81, Wat83, van88]. User-Defined [Wal80, Wal81]. Using [AGT89, Bob80, CGJ^{+}97a, CES86, CH87, DP93, Dil90, DMM01, DJP^{+}16, FLBB89, GSW95, GSO94, HRB90, ISIRS22, JTM98, Kar84, LaL89, Lam84, LM18, LWR21, Mye90, Ode93, Pet83b, PP94, PBR^{+}15, SSFZ^{+}23, SS84, SS96, Sok87, SGL98, TvS82, ACM11, BH99, CSW06, CGS^{+}03, DR05, GS99, GCRN11, KWL09, KSK07, MTSS09, RD03, SZLY21, ST00a, SGL96, TFK^{+}11, VJB12, XA07, YUW02, ZSD09, Pem83]. Utilizing [ES97].

VAL [McG82, Wet82]. Validation [How80, KC01, MOS07a]. Value [HL82, dBH21, HL05, SW97a]. valued [RMH06, SRW02]. Values [DD85, Han92, HSD22, Wet82]. Variable [MS83, MTG80, FMoPS11, GLMM05]. Variables [GSW95, JPP91, Lam88, LH91, Pet83b, Rem81, Sch85, BGP99, HVB^{+}99, NS13, SV96]. Variant [IV06]. variants [FG03]. Variational [CEW14]. Variety [NcS20]. Vector [AK87, Bud84, CBMO19, Fis80, FTJ95, KD94, Per79, KK07]. Verifiable [YB85]. Verifiably [HMY^{+}24]. Verification [App15, BDP14, BCD^{+}15, CDFP89, CES86, CPS93, CHMY19, DMZ^{+}24, Dil90, EGP14, GL94, ITF^{+}22, JBK18, Jon94, JTM98, KKW14, LFF14, LJ99, LCK^{+}22, LS79, MTK21, NBG13, RY88, SZLY21, SSFZ^{+}23, BDL^{+}08, CEI^{+}07, GPF08, GM12, Qia00]. Verified [BFGT08, BKL^{+}97, HL22, HRL^{+}23, JLP^{+}14, DSW11]. Verifying [AS89, BFG08, CGJ97b, DJP^{+}16, GEGP17, LM18, YS10, Mon08]. Version [YR94]. Versions [HPR89]. Versus [DPP22, Pal98, Pip97, UM02]. Vertices [BGH^{+}13]. Very [GLR83]. VHDL [BKL^{+}97]. via [CEI^{+}07, FKW98, GPF08, GSO94, HLH19, HOYY18, ITF^{+}22, MMM^{+}07, PE08, Ras24, RTP17, SRW02, SV20, SCP23, Tra08, WCM00]. View [KBP22, SZBH86, FGM^{+}07a]. view-update [FGM^{+}07a]. Virtual [Jef85, RRB19, CEG07, KN06]. Visibly [CBMO19, JKT23]. Visual [Mye90, BCM99]. vita [MP10a, MP10b]. VLSI [LVV^{+}83]. Volpano [Bur91]. Volume [Ano18, TGT20]. vs [HR02].

W [Tic88]. Wait [Her91]. Wait-Free

[Her91]. Waite [BP82]. Warp [LW93]. way [VHM^{+}01]. Weak [AMT14, KZC15]. weakening [SYYH07]. Weaker [Boo82]. web [BFG08, BLRS12, CHY12, CGP09, CMS03]. Weight [GHS83]. While [Pet83a, BC85b, GM81]. while-Programs [BC85b]. Whole [BDH^{+}16]. Widening [KKW14, VJB12]. win} [Lam90]. Within [FKW98]. Without [Cop94, Ode93, AS89, Cas95, Sto04, VR95]. Witnessing [TA08b]. Workbench [CPS93]. World [GG85, DF11]. World-Model-Based [GG85]. Worst [CFG19, WW95]. Worst-Case [CFG19]. wp [BH99]. write [AE01]. Writing [Pet83a, Win87]. WYSINWYX [BR10].

X [OLH^{+}16, MSM^{+}16]. X-Sensitive [OLH^{+}16]. X10 [GHH^{+}19]. XARK [ATD08]. XML [HVP05, HFC09]. XSL [MOS07a].

Years [Apt81].

ZGC [YW22]. Zones [GMN^{+}21].

References

Ancona:2007:PCT

[AAD^{+}07] D. Ancona, C. Anderson, F. Damiani, S. Drossopoulou, P. Giannini, and E. Zucca. A provenly correct translation of Fickle into Java. ACM Transactions on Programming Languages and Systems, 29(2):13:1– 13:67, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Attie:2004:SFT

[AAE04] Paul C. Attie, Anish Arora, and E. Allen Emerson. Synthesis of fault-tolerant concurrent programs. ACM Transactions on Programming Languages and Systems, 26(1):125–185, January 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ahmed:2010:SFT

[AAR^{+}10] Amal Ahmed, Andrew W. Appel, Christopher D. Richards, Kedar N. Swadi, Gang Tan, and Daniel C. Wang. Semantic foundations for typed assembly languages. ACM Transactions on Programming Languages and Systems, 32(3):7:1– 7:67, March 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anderson:1981:LLC

[AB81] S. O. Anderson and R. C. Backhouse. Locally least-cost error recovery in Early’s algorithm. ACM Transactions on Programming Languages and Systems, 3 (3):318–347, July 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Arbab:1994:SCD

[AB94] Bijan Arbab and Daniel Berry. Some comments on “A denotational semantics for Prolog”. ACM Transactions on Programming Languages and Systems, 16(3):605–606, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177605.html.

Amtoft:2020:TSI

[AB20] Torben Amtoft and Anindya Banerjee. A theory of slicing for imperative probabilistic programs. ACM Transactions on Programming Languages and Systems, 42(2):6:1– 6:71, May 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3372895.

Acar:2009:EAS

[ABB^{+}09] Umut A. Acar, Guy E. Blelloch, Matthias Blume, Robert Harper, and Kanat Tangwongsan. An experimental analysis of selfadjusting computation. ACM Transactions on Programming Languages and Systems, 32(1): 3:1–3:53, October 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Abate:2021:EAT

[ABC^{+}21] Carmine Abate, Roberto Blanco, Stefan Ciobâca, Adrien Durier, Deepak Garg, Catalin Hrittcu, Marco Patrignani, Éric Tanter, and Jérémy Thibault. An extended account of trace-relating compiler correctness and secure compilation. ACM Transactions on Programming Languages and Systems, 43(4):14:1– 14:48, December 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3460860.

Alur:2005:ARS

[ABE^{+}05] Rajeev Alur, Michael Benedikt, Kousha Etessami, Patrice Godefroid, Thomas Reps, and Mihalis Yannakakis. Analysis of recursive state machines. ACM Transactions on Programming Languages and Systems, 27(4): 786–818, July 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ascari:2024:LDD

[ABG24] Flavio Ascari, Roberto Bruni, and Roberta Gori. Limits and difficulties in the design of under-approximation abstract domains. ACM Transactions on Programming Languages and Systems, 46(3):11:1– 11:??, September 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3666014.

Acar:2006:AFP

[ABH06] Umut A. Acar, Guy E. Blelloch, and Robert Harper. Adaptive functional programming. ACM Transactions on Programming Languages and Systems, 28(6): 990–1034, November 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Abadi:2011:STM

[ABHI11] Martín Abadi, Andrew Birrell, Tim Harris, and Michael Isard. Semantics of transactional memory and automatic mutual exclusion. ACM Transactions on Programming Languages and Systems, 33(1):2:1–2:50, January 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Amadio:2003:RDC

[ABL03] Roberto M. Amadio, Gérard Boudol, and Cédric Lhoussaine. The receptive distributed π-calculus. ACM Transactions on Programming Languages and Systems, 25(5):549– 577, September 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Abadi:1993:CAC

[ABLP93] Martín Abadi, Michael Burrows, Butler Lampson, and Gordon Plotkin. A calculus for access control in distributed systems. ACM Transactions on Programming Languages and Systems, 15(4):706– 734, September 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
155225.html.

Afek:1993:LC

[ABM93] Yehuda Afek, Geoffrey Brown, and Michael Merritt. Lazy caching. ACM Transactions on Programming Languages and Systems, 15(1):182– 205, January 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
151651.html.

Apt:1998:AIL

[ABPS98] Krzsztof R. Apt, Jacob Brunekreef, Vincent Partington, and Andrea Schaerf. Alma-O: An imperative language that supports declarative programming. ACM Transactions on Programming Languages and Systems, 20(5):1014– 1066, September 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-5/
p1014-apt/.

Andre:1981:MAC

[ABR81] F. André, J. P. Banatre, and J. P. Routeau. A multiprocessing approach to compiletime symbol resolution. ACM Transactions on Programming Languages and Systems, 3(1): 11–23, January 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ariola:2009:SCA

[ABS09] Zena M. Ariola, Aaron Bohannon, and Amr Sabry. Sequent calculi and abstract machines. ACM Transactions on Programming Languages and Systems, 31 (4):13:1–13:48, May 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Amadio:1993:SRT

[AC93] Roberto M. Amadio and Luca Cardelli. Subtyping recursive types. ACM Transactions on Programming Languages and Systems, 15(4):575– 631, September 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        155231.html.

Ashley:1994:FCP

[AC94] J. Michael Ashley and Charles Consel. Fixpoint computation for polyvariant static analyses of higher-order applicative programs. ACM Transactions on Programming Languages and Systems, 16(5):1431– 1448, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        186037.html.

Abadi:1996:SM

[AC96] Martín Abadi and Luca Cardelli.

On subtyping and matching. ACM Transactions on Programming Languages and Systems, 18(4):401–423, July 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
233563.html.

Attali:1996:NSE

[ACE96] Isabelle Attali, Denis Caromel, and Sidi Ould Ehmety. A natural semantics for Eiffel dynamic binding. ACM Transactions on Programming Languages and Systems, 18(6):711– 729, November 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
235830.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/236118.html.

Alur:2011:SMC

[ACM11] Rajeev Alur, Swarat Chaudhuri, and P. Madhusudan. Software model checking using languages of nested trees. ACM Transactions on Programming Languages and Systems, 33(5):15:1– 15:45, November 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Abadi:1991:DTS

[ACPP91] Martín Abadi, Luca Cardelli, Benjamin Pierce, and Gordon Plotkin. Dynamic typing in a statically typed language. ACM Transactions on Programming Languages and Systems, 13(2): 237–268, April 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
103138.html.

Archer:1984:URR

[ACS84] James E. Archer, Jr., Richard Conway, and Fred B. Schneider. User recovery and reversal in interactive systems. ACM Transactions on Programming Languages and Systems, 6 (1):1–19, January 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Aggarwal:1990:ALP

[ACW90] S. Aggarwal, C. Courcoubetis, and P. Wolper. Adding liveness properties to coupled finitestate machines. ACM Transactions on Programming Languages and Systems, 12(2):303– 339, April 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
78948.html.

Ashley:1998:PFF

[AD98] J. Michael Ashley and R. Kent Dybvig. A practical and flexible flow analysis for higher-order languages. ACM Transactions on Programming Languages and Systems, 20(4):845–868, July 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-4/
p845-ashley/.

Afek:1994:BFF

[ADG^{+}94] Yehuda Afek, Danny Dolev, Eli Gafni, Michael Merritt, and Nir Shavit. A bounded firstin, first-enabled solution to the ℓ-exclusion problem. ACM Transactions on Programming Languages and Systems, 16(3): 939–953, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177731.html.

Alglave:2021:ACF

[ADG^{+}21] Jade Alglave, Will Deacon, Richard Grisenthwaite, Antoine Hacquard, and Luc Maranget. Armed cats: Formal concurrency modelling at Arm. ACM Transactions on Programming Languages and Systems, 43(2): 8:1–8:54, July 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3458926.

Ancona:1991:ECL

[ADGM91] M. Ancona, G. Dodero, V. Gianuzzi, and M. Morgavi. Efficient construction of LR(k) states and tables. ACM Transactions on Programming Languages and Systems, 13(1):150– 178, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        102809.html.

Ager:2006:FPE

[ADR06] Mads Sig Ager, Olivier Danvy, and Henning Korsholm Rohde. Fast partial evaluation of pattern matching in strings. ACM Transactions on Programming Languages and Systems, 28(4): 696–714, July 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Attie:1998:SCS

[AE98] Paul C. Attie and E. Allen Emerson. Synthesis of concurrent systems with many similar processes. ACM Transactions on Programming Languages and Systems, 20(1):51–115, January 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-1/
                                        p51-attie/.

Attie:2001:SCP

[AE01] Paul C. Attie and E. Allen Emerson. Synthesis of concurrent programs for an atomic read/write model of computation. ACM Transactions on Programming Languages and Systems, 23(2): 187–242, March 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Apt:1984:MDT

[AF84] Krzysztof R. Apt and Nissem Francez. Modeling the distributed termination convention of CSP. ACM Transactions on Programming Languages and Systems, 6(3):370– 379, July 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Appel:1994:E

[AF94] Andrew Appel and Charles Fischer. Editorial. ACM Transactions on Programming Languages and Systems, 16(4):1079, July 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Apt:1980:PSC

[AFdR80] Krzysztof R. Apt, Nissim Francez, and Willem P. de Roever. A proof system for communicating sequential processes. ACM Transactions on Programming Languages and Systems, 2 (3):359–385, July 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remarks [Moi83].

Abadi:2006:TSL

[AFF06] Martin Abadi, Cormac Flanagan, and Stephen N. Freund. Types for safe locking: Static race detection for Java. ACM Transactions on Programming Languages and Systems, 28(2): 207–255, March 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Alpuente:1998:PEF

[AFV98] María Alpuente, Moreno Falaschi, and Germán Vidal. Partial evaluation of functional logic programs. ACM Transactions on Programming Languages and Systems, 20(4):768–844, July 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-4/
p768-alpuente/.

Appel:1993:Eb

[AG93] A. W. Appel and S. L. Graham. Editorial. ACM Transactions on Programming Languages and Systems, 15(5):743–744, November 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Alur:2004:MRH

[AG04] Rajeev Alur and Radu Grosu. Modular refinement of hierarchic reactive machines. ACM Transactions on Programming Languages and Systems, 26(2): 339–369, March 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Aho:1989:CGU

[AGT89] Alfred V. Aho, Mahadevan Ganapathi, and Steven W. K. Tjiang. Code generation using tree matching and dynamic programming. ACM Transactions on Programming Languages and Systems, 11(4):491– 516, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

pubs/toc/Abstracts/0164-0925/
75700.html.

Alur:1998:FF

[AH98] Rajeev Alur and Thomas A.

Henzinger. Finitary fairness. ACM Transactions on Programming Languages and Systems, 20 (6):1171–1194, November 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-6/
p1171-alur/.

Apel:2010:CUF

[AH10] Sven Apel and Delesley Hutchins.

A calculus for uniform feature composition. ACM Transactions on Programming Languages and Systems, 32(5):19:1–19:33, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Aung:2014:SS

[AHJR14] Min Aung, Susan Horwitz, Rich Joiner, and Thomas Reps. Specialization slicing. ACM Transactions on Programming Languages and Systems, 36(2):5:1– 5:??, July 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ahmed:2020:ISI

[Ahm20] Amal Ahmed. Introduction to the special issue on ESOP 2018. ACM Transactions on Programming Languages and Systems, 42(1):1, January 2020. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3368252.

Arsac:1982:STR

[AK82] J. Arsac and Y. Kodratoff. Some techniques for recursion removal from recursive functions. ACM Transactions on Programming Languages and Systems, 4 (2):295–322, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Allen:1987:ATF

[AK87] Randy Allen and Ken Kennedy. Automatic translation of Fortran programs to vector form. ACM Transactions on Programming Languages and Systems, 9 (4):491–542, October 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
29875.html.

Ait-Kaci:1989:EIL

[AKBLN89] Hassan A¨ ıt-Kaci, Robert Boyer, Patrick Lincoln, and Roger Nasr. Efficient implementation of lattice operations. ACM Transactions on Programming Languages and Systems, 11(1):115– 146, January 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
59293.html.

Alglave:2017:DSF

[AKNP17] Jade Alglave, Daniel Kroening, Vincent Nimal, and Daniel Poetzl. Don’t sit on the fence: a static analysis approach to automatic fence insertion. ACM Transactions on Programming Languages and Systems, 39(2): 6:1–6:??, May 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ait-Kaci:1994:FPC

[AKP94] Hassan Aït-Kaci and Andreas Podelski. Functions as passive constraints in LIFE. ACM Transactions on Programming Languages and Systems, 16(4): 1279–1318, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
183526.html.

Abadi:1993:CS

[AL93] Martín Abadi and Leslie Lamport. Composing specifications. ACM Transactions on Programming Languages and Systems, 15 (1):73–132, January 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
151649.html.

Abadi:1994:OFR

[AL94] Martín Abadi and Leslie Lamport. An old-fashioned recipe for real time. ACM Transactions on Programming Languages and Systems, 16(5):1543– 1571, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        186058.html.

Abadi:1995:CS

[AL95] Martin Abadi and Leslie Lamport. Conjoining specifications. ACM Transactions on Programming Languages and Systems, 17(3):507–534, May 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        201069.html.

Ancona:2003:JDJ

[ALZ03] Davide Ancona, Giovanni Lagorio, and Elena Zucca. Jam— designing a Java extension with mixins. ACM Transactions on Programming Languages and Systems, 25(5):641– 712, September 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Atkinson:1985:PPD

[AM85] Malcolm P. Atkinson and Ronald Morrison. Procedures as persistent data objects. ACM Transactions on Programming Languages and Systems, 7(4): 539–559, October 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        4477.html.

Appel:2001:IMR

[AM01] Andrew W. Appel and David McAllester. An indexed model of recursive types for foundational proof-carrying code. ACM Transactions on Programming Languages and Systems, 23(5): 657–683, September 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Alglave:2014:HCM

[AMT14] Jade Alglave, Luc Maranget, and Michael Tautschnig. Herding cats: Modelling, simulation, testing, and data mining for weak memory. ACM Transactions on Programming Languages and Systems, 36(2): 7:1–7:??, July 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Andrews:1981:SR

[And81] Gregory R. Andrews. Synchronizing resources. ACM Transactions on Programming Languages and Systems, 3(4):405– 430, October 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anger:1989:LIC

[Ang89] Frank D. Anger. On Lamport’s interprocessor communication model. ACM Transactions on Programming Languages and Systems, 11(3): 404–417, July 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
65982.html.

Anonymous:1982:IA

[Ano82] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 4(3):522–525, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1983:IA

[Ano83] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 5(3):506–509, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1984:IA

[Ano84] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 6(3):444–447, July 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1985:IA

[Ano85] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 7(3):497–500, July 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1986:AI

[Ano86a] Anonymous. 1986 author index. ACM Transactions on Programming Languages and Systems, 8 (4):609–610, October 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1986:IA

[Ano86b] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 8(3):412–415, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1987:IA

[Ano87] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 9(3):467–470, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1988:AI

[Ano88a] Anonymous. 1988 author index.

ACM Transactions on Programming Languages and Systems, 10 (4):641–642, October 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1988:IA

[Ano88b] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 10(3):508–511, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1989:AI

[Ano89a] Anonymous. 1989 author index.

ACM Transactions on Programming Languages and Systems, 11 (4):669–670, October 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1989:IA

[Ano89b] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 11(3):487–490, July 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1990:AI

[Ano90a] Anonymous. 1990 author index. ACM Transactions on Programming Languages and Systems, 12 (4):700–701, October 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1990:IA

[Ano90b] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 12(3):497–500, July 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1991:AI

[Ano91a] Anonymous. 1991 author index. ACM Transactions on Programming Languages and Systems, 13 (4):634–635, October 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1991:IA

[Ano91b] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 13(3):447–450, July 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1992:AI

[Ano92a] Anonymous. 1992 author index.

ACM Transactions on Programming Languages and Systems, 14 (4):617–618, October 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1992:IA

[Ano92b] Anonymous. Information for authors. ACM Transactions on Programming Languages and Systems, 14(3):466–469, July 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1994:AI

[Ano94] Anonymous. 1994 author index.

ACM Transactions on Programming Languages and Systems, 16(6):1904–1907, October 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1995:AI

[Ano95] Anonymous. 1995 author index.

ACM Transactions on Programming Languages and Systems, 17(6):938–940, November 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:1998:AI

[Ano98] Anonymous. 1998 author index. ACM Transactions on Programming Languages and Systems, 20(6):1323–1325, November 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:2002:ADC

[Ano02a] Anonymous. Automatic derivation of compiler machine descriptions. ACM Transactions on Programming Languages and Systems, 24(4):369–408, July 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:2002:LDD

[Ano02b] Anonymous. On loops, dominators, and dominance frontiers. ACM Transactions on Programming Languages and Systems, 24(5):455–490, September 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Anonymous:2018:CCL

[Ano18] Anonymous. Corrigendum to “Cross-Language Interoperability in a Multi-Language Runtime”, by Grimmer et al., ACM Transactions on Programming Languages and Systems (TOPLAS) Volume 40, Issue 2, Article No. 8. ACM Transactions on Programming Languages and Systems, 40(4):18:1– 18:??, December 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3283723.

Arvind:1989:SDS

[ANP89] Arvind, Rishiyur S. Nikhil, and Keshav K. Pingali. Istructures: Data structures for parallel computing. ACM Transactions on Programming Languages and Systems, 11(4):598– 632, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
69562.html.

Anson:1987:GIC

[Ans87] Ed Anson. A generalized iterative construct and its semantics. ACM Transactions on Programming Languages and Systems, 9 (4):567–581, October 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
30391.html.

Andrews:1988:OSL

[AOC^{+}88] Gregory R. Andrews, Ronald A.

Olsson, Michael Coffin, Irving Elshoff, Kelvin Nilsen, Titus Purdin, and Gregg Townsend. An overview of the SR language and implementation. ACM Transactions on Programming Languages and Systems, 10(1): 51–86, January 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
42324.html.

Apt:1994:OCF

[AP94] Krzysztof R. Apt and Alessando Pellegrini. On the occur-check-free Prolog programs. ACM Transactions on Programming Languages and Systems, 16(3): 687–726, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        177673.html.

Abadi:2007:E

[AP07] Martín Abadi and Jens Palsberg. Editorial. ACM Transactions on Programming Languages and Systems, 29(3):14:1, May 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Appel:1993:Ea

[App93] Andrew W. Appel. Editorial. ACM Transactions on Programming Languages and Systems, 15 (2):209–210, April 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Appel:1994:ABG

[App94a] A. W. Appel. Axiomatic bootstrapping: a guide for compiler hackers. ACM Transactions on Programming Languages and Systems, 16(6):1699– 1718, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        197336.html.

Appel:1994:PS

[App94b] Andrew Appel. Purpose and scope. ACM Transactions on Programming Languages and Systems, 16(1):1–2, January 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Appel:2015:VCP

[App15] Andrew W. Appel. Verification of a cryptographic primitive: SHA-256. ACM Transactions on Programming Languages and Systems, 37(2):7:1– 7:??, April 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Apt:1981:TYH

[Apt81] Krzysztof R. Apt. Ten years of Hoare’s logic: a survey — Part I. ACM Transactions on Programming Languages and Systems, 3 (4):431–483, October 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Apt:1986:CPD

[Apt86] Krzysztof R. Apt. Correctness proofs of distributed termination algorithms. ACM Transactions on Programming Languages and Systems, 8(3): 388–405, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
6000.html.

Apt:2000:RCC

[Apt00]  Krzysztof R. Apt.  The

role of commutativity in constraint propagation algorithms. ACM Transactions on Programming Languages and Systems, 22(6):1002–1036, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/

                                        journals/toplas/2000-22-6/
                                        p1002-apt/p1002-apt.pdf;
                                        http://www.acm.org/pubs/citations/
                                        journals/toplas/2000-22-6/
                                        p1002-apt/.

Andrews:1980:AAI

[AR80] Gregory R. Andrews and Richard P. Reitman. An axiomatic approach to information flow in programs. ACM Transactions on Programming Languages and Systems, 2(1):56–76, January 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Appelbe:1984:ECS

[AR84] William F. Appelbe and A. P. Ravn. Encapsulation constructs in systems programming languages. ACM Transactions on Programming Languages and Systems, 6(2):129–158, April 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Arnold:1980:URG

[AS80] D. B. Arnold and M. R. Sleep. Uniform random generation of balanced parenthesis strings. ACM Transactions on Programming Languages and Systems, 2 (1):122–128, January 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Alpern:1989:VTP

[AS89] Bowen Alpern and Fred B. Schneider. Verifying temporal properties without temporal logic. ACM Transactions on Programming Languages and Systems, 11(1):147– 167, January 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
62028.html.

Andersen:2019:FSP

[ASAVF19] Leif Andersen, Vincent St-Amour, Jan Vitek, and Matthias Felleisen. Feature-specific profiling. ACM Transactions on Programming Languages and Systems, 41(1):4:1–4:??, March 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3275519.

Austin:2017:MFD

[ASF17] Thomas H. Austin, Tommy Schmitz, and Cormac Flanagan. Multiple facets for dynamic information flow with exceptions. ACM Transactions on Programming Languages and Systems, 39 (3):10:1–10:??, July 2017. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Assmann:2000:GRS

[Ass00] Uwe Assmann. Graph rewrite systems for program optimization. ACM Transactions on Programming Languages and Systems, 22(4):583–637, 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-4/p583-assmann/.

Arenaz:2008:XEF

[ATD08] Manuel Arenaz, Juan Touriño, and Ramon Doallo. XARK: an extensible framework for automatic recognition of computational kernels. ACM Transactions on Programming Languages and Systems, 30(6):32:1– 32:56, October 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ashcroft:1982:RS

[AW82] E. A. Ashcroft and W. W. Wadge. R_{/} for semantics. ACM Transactions on Programming Languages and Systems, 4(2): 283–294, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Avrunin:1985:DAD

[AW85] George S. Avrunin and Jack C. Wileden. Describing and analyzing distributed software system designs. ACM Transactions on Programming Languages and Systems, 7(3):380– 403, July 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        3989.html.

Aiken:1995:SST

[AWW95] Alexander Aiken, John H. Williams, and Edward L. Wimmers. Safe: a semantic technique for transforming programs in the presence of errors. ACM Transactions on Programming Languages and Systems, 17(1):

63–84, January 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201002.html.

Alur:2001:MCH

[AY01] Rajeev Alur and Mihalis Yannakakis. Model checking of hierarchical state machines. ACM Transactions on Programming Languages and Systems, 23(3): 273–303, May 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ben-Ari:1984:AFG

[BA84] Mordechai Ben-Ari. Algorithms for on-the-fly garbage collection. ACM Transactions on Programming Languages and Systems, 6 (3):333–344, July 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Blume:1999:HM

[BA99] Matthias Blume and Andrew W.

Appel. Hierarchical modularity. ACM Transactions on Programming Languages and Systems, 21(4):813–847, July 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-4/
p813-blume/p813-blume.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-4/p813-blume/.

Ben-Amram:2008:SCT

[BA08] Amir M. Ben-Amram. Sizechange termination with difference constraints. ACM Transactions on Programming Languages and Systems, 30(3):16:1– 16:31, May 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Backhouse:1984:GDF

[Bac84] Roland Backhouse. Global data flow analysis problems arising in locally least-cost error recovery. ACM Transactions on Programming Languages and Systems, 6 (2):192–214, April 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bondhugula:2016:PAP

[BAC16] Uday Bondhugula, Aravind Acharya, and Albert Cohen. The Pluto+ algorithm: a practical approach for parallelization and locality optimization of affine loop nests. ACM Transactions on Programming Languages and Systems, 38(3):12:1–12:??, May 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bagrodia:1989:SAP

[Bag89]  Rajive Bagrodia.  Synchro-

nization of asynchronous processes in CSP. ACM Transactions on Programming Languages and Systems, 11(4):585– 597, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
69561.html.

Ben-Amram:2012:TIL

[BAGM12] Amir M. Ben-Amram, Samir Genaim, and Abu Naser Masud. On the termination of integer loops. ACM Transactions on Programming Languages and Systems, 34(4):16:1– 16:24, December 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Baker:1982:OPA

[Bak82] T. P. Baker. A one-pass algorithm for overload resolution in Ada. ACM Transactions on Programming Languages and Systems, 4(4):601–614, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ball:1994:ECP

[Bal94] Thomas Ball. Efficiently counting program events with support for on-line queries. ACM Transactions on Programming Languages and Systems, 16(5):1399– 1410, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
186027.html.

Ben-Amram:2007:PTA

[BAL07] Amir M. Ben-Amram and Chin Soon Lee. Program termination analysis in polynomial time. ACM Transactions on Programming Languages and Systems, 29(1):5:1–5:37, January 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Brecht:2006:CGC

[BALP06] Tim Brecht, Eshrat Arjomandi, Chang Li, and Hang Pham. Controlling garbage collection and heap growth to reduce the execution time of Java applications. ACM Transactions on Programming Languages and Systems, 28(5):908–941, September 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Banerjee:1987:MSR

[Ban87] Debasish Banerjee. A methodology for synthesis of recursive functional programs. ACM Transactions on Programming Languages and Systems, 9(3): 441–462, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
24071.html.

Banerjee:2011:MFT

[Ban11] Utpal Banerjee. Mathematical foundation of trace scheduling. ACM Transactions on Programming Languages and Systems, 33 (3):10:1–10:24, April 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barnden:1981:NCA

[Bar81] J. A. Barnden. Nonsequentiality and concrete activity phases in discrete-event simulation languages. ACM Transactions on Programming Languages and Systems, 3(3):293–317, July 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barstow:1985:CTD

[Bar85] David Barstow. On convergence toward a database of program transformations. ACM Transactions on Programming Languages and Systems, 7(1): 1–9, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
2364.html.

Beyer:1979:SED

[BB79] Eric Beyer and Peter Buneman.

A space efficient dynamic allocation algorithm for queuing messages. ACM Transactions on Programming Languages and Systems, 1(2):287–294, October 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Breuer:1994:DET

[BB94] Peter T. Breuer and Jonathan P.

Bowen. Decompilation: The enumeration of types and grammers. ACM Transactions on Programming Languages and Systems, 16(5):1613–1647, September 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

                                        pubs/toc/Abstracts/0164-0925/
                                        186093.html.

Bhaskaracharya:2016:ASO

[BBC16] Somashekaracharya G. Bhaskaracharya, Uday Bondhugula, and Albert Cohen. Automatic storage optimization for arrays. ACM Transactions on Programming Languages and Systems, 38(3): 11:1–11:??, May 2016. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bengtson:2011:RTS

[BBF^{+}11] Jesper Bengtson, Karthikeyan Bhargavan, Cédric Fournet, Andrew D. Gordon, and Sergio Maffeis. Refinement types for secure implementations. ACM Transactions on Programming Languages and Systems, 33(2): 8:1–8:45, January 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Biering:2007:BHH

[BBTS07] Bodil Biering, Lars Birkedal, and Noah Torp-Smith. BIhyperdoctrines, higher-order separation logic, and abstraction. ACM Transactions on Programming Languages and Systems, 29(5):24:1–24:35, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barabash:2005:PIM

[BBYG^{+}05] Katherine Barabash, Ori Ben-Yitzhak, Irit Goft, Elliot K. Kolodner, Victor Leikehman, Yoav Ossia, Avi Owshanko, and Erez Petrank. A parallel, incremental, mostly concurrent garbage collector for servers. ACM Transactions on Programming Languages and Systems, 27 (6):1097–1146, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bobrow:1979:CEL

[BC79] Daniel G. Bobrow and Douglas W. Clark. Compact encodings of list structure. ACM Transactions on Programming Languages and Systems, 1(2): 266–286, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bates:1985:PP

[BC85a] Joseph L. Bates and Robert L.

Constable. Proofs as programs. ACM Transactions on Programming Languages and Systems, 7 (1):113–136, January 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
2528.html.

Bergeretti:1985:IFD

[BC85b] Jean-François Bergeretti and Bernard A. Carré. Informationflow and data-flow analysis of while-programs. ACM Transactions on Programming Languages and Systems, 7(1):37– 61, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

                                        pubs/toc/Abstracts/0164-0925/
                                        2366.html.

Brogi:1991:CLS

[BC91] Antonio Brogi and Paolo Ciancarini. The concurrent language, Shared Prolog. ACM Transactions on Programming Languages and Systems, 13(1): 99–123, January 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        102807.html.

Bugliesi:2004:ACM

[BCC04] Michele Bugliesi, Giuseppe Castagna, and Silvia Crafa. Access control for mobile agents: The calculus of boxed ambients. ACM Transactions on Programming Languages and Systems, 26 (1):57–124, January 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bossi:1990:MSL

[BCD90] A. Bossi, N. Cocco, and S. Dulli. A method for specializing logic programs. ACM Transactions on Programming Languages and Systems, 12(2):253– 302, April 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        78947.html.

Betts:2015:DIV

[BCD^{+}15] Adam Betts, Nathan Chong, Alastair F. Donaldson, Jeroen Ketema, Shaz Qadeer, Paul Thomson, and John Wickerson. The design and implementation of a verification technique for GPU kernels. ACM Transactions on Programming Languages and Systems, 37(3): 10:1–10:??, June 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bugliesi:2015:ART

[BCEM15] Michele Bugliesi, Stefano Calzavara, Fabienne Eigner, and Matteo Maffei. Affine refinement types for secure distributed programming. ACM Transactions on Programming Languages and Systems, 37(4):11:1–11:??, August 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Benton:2004:MCA

[BCF04] Nick Benton, Luca Cardelli, and Cédric Fournet. Modern concurrency abstractions for C#. ACM Transactions on Programming Languages and Systems, 26(5):769–804, September 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bruynooghe:2007:TAL

[BCG^{+}07] Maurice Bruynooghe, Michael Codish, John P. Gallagher, Samir Genaim, and Wim Vanhoof. Termination analysis of logic programs through combination of type-based norms. ACM Transactions on Programming Languages and Systems, 29 (2):10:1–10:44, April 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bottoni:1999:SDC

[BCM99] P. Bottoni, M. F. Costabile, and P. Mussio. Specification and dialogue control of visual interaction through visual rewriting systems. ACM Transactions on Programming Languages and Systems, 21 (6):1077–1136, November 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/1999-21-6/
                                        p1077-bottoni/p1077-bottoni.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-6/p1077-bottoni/.

Bhatia:2008:RSE

[BCP08] Sapan Bhatia, Charles Consel, and Calton Pu. Remote specialization for efficient embedded operating systems. ACM Transactions on Programming Languages and Systems, 30(4): 22:1–22:32, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Briggs:1994:IGC

[BCT94] Preston Briggs, Keith D. Cooper, and Linda Torczon. Improvements to graph coloring register allocation. ACM Transactions on Programming Languages and Systems, 16(3): 428–455, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177575.html.

Bergstra:1997:TCT

[BDFH97] J. A. Bergstra, T. B. Dinesh, J. Field, and J. Heering. Toward a complete transformational toolkit for compilers. ACM Transactions on Programming Languages and Systems, 19(5):639–684, September 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-5/
p639-bergstra/.

Bartoletti:2009:LPR

[BDFZ09] Massimo Bartoletti, Pierpaolo Degano, Gian-Luigi Ferrari, and Roberto Zunino. Local policies for resource usage analysis. ACM Transactions on Programming Languages and Systems, 31(6):23:1–23:43, August 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Blackburn:2016:TWT

[BDH^{+}16] Stephen M. Blackburn, Amer Diwan, Matthias Hauswirth, Peter F. Sweeney, José Nelson Amaral, Tim Brecht, Lubomír Bulej, Cliff Click, Lieven Eeckhout, Sebastian Fischmeister, Daniel Frampton, Laurie J. Hendren, Michael Hind, Antony L. Hosking, Richard E. Jones, Tomas Kalibera, Nathan Keynes, Nathaniel Nystrom, and Andreas Zeller. The truth, the whole truth, and nothing but the truth: a pragmatic guide to assessing empirical evaluations. ACM Transactions on Programming Languages and Systems, 38(4):15:1–15:??, October 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Botincan:2013:PDP

[BDJ13] Matko Botincan, Mike Dodds, and Suresh Jagannathan. Proofdirected parallelization synthesis by separation logic. ACM Transactions on Programming Languages and Systems, 35(2): 8:1–8:??, July 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bernardeschi:2008:DBV

[BDL^{+}08] C. Bernardeschi, N. De Francesco, G. Lettieri, L. Martini, and P. Masci. Decomposing bytecode verification by abstract interpretation. ACM Transactions on Programming Languages and Systems, 31(1):3:1–3:63, December 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bueno:1999:EAI

[BdlBH99] Francisco Bueno, María García de la Banda, and Manuel Hermenegildo. Effectiveness of abstract interpretation in automatic parallelization: a case study in logic programming. ACM Transactions on Programming Languages and Systems, 21 (2):189–239, March 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-2/p189-bueno/.

Biernacki:2015:DCP

[BDM15] Dariusz Biernacki, Olivier Danvy, and Kevin Millikin. A dynamic continuation-passing style for dynamic delimited continuations. ACM Transactions on Programming Languages and Systems, 38(1):2:1–2:??, October 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bowman:1993:RAN

[BDP93] Mic Bowman, Saumya K. Debray, and Larry L. Peterson. Reasoning about naming systems. ACM Transactions on Programming Languages and Systems, 15(5):795– 825, November 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
161471.html.

Barthe:2014:FVS

[BDP14] Gilles Barthe, Delphine Demange, and David Pichardie. Formal verification of an SSA-Based middle-end for CompCert. ACM Transactions on Programming Languages and Systems, 36(1):4:1–4:??, March 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bossi:1994:TAP

[BE94] Annalisa Bossi and Sandro Etalle. Transforming acyclic programs. ACM Transactions on Programming Languages and Systems, 16(4):1081– 1096, July 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        183434.html.

Bouajjani:2013:ARP

[BE13] Ahmed Bouajjani and Michael Emmi. Analysis of recursively parallel programs. ACM Transactions on Programming Languages and Systems, 35(3):10:1– 10:49, November 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Beemster:1994:SOG

[Bee94] Marcel Beemster. Strictness optimization for graph reduction machines (Why id might not be strict). ACM Transactions on Programming Languages and Systems, 16(5):1449– 1466, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        186040.html.

Brockschmidt:2016:ARS

[BEF^{+}16] Marc Brockschmidt, Fabian Emmes, Stephan Falke, Carsten Fuhs, and Jürgen Giesl. Analyzing runtime and size complexity of integer programs. ACM Transactions on Programming Languages and Systems, 38(4): 13:1–13:??, October 2016. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bernstein:1980:OGN

[Ber80]  Arthur Bernstein.  Output

guards and nondeterminism in “Communicating Sequential Processes”. ACM Transactions on Programming Languages and Systems, 2(2):234–238, April 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). Bernstein presents a distributed algorithm for CSP output guards based on priority ordering of processes.

Berzins:1994:SMS

[Ber94] V. Berzins. Software merge: Semantics of combining changes to programs. ACM Transactions on Programming Languages and Systems, 16(6):1875– 1903, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
197403.html.

Burke:1987:PML

[BF87] Michael G. Burke and Gerald A. Fisher. A practical method for LR and LL syntactic error diagnosis and recovery. ACM Transactions on Programming Languages and Systems, 9 (2):164–197, April 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        22720.html.

Bhargavan:2008:VPB

[BFG08] Karthikeyan Bhargavan, Cédric Fournet, and Andrew D. Gordon. Verifying policy-based web services security. ACM Transactions on Programming Languages and Systems, 30(6):30:1– 30:59, October 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bhargavan:2008:VII

[BFGT08] Karthikeyan Bhargavan, Cédric Fournet, Andrew D. Gordon, and Stephen Tse. Verified interoperable implementations of security protocols. ACM Transactions on Programming Languages and Systems, 31(1):5:1– 5:57, December 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barbosa:1989:CHL

[BG89a] Valmir Barbosa and Eli Gafni. Concurrency in heavily loaded neighborhood-constrained systems. ACM Transactions on Programming Languages and Systems, 11(4):562–584, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

                                        acm.org/pubs/toc/Abstracts/
                                        0164-0925/69560.html.

Bernstein:1989:SEP

[BG89b] David Bernstein and Izidor Gertner. Scheduling expressions on a pipelined processor with a maximal delay of one cycle. ACM Transactions on Programming Languages and Systems, 11 (1):57–66, January 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
59291.html.

Baillot:2022:TCP

[BG22] Patrick Baillot and Alexis Ghyselen. Types for complexity of parallel computation in pi-calculus. ACM Transactions on Programming Languages and Systems, 44(3):15:1– 15:50, September 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3495529.

Binkley:2013:EIL

[BGH^{+}13] David Binkley, Nicolas Gold, Mark Harman, Syed Islam, Jens Krinke, and Zheng Li. Efficient identification of linchpin vertices in dependence clusters. ACM Transactions on Programming Languages and Systems, 35 (2):7:1–7:??, July 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barthe:2009:CTO

[BGKR09] Gilles Barthe, Benjamin Grégoire, César Kunz, and Tamara Rezk. Certificate translation for optimizing compilers. ACM Transactions on Programming Languages and Systems, 31(5):18:1– 18:45, June 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Barbuti:1993:GFS

[BGL93] Roberto Barbuti, Roberto Giacobazzi, and Giorgio Levi. A general framework for semanticsbased bottom-up abstract interpretation of logic programs. ACM Transactions on Programming Languages and Systems, 15 (1):133–181, January 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        151650.html.

Boruch-Gruszecki:2023:CT

[BGOL^{+}23] Aleksander Boruch-Gruszecki, Martin Odersky, Edward Lee, Ondrej Lhoták, and Jonathan Brachthäuser. Capturing types. ACM Transactions on Programming Languages and Systems, 45 (4):21:1–21:??, December 2023. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3618003.

Bultan:1999:MCC

[BGP99] Tevfik Bultan, Richard Gerber, and William Pugh. Modelchecking concurrent systems with unbounded integer variables: symbolic representations, approximations, and experimental results. ACM Transactions on Programming Languages and Systems, 21(4):747–789, July 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-4/
p747-bultan/p747-bultan.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-4/p747-bultan/.

Butler:1999:RAG

[BH99] Michael Butler and Pieter Hartel. Reasoning about Grover’s quantum search algorithm using probabilistic wp. ACM Transactions on Programming Languages and Systems, 21(3):417–429, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-3/
p417-butler/p417-butler.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-3/p417-butler/.

Back:2005:KJR

[BH05a] Godmar Back and Wilson C.

Hsieh. The KaffeOS Java runtime system. ACM Transactions on Programming Languages and Systems, 27(4):583– 630, July 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Buhr:2005:ISM

[BH05b] Peter A. Buhr and Ashif S.

Harji. Implicit-signal monitors. ACM Transactions on Programming Languages and Systems, 27 (6):1270–1343, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Binkley:2007:ESO

[BHK07] David Binkley, Mark Harman, and Jens Krinke. Empirical study of optimization techniques for massive slicing. ACM Transactions on Programming Languages and Systems, 30(1):3:1– 3:33, November 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Blackburn:2007:PBP

[BHM^{+}07] Stephen M. Blackburn, Matthew Hertz, Kathryn S. Mckinley, J. Eliot B. Moss, and Ting Yang. Profile-based pretenuring. ACM Transactions on Programming Languages and Systems, 29 (1):2:1–2:57, January 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Berger:2019:IPL

[BHM^{+}19] Emery D. Berger, Celeste Hollenbeck, Petr Maj, Olga Vitek, and Jan Vitek. On the impact of programming languages on code quality: a reproduction study. ACM Transactions on Programming Languages and Systems, 41(4):21:1– 21:??, November 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3340571.

Bird:1984:PAS

[Bir84] R. S. Bird. The promotion and accumulation strategies in transformational programming. ACM Transactions on Programming Languages and Systems, 6 (4):487–504, October 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See addendum [Bir85].

Bird:1985:APA

[Bir85] R. S. Bird. Addendum: “The Promotion and Accumulation Strategies in Transformational Programming”. ACM Transactions on Programming Languages and Systems, 7(3):490– 492, July 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Bir84].

Barthe:2011:AMC

[BK11] Gilles Barthe and César Kunz.

An abstract model of certificate translation. ACM Transactions on Programming Languages and Systems, 33(4):13:1– 13:46, July 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Broy:1980:DIA

[BKB80] Manfred Broy and Bernd Krieg-Brückner. Derivation of invariant assertions during program development by transformation. ACM Transactions on Programming Languages and Systems, 2 (3):321–337, July 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Breuer:1997:RCS

[BKL^{+}97] Peter T. Breuer, Carlos Delgado Kloos, Andrés Marín López, Natividad Martínez Madrid, and Luis Sánchez Fernández. A refinement calculus for the synthesis of verified hardware descriptions in VHDL. ACM Transactions on Programming Languages and Systems, 19 (4):586–616, July 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-4/
p586-breuer/.

Barthe:2013:PRR

[BKOZB13] Gilles Barthe, Boris Köpf, Federico Olmedo, and Santiago Zanella-Béguelin. Probabilistic relational reasoning for differential privacy. ACM Transactions on Programming Languages and Systems, 35(3):9:1–9:49, November 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Buchsbaum:1998:NSL

[BKRW98] Adam L. Buchsbaum, Haim Kaplan, Anne Rogers, and Jeffery R. Westbrook. A new, simpler linear-time dominators algorithm. ACM Transactions on Programming Languages and Systems, 20(6):1265– 1296, November 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/

journals/toplas/1998-20-6/
p1265-buchsbaum/. See corri-

gendum [BKRW05].

Buchsbaum:2005:CNS

[BKRW05] Adam L. Buchsbaum, Haim Kaplan, Anne Rogers, and Jeffery R. Westbrook. Corrigendum: “A new, simpler lineartime dominators algorithm”. ACM Transactions on Programming Languages and Systems, 27(3):383–387, May 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [BKRW98].

Back:1988:DCA

[BKS88] R. J. R. Back and R. Kurki-Suonio. Distributed cooperation with action systems. ACM Transactions on Programming Languages and Systems, 10(4): 513–554, October 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
48023.html.

Bic:1987:DDM

[BL87] Lubomir Bic and Craig Lee. A data-driven model for a subset of logic programming. ACM Transactions on Programming Languages and Systems, 9(4): 618–645, October 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
31333.html.

Ball:1994:OPT

[BL94a] Thomas Ball and James R. Larus. Optimally profiling and tracing programs. ACM Transactions on Programming Languages and Systems, 16(4): 1319–1360, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        183527.html.

Bates:1994:RSL

[BL94b] Joseph Bates and Alon Lavie. Recognizing substrings of LR(k) languages in linear time. ACM Transactions on Programming Languages and Systems, 16(3): 1051–1077, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        177768.html.

Blanchet:2003:EAJ

[Bla03] Bruno Blanchet. Escape analysis for Java^{T M}: Theory and practice. ACM Transactions on Programming Languages and Systems, 25(6):713–775, November 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bodden:2012:PEF

[BLH12] Eric Bodden, Patrick Lam, and Laurie Hendren. Partially evaluating finite-state runtime monitors ahead of time. ACM Transactions on Programming Languages and Systems, 34(2):

7:1–7:52, June 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bloss:1994:PAO

[Blo94] Adrienne Bloss. Path analysis and the optimization of nonstrict functional languages. ACM Transactions on Programming Languages and Systems, 16(3):328–369, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177497.html.

Boudol:2012:RAW

[BLRS12] Gérard Boudol, Zhengqin Luo, Tamara Rezk, and Manuel Serrano. Reasoning about Web applications: An operational semantics for HOP. ACM Transactions on Programming Languages and Systems, 34(2):10:1– 10:40, June 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Blume:1999:DAS

[Blu99]  Matthias Blume.  Depen-

dency analysis for Standard ML. ACM Transactions on Programming Languages and Systems, 21(4):790–812, July 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-4/
p790-blume/p790-blume.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-4/p790-blume/.

Brandis:1994:SPG

[BM94] M. M. Brandis and H. Moessenboeck. Single-pass generation of static single-assignment form for structured languages. ACM Transactions on Programming Languages and Systems, 16(6): 1684–1698, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        197331.html.

Brogi:1994:MLP

[BMPT94] Antonio Brogi, Paolo Mancarella, Dino Pedreschi, and Franco Turini. Modular logic programming. ACM Transactions on Programming Languages and Systems, 16(4):1361– 1398, July 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        183528.html.

Bistarelli:2001:SBC

[BMR01] Stefano Bistarelli, Ugo Montanari, and Francesca Rossi. Semiring-based constraint logic programming: syntax and semantics. ACM Transactions on Programming Languages and Systems, 23(1):1–29, January 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ball:2005:PPA

[BMR05] Thomas Ball, Todd Millstein, and Sriram K. Rajamani. Polymorphic predicate abstraction. ACM Transactions on Programming Languages and Systems, 27 (2):314–343, March 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Borstler:1991:TCT

[BMW91] Jürgen Börstler, Ulrich Möncke, and Reinhard Wilhelm. Table compression for tree automata. ACM Transactions on Programming Languages and Systems, 13(3):295–314, July 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
117013.html.

Broy:1994:AFC

[BN94] Manfred Broy and Greg Nelson. Adding fair choice to Dijkstra’s calculus. ACM Transactions on Programming Languages and Systems, 16(3):924– 938, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177727.html.

Bertsch:1999:FPT

[BN99] Eberhard Bertsch and Mark-Jan Nederhof. On failure of the pruning technique in “Error Repair in Shift-Reduce Parsers”. ACM Transactions on Programming Languages and Systems, 21 (1):1–10, January 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-1/p1-bertsch/. See

[MYD95].

Banerjee:2018:LAF

[BNN18] Anindya Banerjee, David A. Naumann, and Mohammad Nikouei. A logical analysis of framing for specifications with pure method calls. ACM Transactions on Programming Languages and Systems, 40(2):6:1– 6:??, June 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Banerjee:2022:RPL

[BNNN22] Anindya Banerjee, Ramana Nagasamudram, David Naumann, and Mohammad Nikouei. A relational program logic with data abstraction and dynamic framing. ACM Transactions on Programming Languages and Systems, 44(4):25:1–25:??, December 2022. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https:

                                        //dl.acm.org/doi/10.1145/
                                        3551497.

Busi:2021:SIE

[BNV^{+}21] Matteo Busi, Job Noorman, Jo Van Bulck, Letterio Galletta, Pierpaolo Degano, Jan Tobias Mühlberg, and Frank Piessens. Securing interruptible enclaved execution on small microprocessors. ACM Transactions on Programming Languages and Systems, 43(3):12:1– 12:77, September 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3470534.

Bohm:1994:TIP

[BO94] A. P. W. Böhm and R. R. Oldehoeft. Two issues in parallel language design. ACM Transactions on Programming Languages and Systems, 16(6):1675– 1683, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
197325.html.

Bobrow:1980:MRS

[Bob80] Daniel G. Bobrow. Managing reentrant structures using reference counts. ACM Transactions on Programming Languages and Systems, 2(3):269– 273, July 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Boehm:1985:SEA

[Boe85] Hans-Juergen Boehm. Side effects and aliasing can have simple axiomatic descriptions. ACM Transactions on Programming Languages and Systems, 7(4): 637–655, October 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4474.html.

Boom:1982:WPL

[Boo82] H. J. Boom. A weaker precondition for loops. ACM Transactions on Programming Languages and Systems, 4(4):668– 677, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Borning:1981:PLA

[Bor81] Alan Borning. The programming language aspects of ThingLab, a constraint-oriented simulation laboratory. ACM Transactions on Programming Languages and Systems, 3(4):353–387, October 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Boute:1988:SSP

[Bou88] Raymond T. Boute. System semantics: Principles, applications, and implementation. ACM Transactions on Programming Languages and Systems, 10 (1):118–155, January 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
45067.html.

Boute:1992:EDF

[Bou92] Raymond T. Boute. The Euclidean definition of the functions div and mod. ACM Transactions on Programming Languages and Systems, 14(2): 127–144, April 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
128862.html.

Boute:2005:FDL

[Bou05] Raymond Boute. Functional declarative language design and predicate calculus: a practical approach. ACM Transactions on Programming Languages and Systems, 27(5):988– 1047, September 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Boute:2006:CSD

[Bou06]  Raymond T. Boute.  Cal-

culational semantics: Deriving programming theories from equations by functional predicate calculus. ACM Transactions on Programming Languages and Systems, 28(4):747– 793, July 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Bar-On:1985:OPG

[BOV85] Ilan Bar-On and Uzi Vishkin.

Optimal parallel generation of a computation tree form. ACM Transactions on Programming Languages and Systems, 7(2): 348–357, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3478.html.

Boyland:1996:CAG

[Boy96] John Tang Boyland. Conditional attribute grammars. ACM Transactions on Programming Languages and Systems, 18(1): 73–108, January 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        225544.html.

Boyland:2010:SFP

[Boy10] John Tang Boyland. Semantics of fractional permissions with nesting. ACM Transactions on Programming Languages and Systems, 32(6):22:1–22:33, August 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Broy:1982:CAA

[BP82] Manfred Broy and Peter Pepper. Combining algebraic and algorithmic reasoning: An approach to the Schorr–Waite algorithm. ACM Transactions on Programming Languages and Systems, 4 (3):362–381, July 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Burns:1989:USS

[BP89] James E. Burns and Jan Pachi. Uniform self-stabilizing rings. ACM Transactions on Programming Languages and Systems, 11 (2):330–344, April 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        63403.html.

Bendersky:2012:SOB

[BP12] Anna Bendersky and Erez Petrank. Space overhead bounds for dynamic memory management with partial compaction.

ACM Transactions on Programming Languages and Systems, 34(3):13:1–13:43, October 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Balabonski:2016:DFM

[BPP16] Thibaut Balabonski, François Pottier, and Jonathan Protzenko. The design and formalization of Mezzo, a permission-based programming language. ACM Transactions on Programming Languages and Systems, 38(4): 14:1–14:??, October 2016. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Basso:2023:OAC

[BPRB23] Matteo Basso, Aleksandar Prokopec, Andrea Rosà, and Walter Binder. Optimization-aware compiler-level event profiling. ACM Transactions on Programming Languages and Systems, 45 (2):10:1–10:??, June 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591473.

Baumgartner:1997:ISC

[BR97] Gerald Baumgartner and Vincent F. Russo. Implementing signatures for C++. ACM Transactions on Programming Languages and Systems, 19(1): 153–187, January 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/

1997-19-1/p153-baumgartner/

.

Balakrishnan:2010:WWY

[BR10] Gogul Balakrishnan and Thomas Reps. WYSINWYX: What you see is not what you eXecute. ACM Transactions on Programming Languages and Systems, 32(6):23:1–23:84, August 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Brent:1989:EIF

[Bre89] R. P. Brent. Efficient implementation of the first-fit strategy for dynamic storage allocation. ACM Transactions on Programming Languages and Systems, 11(3):388–403, July 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
65981.html.

Buckley:1983:EIG

[BS83] G. N. Buckley and Abraham Silberschatz. An effective implementation for the generalized input-output construct of CSP. ACM Transactions on Programming Languages and Systems, 5 (2):223–235, April 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). They present a distributed algorithm for CSP output guards based on priority ordering of processes. Their algorithm has the property that two processes that can communicate and do not establish communication with a third process will communicate within a bounded time.

Bahlke:1986:PSF

[BS86] Rolf Bahlke and Gregor Snelting. The PSG system: From formal language definitions to interactive programming environments. ACM Transactions on Programming Languages and Systems, 8(4):547– 576, October 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
20890.html.

Bermudez:1988:NRB

[BS88] Manuel E. Bermudez and Karl M. Schimpf. On the (non-) relationship between SLR(1) and NQLALR(1) grammars (technical correspondence). ACM Transactions on Programming Languages and Systems, 10(2): 338–342, April 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
42276.html.

Bruce:2003:PTS

[BSvGF03] Kim B. Bruce, Angela Schuett, Robert van Gent, and Adrian Fiech. PolyTOIL: a type-safe polymorphic object-oriented language. ACM Transactions on Programming Languages and Systems, 25(2):225–290, March 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Burke:1993:IOE

[BT93] Michael Burke and Linda Torczon. Interprocedural optimization: Eliminating unnecessary recompilation. ACM Transactions on Programming Languages and Systems, 15(3): 367–399, July 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        169678.html.

Budd:1984:ACV

[Bud84] Timothy A. Budd. An APL compiler for a vector processor. ACM Transactions on Programming Languages and Systems, 6 (3):297–313, July 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Burton:1984:ACP

[Bur84] F. Warren Burton. Annotations to Control Parallelism and Reduction Order in the Distributed Evaluation of Functional Programs. ACM Transactions on Programming Languages and Systems, 6(2):159– 174, April 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Burke:1990:IBA

[Bur90a] Michael Burke. An intervalbased approach to exhaustive and incremental interprocedural data-flow analysis. ACM Transactions on Programming Languages and Systems, 12(3): 341–395, July 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
78963.html.

Burton:1990:TCT

[Bur90b] F. Warren Burton. Technical correspondence: Type extension through polymorphism. ACM Transactions on Programming Languages and Systems, 12(1): 135–138, January 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214515.html;   http://www.
acm.org/pubs/toc/Abstracts/
0164-0925/77560.html. See

[Vol91, Bur91].

Burton:1991:TCA

[Bur91] F. Warren Burton. Technical correspondence: Author’s reply to “Subtypes and Quantification”, by D. M. Volpano. ACM Transactions on Programming Languages and Systems, 13 (4):633, October 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Bur90b, Vol91].

Broy:1987:ADP

[BWP87] Manfred Broy, Martin Wirsing, and Peter Pepper. On the algebraic definition of programming languages. ACM Transactions on Programming Languages and Systems, 9(1):54– 99, January 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        10501.html.

Cameron:1989:EHL

[Cam89]  Robert D. Cameron.  Effi-

cient high-level iteration with accumulators. ACM Transactions on Programming Languages and Systems, 11(2):194– 211, April 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        63401.html.

Carlisle:1995:TCC

[Car95] W. H. Carlisle. Type checking concurrent I/O. ACM Transactions on Programming Languages and Systems, 17(3): 448–460, May 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        203097.html.

Castagna:1995:CCC

[Cas95] Giuseppe Castagna. Covariance and contravariance: Conflict without a cause. ACM Transactions on Programming Languages and Systems, 17(3): 431–447, May 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        203096.html.

Cattell:1980:ADC

[Cat80]  R. G. G. Cattell.  Auto-

matic derivation of code generators from machine descriptions. ACM Transactions on Programming Languages and Systems, 2 (2):173–190, April 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Casanova:1980:FSR

[CB80] Marco R. Casanova and Phillip A.

Bernstein. A formal system for reasoning about programs accessing a relational database. ACM Transactions on Programming Languages and Systems, 2 (3):386–414, July 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Charron-Bost:1995:LTP

[CBDGF95] Bernadette Charron-Bost, Carole Delporte-Gallet, and Hugues Fauconnier. Local and temporal predicates in distributed systems. ACM Transactions on Programming Languages and Systems, 17(1):157– 179, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201005.html.

Cotton-Barratt:2019:MVP

[CBMO19] Conrad Cotton-Barratt, Andrzej S. Murawski, and C.- H. Luke Ong. ML, visibly pushdown class memory automata, and extended branching vector addition systems with states.

ACM Transactions on Programming Languages and Systems, 41 (2):11:1–11:??, June 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3310338.

Click:1995:CAC

[CC95] Cliff Click and Keith D. Cooper. Combining analyses, combining optimizations. ACM Transactions on Programming Languages and Systems, 17(2):181– 196, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201061.html.

Clarke:1997:URE

[CC97] Charles L. A. Clarke and Gordon V. Cormack. On the use of regular expressions for searching text. ACM Transactions on Programming Languages and Systems, 19(3):413– 426, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-3/p413-clarke/.

Chargueraud:2023:OSH

[CCEG23] Arthur Charguéraud, Adam Chlipala, Andres Erbsen, and Samuel Gruetter. Omnisemantics: Smooth handling of nondeterminism. ACM Transactions on Programming Languages and Systems, 45(1):

5:1–5:??, March 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579834.

Constable:1979:HAF

[CD79] Robert L. Constable and James E.

Donahue. A hierarchical approach to formal semantics with application to the definition of PL/CS. ACM Transactions on Programming Languages and Systems, 1(1):98–114, July 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Carchiolo:1989:ELT

[CDFP89] Vincenza Carchiolo, Antonella Di Stefano, Alberto Faro, and Giuseppe Pappalardo. ECCS and LIPS: Two languages for OSI systems specification and verification. ACM Transactions on Programming Languages and Systems, 11(2):284– 329, April 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
63402.html.

Chen:2018:BPP

[CDK^{+}18] Hong-Yi Chen, Cristina David, Daniel Kroening, Peter Schrammel, and Björn Wachter. Bitprecise procedure-modular termination analysis. ACM Transactions on Programming Languages and Systems, 40(1):1:1– 1:??, January 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Casey:2007:OIB

[CEG07] Kevin Casey, M. Anton Ertl, and David Gregg. Optimizing indirect branch prediction accuracy in virtual machine interpreters. ACM Transactions on Programming Languages and Systems, 29(6):37:1–37:36, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chander:2007:ERB

[CEI^{+}07] Ajay Chander, David Espinosa, Nayeem Islam, Peter Lee, and George C. Necula. Enforcing resource bounds via static verification of dynamic checks. ACM Transactions on Programming Languages and Systems, 29(5): 28:1–28:18, August 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Clarke:1986:AVF

[CES86] E. M. Clarke, E. A. Emerson, and A. P. Sistla. Automatic verification of finite-state concurrent systems using temporal logic specifications. ACM Transactions on Programming Languages and Systems, 8(2): 244–263, April 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/

5399.html. An early version appeared in Proc. 10th ACM Symposium on Principles of Programming Languages, 1983.

Chen:2014:ETI

[CEW14] Sheng Chen, Martin Erwig, and Eric Walkingshaw. Extending type inference to variational programs. ACM Transactions on Programming Languages and Systems, 36(1):1:1– 1:??, March 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Choi:1994:SSP

[CF94] Jong-Deok Choi and Jeanne Ferrante. Static slicing in the presence of goto statements. ACM Transactions on Programming Languages and Systems, 16(4): 1097–1113, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
183438.html.

Cytron:1995:ECN

[CF95] Ron K. Cytron and Jeanne Ferrante. Efficiently computing ϕ-nodes on-the-fly. ACM Transactions on Programming Languages and Systems, 17(3): 487–506, May 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
203099.html.

Clements:2004:TRM

[CF04] John Clements and Matthias Felleisen. A tail-recursive machine with stack inspection.

ACM Transactions on Programming Languages and Systems, 26 (6):1029–1052, November 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cortesi:1997:CAI

[CFG^{+}97] Agostino Cortesi, Gilberto File, Roberto Giacobazzi, Catuscia Palamidessi, and Francesco Ranzato. Complementation in abstract interpretation. ACM Transactions on Programming Languages and Systems, 19(1): 7–47, January 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-1/p7-cortesi/.

Chatterjee:2019:NPW

[CFG19] Krishnendu Chatterjee, Hongfei Fu, and Amir Kafshdar Goharshady. Non-polynomial worst-case analysis of recursive programs. ACM Transactions on Programming Languages and Systems, 41(4):20:1– 20:??, November 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3339984.

Codish:1994:SAC

[CFM94] Michael Codish, Moreno Falaschi, and Kim Marriott. Suspension analyses for concurrent logic programs. ACM Transactions on Programming Languages and Systems, 16(3):

649–686, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177656.html.

Chatterjee:2018:AAQ

[CFNH18] Krishnendu Chatterjee, Hongfei Fu, Petr Novotný, and Rouzbeh Hasheminezhad. Algorithmic analysis of qualitative and quantitative termination problems for affine probabilistic programs. ACM Transactions on Programming Languages and Systems, 40 (2):7:1–7:??, June 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cortes:2004:HLA

[CFP^{+}04] Corinna Cortes, Kathleen Fisher, Daryl Pregibon, Anne Rogers, and Frederick Smith. Hancock: a language for analyzing transactional data streams. ACM Transactions on Programming Languages and Systems, 26(2): 301–338, March 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cytron:1991:ECS

[CFR^{+}91] Ron Cytron, Jeanne Ferrante, Barry K. Rosen, Mark N. Wegman, and F. Kenneth Zadeck. Efficiently computing static single assignment form and the control dependence graph. ACM Transactions on Programming Languages and Systems, 13(4): 451–490, October 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        115320.html.

Clark:1986:PPP

[CG86] Keith Clark and Steve Gregory. Parlog: parallel programming in logic. ACM Transactions on Programming Languages and Systems, 8(1):1– 49, January 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        5390.html.

Chin:1995:ROA

[CG95] Wei-Ngan Chin and Eak-Khoon Goh. A reexamination of “Optimization of Array Subscript Range Checks”. ACM Transactions on Programming Languages and Systems, 17(2):217– 227, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        201063.html.

Christensen:2004:OPE

[CG04] Niels H. Christensen and Robert Glück. Offline partial evaluation can be as accurate as online partial evaluation. ACM Transactions on Programming Languages and Systems, 26(1): 191–220, January 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chatterjee:2019:FAD

[CGG^{+}19] Krishnendu Chatterjee, Amir Kafshdar Goharshady, Prateesh Goyal, Rasmus Ibsen-Jensen, and Andreas Pavlogiannis. Faster algorithms for dynamic algebraic queries in basic RSMs with constant treewidth. ACM Transactions on Programming Languages and Systems, 41(4):23:1– 23:??, November 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3363525.

Calder:1997:EBS

[CGJ^{+}97a] Brad Calder, Dirk Grunwald, Michael Jones, Donald Lindsay, James Martin, Michael Mozer, and Benjamin Zorn. Evidencebased static branch prediction using machine learning. ACM Transactions on Programming Languages and Systems, 19(1): 188–222, January 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-1/p188-calder/.

Clarke:1997:VPN

[CGJ97b] E. M. Clarke, O. Grumberg, and S. Jha. Verifying parameterized networks. ACM Transactions on Programming Languages and Systems, 19(5):726– 750, September 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/

                                        journals/toplas/1997-19-5/
                                        p726-clarke/.

Clarke:1994:MCA

[CGL94] Edmund M. Clarke, Orna Grumberg, and David E. Long. Model checking and abstraction. ACM Transactions on Programming Languages and Systems, 16(5):1512–1542, September 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        186051.html.

Castagna:2009:TCW

[CGP09] Giuseppe Castagna, Nils Gesbert, and Luca Padovani. A theory of contracts for Web services. ACM Transactions on Programming Languages and Systems, 31 (5):19:1–19:61, June 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Choi:2003:SAS

[CGS^{+}03] Jong-Deok Choi, Manish Gupta, Mauricio J. Serrano, Vugranam C. Sreedhar, and Samuel P. Midkiff. Stack allocation and synchronization optimizations for Java using escape analysis. ACM Transactions on Programming Languages and Systems, 25(6): 876–910, November 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chatterjee:1995:OEA

[CGST95] Siddhartha Chatterjee, John R. Gilbert, Robert Schreiber, and Shang-Hua Teng. Optimal evaluation of array expressions on massively parallel machines. ACM Transactions on Programming Languages and Systems, 17 (1):123–156, January 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201004.html.

Cohen:1987:PCU

[CH87] Jacques Cohen and Timothy J.

Hickey. Parsing and compiling using Prolog. ACM Transactions on Programming Languages and Systems, 9(2):125– 163, April 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
22946.html.

Chow:1990:PBC

[CH90] Fred C. Chow and John L.

Hennessy. The priority-based coloring approach to register allocation. ACM Transactions on Programming Languages and Systems, 12(4):501– 536, October 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
88621.html.

Charlesworth:1987:MR

[Cha87]  Arthur Charlesworth.  The

multiway rendezvous. ACM Transactions on Programming Languages and Systems, 9(3):

350–366, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        24050.html.

Chatterjee:1993:CND

[Cha93] Siddhartha Chatterjee. Compiling nested data-parallel programs for shared-memory multiprocessors. ACM Transactions on Programming Languages and Systems, 15(3): 400–462, July 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174152.html.

Charlesworth:2002:UAC

[Cha02] Arthur Charlesworth. The undecidability of associativity and commutativity analysis. ACM Transactions on Programming Languages and Systems, 24(5): 554–565, September 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chitil:2005:PPL

[Chi05] Olaf Chitil. Pretty printing with lazy dequeues. ACM Transactions on Programming Languages and Systems, 27(1): 163–184, January 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cogumbreiro:2019:DDV

[CHMY19] Tiago Cogumbreiro, Raymond Hu, Francisco Martins, and Nobuko Yoshida. Dynamic deadlock verification for general barrier synchronisation. ACM Transactions on Programming Languages and Systems, 41(1): 1:1–1:??, March 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3229060.

Carbone:2012:SCC

[CHY12] Marco Carbone, Kohei Honda, and Nobuko Yoshida. Structured communication-centered programming for Web services. ACM Transactions on Programming Languages and Systems, 34 (2):8:1–8:78, June 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cameron:1984:GBD

[CI84] Robert D. Cameron and M. Robert Ito. Grammar-based definition of metaprogramming systems. ACM Transactions on Programming Languages and Systems, 6 (1):20–54, January 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chatterjee:2018:AAP

[CIJGP18] Krishnendu Chatterjee, Rasmus Ibsen-Jensen, Amir Kafshdar Goharshady, and Andreas Pavlogiannis. Algorithms for algebraic path properties in concurrent systems of constant treewidth components. ACM Transactions on Programming Languages and Systems, 40(3):9:1–9:??, August 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cejtin:1995:HOD

[CJK95] Henry Cejtin, Suresh Jagannathan, and Richard Kelsey. Higher-order distributed objects. ACM Transactions on Programming Languages and Systems, 17(5):704–739, September 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        213986.html.

Consel:1993:PPE

[CK93] Charles Consel and Siau Cheng Khoo. Parameterized partial evaluation. ACM Transactions on Programming Languages and Systems, 15(3): 463–493, July 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174155.html.

Carr:1994:IRM

[CK94] S. Carr and K. Kennedy. Improving the ratio of memory operations in floating-point operations in loops. ACM Transactions on Programming Languages and Systems, 16(6):1768– 1810, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        197366.html.

Cooper:1986:IIA

[CKT86] Keith D. Cooper, Ken Kennedy, and Linda Torczon. The impact of interprocedural analysis and optimizations in the R(n) programming environment. ACM Transactions on Programming Languages and Systems, 8(4): 491–523, October 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
6489.html.

Crowl:1994:PPC

[CL94] Lawrence A. Crowl and Thomas J.

LeBlanc. Parallel programming with control abstraction. ACM Transactions on Programming Languages and Systems, 16(3):524–576, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177584.html.

Chambers:1995:TMM

[CL95] Craig Chambers and Gary T.

Leavens. Typechecking and modules for multimethods. ACM Transactions on Programming Languages and Systems, 17(6): 805–843, November 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
218571.html.

Clarke:1980:SRI

[Cla80] Edmund Melson Clarke, Jr. Synthesis of resource invariants for concurrent programs. ACM Transactions on Programming Languages and Systems, 2(3): 338–358, July 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chandy:1984:DPP

[CM84] K. M. Chandy and Jayadev Misra. The drinking philosopher’s problem. ACM Transactions on Programming Languages and Systems, 6(4):632– 646, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Chandy:1986:ESR

[CM86a] K. M. Chandy and Jayadev Misra. An example of stepwise refinement of distributed programs: Quiescence detection. ACM Transactions on Programming Languages and Systems, 8 (3):326–343, July 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5958.html.

Chirica:1986:TCI

[CM86b] Laurian M. Chirica and David F. Martin. Toward compiler implementation correctness proofs. ACM Transactions on Programming Languages and Systems, 8 (2):185–214, April 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
30847.html.

Copperman:1993:TCF

[CM93] Max Copperman and Charles E.

McDowell. Technical correspondence: a further note on Hennessy’s “Symbolic Debugging of Optimized Code”. ACM Transactions on Programming Languages and Systems, 15(2): 357–365, April 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

214526.html. See [Hen82].

Codish:1995:IAI

[CMB^{+}95] Michael Codish, Anne Mulkers, Maurice Bruynooghe, Maria Garcia de la Banda, and Manuel Hermenegildo. Improving abstract interpretations by combining domains. ACM Transactions on Programming Languages and Systems, 17(1):28– 44, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
200998.html.

Clifton:2006:MDR

[CMLC06] Curtis Clifton, Todd Millstein, Gary T. Leavens, and Craig Chambers. MultiJava: Design rationale, compiler implementation, and applications. ACM Transactions on Programming Languages and Systems, 28(3): 517–575, May 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Choi:1991:TDP

[CMN91] Jong-Deok Choi, Barton P. Miller, and Robert H. B. Netzer. Techniques for debugging parallel programs with flowback analysis. ACM Transactions on Programming Languages and Systems, 13(4):491– 530, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        115324.html.

Chiari:2023:MCO

[CMPP23] Michele Chiari, Dino Mandrioli, Francesco Pontiggia, and Matteo Pradella. A model checker for operator precedence languages. ACM Transactions on Programming Languages and Systems, 45 (3):19:1–19:??, September 2023. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3608443.

Christensen:2003:EJH

[CMS03] Aske Simon Christensen, Anders Møller, and Michael I. Schwartzbach. Extending Java for high-level Web service construction. ACM Transactions on Programming Languages and Systems, 25(6):814–875, November 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cohen:1983:CCA

[CN83] Jacques Cohen and Alexandru Nicolau. Comparison of compacting algorithms for garbage collection. ACM Transactions on Programming Languages and Systems, 5(4):532–553, October 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Clemm:1990:MEI

[CO90] Geoffrey Clemm and Leon Osterweil. A mechanism for environment integration. ACM Transactions on Programming Languages and Systems, 12(1): 1–25, January 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
77607.html.

Coblenz:2020:OTA

[COE^{+}20] Michael Coblenz, Reed Oei, Tyler Etzel, Paulette Koronkevich, Miles Baker, Yannick Bloem, Brad A. Myers, Joshua Sunshine, and Jonathan Aldrich. Obsidian: Typestate and assets for safer blockchain programming. ACM Transactions on Programming Languages and Systems, 42(3):14:1– 14:82, December 2020. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3417516.

Cohen:1983:ERR

[Coh83] Norman H. Cohen. Eliminating redundant recursive calls. ACM Transactions on Programming Languages and Systems, 5 (3):265–299, July 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cohen:1985:NCE

[Coh85] Norman H. Cohen. A note on Cohen’s “Eliminating Redundant Recursive Calls”. ACM Transactions on Programming Languages and Systems, 7(4): 680–685, October 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        215006.html.

Cohen:1991:TCT

[Coh91] Norman H. Cohen. Technical correspondence: Type-extension type tests can be performed in constant time. ACM Transactions on Programming Languages and Systems, 13(4):626– 629, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        115297.html.  See [Wir88,

Wir91].

Colussi:1984:RES

[Col84] L. Colussi. Recursion as an effective step in program development. ACM Transactions on Programming Languages and Systems, 6(1):55–67, January 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Comer:1980:NMS

[Com80] Douglas Comer. A note on median split trees. ACM Transactions on Programming Languages and Systems, 2(1):129– 133, January 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Copperman:1994:DOC

[Cop94] Max Copperman. Debugging optimized code without being misled. ACM Transactions on Programming Languages and Systems, 16(3):387– 427, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177517.html.

Carle:1995:MBI

[CP95] Alan Carle and Lori Pollock. Matching-based incremental evaluators for hierarchical attribute grammar dialects. ACM Transactions on Programming Languages and Systems, 17 (2):394–429, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201071.html.

Carle:1996:OCP

[CP96] Alan Carle and Lori Pollock. On the optimality of change propagation for incremental evaluation of hierarchical attribute grammars. ACM Transactions on Programming Languages and Systems, 18(1):16– 29, January 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
225542.html.

Cohen:2017:LPC

[CP17] Nachshon Cohen and Erez Petrank. Limitations of partial compaction: Towards practical bounds. ACM Transactions on Programming Languages and Systems, 39(1):2:1– 2:??, March 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Corchuelo:2002:RSE

[CPRT02] Rafael Corchuelo, José A. Pérez, Antonio Ruiz, and Miguel Toro. Repairing syntax errors in LR parsers. ACM Transactions on Programming Languages and Systems, 24(6):698–710, November 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cleaveland:1993:CWS

[CPS93] Rance Cleaveland, Joachim Parrow, and Bernhard Steffen. The Concurrency Workbench: a semantics-based tool for the verification of concurrent systems. ACM Transactions on Programming Languages and Systems, 15 (1):36–72, January 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
151648.html.

Carson:1987:GSP

[CR87] Scott D. Carson and Paul F.

Reynolds, Jr. The geometry of semaphore programs. ACM Transactions on Programming Languages and Systems, 9(1): 25–53, January 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
9759.html.

Cooke:2008:NTD

[CRN^{+}08] Daniel E. Cooke, J. Nelson Rushton, Brad Nemanich, Robert G. Watson, and Per Andersen. Normalize, transpose, and distribute: an automatic approach for handling nonscalars. ACM Transactions on Programming Languages and Systems, 30 (2):9:1–9:49, March 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cuny:1987:CDD

[CS87] Janice E. Cuny and Lawrence Snyder. Conversion from datadriven to synchronous execution in loop programs. ACM Transactions on Programming Languages and Systems, 9(4): 599–617, October 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
31334.html.

Choy:1995:EFT

[CS95] Manhoi Choy and Ambuj K.

Singh. Efficient fault-tolerant algorithms for distributed resource allocation. ACM Transactions on Programming Languages and Systems, 17(3):535– 559, May 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
203101.html.

Chen:2004:LGS

[CS04] Yifeng Chen and J. W. Sanders. Logic of global synchrony. ACM Transactions on Programming Languages and Systems, 26(2): 221–262, March 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Clausen:2000:JBC

[CSCM00] Lars Ræder Clausen, Ulrik Pagh Schultz, Charles Consel, and Gilles Muller. Java bytecode compression for low-end embedded systems. ACM Transactions on Programming Languages and Systems, 22(3):471– 489, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-3/
p471-clausen/p471-clausen.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-3/p471-clausen/.

Codish:1999:SGD

[CSS99] Michael Codish, Harald Søndergaard, and Peter J. Stuckey. Sharing and groundness dependencies in logic programs. ACM Transactions on Programming Languages and Systems, 21 (5):948–976, September 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-5/
p948-codish/p948-codish.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-5/p948-codish/.

Cooper:2001:OSR

[CSV01] Keith D. Cooper, L. Taylor Simpson, and Christopher A. Vick. Operator strength reduction. ACM Transactions on Programming Languages and Systems, 23(5):603–625, September 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Carlsson:2006:MAC

[CSW06] Richard Carlsson, Konstantinos Sagonas, and Jesper Wilhelmsson. Message analysis for concurrent programs using message passing. ACM Transactions on Programming Languages and Systems, 28(4):715– 746, July 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Collberg:2007:DGB

[CTT07] Christian S. Collberg, Clark Thomborson, and Gregg M. Townsend. Dynamic graphbased software fingerprinting. ACM Transactions on Programming Languages and Systems, 29(6):35:1–35:67, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Cheney:2008:NLP

[CU08] James Cheney and Christian Urban. Nominal logic programming. ACM Transactions on Programming Languages and Systems, 30(5):26:1–26:47, August 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Constable:1984:TTP

[CZ84] Robert L. Constable and Daniel R. Zlatin. The type theory of PL/CV3. ACM Transactions on Programming Languages and Systems, 6(1):94–117, January 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Damiani:2003:RIT

[Dam03] Ferruccio Damiani. Rank 2 intersection types for local definitions and conditional expressions. ACM Transactions on Programming Languages and Systems, 25(4):401–451, July 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Danvy:2023:THA

[Dan23] Olivier Danvy. The tortoise and the hare algorithm for finite lists, compositionally. ACM Transactions on Programming Languages and Systems, 45(1):

1:1–1:??, March 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3564619.

Darlington:1990:SDG

[Dar90] Jared L. Darlington. Search direction by goal failure in goaloriented programming. ACM Transactions on Programming Languages and Systems, 12(2): 224–252, April 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
78946.html.

Dujardin:1998:FAC

[DAS98] Eric Dujardin, Eric Amiel, and Eric Simon. Fast algorithms for compressed multimethod dispatch table generation. ACM Transactions on Programming Languages and Systems, 20(1): 116–165, January 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-1/
p116-dujardin/.

Dillon:1988:CET

[DAW88] Laura K. Dillon, George S.

Avrunin, and Jack C. Wiledon. Constrained expressions: Toward broad applicability of analysis methods for distributed software systems. ACM Transactions on Programming Languages and Systems, 10(3):

374–402, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
44502.html.

Dunlop:1985:GSU

[DB85] Douglas D. Dunlop and Victor R. Basili. Generalizing specifications for uniformly implemented loops. ACM Transactions on Programming Languages and Systems, 7(1):137– 158, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
2708.html.

deBruin:1985:DSD

[dBB85] Arie de Bruin and Wim Böhm. The denotational semantics of dynamic networks of processes. ACM Transactions on Programming Languages and Systems, 7 (4):656–679, October 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4473.html.

deBoer:2021:CCR

[dBH21] Frank S. de Boer and Hans-Dieter A. Hiep. Completeness and complexity of reasoning about call-by-value in Hoare logic. ACM Transactions on Programming Languages and Systems, 43(4):17:1– 17:35, December 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477143.

deBoer:2024:PCP

[dBJPT24] Frank de Boer, Einar Broch Johnsen, Violet Ka I. Pun, and Silvia Lizeth Tapia Tarifa. Proving correctness of parallel implementations of transition system models. ACM Transactions on Programming Languages and Systems, 46(3):9:1– 9:??, September 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3660630.

Deng:2022:SDR

[DC22] Chaoqiang Deng and Patrick Cousot. The systematic design of responsibility analysis by abstract interpretation. ACM Transactions on Programming Languages and Systems, 44(1): 3:1–3:90, March 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3484938.

Donahue:1985:DTV

[DD85] James Donahue and Alan Demers. Data types are values. ACM Transactions on Programming Languages and Systems, 7 (3):426–445, July 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3987.html.

DeSutter:2005:LTB

[DDD05] Bjorn De Sutter, Bruno De Bus, and Koen De Bosschere. Linktime binary rewriting techniques for program compaction. ACM Transactions on Programming Languages and Systems, 27(5): 882–945, September 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Drossopoulou:2002:MDO

[DDDCG02] Sophia Drossopoulou, Ferruccio Damiani, Mariangiola Dezani-Ciancaglini, and Paola Giannini. More dynamic object reclassification: Fickle_{∥}. ACM Transactions on Programming Languages and Systems, 24(2): 153–191, March 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dencker:1984:OPT

[DDH84] Peter Dencker, Karl Dürre, and Johannes Heuft. Optimization of parser tables for portable compilers. ACM Transactions on Programming Languages and Systems, 6(4):546–572, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dietl:2011:SOT

[DDM11] Werner Dietl, Sophia Drossopoulou, and Peter Müller. Separating ownership topology and encapsulation with generic universe types. ACM Transactions on Programming Languages and Systems, 33(6):20:1– 20:62, December 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Das:2022:NST

[DDMP22] Ankush Das, Henry Deyoung, Andreia Mordido, and Frank Pfenning. Nested session types. ACM Transactions on Programming Languages and Systems, 44 (3):19:1–19:45, September 2022. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3539656.

Decorte:1999:CBT

[DDV99] Stefaan Decorte, Danny De Schreye, and Henk Vandecasteele. Constraint-based termination analysis of logic programs. ACM Transactions on Programming Languages and Systems, 21(6):1137–1195, November 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-6/p1137-decorte/.

Debray:1989:SIM

[Deb89] Saumya K. Debray. Static inference of modes and data dependencies in logic programs. ACM Transactions on Programming Languages and Systems, 11(3):418–450, July 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
65983.html.

Debray:1995:CDA

[Deb95] Saumya K. Debray. On the complexity of dataflow analysis of logic programs. ACM Transactions on Programming Languages and Systems, 17(2): 331–365, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        201068.html.

DeMillo:1983:GEI

[DeM83] Richard A. DeMillo. Guest Editor’s introduction. ACM Transactions on Programming Languages and Systems, 5(3):355, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Debray:2000:CTC

[DEMD00] Saumya K. Debray, William Evans, Robert Muth, and Bjorn De Sutter. Compiler techniques for code compaction. ACM Transactions on Programming Languages and Systems, 22(2): 378–415, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-2/p378-debray/.

Dershowitz:1985:PAI

[Der85] Nachum Dershowitz. Program abstraction and instantiation. ACM Transactions on Programming Languages and Systems, 7 (3):446–477, July 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3986.html.

DeFraine:2012:EAC

[DES12] Bruno De Fraine, Erik Ernst, and Mario Südholt. Essential AOP: The A calculus. ACM Transactions on Programming Languages and Systems, 34(3): 12:1–12:43, October 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Davidson:1980:DAR

[DF80] Jack W. Davidson and Christopher W. Fraser. The design and application of a retargetable peephole optimizer. ACM Transactions on Programming Languages and Systems, 2 (2):191–202, April 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See also corrigendum [DF81].

Davidson:1981:CDA

[DF81] Jack W. Davidson and Christopher W. Fraser. Corrigendum: “The Design and Application of a Retargetable Peephole Optimizer”. ACM Transactions on Programming Languages and Systems, 3(1):110, January 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [DF80].

Davidson:1984:CST

[DF84] Jack W. Davidson and Christopher W. Fraser. Code selection through object code optimization. ACM Transactions on Programming Languages and Systems, 6(4):505–526, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Douence:1998:SSF

[DF98] Rémi Douence and Pascal Fradet. A systematic study of functional language implementations. ACM Transactions on Programming Languages and Systems, 20(2):344–387, March 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-2/
                                        p344-douence/.

Dimoulas:2011:CSH

[DF11] Christos Dimoulas and Matthias Felleisen. On contract satisfaction in a higher-order world. ACM Transactions on Programming Languages and Systems, 33 (5):16:1–16:29, November 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Demetrescu:2015:RIP

[DFR15] Camil Demetrescu, Irene Finocchi, and Andrea Ribichini. Reactive imperative programming with dataflow constraints. ACM Transactions on Programming Languages and Systems, 37(1): 3:1–3:??, January 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

DalLago:2019:PTM

[DG19] Ugo Dal Lago and Charles Grellois. Probabilistic termination by monadic affine sized typing. ACM Transactions on Programming Languages and Systems, 41 (2):10:1–10:??, June 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3293605.

Dams:1997:AIR

[DGG97] Dennis Dams, Rob Gerth, and Orna Grumberg. Abstract interpretation of reactive systems. ACM Transactions on Programming Languages and Systems, 19 (2):253–291, March 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-2/p253-dams/.

Dewar:1979:PRE

[DGL^{+}79] Robert B. K. Dewar, Arthur Grand, Ssu-Cheng Liu, Jacob T. Schwartz, and Edmond Schonberg. Programming by refinement, as exemplified by the SETL representation sublanguage. ACM Transactions on Programming Languages and Systems, 1(1):27–49, July 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

DeBoer:1997:PCC

[DGMP97] Frank S. De Boer, Maurizio Gabbrielli, Elena Marchiori, and Catuscia Palamidessi. Proving concurrent constraint programs correct. ACM Transactions on Programming Languages and Systems, 19(5):685– 725, September 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-5/
p685-de_boer/.

Duesterwald:1997:PFD

[DGS97] Evelyn Duesterwald, Rajiv Gupta, and Mary Lou Soffa. A practical framework for demanddriven interprocedural data flow analysis. ACM Transactions on Programming Languages and Systems, 19(6):992– 1030, November 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-6/
p992-duesterwald/.

Dhamdhere:1991:PAG

[Dha91] Dhananjay M. Dhamdhere. Practical adaptation of the global optimization algorithm of Morel and Renvoise. ACM Transactions on Programming Languages and Systems, 13(2): 291–294, April 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
103161.html;   http://www.

acm.org/pubs/toc/Abstracts/
0164-0925/214520.html.

delaBanda:1996:GAC

[dHB^{+}96] M. Garcia de la Banda, M. Hermenegildo, M. Bruynooghe, V. Dumortier, G. Janssens, and W. Simoens. Global analysis of constraint logic programs. ACM Transactions on Programming Languages and Systems, 18(5): 564–614, September 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
232734.html.

Din:2024:LAG

[DHH^{+}24] Crystal Chang Din, Reiner Hähnle, Ludovic Henrio, Einar Broch Johnsen, Violet Ka I. Pun, and S. Lizeth Tapia Tarifa. Locally abstract, globally concrete semantics of concurrent programming languages. ACM Transactions on Programming Languages and Systems, 46(1): 3:1–3:??, March 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3648439.

DeLaBanda:2000:ICL

[DHM00] Marí García De La Banda, Manuel Hermengildo, and Kim Marriott. Independence in CLP languages. ACM Transactions on Programming Languages and Systems, 22(2):296– 339, March 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-2/p296-garcia_de_
                                        la_banda/.

Dolby:2012:DCA

[DHM^{+}12] Julian Dolby, Christian Hammer, Daniel Marino, Frank Tip, Mandana Vaziri, and Jan Vitek. A data-centric approach to synchronization. ACM Transactions on Programming Languages and Systems, 34(1):4:1– 4:48, April 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dolev:2009:SSP

[DHS09] Shlomi Dolev, Yinnon Haviv, and Mooly Sagiv. Selfstabilization preserving compiler. ACM Transactions on Programming Languages and Systems, 31(6):22:1–22:22, August 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

DeMoura:2009:RC

[DI09] Ana Lúcia De Moura and Roberto Ierusalimschy. Revisiting coroutines. ACM Transactions on Programming Languages and Systems, 31(2):6:1– 6:31, February 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dillon:1990:USE

[Dil90] Laura K. Dillon. Using symbolic execution for verification of Ada tasking programs. ACM Transactions on Programming Languages and Systems, 12(4): 643–669, October 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
96551.html.

deJonge:2012:NFE

[dJKVS12] Maartje de Jonge, Lennart C. L.

Kats, Eelco Visser, and Emma Söderberg. Natural and flexible error recovery for generated modular language environments. ACM Transactions on Programming Languages and Systems, 34 (4):15:1–15:50, December 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dodds:2016:VCS

[DJP^{+}16] Mike Dodds, Suresh Jagannathan, Matthew J. Parkinson, Kasper Svendsen, and Lars Birkedal. Verifying custom synchronization constructs using higher-order separation logic. ACM Transactions on Programming Languages and Systems, 38 (2):4:1–4:??, January 2016. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Darulova:2017:TCR

[DK17] Eva Darulova and Viktor Kuncak. Towards a compiler for reals. ACM Transactions on Programming Languages and Systems, 39(2):8:1–8:??, May 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

David:2018:PSP

[DKKL18] Cristina David, Pascal Kesseli, Daniel Kroening, and Matt Lewis. Program synthesis for program analysis. ACM Transactions on Programming Languages and Systems, 40(2):5:1– 5:??, June 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Drinic:2007:PPC

[DKV07] Milenko Drinić, Darko Kirovski, and Hoi Vo. PPMexe: Program compression. ACM Transactions on Programming Languages and Systems, 29(1):3:1–3:31, January 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Debray:1993:CAL

[DL93] Saumya K. Debray and Nai-Wei Lin. Cost analysis of logic programs. ACM Transactions on Programming Languages and Systems, 15(5):826– 875, November 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        161472.html.

Danicic:2018:SBS

[DL18] Sebastian Danicic and Michael R. Laurence. Static backward slicing of non-deterministic programs and systems. ACM Transactions on Programming Languages and Systems, 40(3): 11:1–11:??, August 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See Editor’s foreword [Mye18].

Dissegna:2016:AIB

[DLR16] Stefano Dissegna, Francesco Logozzo, and Francesco Ranzato. An abstract interpretationbased model of tracing just-in-time compilation. ACM Transactions on Programming Languages and Systems, 38(2):7:1– 7:??, January 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Degano:1988:EIL

[DMM88] Pierpaolo Degano, Stefano Mannucci, and Bruno Mojana. Efficient incremental LR parsing for syntax-directed editors. ACM Transactions on Programming Languages and Systems, 10(3):345–373, July 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214503.html.

Diwan:2001:UTA

[DMM01] Amer Diwan, Kathryn S.

McKinley, and J. Eliot B. Moss. Using types to analyze and optimize object-oriented programs. ACM Transactions on Programming Languages and Systems, 23 (1):30–72, January 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Danvy:1996:EED

[DMP96] Olivier Danvy, Karoline Malmkjær, and Jens Palsberg. Etaexpansion does the trick. ACM Transactions on Programming Languages and Systems, 18(6): 730–751, November 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
235831.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/236119.html.

DiVincenzo:2024:GCS

[DMZ^{+}24] Jenna DiVincenzo, Ian McCormack, Conrad Zimmerman, Hemant Gouni, Jacob Gorenburg, Jan-Paul Ramos-Dávila, Mona Zhang, Joshua Sunshine, Éric Tanter, and Jonathan Aldrich. Gradual C0: Symbolic execution for gradual verification. ACM Transactions on Programming Languages and Systems, 46(4): 14:1–14:??, December 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3704808.

Ducasse:2006:TMF

[DNS^{+}06] Stéphane Ducasse, Oscar Nierstrasz, Nathanael Schärli, Roel Wuyts, and Andrew P. Black. Traits: a mechanism for finegrained reuse. ACM Transactions on Programming Languages and Systems, 28(2):331– 388, March 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

DeRemer:1982:ECL

[DP82] Frank DeRemer and Thomas Pennello. Efficient computation of LALR(1) look-ahead sets. ACM Transactions on Programming Languages and Systems, 4 (4):615–649, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dhamdhere:1993:EAB

[DP93] Dhananjay M. Dhamdhere and Harish Patil. An elimination algorithm for bidirectional data flow problems using edge placement. ACM Transactions on Programming Languages and Systems, 15(2):312– 336, April 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
169684.html.

Debray:1997:ICF

[DP97] Saumya K. Debray and Todd A.

Proebsting. Interprocedural control flow analysis of first-order programs with tail-call optimization. ACM Transactions on Programming Languages and Systems, 19(4):568–585, July 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-4/
p568-debray/.

DeRose:1999:TTM

[DP99] Luiz De Rose and David Padua.

Techniques for the translation of MATLAB programs into Fortran 90. ACM Transactions on Programming Languages and Systems, 21(2):286– 323, March 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-2/p286-de_rose/.

Devriese:2022:TPV

[DPP22] Dominique Devriese, Marco Patrignani, and Frank Piessens. Two parametricities versus three universal types. ACM Transactions on Programming Languages and Systems, 44(4):23:1– 23:??, December 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3539657.

Dovier:2000:SCL

[DPPR00] Agostino Dovier, Carla Piazza, Enrico Pontelli, and Gianfranco Rossi. Sets and constraint logic programming. ACM Transactions on Programming Languages and Systems, 22(5):861– 931, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/2000-22-5/
                                        p861-dovier/p861-dovier.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/toplas/
                                        2000-22-5/p861-dovier/.

Das:2005:PFI

[DR05] Dibyendu Das and U. Ramakrishna. A practical and fast iterative algorithm for ϕ-function computation using DJ graphs.

ACM Transactions on Programming Languages and Systems, 27 (3):426–440, May 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dawson:1996:PPU

[DRSS96] Steven Dawson, C. R. Ramakrishnan, Steven Skiena, and Terrance Swift. Principles and practice of unification factoring. ACM Transactions on Programming Languages and Systems, 18(5):528– 563, September 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
232722.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/232722.html;   http:
//www.acm.org/pubs/toc/Abstracts/
toplas/235455.html.

Dekel:1983:PGP

[DS83] Eliezer Dekel and Sartaj Sahni.

Parallel generation of postfix and tree forms. ACM Transactions on Programming Languages and Systems, 5(3):300– 317, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Drechsler:1988:TCS

[DS88] Karl-Heinz Drechsler and Manfred P. Stadel. Technical correspondence: a solution to a problem with Morel and Renvoise’s “Global Optimization by Suppression of Partial Redundancies”. ACM Transactions on Programming Languages and Systems, 10(4):635– 640, October 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        214509.html.  See remark

[Sor89].

Dewan:1990:ASA

[DS90] Prasun Dewan and Marvin Solomon. An approach to support automatic generation of user interfaces. ACM Transactions on Programming Languages and Systems, 12(4):566– 609, October 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        214518.html.

Dhamdhere:1998:DCD

[DS98] D. M. Dhamdhere and K. V. Sankaranarayanan. Dynamic currency determination in optimized programs. ACM Transactions on Programming Languages and Systems, 20(6):1111– 1130, November 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-6/
                                        p1111-dhamdhere/.

DOsualdo:2021:TLC

[DSFG21] Emanuele D’Osualdo, Julian Sutherland, Azadeh Farzan, and Philippa Gardner. TaDA live:

Compositional reasoning for termination of fine-grained concurrent programs. ACM Transactions on Programming Languages and Systems, 43(4):16:1– 16:134, December 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3477082.

Dewar:1982:TDG

[DSW82] Robert B. K. Dewar, Micha Sharir, and Elia Weixelbaum. Transformational derivation of a garbage collection algorithm. ACM Transactions on Programming Languages and Systems, 4 (4):650–667, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Derrick:2011:MVP

[DSW11] John Derrick, Gerhard Schellhorn, and Heike Wehrheim. Mechanically verified proof obligations for linearizability. ACM Transactions on Programming Languages and Systems, 33(1): 4:1–4:43, January 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ducournau:2008:PHA

[Duc08] Roland Ducournau. Perfect hashing as an almost perfect subtype test. ACM Transactions on Programming Languages and Systems, 30(6):33:1–33:56, October 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Duggan:1999:DTD

[Dug99]  Dominic Duggan.  Dynamic

typing for distributed programming in polymorphic languages. ACM Transactions on Programming Languages and Systems, 21 (1):11–45, January 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-1/p11-duggan/.

Duggan:2002:TSL

[Dug02] Dominic Duggan. Type-safe linking with recursive DLLs and shared libraries. ACM Transactions on Programming Languages and Systems, 24(6):711– 804, November 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

DeSutter:2007:PID

[DVD07] Bjorn De Sutter, Ludo Van Put, and Koen De Bosschere. A practical interprocedural dominance algorithm. ACM Transactions on Programming Languages and Systems, 29(4):19:1–19:44, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Dantoni:2015:FTB

[DVLM15] Loris D’antoni, Margus Veanes, Benjamin Livshits, and David Molnar. Fast: a transducerbased language for tree manipulation. ACM Transactions on Programming Languages and Systems, 38(1):1:1–1:??, October 2015. CODEN ATPSDT.

ISSN 0164-0925 (print), 1558-4593 (electronic).

Debray:1989:FCL

[DW89] Saumya K. Debray and David S.

Warren. Functional computations in logic programs. ACM Transactions on Programming Languages and Systems, 11(3): 451–481, July 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
65984.html.

Dantas:2008:APA

[DWWW08] Daniel S. Dantas, David Walker, Geoffrey Washburn, and Stephanie Weirich. AspectML: a polymorphic aspect-oriented functional programming language. ACM Transactions on Programming Languages and Systems, 30(3): 14:1–14:60, May 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Etalle:2001:TCP

[EGM01] Sandro Etalle, Maurizio Gabbrielli, and Maria Chiara Meo. Transformations of CCP programs. ACM Transactions on Programming Languages and Systems, 23(3):304–395, May 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Esparza:2014:PBV

[EGP14] Javier Esparza, Pierre Ganty, and Tomás Poch. Patternbased verification for multithreaded programs. ACM Transactions on Programming Languages and Systems, 36(3):9:1– 9:??, September 2014. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Economou:2023:FRT

[EKD23] Dimitrios J. Economou, Neel Krishnaswami, and Jana Dunfield. Focusing on refinement typing. ACM Transactions on Programming Languages and Systems, 45 (4):22:1–22:??, December 2023. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3610408.

Ellis:1982:TCS

[Ell82] John R. Ellis. Technical correspondence: On Steensgaard-Madsen’s “A Statement-Oriented Approach to Data Abstraction”. ACM Transactions on Programming Languages and Systems, 4 (1):120–122, January 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [SM81, SM82].

Elder:2014:ADA

[ELS^{+}14] Matt Elder, Junghee Lim, Tushar Sharma, Tycho Andersen, and Thomas Reps. Abstract domains of affine relations. ACM Transactions on Programming Languages and Systems, 36(4): 11:1–11:??, October 2014. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Eilers:2020:MPP

[EMH20] Marco Eilers, Peter Müller, and Samuel Hitz. Modular product programs. ACM Transactions on Programming Languages and Systems, 42(1):3:1– 3:??, January 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3324783.

Ernst:1980:SAD

[EO80] George W. Ernst and William F.

Ogden. Specification of abstract data types in Modula. ACM Transactions on Programming Languages and Systems, 2 (4):522–543, October 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Emerson:1997:USW

[ES97] E. A. Emerson and A. P. Sistla.

Utilizing symmetry when modelchecking under fairness assumptions: An automata-theoretic approach. ACM Transactions on Programming Languages and Systems, 19(4):617–638, July 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1997-19-4/
p617-emerson/.

Eugster:2007:TBP

[Eug07] Patrick Eugster. Type-based publish/subscribe: Concepts and experiences. ACM Transactions on Programming Languages and Systems, 29(1):6:1– 6:50, January 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Finlay:1993:TCC

[FA93] Alan Finlay and Lloyd Allison. Technical correspondence: a correction to the denotational semantics for the Prolog of Nicholson and Foo. ACM Transactions on Programming Languages and Systems, 15(1):206– 208, January 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        151652.html. See [NF89].

Fateman:1982:HLL

[Fat82] Richard J. Fateman. Highlevel language implications of the proposed IEEE floatingpoint standard. ACM Transactions on Programming Languages and Systems, 4(2):239– 257, April 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Feng:2012:BQP

[FDY12] Yuan Feng, Runyao Duan, and Mingsheng Ying. Bisimulation for quantum processes. ACM Transactions on Programming Languages and Systems, 34(4): 17:1–17:43, December 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Feather:1982:SAP

[Fea82] Martin S. Feather. A system for assisting program transformation. ACM Transactions on Programming Languages and Systems, 4(1):1–20, January 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Feather:1987:LSS

[Fea87] Martin S. Feather. Language support for the specification and development of composite systems. ACM Transactions on Programming Languages and Systems, 9(2):198– 234, April 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
22947.html.

Flanagan:1999:CSB

[FF99] Cormac Flanagan and Matthias Felleisen. Componential setbased analysis. ACM Transactions on Programming Languages and Systems, 21(2):370– 416, March 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-2/p370-flanagan/.

Furr:2008:CTS

[FF08] Michael Furr and Jeffrey S. Foster. Checking type safety of foreign function calls. ACM Transactions on Programming Languages and Systems, 30(4):

18:1–18:63, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Florence:2018:PPP

[FFF^{+}18] Spencer P. Florence, Burke Fetscher, Matthew Flatt, William H. Temps, Vincent St-Amour, Tina Kiguradze, Dennis P. West, Charlotte Niznik, Paul R. Yarnold, Robert Bruce Findler, and Steven M. Belknap. POP-PL: a patient-oriented prescription programming language. ACM Transactions on Programming Languages and Systems, 40(3):10:1–10:??, August 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Flanagan:2008:TAS

[FFLQ08] Cormac Flanagan, Stephen N. Freund, Marina Lifshin, and Shaz Qadeer. Types for atomicity: Static checking and inference for Java. ACM Transactions on Programming Languages and Systems, 30(4):20:1– 20:52, July 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fournet:2003:SIT

[FG03] Cedric Fournet and Andrew D. Gordon. Stack inspection: Theory and variants. ACM Transactions on Programming Languages and Systems, 25(3):360– 399, May 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Freudenberger:1994:ASC

[FGL94] Stefan Freudenberger, Thomas R.

Gross, and P. Geoffrey Lowney. Avoidance and suppression of compensation code in a trace scheduling compiler. ACM Transactions on Programming Languages and Systems, 16(4): 1156–1214, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
183446.html.

Foster:2007:CBT

[FGM^{+}07a] J. Nathan Foster, Michael B.

Greenwald, Jonathan T. Moore, Benjamin C. Pierce, and Alan Schmitt. Combinators for bidirectional tree transformations: a linguistic approach to the view-update problem. ACM Transactions on Programming Languages and Systems, 29(3): 17:1–17:65, May 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fournet:2007:TDA

[FGM07b] Cédric Fournet, Andrew D. Gordon, and Sergio Maffeis. A type discipline for authorization policies. ACM Transactions on Programming Languages and Systems, 29(5):25:1–25:37, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fernandez:2004:ICS

[FH04] Antonio J. Fernández and Patricia M. Hill. An interval constraint system for lattice domains. ACM Transactions on Programming Languages and Systems, 26(1):1– 46, January 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fidge:1993:FDP

[Fid93] C. J. Fidge. A formal definition of priority in CSP. ACM Transactions on Programming Languages and Systems, 15(4): 681–705, September 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
155221.html.

Fischer:1980:PCA

[Fis80] Charles N. Fischer. On parsing and compiling arithmetic expressions on vector computers. ACM Transactions on Programming Languages and Systems, 2 (2):203–224, April 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Forejt:2017:PPA

[FJK^{+}17] Vojtăch Forejt, Saurabh Joshi, Daniel Kroening, Ganesh Narayanaswamy, and Subodh Sharma. Precise predictive analysis for discovering communication deadlocks in MPI programs. ACM Transactions on Programming Languages and Systems, 39(4):15:1– 15:??, September 2017. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Foster:2006:FIT

[FJKA06] Jeffrey S. Foster, Robert Johnson, John Kodumal, and Alex Aiken. Flow-insensitive type qualifiers. ACM Transactions on Programming Languages and Systems, 28(6):1035– 1087, November 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fuchs:1985:OPF

[FK85] David R. Fuchs and Donald E.

Knuth. Optimal prepaging and font caching. ACM Transactions on Programming Languages and Systems, 7(1):62– 79, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
2367.html.

Fokkink:1998:WAR

[FKW98] Wan Fokkink, Jasper Kamperman, and Pum Walters. Within ARM’s reach: Compilation of left-linear rewrite systems via minimal rewrite systems. ACM Transactions on Programming Languages and Systems, 20(3): 679–706, May 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-3/
p679-fokkink/.

Fokkink:2000:LRE

[FKW00] Wan Fokkink, Jasper Kamperman, and Pum Walters. Lazy rewriting on eager machinery. ACM Transactions on Programming Languages and Systems, 22 (1):45–86, January 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-1/p45-fokkink/.

Fraser:1981:EDS

[FL81] Christopher W. Fraser and A. A. Lopez. Editing data structures. ACM Transactions on Programming Languages and Systems, 3 (2):115–125, April 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fradet:1991:CFL

[FL91] Pascal Fradet and Daniel Le Métayer. Compilation of functional languages by program transformation. ACM Transactions on Programming Languages and Systems, 13(1):21– 51, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
102805.html.

Frechtling:2015:MMS

[FL15] Michael Frechtling and Philip H. W. Leong. MCALIB: Measuring sensitivity to rounding error with Monte Carlo programming. ACM Transactions on Programming Languages and Systems, 37 (2):5:1–5:??, April 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Fischer:1989:DFA

[FLBB89] Michael J. Fischer, Nancy A.

Lynch, James E. Burns, and Allan Borodin. Distributed FIFO allocation of identical resources using small shared space. ACM Transactions on Programming Languages and Systems, 11(1): 90–114, January 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
59292.html.

Finkel:1987:DDI

[FM87a] Raphael Finkel and Udi Manber. DIB — a distributed implementation of backtracking. ACM Transactions on Programming Languages and Systems, 9 (2):235–256, April 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
24067.html.

Fraser:1987:ERC

[FM87b] Christopher W. Fraser and Eugene W. Myers. An editor for revision control. ACM Transactions on Programming Languages and Systems, 9(2):277– 295, April 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
22948.html.

Freund:1999:TSO

[FM99] Stephen N. Freund and John C.

Mitchell. The type system for object initialization in the Java bytecode language. ACM Transactions on Programming Languages and Systems, 21(6):1196– 1250, November 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-6/p1196-freund/.

Flexeder:2011:FIL

[FMoPS11] Andrea Flexeder, Markus Müllerolm, Michael Petter, and Helmut Seidl. Fast interprocedural linear two-variable equalities. ACM Transactions on Programming Languages and Systems, 33(6): 21:1–21:33, December 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Frohn:2020:ILR

[FNBG20] Florian Frohn, Matthias Naaf, Marc Brockschmidt, and Jürgen Giesl. Inferring lower runtime bounds for integer programs. ACM Transactions on Programming Languages and Systems, 42 (3):13:1–13:50, December 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3410331.

Foster:1996:CPP

[Fos96] Ian Foster. Compositional parallel programming languages. ACM Transactions on Programming Languages and Systems, 18(4):454–476, July 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
233565.html.

Ferrante:1987:PDG

[FOW87] Jeanne Ferrante, Karl J. Ottenstein, and Joe D. Warren. The program dependence graph and its use in optimization. ACM Transactions on Programming Languages and Systems, 9 (3):319–349, July 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
24041.html.

Fisher:2002:GE

[FP02] Kathleen Fisher and Benjamin C. Pierce. Guest editorial. ACM Transactions on Programming Languages and Systems, 24(2):111, March 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Facchinetti:2019:HOD

[FPS19] Leandro Facchinetti, Zachary Palmer, and Scott Smith. Higher-order demand-driven program analysis. ACM Transactions on Programming Languages and Systems, 41(3):14:1– 14:??, July 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3310340.

Francez:1980:CDT

[Fra80a]  Nissim Francez.  Corrigen-

dum: “Distributed Termination”. ACM Transactions on Programming Languages and Systems, 2(3):463, July 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Fra80b, Moh81, Fra81].

Francez:1980:DT

[Fra80b] Nissim Francez. Distributed termination. ACM Transactions on Programming Languages and Systems, 2(1):42–55, January 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See also corrigendum [Fra80a] and remarks [Moh81, Fra81].

Francez:1981:TCR

[Fra81] N. Francez. Technical correspondence: Reply from Francez. ACM Transactions on Programming Languages and Systems, 3 (1):112–113, January 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Fra80b, Moh81].

Farmer:1990:CPC

[FRW90] William M. Farmer, John D. Ramsdell, and Ronald J. Watro. A correctness proof for combinator reduction with cycles. ACM Transactions on Programming Languages and Systems, 12 (1):123–134, January 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
77612.html.

Fan:2023:SMO

[FSH23] Hongyu Fan, Zhihang Sun, and Fei He. Satisfiability modulo ordering consistency theory for SC, TSO, and PSO memory models. ACM Transactions on Programming Languages and Systems, 45 (1):6:1–6:??, March 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579835.

Freudenberger:1983:ESO

[FSS83] Stefan M. Freudenberger, Jacob T. Schwartz, and Micha Sharir. Experience with the SETL optimizer. ACM Transactions on Programming Languages and Systems, 5(1):26– 45, January 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Foster:1994:CAS

[FT94] Ian Foster and Stephen Taylor.

A compiler approach to scalable concurrent-program design. ACM Transactions on Programming Languages and Systems, 16(3):577–604, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177612.html.

Fricker:1995:ICI

[FTJ95] Christine Fricker, Olivier Temam, and William Jalby. Influence of cross-interferences on blocked loops: a case study with matrix-vector multiply. ACM Transactions on Programming Languages and Systems, 17(4): 561–575, July 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
210185.html.

Francez:1985:SIC

[FY85] Nissim Francez and Shaula A. Yemini. Symmetric intertask communication. ACM Transactions on Programming Languages and Systems, 7(4):622– 636, October 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4475.html.

George:1996:IRC

[GA96] Lal George and Andrew W. Appel. Iterated register coalescing. ACM Transactions on Programming Languages and Systems, 18(3):300–324, May 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
229546.html.

Gazinger:1983:PSP

[Gaz83] Harald Gazinger. Parameterized specifications: Parameter passing and implementation with respect to observability. ACM Transactions on Programming Languages and Systems, 5(3): 318–354, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Greiner:1999:PTE

[GB99] John Greiner and Guy E. Blelloch. A provably time-efficient parallel implementation of full speculation. ACM Transactions on Programming Languages and Systems, 21(2):240– 285, March 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-2/p240-greiner/.

Gouda:1986:PLN

[GC86] Mohamed G. Gouda and Chung-Kou Chang. Proving liveness for networks of communicating finite state machines. ACM Transactions on Programming Languages and Systems, 8(1): 154–182, January 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5002.html.

Grove:2001:FCG

[GC01] David Grove and Craig Chambers. A framework for call graph construction algorithms. ACM Transactions on Programming Languages and Systems, 23(6):685–746, November 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gulavani:2011:BSA

[GCRN11] Bhargav S. Gulavani, Supratik Chakraborty, G. Ramalingam, and Aditya V. Nori. Bottomup shape analysis using LISF. ACM Transactions on Programming Languages and Systems, 33 (5):17:1–17:41, November 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gergeron:1982:SAS

[GD82] J. Gergeron and A. Dubuque. A structured APL system. ACM Transactions on Programming Languages and Systems, 4(4): 585–600, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Greenman:2023:TUI

[GDF23] Ben Greenman, Christos Dimoulas, and Matthias Felleisen. Typed-untyped interactions: a comparative analysis. ACM Transactions on Programming Languages and Systems, 45(1): 4:1–4:??, March 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3579833.

Gordon:2017:VIL

[GEGP17] Colin S. Gordon, Michael D. Ernst, Dan Grossman, and Matthew J. Parkinson. Verifying invariants of lock-free data structures with rely-guarantee and refinement types. ACM Transactions on Programming Languages and Systems, 39(3): 11:1–11:??, July 2017. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gelernter:1985:GCL

[Gel85] David Gelernter. Generative communication in Linda. ACM Transactions on Programming Languages and Systems, 7(1): 80–112, January 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
2433.html.

Georgeff:1984:TRS

[Geo84] Michael Georgeff. Transformations and reduction strategies for typed lambda expressions. ACM Transactions on Programming Languages and Systems, 6 (4):603–631, October 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remark [NN86].

Ganapathi:1985:AGD

[GF85] Mahadevan Ganapathi and Charles N. Fischer. Affix grammar driven code generation. ACM Transactions on Programming Languages and Systems, 7 (4):560–599, October 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4486.html.

Gini:1985:DWM

[GG85] Giuseppina C. Gini and Maria L.

Gini. Dealing with worldmodel-based programs. ACM Transactions on Programming Languages and Systems, 7(2):

334–347, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3479.html.

Gesbert:2015:LAD

[GGL15] Nils Gesbert, Pierre Genevès, and Nabil Layaïda. A logical approach to deciding semantic subtyping. ACM Transactions on Programming Languages and Systems, 38(1):3:1–3:??, October 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gorinova:2022:CIT

[GGSV22] Maria I. Gorinova, Andrew D. Gordon, Charles Sutton, and Matthijs Vákár. Conditional independence by typing. ACM Transactions on Programming Languages and Systems, 44(1): 4:1–4:54, March 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3490421.

Griswold:1980:AUP

[GH80] Ralph E. Griswold and David R. Hanson. An alternative to the use of patterns in string processing. ACM Transactions on Programming Languages and Systems, 2(2):153–172, April 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gerber:1997:SRT

[GH97] Richard Gerber and Seongsoo Hong. Slicing realtime programs for enhanced schedulability. ACM Transactions on Programming Languages and Systems, 19(3):525– 555, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-3/p525-gerber/.

GarciaDeLaBanda:1996:GAC

[GHB^{+}96] M. Garcia De La Banda, M. Hermenegildo, M. Bruynooghe, V. Dumortier, G. Janssens, and W. Simoens. Global analysis of constraint logic programs. ACM Transactions on Programming Languages and Systems, 18(5): 564–614, September 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
235456.html.

Grove:2019:FRR

[GHH^{+}19] David Grove, Sara S. Hamouda, Benjamin Herta, Arun Iyengar, Kiyokuni Kawachiya, Josh Milthorpe, Vijay Saraswat, Avraham Shinnar, Mikio Takeuchi, and Olivier Tardieu. Failure recovery in resilient X10. ACM Transactions on Programming Languages and Systems, 41(3): 15:1–15:??, July 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3332372.

Griswold:1981:GI

[GHK81] Ralph E. Griswold, David R. Hanson, and John T. Korb. Generators in Icon. ACM Transactions on Programming Languages and Systems, 3(2):144– 161, April 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ghosh:1993:ASP

[Gho93] Sukumar Ghosh. An alternative solution to a problem on self-stabilization. ACM Transactions on Programming Languages and Systems, 15(4):735– 742, September 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        155228.html.

Graham:1980:ICF

[GHR80] Susan L. Graham, Michael A. Harrison, and Walter L. Ruzzo. An improved context-free recognizer. ACM Transactions on Programming Languages and Systems, 2(3):415–462, July 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gallager:1983:DAM

[GHS83] R. G. Gallager, P. A. Humblet, and P. M. Spira. A distributed algorithm for minimumweight spanning trees. ACM Transactions on Programming Languages and Systems, 5(1): 66–77, January 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Giegerich:1983:FFD

[Gie83] Robert Giegerich. A formal framework for the derivation of machine-specific optimizers. ACM Transactions on Programming Languages and Systems, 5 (3):478–498, July 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gupta:1993:APE

[GJ93] Gopal Gupta and Bharat Jayaraman. Analysis of Or-parallel execution models. ACM Transactions on Programming Languages and Systems, 15(4):659– 680, September 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
155220.html.

Glenstrup:2005:TAS

[GJ05] Arne John Glenstrup and Neil D. Jones. Termination analysis and specialization-point insertion in offline partial evaluation. ACM Transactions on Programming Languages and Systems, 27(6):1147–1215, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Garlan:1994:TAM

[GKL94] David Garlan, Charles W.

Krueger, and Barbara S. Lerner.

TransformGen: Automating the maintenance of structureoriented environments. ACM Transactions on Programming Languages and Systems, 16(3): 727–774, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177697.html.

Gharat:2020:GPG

[GKM20] Pritam M. Gharat, Uday P. Khedker, and Alan Mycroft. Generalized points-to graphs: a precise and scalable abstraction for points-to analysis. ACM Transactions on Programming Languages and Systems, 42(2): 8:1–8:78, May 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3382092.

Gries:1980:APC

[GL80] David Gries and Gary Levin. Assignment and procedure call proof rules. ACM Transactions on Programming Languages and Systems, 2(4):564–579, October 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Grumberg:1994:MCM

[GL94] Orna Grumberg and David E. Long. Model checking and modular verification. ACM Transactions on Programming Languages and Systems, 16(3): 843–871, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177725.html.

Gavanelli:2005:DIK

[GLMM05] Marco Gavanelli, Evelina Lamma, Paola Mello, and Michela Milano. Dealing with incomplete knowledge on CLP(FD) variable domains. ACM Transactions on Programming Languages and Systems, 27(2):236–263, March 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Greenberg:1988:SEA

[GLO88] Albert G. Greenberg, Boris D.

Lubachevsky, and Andrew M. Odlyzko. Simple, efficient, asynchronous parallel algorithms for maximization. ACM Transactions on Programming Languages and Systems, 10(2):313– 337, April 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
42278.html.

Gottlieb:1983:BTE

[GLR83] Allan Gottlieb, Boris D. Lubachevsky, and Larry Rudolph. Basic techniques for the efficient coordination of very large numbers of cooperating sequential processors. ACM Transactions on Programming Languages and Systems, 5(2):164– 189, April 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ghezzi:1979:IP

[GM79] Carlo Ghezzi and Dino Mandrioli. Incremental parsing. ACM Transactions on Programming Languages and Systems, 1(1):58–70, July 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Greif:1981:SSW

[GM81] Irene Greif and Albert R. Meyer. Specifying the semantics of while programs: a tutorial and critique of a paper by Hoare and Lauer. ACM Transactions on Programming Languages and Systems, 3 (4):484–507, October 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ganty:2012:AVA

[GM12] Pierre Ganty and Rupak Majumdar. Algorithmic verification of asynchronous programs. ACM Transactions on Programming Languages and Systems, 34 (1):6:1–6:48, April 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gannon:1981:DAI

[GMH81] John Gannon, Paul McMullin, and Richard Hamlet. Data abstraction, implementation, specification, and testing. ACM Transactions on Programming Languages and Systems, 3(3): 211–223, July 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ghosh:1999:CME

[GMM99] Somnath Ghosh, Margaret Martonosi, and Sharad Malik. Cache miss equations: a compiler framework for analyzing and tuning memory behavior. ACM Transactions on Programming Languages and Systems, 21(4):703–746, July 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-4/
p703-ghosh/p703-ghosh.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-4/p703-ghosh/.

Gange:2021:FLZ

[GMN^{+}21] Graeme Gange, Zequn Ma, Jorge A. Navas, Peter Schachte, Harald Søndergaard, and Peter J. Stuckey. A fresh look at zones and octagons. ACM Transactions on Programming Languages and Systems, 43(3): 11:1–11:51, September 2021. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3457885.

Grant:2000:BCD

[GMP^{+}00] Brian Grant, Markus Mock, Matthai Philipose, Craig Chambers, and Susan J. Eggers. The benefits and costs of DyC’s runtime optimizations. ACM Transactions on Programming Languages and Systems, 22(5):932– 972, 2000. CODEN ATPSDT.

ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/2000-22-5/
                                        p932-grant/p932-grant.pdf;
                                         http://www.acm.org/pubs/
                                        citations/journals/toplas/
                                        2000-22-5/p932-grant/.

Giacobazzi:2024:AAI

[GMP24a] Roberto Giacobazzi, Isabella Mastroeni, and Elia Perantoni. Adversities in abstract interpretation — accommodating robustness by abstract interpretation. ACM Transactions on Programming Languages and Systems, 46(2):5:1–5:??, June 2024. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649309.

Giallorenzo:2024:COO

[GMP24b] Saverio Giallorenzo, Fabrizio Montesi, and Marco Peressotti. Choral: Object-oriented choreographic programming. ACM Transactions on Programming Languages and Systems, 46(1): 1:1–1:??, March 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3632398.

Grossman:2000:STA

[GMZ00] Dan Grossman, Greg Morrisett, and Steve Zdancewic. Syntactic type abstraction. ACM Transactions on Programming Languages and Systems, 22(6):1037– 1080, 2000. CODEN ATPSDT.

ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-6/
p1037-grossman/p1037-grossman.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-6/p1037-grossman/.

Gange:2015:IAM

[GNS^{+}15] Graeme Gange, Jorge A.

Navas, Peter Schachte, Harald Søndergaard, and Peter J. Stuckey. Interval analysis and machine arithmetic: Why signedness ignorance is bliss. ACM Transactions on Programming Languages and Systems, 37 (1):1:1–1:??, January 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gomard:1992:SAP

[Gom92] Carsten K. Gomard. A selfapplicable partial evaluator for the lambda calculus: Correctness and pragmatics. ACM Transactions on Programming Languages and Systems, 14(2): 147–172, April 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
128864.html.

Gorlatch:2004:SRC

[Gor04] Sergei Gorlatch. Send-receive considered harmful: Myths and realities of message passing. ACM Transactions on Programming Languages and Systems, 26 (1):47–56, January 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gordon:2021:PIS

[Gor21] Colin S. Gordon. Polymorphic iterable sequential effect systems. ACM Transactions on Programming Languages and Systems, 43(1): 4:1–4:79, April 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3450272.

Grit:1981:DIT

[GP81] Dale H. Grit and Rex L. Page. Deleting irrelevant tasks in an expression-oriented multiprocessor system. ACM Transactions on Programming Languages and Systems, 3(1):49–59, January 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Girkar:1995:ETL

[GP95] Milind Girkar and Constantine D. Polychronopoulos. Extracting task-level parallelism. ACM Transactions on Programming Languages and Systems, 17(4):600–634, July 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        210189.html.

Gupta:2001:PEP

[GPA^{+}01] Gopal Gupta, Enrico Pontelli, Khayri A. M. Ali, Mats Carlsson, and Manuel V. Hermenegildo. Parallel execution of Prolog programs: a survey. ACM Transactions on Programming Languages and Systems, 23 (4):472–602, July 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gal:2008:JBV

[GPF08] Andreas Gal, Christian W.

Probst, and Michael Franz. Java bytecode verification via static single assignment form. ACM Transactions on Programming Languages and Systems, 30(4): 21:1–21:21, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Grothoff:2007:EOC

[GPV07] Christian Grothoff, Jens Palsberg, and Jan Vitek. Encapsulating objects with confined types. ACM Transactions on Programming Languages and Systems, 29(6):32:1–32:41, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gil:2008:TDB

[GPWZ08] Joseph (Yossi) Gil, William Pugh, Grant E. Weddell, and Yoav Zibin. Two-dimensional bidirectional object layout. ACM Transactions on Programming Languages and Systems, 30(5):28:1–28:38, August 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gries:1979:SEB

[Gri79] David Gries. Is sometimes ever better than alway? ACM Transactions on Programming Languages and Systems, 1(2): 258–265, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Griswold:1982:EEI

[Gri82] Ralph E. Griswold. The evaluation of expressions in Icon. ACM Transactions on Programming Languages and Systems, 4 (4):563–584, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Grossman:2006:QTI

[Gro06]  Dan Grossman.  Quantified

types in an imperative language. ACM Transactions on Programming Languages and Systems, 28 (3):429–475, May 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Giesl:2011:ATP

[GRSK^{+}11] Jürgen Giesl, Matthias Raffelsieper, Peter Schneider-Kamp, Stephan Swiderski, and René Thiemann. Automated termination proofs for Haskell by term rewriting. ACM Transactions on Programming Languages and Systems, 33(2):7:1–7:39, January 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Giacobazzi:1998:LMR

[GS98] Roberto Giacobazzi and Francesca Scozzari. A logical model for relational abstract domains. ACM Transactions on Programming Languages and Systems, 20 (5):1067–1109, September 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-5/
p1067-giacobazzi/.

Gloy:1999:PPU

[GS99] Nikolas Gloy and Michael D.

Smith. Procedure placement using temporal-ordering information. ACM Transactions on Programming Languages and Systems, 21(5):977–1027, September 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-5/
p977-gloy/p977-gloy.pdf;
http://www.acm.org/pubs/citations/
journals/toplas/1999-21-5/
p977-gloy/.

Gawlitza:2011:SSR

[GS11] Thomas Martin Gawlitza and Helmut Seidl. Solving systems of rational equations through strategy iteration. ACM Transactions on Programming Languages and Systems, 33(3):11:1– 11:48, April 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gupta:1994:ERA

[GSO94] Rajiv Gupta, Mary Lou Soffa, and Denise Ombres. Efficient register allocation via coloring using clique separators. ACM Transactions on Programming Languages and Systems, 16(3): 370–386, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177499.html.

Grimmer:2018:CLI

[GSS^{+}18] Matthias Grimmer, Roland Schatz, Chris Seaton, Thomas Würthinger, and Mikel Luján. Cross-language interoperability in a multi-language runtime. ACM Transactions on Programming Languages and Systems, 40 (2):8:1–8:??, June 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gerlek:1995:BIV

[GSW95] Michael P. Gerlek, Eric Stoltz, and Michael Wolfe. Beyond induction variables: Detecting and classifying sequences using a demand-driven SSA form. ACM Transactions on Programming Languages and Systems, 17(1): 85–122, January 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201003.html.

Garcia:2014:FTO

[GTWA14] Ronald Garcia, Éric Tanter, Roger Wolff, and Jonathan Aldrich. Foundations of typestate-oriented programming. ACM Transactions on Programming Languages and Systems, 36(4):12:1–12:??, October 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gudeman:1992:DSG

[Gud92] David A. Gudeman. Denotational semantics of a goaldirected language. ACM Transactions on Programming Languages and Systems, 14(1):107– 125, January 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
104659.html.

Grosser:2015:PAG

[GVC15] Tobias Grosser, Sven Verdoolaege, and Albert Cohen. Polyhedral AST generation is more than scanning polyhedra. ACM Transactions on Programming Languages and Systems, 37(4):12:1–12:??, August 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gudjonsson:1999:CTM

[GW99] Gudjón Gudjónsson and William H.

Winsborough. Compile-time memory reuse in logic programming languages through update in place. ACM Transactions on Programming Languages and Systems, 21(3):430–501, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-3/

                                        p430-gudjonsson/p430-gudjonsson.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-3/p430-gudjonsson/.

Glesner:2004:NSS

[GZ04] Sabine Glesner and Wolf Zimmermann. Natural semantics as a static program analysis framework. ACM Transactions on Programming Languages and Systems, 26(3):510– 577, May 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gil:2005:EST

[GZ05] Joseph (Yossi) Gil and Yoav Zibin. Efficient subtyping tests with PQ-encoding. ACM Transactions on Programming Languages and Systems, 27(5):819– 856, September 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Gil:2007:EDD

[GZ07] Joseph (Yossi) Gil and Yoav Zibin. Efficient dynamic dispatching with type slicing. ACM Transactions on Programming Languages and Systems, 30(1): 5:1–5:53, November 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hoffmann:2012:MAR

[HAH12] Jan Hoffmann, Klaus Aehlig, and Martin Hofmann. Multivariate amortized resource analysis. ACM Transactions on Programming Languages and Systems, 34(3):14:1–14:62, October 2012.

CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hailperin:1998:COC

[Hai98] Max Hailperin. Cost-optimal code motion. ACM Transactions on Programming Languages and Systems, 20(6):1297– 1322, November 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-6/
p1297-hailperin/.

Hailperin:2005:CCC

[Hai05]  Max Hailperin.  Comparing

conservative coalescing criteria. ACM Transactions on Programming Languages and Systems, 27 (3):571–582, May 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Halstead:1985:MLC

[Hal85] Robert H. Halstead, Jr. Multilisp: A Language for Concurrent Symbolic Computation. ACM Transactions on Programming Languages and Systems, 7 (4):501–538, October 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4478.html.

Hall:2005:IPA

[HAM^{+}05] Mary W. Hall, Saman P. Amarasinghe, Brian R. Murphy, Shih-Wei Liao, and Monica S.

Lam. Interprocedural parallelization analysis in SUIF. ACM Transactions on Programming Languages and Systems, 27(4): 662–731, July 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hansen:1981:CMI

[Han81a] Wilfred J. Hansen. A cost model for the internal organization of B^{+}-tree nodes. ACM Transactions on Programming Languages and Systems, 3(4): 508–532, October 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hanson:1981:APP

[Han81b] David R. Hanson. Algorithm 568. PDS — a portable directory system. ACM Transactions on Programming Languages and Systems, 3(2):162– 167, April 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://doi.acm.org/10.
1145/357133.357137.

Hansen:1992:SRF

[Han92]  Wilfred J. Hansen.  Subse-

quence references: First-class values for substrings. ACM Transactions on Programming Languages and Systems, 14(4): 471–489, October 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
133234.html.

Hannan:1994:OSD

[Han94]  John Hannan.  Operational

semantics-directed compilers and machine architectures. ACM Transactions on Programming Languages and Systems, 16(4): 1215–1247, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
183458.html.

Harel:1980:PNA

[Har80] David Harel. And/Or programs: a new approach to structured programming. ACM Transactions on Programming Languages and Systems, 2(1): 1–17, January 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hauser:1996:HFP

[Hau96] John R. Hauser. Handling floating-point exceptions in numeric programs. ACM Transactions on Programming Languages and Systems, 18(2):139– 174, March 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
227701.html.

Havlak:1997:NRI

[Hav97] Paul Havlak. Nesting of reducible and irreducible loops. ACM Transactions on Programming Languages and Systems, 19(4):557–567, July 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-4/
p557-havlak/.

Hind:1999:IPA

[HBCC99] Michael Hind, Michael Burke, Paul Carini, and Jong-Deok Choi. Interprocedural pointer alias analysis. ACM Transactions on Programming Languages and Systems, 21(4): 848–894, July 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/articles/journals/toplas/
1999-21-4/p848-hind/p848-
hind.pdf;   http://www.acm.
org/pubs/citations/journals/
toplas/1999-21-4/p848-hind/

.

Harman:2009:DCS

[HBG^{+}09] Mark Harman, David Binkley, Keith Gallagher, Nicolas Gold, and Jens Krinke. Dependence clusters in source code. ACM Transactions on Programming Languages and Systems, 32(1): 1:1–1:33, October 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hassen:1998:TDP

[HBJ98] Saniya Ben Hassen, Henri E. Bal, and Ceriel J. H. Jacobs. A task- and data-parallel programming language based on shared objects. ACM Transactions on Programming Languages and Systems, 20(6):1131– 1170, November 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-6/
p1131-ben_hassen/.

Hertz:2006:GOL

[HBM^{+}06] Matthew Hertz, Stephen M.

Blackburn, J. Eliot B. Moss, Kathryn S. McKinley, and Darko Stefanović. Generating object lifetime traces with Merlin. ACM Transactions on Programming Languages and Systems, 28(3):476–516, May 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hecker:2022:TSC

[HBS22] Martin Hecker, Simon Bischof, and Gregor Snelting. On timesensitive control dependencies. ACM Transactions on Programming Languages and Systems, 44 (1):2:1–2:37, March 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3486003.

Hickey:1992:CAM

[HCHP92] Timothy J. Hickey, Jacques Cohen, Hitofumi Hotta, and Thierry PetitJean. Computerassisted microanalysis of parallel programs. ACM Transactions on Programming Languages and Systems, 14(1):54– 106, January 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        126699.html.

Huang:2010:DBR

[HCS10] Yuqiang Huang, Bruce R. Childers, and Mary Lou Soffa. Detecting bugs in register allocation. ACM Transactions on Programming Languages and Systems, 32(4):15:1–15:36, April 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Holt:1982:ISS

[HCW82] Richard C. Holt, J. R. Cordy, and David B. Wortman. An introduction to S/SL: Syntax/ Semantic Language. ACM Transactions on Programming Languages and Systems, 4(2): 149–178, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hirzel:2002:UTL

[HDH02] Martin Hirzel, Amer Diwan, and Johannes Henkel. On the usefulness of type and liveness accuracy for garbage collection and leak detection. ACM Transactions on Programming Languages and Systems, 24(6):593– 624, November 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hennessy:1982:SDO

[Hen82] John L. Hennessy. Symbolic debugging of optimized code. ACM Transactions on Programming Languages and Systems, 4 (3):323–344, July 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remark [CM93].

Henderson:1983:TCL

[Hen83] Peter B. Henderson. Technical correspondence: On LaLonde and des Rivieres’ “Handling Operator Precedence in Arithmetic Expressions with Tree Transformations”. ACM Transactions on Programming Languages and Systems, 5(1):122, January 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [LdR81, LaL83].

Hennessy:1986:PSS

[Hen86] Matthew Hennessy. Proving systolic systems correct. ACM Transactions on Programming Languages and Systems, 8(3): 344–387, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5999.html.

Henglein:1993:TIP

[Hen93] Fritz Henglein. Type inference with polymorphic recursion. ACM Transactions on Programming Languages and Systems, 15 (2):253–289, April 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
169692.html.

Herlihy:1991:WFS

[Her91]  Maurice Herlihy.  Wait-free

synchronization. ACM Transactions on Programming Languages and Systems, 13(1):124– 149, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        102808.html.

Herlihy:1993:MIH

[Her93] Maurice Herlihy. A methodology for implementing highly concurrent data objects. ACM Transactions on Programming Languages and Systems, 15(5): 745–770, November 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        161469.html.

Hesselink:1988:MAN

[Hes88] W. H. Hesselink. A mathematical approach to nondeterminism in data types. ACM Transactions on Programming Languages and Systems, 10(1): 87–117, January 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        42194.html.

Haynes:1987:ECP

[HF87] Christopher T. Haynes and Daniel P. Friedman. Embedding continuations in procedural objects. ACM Transactions on Programming Languages and Systems, 9(4):582– 598, October 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
30392.html.

Hosoya:2009:PPX

[HFC09] Haruo Hosoya, Alain Frisch, and Giuseppe Castagna. Parametric polymorphism for XML. ACM Transactions on Programming Languages and Systems, 32(1): 2:1–2:56, October 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hennessy:1983:PCO

[HG83] John L. Hennessy and Thomas Gross. Postpass code optimization of pipeline constraints. ACM Transactions on Programming Languages and Systems, 5 (3):422–448, July 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hall:1996:TCH

[HHPW96] Cordelia V. Hall, Kevin Hammond, Simon L. Peyton Jones, and Philip L. Wadler. Type classes in Haskell. ACM Transactions on Programming Languages and Systems, 18(2):109– 138, March 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
227700.html.

Hilfinger:1988:APD

[Hil88] Paul N. Hilfinger. An Ada package for dimensional analysis. ACM Transactions on Programming Languages and Systems, 10 (2):189–203, April 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        42346.html.

Hu:1997:FDE

[HIT97] Zhenjiang Hu, Hideya Iwasaki, and Masato Takechi. Formal derivation of efficient parallel programs by construction of list homomorphisms. ACM Transactions on Programming Languages and Systems, 19(3): 444–461, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1997-19-3/p444-hu/.

Heering:1985:TMP

[HK85] Jan Heering and Paul Klint. Towards monolingual programming environments. ACM Transactions on Programming Languages and Systems, 7(2): 183–213, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        3321.html.

Henzinger:2007:EMP

[HK07] Thomas A. Henzinger and Christoph M. Kirsch. The embedded machine: Predictable, portable real-time code. ACM Transactions on Programming Languages and Systems, 29(6): 33:1–33:29, October 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Haines:1994:CFC

[HKMN94] N. Haines, D. Kindred, J. G.

Morrisett, and S. M. Nettles. Composing first-class transactions. ACM Transactions on Programming Languages and Systems, 16(6):1719– 1736, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
197346.html.

Heering:1992:IGL

[HKR92] J. Heering, P. Klint, and J. Rekers. Incremental generation of lexical scanners. ACM Transactions on Programming Languages and Systems, 14(4):490– 520, October 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
133240.html.

Heering:1994:LIP

[HKR94] J. Heering, P. Klint, and J. Rekers. Lazy and incremental program generation. ACM Transactions on Programming Languages and Systems, 16(3):1010– 1023, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        177750.html.

Herlihy:1982:VTM

[HL82] Maurice P. Herlihy and Barbara Liskov. A value transmission method for abstract data types. ACM Transactions on Programming Languages and Systems, 4 (4):527–551, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hirschowitz:2005:MMC

[HL05] Tom Hirschowitz and Xavier Leroy. Mixin modules in a callby-value setting. ACM Transactions on Programming Languages and Systems, 27(5):857– 881, September 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Haslbeck:2022:FDM

[HL22] Maximilian P. L. Haslbeck and Peter Lammich. For a few dollars more: Verified fine-grained algorithm analysis down to LLVM. ACM Transactions on Programming Languages and Systems, 44 (3):14:1–14:36, September 2022. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3486169.

Hague:2019:CMC

[HLH19] Matthew Hague, Anthony W. Lin, and Chih-Duo Hong. CSS minification via constraint solving. ACM Transactions on Programming Languages and Systems, 41(2):12:1–12:??, June 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https://dl.

acm.org/ft_gateway.cfm?id=
3310337.

Hu:2023:TPO

[HLH^{+}23] Jingmei Hu, Eric Lu, David A.

Holland, Ming Kawaguchi, Stephen Chong, and Margo Seltzer. Towards porting operating systems with program synthesis. ACM Transactions on Programming Languages and Systems, 45(1): 2:1–2:??, March 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3563943.

Hull:1984:CSP

[HM84] M. Elizabeth C. Hull and R. M.

McKeag. Communicating sequential processes for centralized and distributed operating system design. ACM Transactions on Programming Languages and Systems, 6(2):175– 191, April 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Harper:1993:TSS

[HM93] Robert Harper and John C.

Mitchell. On the type structure of standard ML. ACM Transactions on Programming Languages and Systems, 15(2): 211–252, April 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

                                        pubs/toc/Abstracts/0164-0925/

169696.html. Earlier version appears as “The Essence of ML” in Proc. 15th ACM Symp. on Principles of Programming Languages, 1988, pp. 28–46.

Hamlen:2006:CCE

[HMS06] Kevin W. Hamlen, Greg Morrisett, and Fred B. Schneider. Computability classes for enforcement mechanisms. ACM Transactions on Programming Languages and Systems, 28(1): 175–205, January 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Haas:2024:LPM

[HMY^{+}24] Julian Haas, Ragnar Mogk, Elena Yanakieva, Annette Bieniusa, and Mira Mezini. LoRe: a programming model for verifiably safe local-first software. ACM Transactions on Programming Languages and Systems, 46 (1):2:1–2:??, March 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3633769.

Hicks:2005:DSU

[HN05] Michael Hicks and Scott Nettles. Dynamic software updating. ACM Transactions on Programming Languages and Systems, 27(6):1049–1096, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hoffman:1982:PE

[HO82] Christoph M. Hoffman and Michael J. O’Donnell. Programming with equations. ACM Transactions on Programming Languages and Systems, 4(1): 83–112, January 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Higuchi:2007:STS

[HO07] Tomoyuki Higuchi and Atsushi Ohori. A static type system for JVM access control. ACM Transactions on Programming Languages and Systems, 29(1): 4:1–4:42, January 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hobson:1984:DEE

[Hob84] Richard F. Hobson. A directly executable encoding for APL. ACM Transactions on Programming Languages and Systems, 6 (3):314–332, July 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Holt:1987:DDC

[Hol87] Richard C. Holt. Data descriptors: a compile-time model of data and addressing. ACM Transactions on Programming Languages and Systems, 9(3): 367–389, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
24051.html.

Horwitz:1997:PFI

[Hor97] Susan Horwitz. Precise flowinsensitive may-alias analysis is NP-Hard. ACM Transactions on Programming Languages and Systems, 19(1):1– 6, January 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-1/p1-horwitz/.

Howden:1980:ASV

[How80] W. E. Howden. Applicability of software validation techniques to scientific programs. ACM Transactions on Programming Languages and Systems, 2 (3):307–320, July 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Heo:2018:ASA

[HOYY18] Kihong Heo, Hakjoo Oh, Hongseok Yang, and Kwangkeun Yi. Adaptive static analysis via learning with Bayesian optimization. ACM Transactions on Programming Languages and Systems, 40(4):14:1– 14:??, December 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3121135.

Haghighat:1996:SAP

[HP96] Mohammad R. Haghighat and Constantine D. Polychronopoulos. Symbolic analysis for parallelizing compilers. ACM Transactions on Programming Languages and Systems, 18(4): 477–518, July 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
233568.html.

Hu:2024:LAI

[HP24] Jason Z. S. Hu and Brigitte Pientka. A layered approach to intensional analysis in type theory. ACM Transactions on Programming Languages and Systems, 46 (4):15:1–15:??, December 2024. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3707203.

Hermenegildo:2000:IAC

[HPMS00] Manuel Hermenegildo, German Puebla, Kim Marriott, and Peter J. Stuckey. Incremental analysis of constraint logic programs. ACM Transactions on Programming Languages and Systems, 22 (2):187–223, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-2/p187-hermenegildo/

.

Horwitz:1989:INV

[HPR89] Susan Horwitz, Jan Prins, and Thomas Reps. Integrating noninterfering versions of programs. ACM Transactions on Programming Languages and Systems, 11(3):345–387, July 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        65980.html.

Henzinger:2002:AGR

[HQRT02] Thomas A. Henzinger, Shaz Qadeer, Sriram K. Rajamani, and Serdar Tasiran. An assumeguarantee rule for checking simulation. ACM Transactions on Programming Languages and Systems, 24(1):51–64, January 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hennessy:2002:IFV

[HR02] Matthew Hennessy and James Riely. Information flow vs. resource access in the asynchronous pi-calculus. ACM Transactions on Programming Languages and Systems, 24(5): 566–591, September 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Horwitz:1990:ISU

[HRB90] Susan Horwitz, Thomas Reps, and David Binkley. Interprocedural slicing using dependence graphs. ACM Transactions on Programming Languages and Systems, 12(1):26– 60, January 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        77608.html.

Hietala:2023:VOQ

[HRL^{+}23] Kesha Hietala, Robert Rand, Liyi Li, Shih-Han Hung, Xiaodi Wu, and Michael Hicks. A verified optimizer for quantum circuits. ACM Transactions on Programming Languages and Systems, 45(3):18:1– 18:??, September 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604630.

Haselwarter:2023:SFF

[HRV^{+}23] Philipp G. Haselwarter, Exequiel Rivas, Antoine Van Muylder, Théo Winterhalter, Carmine Abate, Nikolaj Sidorenco, Catalin Hritcu, Kenji Maillard, and Bas Spitters. SSProve: a foundational framework for modular cryptographic proofs in Coq. ACM Transactions on Programming Languages and Systems, 45(3):15:1– 15:??, September 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3594735.

Harrold:1994:ECI

[HS94] Mary Jean Harrold and Mary Lou Soffa. Efficient computation of interprocedural definitionuse chains. ACM Transactions on Programming Languages and Systems, 16(2):175– 204, March 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174663.html.

Huang:2011:MSS

[HS11] Shan Shan Huang and Yannis Smaragdakis. Morphing: Structurally shaping a class by reflecting on others. ACM Transactions on Programming Languages and Systems, 33(2):6:1– 6:44, January 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hu:2022:SPS

[HSD22] Qinheping Hu, Rishabh Singh, and Loris D’Antoni. Solving program sketches with large integer values. ACM Transactions on Programming Languages and Systems, 44(2): 9:1–9:28, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3532849.

Hirzel:2017:SEL

[HSG17] Martin Hirzel, Scott Schneider, and Bugra Gedik. SPL: an extensible language for distributed stream processing. ACM Transactions on Programming Languages and Systems, 39(1):5:1– 5:??, March 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hart:1983:TPC

[HSP83] Sergiu Hart, Micha Sharir, and Amir Pnueli. Termination of probabilistic concurrent program. ACM Transactions on Programming Languages and Systems, 5(3):356–380, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hayden:2014:KEG

[HSS^{+}14] Christopher M. Hayden, Karla Saur, Edward K. Smith, Michael Hicks, and Jeffrey S. Foster. Kitsune: Efficient, general-purpose dynamic software updating for C. ACM Transactions on Programming Languages and Systems, 36(4):13:1–13:??, October 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Horwitz:1986:GEE

[HT86] Susan Horwitz and Tim Teitelbaum. Generating editing environments based on relations and attributes. ACM Transactions on Programming Languages and Systems, 8(4):577– 608, October 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
6512.html.

Helsen:2004:PSM

[HT04] Simon Helsen and Peter Thiemann. Polymorphic specialization for ML. ACM Transactions on Programming Languages and Systems, 26(4):652– 701, July 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Holzle:1996:RRP

[HU96] Urs Hölzle and David Ungar. Reconciling responsiveness with performance in pure object-oriented languages. ACM Transactions on Programming Languages and Systems, 18(4): 355–400, July 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/toplas/
                                        233562.html.

Huang:1990:DDD

[Hua90] Shing-Tsaan Huang. A distributed deadlock detection algorithm for CSP-like communication. ACM Transactions on Programming Languages and Systems, 12(1):102– 122, January 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        77611.html.

Huang:1993:LEU

[Hua93] Shing-Tsaan Huang. Leader election in uniform rings. ACM Transactions on Programming Languages and Systems, 15(3): 563–573, July 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174161.html.

Hudson:1991:IAE

[Hud91] Scott E. Hudson. Incremental attribute evaluation: a flexible algorithm for lazy update. ACM Transactions on Programming Languages and Systems, 13(3):315–341, July 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
117012.html.

Haridi:1999:ELV

[HVB^{+}99] Seif Haridi, Peter Van Roy, Per Brand, Michael Mehl, Ralf Scheidhauer, and Gert Smolka. Efficient logic variables for distributed computing. ACM Transactions on Programming Languages and Systems, 21(3):569–626, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-3/
p569-haridi/p569-haridi.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-3/p569-haridi/.

Hirzel:2007:FOP

[HVDH07] Martin Hirzel, Daniel Von Dincklage, Amer Diwan, and Michael Hind. Fast online pointer analysis. ACM Transactions on Programming Languages and Systems, 29(2):11:1– 11:55, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Hosoya:2005:RET

[HVP05] Haruo Hosoya, Jérôme Vouillon, and Benjamin C. Pierce. Regular expression types for XML.

ACM Transactions on Programming Languages and Systems, 27 (1):46–90, January 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Holt:1982:MIE

[HW82] Richard C. Holt and David B. Wortman. A model for implementing Euclid modules and prototypes. ACM Transactions on Programming Languages and Systems, 4(4):552–562, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Herlihy:1990:LCC

[HW90] Maurice P. Herlihy and Jeannette M. Wing. Linearizability: a correctness condition for concurrent objects. ACM Transactions on Programming Languages and Systems, 12(3): 463–492, July 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        78972.html.

Hudak:1991:CIE

[HY91] Paul Hudak and Jonathan Young. Collecting interpretations of expressions. ACM Transactions on Programming Languages and Systems, 13(2): 269–290, April 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        103139.html.

Honda:2007:UTS

[HY07] Kohei Honda and Nobuko Yoshida. A uniform type structure for secure information flow. ACM Transactions on Programming Languages and Systems, 29 (6):31:1–31:100, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Igarashi:2005:RUA

[IK05] Atsushi Igarashi and Naoki Kobayashi. Resource usage analysis. ACM Transactions on Programming Languages and Systems, 27(2):264–313, March 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Igarashi:2001:FJM

[IPW01] Atsushi Igarashi, Benjamin C.

Pierce, and Philip Wadler. Featherweight Java: a minimal core calculus for Java and GJ. ACM Transactions on Programming Languages and Systems, 23(3):396–450, May 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ish-Shalom:2022:RCB

[ISIRS22] Oren Ish-Shalom, Shachar Itzhaky, Noam Rinetzky, and Sharon Shoham. Runtime complexity bounds using squeezers. ACM Transactions on Programming Languages and Systems, 44(3): 17:1–17:36, September 2022. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527632.

Inoue:1988:AFP

[ISY88] Katsuro Inoue, Hiroyuki Seki, and Hikaru Yagi. Analysis of functional programs to detect run-time garbage cells. ACM Transactions on Programming Languages and Systems, 10(4): 555–578, October 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
48025.html.

Inverso:2022:BVM

[ITF^{+}22] Omar Inverso, Ermenegildo Tomasco, Bernd Fischer, Salvatore La Torre, and Gennaro Parlato. Bounded verification of multi-threaded programs via lazy sequentialization. ACM Transactions on Programming Languages and Systems, 44(1): 1:1–1:50, March 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3478536.

Igarashi:2006:VPT

[IV06] Atsushi Igarashi and Mirko Viroli. Variant parametric types: a flexible subtyping scheme for generics. ACM Transactions on Programming Languages and Systems, 28(5):795– 847, September 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Iverson:1979:O

[Ive79] Kenneth E. Iverson. Operators.

ACM Transactions on Programming Languages and Systems, 1 (2):161–176, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jagannathan:1994:MBB

[Jag94]  Suresh Jagannathan.  Met-

alevel building blocks for modular systems. ACM Transactions on Programming Languages and Systems, 16(3):456– 492, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177578.html.

Jay:2004:PC

[Jay04] C. Barry Jay. The pattern calculus. ACM Transactions on Programming Languages and Systems, 26(6):911–937, November 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joisha:2006:AAS

[JB06] Pramod G. Joisha and Prithviraj Banerjee. An algebraic array shape inference system for MATLAB(R). ACM Transactions on Programming Languages and Systems, 28(5):848– 907, September 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jangda:2020:EFT

[JB20] Abhinav Jangda and Uday Bondhugula. An effective fusion and tile size model for PolyMage. ACM Transactions on Programming Languages and Systems, 42 (3):12:1–12:27, December 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3404846.

Jacobs:2018:MTV

[JBK18] Bart Jacobs, Dragan Bosnacki, and Ruurd Kuiper. Modular termination verification of singlethreaded and multithreaded programs. ACM Transactions on Programming Languages and Systems, 40(3):12:1–12:??, August 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Janssen:1997:MGR

[JC97] Johan Janssen and Henk Corporaal. Making graphs reducible with controlled node splitting. ACM Transactions on Programming Languages and Systems, 19 (6):1031–1052, November 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-6/
p1031-janssen/.

Jacek:2019:OCW

[JCMM19] Nicholas Jacek, Meng-Chieh Chiu, Benjamin M. Marlin, and J. Eliot B. Moss. Optimal choice of when to garbage collect. ACM Transactions on Programming Languages and Systems, 41(1): 3:1–3:??, March 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jefferson:1985:VT

[Jef85] David R. Jefferson. Virtual time.

ACM Transactions on Programming Languages and Systems, 7 (3):404–425, July 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3988.html.

Jeffery:2003:GLS

[Jef03] Clinton L. Jeffery. Generating LR syntax error messages from examples. ACM Transactions on Programming Languages and Systems, 25(5):631– 640, September 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jensen:1997:DPA

[Jen97] Thomas Jensen. Disjunctive program analysis for algebraic data types. ACM Transactions on Programming Languages and Systems, 19(5):751– 803, September 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-5/
p751-jensen/.

Juelich:1981:CAS

[JF81] Otto C. Juelich and Clinton R. Foulk. Compilation of acyclic smooth programs for parallel execution. ACM Transactions on Programming Languages and Systems, 3(1):24–48, January 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jeon:2019:MLA

[JJCO19] Minseok Jeon, Sehun Jeong, Sungdeok Cha, and Hakjoo Oh. A machine-learning algorithm with disjunctive model for datadriven program analysis. ACM Transactions on Programming Languages and Systems, 41(2): 13:1–13:??, June 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3293607.

Jackson:1998:IFM

[JJD98] Daniel Jackson, Somesh Jha, and Craig A. Damon. Isomorphfree model enumeration: a new method for checking relational specifications. ACM Transactions on Programming Languages and Systems, 20(2): 302–343, March 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-2/
                                        p302-jackson/.

Jia:2023:DBP

[JKT23] Xiaodong Jia, Ashish Kumar, and Gang Tan. A derivativebased parser generator for visibly pushdown grammars. ACM Transactions on Programming Languages and Systems, 45(2): 9:1–9:??, June 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591472.

Jimenez:2002:RTN

[JLF02] Marta Jiménez, José M. Llabería, and Agustín Fernández. Register tiling in nonrectangular iteration spaces. ACM Transactions on Programming Languages and Systems, 24(4):409–453, July 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jagannathan:2014:ARV

[JLP^{+}14] Suresh Jagannathan, Vincent Laporte, Gustavo Petri, David Pichardie, and Jan Vitek. Atomicity refinement for verified compilation. ACM Transactions on Programming Languages and Systems, 36(2):6:1–6:??, July 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jeannet:2010:RAI

[JLRS10] Bertrand Jeannet, Alexey Loginov, Thomas Reps, and Mooly Sagiv. A relational approach to interprocedural shape analysis. ACM Transactions on Programming Languages and Systems, 32 (2):5:1–5:52, January 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jaffar:1992:CLS

[JMSY92] Joxan Jaffar, Spiro Michaylov, Peter J. Stuckey, and Roland H. C. Yap. The CLP(R) language and system. ACM Transactions on Programming Languages and Systems, 14(3): 339–395, July 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        129398.html.

Jeffrey:2010:ESA

[JNGG10] Dennis Jeffrey, Vijay Nagarajan, Rajiv Gupta, and Neelam Gupta. Execution suppression: an automated iterative technique for locating memory errors. ACM Transactions on Programming Languages and Systems, 32(5):17:1– 17:36, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joshi:2006:DPA

[JNZ06] Rajeev Joshi, Greg Nelson, and Yunhong Zhou. Denali: a practical algorithm for generating optimal code. ACM Transactions on Programming Languages and Systems, 28(6):967–989, November 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jones:1983:TST

[Jon83] C. B. Jones. Tentative steps toward a development method for interfering programs. ACM Transactions on Programming Languages and Systems, 5(4): 596–619, October 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jones:1990:EEC

[Jon90]  Larry G. Jones.  Efficient

evaluation of circular attribute grammars. ACM Transactions on Programming Languages and Systems, 12(3): 429–462, July 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
78971.html.

Jonsson:1994:CSV

[Jon94] Bengt Jonsson. Compositional specification and verification of distributed systems. ACM Transactions on Programming Languages and Systems, 16(2): 259–303, March 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
174665.html.

Jazayeri:1981:SES

[JP81] Medhi Jazayeri and Diane Pozefsky. Space-efficient storage management in an attribute grammar evaluator. ACM Transactions on Programming Languages and Systems, 3(4):388– 404, October 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jourdan:2017:SPC

[JP17] Jacques-Henri Jourdan and François Pottier. A simple, possibly correct LR parser for C11. ACM Transactions on Programming Languages and Systems, 39 (4):14:1–14:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Jagadeesan:1991:FAS

[JPP91] Radha Jagadeesan, Keshav Pingali, and Prakash Panangaden. A fully abstract semantics for a first-order functional language with logic variables. ACM Transactions on Programming Languages and Systems, 13(4): 577–625, October 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        115371.html.

Jacobs:2008:PMC

[JPS^{+}08] Bart Jacobs, Frank Piessens, Jan Smans, K. Rustan M. Leino, and Wolfram Schulte. A programming model for concurrent object-oriented programs. ACM Transactions on Programming Languages and Systems, 31(1): 1:1–1:48, December 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Joung:1994:CFO

[JS94] Yuh-Jzer Joung and Scott A.

Smolka. Coordinating firstorder multiparty interactions. ACM Transactions on Programming Languages and Systems, 16(3):954–985, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177739.html.

Joisha:2012:TTE

[JSB^{+}12] Pramod G. Joisha, Robert S.

Schreiber, Prithviraj Banerjee, Hans-J. Boehm, and Dhruva R. Chakrabarti. On a technique for transparently empowering classical compiler optimizations on multithreaded code. ACM Transactions on Programming Languages and Systems, 34(2): 9:1–9:42, June 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Juan:1998:CVC

[JTM98] Eric Y. T. Juan, Jeffrey J. P.

Tsai, and Tadao Murata. Compositional verification of concurrent systems using Petri-net-based condensation rules. ACM Transactions on Programming Languages and Systems, 20(5): 917–979, September 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-5/
p917-juan/.

Jakobs:2017:PPF

[JW17] Marie-Christine Jakobs and Heike Wehrheim. Programs from proofs: a framework for the safe execution of untrusted software. ACM Transactions on Programming Languages and Systems, 39(2):7:1–7:??, May 2017. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ji:2024:DBS

[JZX^{+}24] Ruyi Ji, Yuwei Zhao, Yingfei Xiong, Di Wang, Lu Zhang, and Zhenjiang Hu. Decompositionbased synthesis for applying divide-and-conquer-like algorithmic paradigms. ACM Transactions on Programming Languages and Systems, 46(2): 8:1–8:??, June 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3648440.

Kaiser:1989:IDS

[Kai89]  Gail E. Kaiser.  Incremen-

tal dynamic semantics for language-based programming environments. ACM Transactions on Programming Languages and Systems, 11(2):169– 193, April 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        63400.html.

Kamin:1983:FDT

[Kam83] Samuel Kamin. Final data types and their specification. ACM Transactions on Programming Languages and Systems, 5(1): 97–121, January 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Karp:1984:PFF

[Kar84] Richard Alan Karp. Proving failure-free properties of concurrent systems using temporal logic. ACM Transactions on Programming Languages and Systems, 6(2):239–253, April 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Katayama:1984:TAG

[Kat84]  Takuya Katayama.  Trans-

lation of attribute grammars into procedures. ACM Transactions on Programming Languages and Systems, 6(3):345– 369, July 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Katz:1993:SCC

[Kat93] Shmuel Katz. A superimposition control construct for distributed systems. ACM Transactions on Programming Languages and Systems, 15(2):337– 356, April 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
169682.html.

Kaufman:1984:TLR

[Kau84] Arie Kaufman. Tailored-list and recombination-delaying buddy systems. ACM Transactions on Programming Languages and Systems, 6(1):118–125, January 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kandemir:1999:GCO

[KBC^{+}99] M. Kandemir, P. Banerjee, A. Choudhary, J. Ramanujam, and N. Shenoy. A global communication optimization technique based on data-flow analysis and linear algebra. ACM Transactions on Programming Languages and Systems, 21(6): 1251–1297, November 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-6/p1251-kandemir/.

Keizer:2022:SCC

[KBP22] Alex C. Keizer, Henning Basold, and Jorge A. Pérez. Session coalgebras: a coalgebraic view on regular and contextfree session types. ACM Transactions on Programming Languages and Systems, 44(3):18:1– 18:45, September 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527633.

Kim:2001:ERV

[KC01] Ik-Soon Kim and Kwang-Moo Choe. Error repair with validation in LR-based parsing. ACM Transactions on Programming Languages and Systems, 23 (4):451–471, July 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kennedy:1999:PRE

[KCL^{+}99] Robert Kennedy, Sun Chan, Shin-Ming Liu, Raymond Lo, Peng Tu, and Fred Chow. Partial redundancy elimination in SSA form. ACM Transactions on Programming Languages and Systems, 21(3):627–676, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-3/
p627-kennedy/p627-kennedy.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-3/p627-kennedy/.

Khedker:1994:GTB

[KD94] Uday P. Khedker and Dhananjay M. Dhamdhere. A generalized theory of bit vector data flow analysis. ACM Transactions on Programming Languages and Systems, 16(5):1472– 1511, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
186043.html.

Kolokasis:2024:TEF

[KEA^{+}24] Iacovos G. Kolokasis, Giannos Evdorou, Shoaib Akram, Christos Kozanitis, Anastasios Papagiannis, Foivos S. Zakkak, Polyvios Pratikakis, and Angelos Bilas. TeraHeap: Exploiting flash storage for mitigating DRAM pressure in managed big data frameworks. ACM Transactions on Programming Languages and Systems, 46(4):12:1– 12:??, December 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3700593.

Kistler:2000:ADM

[KF00] Thomas Kistler and Michael Franz. Automated data-member layout of heap objects to improve memory-hierarchy performance. ACM Transactions on Programming Languages and Systems, 22(3):490– 505, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/2000-22-3/
                                        p490-kistler/p490-kistler.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-3/p490-kistler/.

Kistler:2003:CPO

[KF03] Thomas Kistler and Michael Franz. Continuous program optimization: a case study. ACM Transactions on Programming Languages and Systems, 25 (4):500–548, July 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Knowles:2010:HTC

[KF10] Kenneth Knowles and Cormac Flanagan. Hybrid type checking. ACM Transactions on Programming Languages and Systems, 32 (2):6:1–6:34, January 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Keen:2004:JFD

[KGMO04] Aaron W. Keen, Tingjian Ge, Justin T. Maris, and Ronald A. Olsson. JR: Flexible distributed programming in an extended Java. ACM Transactions on Programming Languages and Systems, 26(3):578–608, May 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kaiser:1992:OBP

[KH92] Gail E. Kaiser and Brent Hailpern. An object-based programming model for shared data. ACM Transactions on Programming Languages and Systems, 14(2):201–264, April 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org/pubs/toc/Abstracts/
0164-0925/128866.html.

Kennedy:1998:ADL

[KK98] Ken Kennedy and Ulrich Kremer. Automatic data layout for distributed-memory machines. ACM Transactions on Programming Languages and Systems, 20(4):869–916, July 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-4/
p869-kennedy/.

Karkare:2007:IBC

[KK07] Bageshri Karkare and Uday P. Khedker. An improved bound for call strings based interprocedural analysis of bit vector frameworks. ACM Transactions on Programming Languages and Systems, 29(6):38:1–38:13, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Korach:1990:MTD

[KKM90] E. Korach, S. Kutten, and S. Moran. A modular technique for the design of efficient distributed leader finding algorithms. ACM Transactions on Programming Languages and Systems, 12(1):84– 101, January 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        77610.html.

Kawahito:2006:ESE

[KKN06] Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Effective sign extension elimination for Java. ACM Transactions on Programming Languages and Systems, 28(1):106–133, January 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kennaway:1994:AGR

[KKSD94] J. R. Kennaway, J. W. Klop, M. R. Sleep, and F. J. De Vries. On the adequacy of graph rewriting for simulating term rewriting. ACM Transactions on Programming Languages and Systems, 16(3):493– 523, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177577.html.

Kaiser:2014:WAM

[KKW14] Alexander Kaiser, Daniel Kroening, and Thomas Wahl. A widening approach to multithreaded program verification. ACM Transactions on Programming Languages and Systems, 36(4): 14:1–14:??, October 2014. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Koopman:1992:CBC

[KLS92] Philip J. Koopman, Jr., Peter Lee, and Daniel P. Siewiorek. Cache behavior of combinator graph reduction. ACM Transactions on Programming Languages and Systems, 14(2):265– 297, April 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

128867.html. Also see [KLS92].

Kristensen:1981:MCL

[KM81] Bent Bruun Kristensen and Ole Lehrmann Madsen. Methods for computing LALR(k) lookahead. ACM Transactions on Programming Languages and Systems, 3(1):60–82, January 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kelly:1998:OCC

[KMM^{+}98] Andrew D. Kelly, Kim Marriott, Andrew MacDonald, Peter J. Stuckey, and Roland Yap. Optimizing compilation of CLP(R). ACM Transactions on Programming Languages and Systems, 20 (6):1223–1250, November 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-6/
                                        p1223-kelly/.

Klein:2006:MCM

[KN06] Gerwin Klein and Tobias Nipkow. A machine-checked model for a Java-like language, virtual machine, and compiler. ACM Transactions on Programming Languages and Systems, 28 (4):619–695, July 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Knapp:1990:EFD

[Kna90] Edgar Knapp. An exercise in the formal derivation of parallel programs: Maximum flows in graphs. ACM Transactions on Programming Languages and Systems, 12(2):203– 223, April 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        78945.html.

Kobayashi:1998:PDF

[Kob98] Naoki Kobayashi. A partially deadlock-free typed process calculus. ACM Transactions on Programming Languages and Systems, 20(2):436–482, March 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-2/
p436-kobayashi/.

Kim:2006:ERI

[KOE^{+}06] Seon Wook Kim, Chong-Liang Ooi, Rudolf Eigenmann, Babak Falsafi, and T. N. Vijaykumar. Exploiting reference idempotency to reduce speculative storage overflow. ACM Transactions on Programming Languages and Systems, 28(5):942– 965, September 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kozen:1997:KAT

[Koz97] Dexter Kozen. Kleene algebra with tests. ACM Transactions on Programming Languages and Systems, 19(3):427– 443, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-3/p427-kozen/.

Kurlander:1995:EIS

[KPF95] Steven M. Kurlander, Todd A.

Proebsting, and Charles N. Fischer. Efficient instruction scheduling for delayed-load architectures. ACM Transactions on Programming Languages and Systems, 17(5):740– 776, September 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
213987.html.

Katzenelson:1992:TMT

[KPS92] Jacob Katzenelson, Shlomit S. Pinter, and Eugen Schenfeld. Type matching, type-graphs, and the Schanuel conjecture. ACM Transactions on Programming Languages and Systems, 14 (4):574–588, October 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
133247.html.

Kobayashi:1999:LPC

[KPT99] Naoki Kobayashi, Benjamin C. Pierce, and David N. Turner. Linearity and the Pi-Calculus. ACM Transactions on Programming Languages and Systems, 21(5):914–947, September 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-5/
p914-kobayashi/p914-kobayashi.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-5/p914-kobayashi/.

Kennedy:1979:DAG

[KR79] Ken Kennedy and Jayashree Ramanathan. A deterministic attribute grammar evaluator based on dynamic scheduling. ACM Transactions on Programming Languages and Systems, 1 (1):142–160, July 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Knoblock:2001:TES

[KR01] Todd B. Knoblock and Jakob Rehof. Type elaboration and subtype completion for Java bytecode. ACM Transactions on Programming Languages and Systems, 23(2):243–272, March 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1982:AAP

[Kro82] Fred T. Krogh. ACM algorithms policy. ACM Transactions on Programming Languages and Systems, 4(3):518– 521, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1983:AAP

[Kro83]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 5(3):502– 505, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1984:AAP

[Kro84]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 6(3):440– 443, July 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1985:AAP

[Kro85]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 7(3):493– 496, July 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1986:AAP

[Kro86]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 8(3):408– 411, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1987:AAP

[Kro87]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 9(3):463– 466, July 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1988:AAP

[Kro88]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 10(3):504– 507, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1989:AAP

[Kro89]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 11(3):483– 486, July 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1990:AAP

[Kro90]  F. T. Krogh.  ACM algo-

rithms policy. ACM Transactions on Programming Languages and Systems, 12(3):493– 496, July 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1991:AAP

[Kro91] Fred T. Krogh. ACM algorithms policy. ACM Transactions on Programming Languages and Systems, 13(3):443– 446, July 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Krogh:1992:AAP

[Kro92] Fred T. Krogh. ACM algorithms policy. ACM Transactions on Programming Languages and Systems, 14(3):462– 465, July 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kim:2018:TFS

[KRR18] Se-Won Kim, Xavier Rival, and Sukyoung Ryu. A theoretical foundation of sensitivity in an abstract interpretation framework. ACM Transactions on Programming Languages and Systems, 40(3):13:1–13:??, August 2018. CODEN ATPSDT.

ISSN 0164-0925 (print), 1558-4593 (electronic).

Korach:1984:DAF

[KRS84] E. Korach, D. Rotem, and N. Santoro. Distributed algorithms for finding centers and medians in networks. ACM Transactions on Programming Languages and Systems, 6(3): 380–401, July 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kruskal:1988:ESM

[KRS88] Clyde P. Kruskal, Larry Rudolph, and Marc Snir. Efficient synchronization on multiprocessors with shared memory. ACM Transactions on Programming Languages and Systems, 10 (4):579–601, October 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
48024.html.

Knoop:1994:OCM

[KRS94] Jens Knoop, Oliver Rüthing, and Bernhard Steffen. Optimal code motion: Theory and practice. ACM Transactions on Programming Languages and Systems, 16(4):1117– 1155, July 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
183443.html.

Kieburtz:1979:CCS

[KS79] Richard B. Kieburtz and Abraham Silberschatz. Comments on “Communicating sequential processes”. ACM Transactions on Programming Languages and Systems, 1(2):218–225, October 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kieburtz:1983:ARE

[KS83] Richard B. Kieburtz and Abraham Silberschatz. Access-right expressions. ACM Transactions on Programming Languages and Systems, 5(1):78–96, January 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Keller:1986:AC

[KS86] Robert M. Keller and M. Ronan Sleep. Applicative caching. ACM Transactions on Programming Languages and Systems, 8 (1):88–108, January 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5004.html.

Kennaway:1988:DSC

[KS88] Richard Kennaway and Ronan Sleep. Director strings as combinators. ACM Transactions on Programming Languages and Systems, 10(4):602– 626, October 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/

48026.html. See corrigendum [KS89].

Kennaway:1989:CDS

[KS89] Richard Kennaway and Ronan Sleep. Corrigendum: “Director Strings as Combinators”. ACM Transactions on Programming Languages and Systems, 11(3):482, July 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [KS88].

Kobayashi:2010:HTS

[KS10] Naoki Kobayashi and Davide Sangiorgi. A hybrid type system for lock-freedom of mobile processes. ACM Transactions on Programming Languages and Systems, 32(5):16:1– 16:49, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Khedker:2007:HRA

[KSK07] Uday P. Khedker, Amitabha Sanyal, and Amey Karkare. Heap reference analysis using access graphs. ACM Transactions on Programming Languages and Systems, 30(1):1:1–1:41, November 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Knoop:1996:PFE

[KSV96] Jens Knoop, Bernhard Steffen, and Jürgen Vollmer. Parallelism for free: efficient and optimal bitvector analyses for parallel programs. ACM Transactions on Programming Languages and Systems, 18(3):268– 299, May 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
229545.html.

Kfoury:1993:TRP

[KTU93] A. J. Kfoury, J. Tiuryn, and P. Urzyczyn. Type recursion in the presence of polymorphic recursion. ACM Transactions on Programming Languages and Systems, 15(2):290– 311, April 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
169687.html.

Kuperman:2000:ATA

[KV00] Orna Kuperman and Moshe Y.

Vardi. An automata-theoretic approach to modular model checking. ACM Transactions on Programming Languages and Systems, 22(1):87– 128, January 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-1/p87-kuperman/.

Kalvala:2009:PTU

[KWL09] Sara Kalvala, Richard Warburton, and David Lacey. Program transformations using temporal logic side conditions. ACM Transactions on Programming Languages and Systems, 31(4): 14:1–14:48, May 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Kasikci:2015:ACD

[KZC15] Baris Kasikci, Cristian Zamfir, and George Candea. Automated classification of data races under both strong and weak memory models. ACM Transactions on Programming Languages and Systems, 37(3): 8:1–8:??, June 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

LaLonde:1981:CSC

[LaL81] Wilf R. LaLonde. The construction of stack-controlling LR parsers for regular right part grammars. ACM Transactions on Programming Languages and Systems, 3(2):168–206, April 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

LaLonde:1983:TCL

[LaL83] Wilf R. LaLonde. Technical correspondence: LaLonde’s reply. ACM Transactions on Programming Languages and Systems, 5 (1):122, January 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [LdR81, Hen83].

LaLonde:1984:TCC

[LaL84] Wilf R. LaLonde. Technical correspondence: Comments on Soisalon-Soininen’s “Inessential Error Entries and Their Use in LR Parser Optimization”. ACM Transactions on Programming Languages and Systems, 6 (3):432–439, July 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [SS82].

LaLonde:1989:DFD

[LaL89]  Wilf R. LaLonde.  Design-

ing families of data types using exemplars. ACM Transactions on Programming Languages and Systems, 11(2):212– 248, April 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
63265.html.

Lamport:1979:NAP

[Lam79] Leslie Lamport. A new approach to proving the correctness of multiprocess programs. ACM Transactions on Programming Languages and Systems, 1(1):84–97, July 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See also corrigendum [Lam80].

Lamport:1980:CNA

[Lam80] Leslie Lamport. Corrigendum:

“A New Approach to Proving the Correctness of Multiprocess Programs”. ACM Transactions on Programming Languages and Systems, 2(1):134, January 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Lam79].

Lamport:1983:SCP

[Lam83] Leslie Lamport. Specifying concurrent program modules. ACM Transactions on Programming Languages and Systems, 5(2): 190–222, April 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lamport:1984:UTI

[Lam84] Leslie Lamport. Using time instead of timeout for faulttolerant distributed systems. ACM Transactions on Programming Languages and Systems, 6 (2):254–280, April 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lamb:1987:ISI

[Lam87] David Alex Lamb. IDL: Sharing intermediate representations. ACM Transactions on Programming Languages and Systems, 9 (3):297–318, July 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        24040.html.

Lamport:1988:CPB

[Lam88] Leslie Lamport. Control predicates are better than dummy variables for reasoning about program control. ACM Transactions on Programming Languages and Systems, 10(2):267– 281, April 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/

pubs/toc/Abstracts/0164-0925/
42348.html.

Lamport:1990:WSP

[Lam90]  Leslie Lamport.  win and

sin: Predicate transformers for concurrency. ACM Transactions on Programming Languages and Systems, 12(3): 396–428, July 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
78970.html.

Lamport:1994:TLA

[Lam94] Leslie Lamport. The temporal logic of actions. ACM Transactions on Programming Languages and Systems, 16(3): 872–923, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177726.html.

Landwehr:1980:ATS

[Lan80] Carl E. Landwehr. An abstract type for statistics collection in Simula. ACM Transactions on Programming Languages and Systems, 2(4):544– 563, October 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Larcheveque:1995:OIP

[Lar95] J.-M. Larchevêque. Optimal incremental parsing. ACM Transactions on Programming Languages and Systems, 17(1):1– 15, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
200996.html.

Lahav:2022:WDA

[LB22] Ori Lahav and Udi Boker. What’s decidable about causally consistent shared memory? ACM Transactions on Programming Languages and Systems, 44 (2):8:1–8:55, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3505273.

Lennon-Bertrand:2022:GCI

[LBMTT22] Meven Lennon-Bertrand, Kenji Maillard, Nicolas Tabareau, and Éric Tanter. Gradualizing the calculus of inductive constructions. ACM Transactions on Programming Languages and Systems, 44(2): 7:1–7:82, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3495528.

Ligatti:2017:SRC

[LBN17] Jay Ligatti, Jeremy Blackburn, and Michael Nachtigal. On subtyping-relation completeness, with an application to isorecursive types. ACM Transactions on Programming Languages and Systems, 39(1):4:1– 4:??, March 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lozano:2019:CRA

[LCBS19] Roberto Castañeda Lozano, Mats Carlsson, Gabriel Hjort Blindell, and Christian Schulte. Combinatorial register allocation and instruction scheduling. ACM Transactions on Programming Languages and Systems, 41 (3):17:1–17:??, July 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3332373.

Lorch:2022:AAV

[LCK^{+}22] Jacob R. Lorch, Yixuan Chen, Manos Kapritsos, Haojun Ma, Bryan Parno, Shaz Qadeer, Upamanyu Sharma, James R. Wilcox, and Xueyuan Zhao. Armada: Automated verification of concurrent code with sound semantic extensibility. ACM Transactions on Programming Languages and Systems, 44(2): 12:1–12:39, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3502491.

Liao:1996:SAD

[LDK^{+}96] Stan Liao, Srinivas Devadas, Kurt Keutzer, Steven Tjiang, and Albert Wang. Storage assignment to decrease code size. ACM Transactions on Programming Languages and Systems, 18(3):235–253, May 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

                                        pubs/toc/Abstracts/0164-0925/
                                        229543.html.

Lee:2007:DIE

[LDM07] Han B. Lee, Amer Diwan, and J. Eliot B. Moss. Design, implementation, and evaluation of a compilation server. ACM Transactions on Programming Languages and Systems, 29(4): 18:1–18:40, August 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

LaLonde:1981:HOP

[LdR81] Wilf R. LaLonde and Jim des Rivieres. Handling operator precedence in arithmetic expressions with tree transformations. ACM Transactions on Programming Languages and Systems, 3 (1):83–103, January 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remarks [Hen83, LaL83].

LeMetayer:1988:AAC

[Le 88] Daniel Le Métayer. ACE: An automatic complexity evaluator. ACM Transactions on Programming Languages and Systems, 10 (2):248–266, April 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        42347.html.

Leeman:1986:FAU

[Lee86] George B. Leeman, Jr. A formal approach to undo operations in programming languages. ACM Transactions on Programming Languages and Systems, 8 (1):50–87, January 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5005.html.

Lee:2009:RFS

[Lee09] Chin Soon Lee. Ranking functions for size-change termination. ACM Transactions on Programming Languages and Systems, 31(3):10:1–10:42, April 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Leiss:1990:KME

[Lei90] Hans Leiss. On Kilbury’s modification of Earley’s algorithm. ACM Transactions on Programming Languages and Systems, 12 (4):610–640, October 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
88637.html.

Leuschel:2004:FIP

[Leu04] Michael Leuschel. A framework for the integration of partial evaluation and abstract interpretation of logic programs. ACM Transactions on Programming Languages and Systems, 26 (3):413–463, May 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liang:2014:RGB

[LFF14] Hongjin Liang, Xinyu Feng, and Ming Fu. Rely-guarantee-based simulation for compositional verification of concurrent program transformations. ACM Transactions on Programming Languages and Systems, 36(1):3:1– 3:??, March 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lueh:2000:FBR

[LGAT00] Guei-Yuan Lueh, Thomas Gross, and Ali-Reza Adl-Tabatabai. Fusion-based register allocation. ACM Transactions on Programming Languages and Systems, 22(3):431–470, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-3/
p431-lueh/p431-lueh.pdf;
http://www.acm.org/pubs/citations/
journals/toplas/2000-22-3/
p431-lueh/.

Lycklama:1991:FCF

[LH91] Edward A. Lycklama and Vassos Hadzilacos. A first-come-first-served mutual-exclusion algorithm with small communication variables. ACM Transactions on Programming Languages and Systems, 13(4):558– 576, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
115370.html.

Lhotak:2008:RAB

[LH08] Ondřej Lhoták and Laurie Hendren. Relations as an abstraction for BDD-based program analysis. ACM Transactions on Programming Languages and Systems, 30(4):19:1– 19:63, July 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liu:2019:RIP

[LHR19] Bozhen Liu, Jeff Huang, and Lawrence Rauchwerger. Rethinking incremental and parallel pointer analysis. ACM Transactions on Programming Languages and Systems, 41(1): 6:1–6:??, March 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lindstrom:1979:BGC

[Lin79] Gary Lindstrom. Backtracking in a generalized control setting. ACM Transactions on Programming Languages and Systems, 1(1):8–26, July 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lin:1993:PIA

[Lin93] Huimin Lin. Procedural implementation of algebraic specification. ACM Transactions on Programming Languages and Systems, 15(5):876– 895, November 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
161473.html.

Liu:1999:SVF

[LJ99] Zhiming Liu and Mathai Joseph.

Specification and verification of fault-tolerance, timing, and scheduling. ACM Transactions on Programming Languages and Systems, 21(1):46– 89, January 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-1/p46-liu/.

Lee:2002:ADC

[LK02] Peizong Lee and Zvi Meir Kedem. Automatic data and computation decomposition on distributed memory parallel computers. ACM Transactions on Programming Languages and Systems, 24(1):1–50, January 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lee:2017:SNS

[LLK^{+}17] Woosuk Lee, Wonchan Lee, Dongok Kang, Kihong Heo, Hakjoo Oh, and Kwangkeun Yi. Sound non-statistical clustering of static analysis alarms. ACM Transactions on Programming Languages and Systems, 39 (4):16:1–16:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lee:2023:OHE

[LLOY23] Dongkwon Lee, Woosuk Lee, Hakjoo Oh, and Kwangkeun Yi. Optimizing homomorphic evaluation circuits by program synthesis and time-bounded exhaustive search. ACM Transactions on Programming Languages and Systems, 45(3):16:1– 16:??, September 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3591622.

Lidman:2018:VRP

[LM18] Jacob Lidman and Sally A. Mckee. Verifying reliability properties using the hyperball abstract domain. ACM Transactions on Programming Languages and Systems, 40(1):3:1–3:??, January 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lamport:2022:PMS

[LM22] Leslie Lamport and Stephan Merz. Prophecy made simple. ACM Transactions on Programming Languages and Systems, 44 (2):6:1–6:27, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3492545.

Leuschel:1998:CGP

[LMD98] Michael Leuschel, Bern Martens, and Danny De Schreye. Controlling generalization and polyvariance in partial deduction of normal logic programs. ACM Transactions on Programming Languages and Systems, 20(1): 208–258, January 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/

                                        journals/toplas/1998-20-1/
                                        p208-leuschel/.

Liu:2021:SDC

[LMM21] Lun Liu, Todd Millstein, and Madanlal Musuvathi. Safe-by-default concurrency for modern programming languages. ACM Transactions on Programming Languages and Systems, 43(3): 10:1–10:50, September 2021. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3462206.

Leino:2002:DAI

[LN02] K. Rustan M. Leino and Greg Nelson. Data abstraction and information hiding. ACM Transactions on Programming Languages and Systems, 24(5):491– 553, September 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Leavens:2015:BSS

[LN15] Gary T. Leavens and David A. Naumann. Behavioral subtyping, specification inheritance, and modular reasoning. ACM Transactions on Programming Languages and Systems, 37(4): 13:1–13:??, August 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Laufer:1994:PTI

[LO94] Konstantin Läufer and Martin Odersky. Polymorphic type inference and abstract data types. ACM Transactions on Programming Languages and Systems, 16(5):1411– 1430, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
186031.html.

Lochbihler:2013:MJM

[Loc13] Andreas Lochbihler. Making the Java memory model safe. ACM Transactions on Programming Languages and Systems, 35 (4):12:1–12:??, December 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Loeckx:1987:ASC

[Loe87] Jacques Loeckx. Algorithmic specifications: a constructive specification method for abstract data types. ACM Transactions on Programming Languages and Systems, 9(4):646– 685, October 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
30399.html.

Luckham:1980:AEH

[LP80] David C. Luckham and W. Polak. Ada exception handling: an axiomatic approach. ACM Transactions on Programming Languages and Systems, 2(2): 225–233, April 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lamport:1999:SYS

[LP99] Leslie Lamport and Lawrence C. Paulson. Should your specification language be typed. ACM Transactions on Programming Languages and Systems, 21(3):502–526, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/1999-21-3/
                                        p502-lamport/p502-lamport.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-3/p502-lamport/.

Leroy:2000:TBA

[LP00] Xavier Leroy and François Pessaux. Type-based analysis of uncaught exceptions. ACM Transactions on Programming Languages and Systems, 22(2): 340–377, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-2/p340-leroy/.

Levanoni:2006:FRC

[LP06] Yossi Levanoni and Erez Petrank. An on-the-fly referencecounting garbage collector for Java. ACM Transactions on Programming Languages and Systems, 28(1):1–69, January 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Leung:2001:STC

[LPP01] Allen Leung, Krishna V. Palem, and Amir Pnueli. Scheduling time-constrained instructions on pipelined processors. ACM Transactions on Programming Languages and Systems, 23(1): 73–103, January 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lieberherr:2004:TOS

[LPSO04] Karl Lieberherr, Boaz Patt-Shamir, and Doug Orleans. Traversals of object structures: Specification and efficient implementation. ACM Transactions on Programming Languages and Systems, 26(2):370–412, March 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lim:2013:TSG

[LR13] Junghee Lim and Thomas Reps.

TSL: a system for generating abstract interpreters and its application to machine-code analysis. ACM Transactions on Programming Languages and Systems, 35 (1):4:1–4:??, April 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lepigre:2019:PSC

[LR19] Rodolphe Lepigre and Christophe Raffalli. Practical subtyping for Curry-style languages. ACM Transactions on Programming Languages and Systems, 41(1): 5:1–5:??, March 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Luckham:1979:VAR

[LS79] David C. Luckham and Norihisa Suzuki. Verification of array, record, and pointer operations in Pascal. ACM Transactions on Programming Languages and Systems, 1(2):226–244, October 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Leverett:1980:CSD

[LS80] Bruce W. Leverett and Thomas G. Szymanski. Chaining spandependent jump instructions. ACM Transactions on Programming Languages and Systems, 2 (3):274–289, July 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lindstrom:1981:RRB

[LS81] Gary Lindstrom and Mary Lou Soffa. Referencing and retention in block-structured coroutines. ACM Transactions on Programming Languages and Systems, 3 (3):263–292, July 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liskov:1983:GAL

[LS83] Barbara Liskov and Robert Scheifler. Guardians and actions: Linguistic support for robust, distributed programs. ACM Transactions on Programming Languages and Systems, 5 (3):381–404, July 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lamport:1984:HLC

[LS84] Leslie Lamport and Fred B.

Schneider. The “Hoare logic” of CSP, and all that. ACM Transactions on Programming Languages and Systems, 6(2): 281–296, April 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lang:1998:SAE

[LS98] Jun Lang and David B. Stewart. A study of the applicability of existing exception-handling techniques to component-base real-time software technology. ACM Transactions on Programming Languages and Systems, 20 (2):274–301, March 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-2/
p274-lang/.

Levi:2003:MSA

[LS03] Francesca Levi and Davide Sangiorgi. Mobile safe ambients. ACM Transactions on Programming Languages and Systems, 25 (1):1–69, January 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Li:2004:ATI

[LS04] Zhiyuan Li and Yonghong Song.

Automatic tiling of iterative stencil loops. ACM Transactions on Programming Languages and Systems, 26(6):975– 1028, November 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liquori:2008:FME

[LS08] Luigi Liquori and Arnaud Spiwack. FeatherTrait: a modest extension of Featherweight Java. ACM Transactions on Programming Languages and Systems, 30 (2):11:1–11:32, March 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liu:2009:DRE

[LS09] Yanhong A. Liu and Scott D. Stoller. From datalog rules to efficient programs with time and space guarantees. ACM Transactions on Programming Languages and Systems, 31(6): 21:1–21:38, August 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Liu:2005:OAA

[LSLR05] Yanhong A. Liu, Scott D. Stoller, Ning Li, and Tom Rothamel. Optimizing aggregate array computations in loops. ACM Transactions on Programming Languages and Systems, 27 (1):91–125, January 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lamport:1982:BGP

[LSP82] Leslie Lamport, Robert Shostak, and Marshall Pease. The Byzantine generals problem. ACM Transactions on Programming Languages and Systems, 4(3): 382–401, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). They proved that Byzantine agreement (the subject of Section ??) cannot be reached unless fewer than one-third of the processes are faulty. This result assumes that authentication, i.e., the crypting of messages to make them unforgeable, is not used. With unforgeable messages, they show that the problem is solvable for any n ≥ t > 0, where n is the total number of processes and t is the number of faulty processes.

Liu:1998:SCI

[LST98] Yanhong A. Liu, Scott D.

Stoller, and Tim Teitelbaum. Static caching for incremental computation. ACM Transactions on Programming Languages and Systems, 20(3): 546–585, May 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-3/
p546-liu/.

League:2002:TPC

[LST02] Christopher League, Zhong Shao, and Valery Trifonov. Type-preserving compilation of Featherweight Java. ACM Transactions on Programming Languages and Systems, 24(2): 112–152, March 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Lengauer:1979:FAF

[LT79] Thomas Lengauer and Robert Endre Tarjan. A fast algorithm for finding dominators in a flowgraph. ACM Transactions on Programming Languages and Systems, 1(1):121–141, July 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Li:2020:PAS

[LTMS20] Yue Li, Tian Tan, Anders Møller, and Yannis Smaragdakis. A principled approach to selective context sensitivity for pointer analysis. ACM Transactions on Programming Languages and Systems, 42(2): 10:1–10:40, May 2020. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  https://dl.acm.org/
                                        doi/abs/10.1145/3381915.

LeCharlier:1994:EEG

[LV94] Baudouin Le Charlier and Pascal Van Hentenryck. Experimental evaluation of a generic abstract interpretation algorithm for PROLOG. ACM Transactions on Programming Languages and Systems, 16(1):35– 101, January 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174627.html.

Lobo-Vesga:2021:PLD

[LVRG21] Elisabet Lobo-Vesga, Alejandro Russo, and Marco Gaboardi.

A programming language for data privacy with accuracy estimations. ACM Transactions on Programming Languages and Systems, 43(2): 6:1–6:42, July 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3452096.

Lipton:1983:VLP

[LVV^{+}83] Richard J. Lipton, Jacobo Valdes, Gopalakrishnan Vijayen, Stephen C. Notth, and Robert Sedgewick. VLSI layout as programming. ACM Transactions on Programming Languages and Systems, 5(3):405–421, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Leivent:1993:MFT

[LW93] Jonathan I. Leivent and Ronald J.

Watro. Mathematical foundations for Time Warp systems. ACM Transactions on Programming Languages and Systems, 15(5):771–794, November 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
161470.html.

Liskov:1994:BNS

[LW94] B. H. Liskov and J. M.

Wing. A behavioral notion of subtyping. ACM Transactions on Programming Languages and Systems, 16(6):1811– 1841, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        197383.html.

Liu:2021:ICU

[LWR21] Peizun Liu, Thomas Wahl, and Thomas Reps. Interprocedural context-unbounded program analysis using observation sequences. ACM Transactions on Programming Languages and Systems, 42(4):16:1– 16:34, February 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3418583.

Lee:1998:PAF

[LY98] Oukseh Lee and Kwangkeun Yi. Proofs about a folklore letpolymorphic type inference algorithm. ACM Transactions on Programming Languages and Systems, 20(4):707–723, July 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-4/
                                        p707-lee/.

Li:2022:FGS

[LZR22] Yuanbo Li, Qirun Zhang, and Thomas Reps. Fast graph simplification for interleaved-Dyck reachability. ACM Transactions on Programming Languages and Systems, 44(2): 11:1–11:28, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3492428.

Mallgren:1982:FSG

[Mal82] William R. Mallgren. Formal specification of graphic data types. ACM Transactions on Programming Languages and Systems, 4(4):687–710, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Merlin:1983:CSS

[MB83] Philip Merlin and Gregor V.

Bochmann. On the construction of submodule specifications and communication protocols. ACM Transactions on Programming Languages and Systems, 5 (1):1–25, January 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Morris:1999:SF

[MB99] J. M. Morris and A. Bunkenburg. Specificational functions. ACM Transactions on Programming Languages and Systems, 21(3):677–701, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-3/
p677-morris/p677-morris.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-3/p677-morris/.

Millstein:2004:MTH

[MBC04] Todd Millstein, Colin Bleckner, and Craig Chambers. Modular typechecking for hierarchically extensible datatypes and functions. ACM Transactions on Programming Languages and Systems, 26(5):836– 889, September 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Morris:2009:TTN

[MBT09] Joseph M. Morris, Alexander Bunkenburg, and Malcolm Tyrrell. Term transformers: a new approach to state. ACM Transactions on Programming Languages and Systems, 31(4): 16:1–16:42, May 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Misra:1982:DGA

[MC82a] Jayadev Misra and K. M. Chandy. A distributed graph algorithm: Knot detection. ACM Transactions on Programming Languages and Systems, 4(4): 678–686, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Misra:1982:TDD

[MC82b] Jayadev Misra and K. M. Chandy. Termination detection of diffusing computations in communicating sequential processes. ACM Transactions on Programming Languages and Systems, 4(1):37–43, January 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

McGraw:1982:VLD

[McG82] James R. McGraw. The VAL language: Description and analysis. ACM Transactions on Programming Languages and Systems, 4(1):44–82, January 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

McKinley:1996:IDL

[MCT96] Kathryn S. McKinley, Steve Carr, and Chau-Wen Tseng. Improving data locality with loop transformations. ACM Transactions on Programming Languages and Systems, 18(4): 424–453, July 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
233564.html.

Morrison:1991:AHA

[MDCB91] R. Morrison, A. Dearle, R. C. H.

Connor, and A. L. Brown. An ad hoc approach to the implementation of polymorphism. ACM Transactions on Programming Languages and Systems, 13(3):342–371, July 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
117017.html.

Moreau:2005:BDR

[MDJ05] Luc Moreau, Peter Dickman, and Richard Jones. Birrell’s distributed reference listing revisited. ACM Transactions on Programming Languages and Systems, 27(6):1344–1395, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Moon:1997:PNC

[ME97] Soo-Mook Moon and Kemal Ebcioğlu. Parallelizing nonnumerical code with selective scheduling and software pipelining. ACM Transactions on Programming Languages and Systems, 19(6):853–898, November 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1997-19-6/
p853-moon/.

Mauney:1988:DEL

[MF88] Jon Mauney and Charles N. Fischer. Determining the extent of lookahead in syntactic error repair. ACM Transactions on Programming Languages and Systems, 10(3): 456–469, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
44505.html.

Matthews:2009:OSM

[MF09] Jacob Matthews and Robert Bruce Findler. Operational semantics for multi-language programs. ACM Transactions on Programming Languages and Systems, 31 (3):12:1–12:44, April 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Millstein:2009:EMP

[MFRW09] Todd Millstein, Christopher Frost, Jason Ryder, and Alessandro Warth. Expressive and modular predicate dispatch for Java. ACM Transactions on Programming Languages and Systems, 31(2):7:1–7:54, February 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Moriconi:1986:PSP

[MH86] Mark Moriconi and Dwight Hare. The PegaSys system: Pictures as formal documentation of large programs. ACM Transactions on Programming Languages and Systems, 8(4): 524–546, October 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
6478.html.

Mirani:2004:FCM

[MH04] Rajiv Mirani and Paul Hudak.

First-class monadic schedules. ACM Transactions on Programming Languages and Systems, 26 (4):609–651, July 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Merro:2006:BBS

[MH06] Massimo Merro and Matthew Hennessy. A bisimulation-based semantic theory of Safe Ambients. ACM Transactions on Programming Languages and Systems, 28(2):290–330, March 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Milne:1985:CRC

[Mil85] George J. Milne. CIRCAL and the representation of communication, concurrency, and time. ACM Transactions on Programming Languages and Systems, 7 (2):270–298, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        3322.html.

Minsky:1984:SLC

[Min84] Naftaly H. Minsky. Selective and locally controlled transport of privileges. ACM Transactions on Programming Languages and Systems, 6(4):573–602, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Miquey:2019:CSC

[Miq19]  Étienne Miquey.  A classi-

cal sequent calculus with dependent types. ACM Transactions on Programming Languages and Systems, 41(2):8:1– 8:??, June 2019. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3230625.

Misra:1981:EPE

[Mis81] Jayadev Misra. An exercise in program explanation. ACM Transactions on Programming Languages and Systems, 3(1): 104–109, January 1981. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Misra:1986:AMA

[Mis86] Jayadev Misra. Axioms for memory access in asynchronous hardware systems. ACM Transactions on Programming Languages and Systems, 8(1):142– 153, January 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5007.html.

Misra:1994:PSP

[Mis94] J. Misra. Powerlist: a structure for parallel recursion. ACM Transactions on Programming Languages and Systems, 16(6): 1737–1767, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
197356.html.

Micallef:1994:EAG

[MK94] Josephine Micallef and Gail E.

Kaiser. Extending attribute grammers to support programmingin-the-large. ACM Transactions on Programming Languages and Systems, 16(5):1572– 1612, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        186091.html.

Ma:1980:DMI

[ML80] Perng-Ti Ma and T. G. Lewis. Design of a machine-independent optimizing system for emulator development. ACM Transactions on Programming Languages and Systems, 2(2):239– 262, April 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Masud:2021:SCD

[ML21] Abu Naser Masud and Björn Lisper. Semantic correctness of dependence-based slicing for interprocedural, possibly nonterminating programs. ACM Transactions on Programming Languages and Systems, 42(4): 19:1–19:56, February 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3434489.

Martelli:1982:EUA

[MM82] Alberto Martelli and Ugo Montanari. An efficient unification algorithm. ACM Transactions on Programming Languages and Systems, 4(2):258–282, February 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Myers:1989:RRA

[MM89] Eugene W. Myers and Webb Miller. Row replacement algorithms for screen editors. ACM Transactions on Programming Languages and Systems, 11(1): 33–56, January 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
59290.html.

Markstrum:2010:JDP

[MME^{+}10] Shane Markstrum, Daniel Marino, Matthew Esquivel, Todd Millstein, Chris Andreae, and James Noble. JavaCOP: Declarative pluggable types for Java. ACM Transactions on Programming Languages and Systems, 32(2): 4:1–4:37, January 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Morzenti:1992:MPR

[MMG92] Angelo Morzenti, Dino Mandrioli, and Carlo Ghezzi. A model parametric real-time logic. ACM Transactions on Programming Languages and Systems, 14(4): 521–573, October 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
129397.html.

Moreira:2000:FMJ

[MMG00] José E. Moreira, Samuel P.

Midkiff, and Manish Gupta. From flop to megaflops: Java for technical computing. ACM Transactions on Programming Languages and Systems, 22(2): 265–295, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-2/p265-moreira/.

Marathe:2007:MMT

[MMM^{+}07] Jaydeep Marathe, Frank Mueller, Tushar Mohan, Sally A. Mckee, Bronis R. De Supinski, and Andy Yoo. METRIC: Memory tracing via dynamic binary rewriting to identify cache inefficiencies. ACM Transactions on Programming Languages and Systems, 29(2):12:1–12:36, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Masticola:1995:LFM

[MMR95] Stephen P. Masticola, Thomas J. Marlowe, and Barbara G. Ryder. Lattice frameworks for multisource and bidirectional data flow problems. ACM Transactions on Programming Languages and Systems, 17(5):777– 803, September 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        213989.html.

Morgan:1996:PPT

[MMS96] Carroll Morgan, Annabelle McIver, and Karen Seidel. Probabilistic predicate transformers. ACM Transactions on Programming Languages and Systems, 18(3):325–353, May 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/

pubs/toc/Abstracts/0164-0925/
229547.html.

Mohan:1981:TCF

[Moh81] C. Mohan. Technical correspondence: On Francez’s “Distributed Termination”. ACM Transactions on Programming Languages and Systems, 3(1): 112, January 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Fra80b, Fra81].

Moitra:1983:TCA

[Moi83] Abha Moitra. Technical correspondence: On Apt, Francez, and de Roever’s “A Proof System for Communicating Sequential Processes”. ACM Transactions on Programming Languages and Systems, 5(3):500– 501, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [AFdR80].

Monniaux:2008:PVF

[Mon08] David Monniaux. The pitfalls of verifying floating-point computations. ACM Transactions on Programming Languages and Systems, 30(3):12:1– 12:41, May 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Morgan:1988:SS

[Mor88] Carroll Morgan. The specification statement. ACM Transactions on Programming Languages and Systems, 10(3): 403–419, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
44503.html.

Moller:2007:SVX

[MOS07a] Anders Møller, Mads Østerby Olesen, and Michael I. Schwartzbach. Static validation of XSL transformations. ACM Transactions on Programming Languages and Systems, 29(4):21:1–21:47, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Muller-Olm:2007:AMA

[MOS07b] Markus Müller-Olm and Helmut Seidl. Analysis of modular arithmetic. ACM Transactions on Programming Languages and Systems, 29(5):29:1–29:27, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Murer:1996:IAS

[MOSS96] Stephan Murer, Stephen Omohundro, David Stoutamire, and Clemens Szyperski. Iteration abstraction in Sather. ACM Transactions on Programming Languages and Systems, 18(1): 1–15, January 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
225541.html.

Mitchell:1988:ATE

[MP88] John C. Mitchell and Gordon D. Plotkin. Abstract types have existential type. ACM Transactions on Programming Languages and Systems, 10(3): 470–502, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

45065.html. Preliminary version appeared in Proc. 12th ACM Symp. on Principles of Programming Languages, 1985.

Moore:2002:AC

[MP02] J. Strother Moore and George Porter. The apprentice challenge. ACM Transactions on Programming Languages and Systems, 24(3):193–216, May 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

McKinley:2007:ECG

[MP07] Kathryn S. McKinley and Keshav Pingali. Editorial: a changing of the guard. ACM Transactions on Programming Languages and Systems, 29(6): 30:1–30:2, October 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Mckinley:2010:DVT

[MP10a] Kathryn S. Mckinley and Keshav Pingali. La dolce vita at TOPLAS. ACM Transactions on Programming Languages and Systems, 32(4):10:1– 10:6, April 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Mckinley:2010:PVT

[MP10b] Kathryn S. Mckinley and Keshav Pingali. La prossima vita at TOPLAS. ACM Transactions on Programming Languages and Systems, 32(6):20:1, August 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Murali:2023:FOL

[MPLM23] Adithya Murali, Lucas Peña, Christof Löding, and P. Madhusudan. A first-order logic with frames. ACM Transactions on Programming Languages and Systems, 45(2): 7:1–7:??, June 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3583057.

Menon:2003:FSA

[MPM03] Vijay Menon, Keshav Pingali, and Nikolay Mateev. Fractal symbolic analysis. ACM Transactions on Programming Languages and Systems, 25(6):776– 813, November 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Moreau:2005:RAP

[MQ05] Luc Moreau and Christian Queinnec. Resource aware programming. ACM Transactions on Programming Languages and Systems, 27(3):441–476, May 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Morgan:1988:RC

[MRG88] C. C. Morgan, K. A. Robinson, and P. H. B. Gardiner. On the refinement calculus. Technical Monograph PRG-70, Oxford University Computing Laboratory, Wolfson Building, Parks Road, Oxford, UK, October 1988. ISBN 0-902928-52-X.

Melo:2020:TIC

[MRGP20] Leandro T. C. Melo, Rodrigo G.

Ribeiro, Breno C. F. Guimarães, and Fernando Magno Quintão Pereira. Type inference for C: Applications to the static analysis of incomplete programs. ACM Transactions on Programming Languages and Systems, 42 (3):15:1–15:71, December 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3421472.

Maher:1983:API

[MS83] B. Maher and D. H. Sleeman.

Automatic program improvement: Variable usage transformations. ACM Transactions on Programming Languages and Systems, 5(2):236–264, April 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Murphy:1988:NDP

[MS88] Sandra L. Murphy and A. Udaya Shankar. A note on the drinking philosophers problem (technical correspondence). ACM Transactions on Programming Languages and Systems, 10(1):

178–188, January 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Marriott:1994:DAI

[MSJ94] Kim Marriott, Harald Søndergaard, and Neil D. Jones. Denotational abstract interpretation of logic programs. ACM Transactions on Programming Languages and Systems, 16(3):607– 648, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177650.html.

Marino:2016:DXU

[MSM^{+}16] Daniel Marino, Abhayendra Singh, Todd Millstein, Madanlal Musuvathi, and Satish Narayanasamy. drf x: an understandable, high performance, and flexible memory model for concurrent languages. ACM Transactions on Programming Languages and Systems, 38(4): 16:1–16:??, October 2016. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Madhavan:2000:EGG

[MSRR00] Maya Madhavan, Priti Shankar, Siddhartha Rai, and U. Ramakrishna. Extending Graham-Glanville techniques for optimal code generation. ACM Transactions on Programming Languages and Systems, 22(6):973– 1001, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-6/
p973-madhavan/p973-madhavan.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-6/p973-madhavan/.

Morris:2008:DNF

[MT08] Joseph M. Morris and Malcolm Tyrrell. Dually nondeterministic functions. ACM Transactions on Programming Languages and Systems, 30(6):34:1–34:34, October 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Moret:1980:AVR

[MTG80] B. M. E. Moret, M. G. Thomason, and R. C. Gonzalez. The activity of a variable and its relation to decision trees. ACM Transactions on Programming Languages and Systems, 2(4): 580–595, October 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Matsushita:2021:RCB

[MTK21] Yusuke Matsushita, Takeshi Tsukada, and Naoki Kobayashi. RustHorn: CHC-based verification for Rust programs. ACM Transactions on Programming Languages and Systems, 43(4): 15:1–15:54, December 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3462205.

MacDonald:2009:DDP

[MTSS09] Steve MacDonald, Kai Tan, Jonathan Schaeffer, and Duane Szafron. Deferring design pattern decisions and automating structural pattern changes using a design-pattern-based programming system. ACM Transactions on Programming Languages and Systems, 31(3):9:1– 9:49, April 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Muller:1992:MLR

[Mul92] Robert Muller. M-LISP: a representation-independent dialect of LISP with reduction semantics. ACM Transactions on Programming Languages and Systems, 14(4):589– 616, October 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
133254.html.

Muller:2021:ISS

[Mül21] Peter Müller. Introduction to the special section on ESOP 2020. ACM Transactions on Programming Languages and Systems, 43(4):13:1, December 2021. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL https:

//dl.acm.org/doi/10.1145/
3484490.

Murtagh:1991:ISM

[Mur91]  Thomas P. Murtagh.  An

improved storage management scheme for block structured languages. ACM Transactions on Programming Languages and Systems, 13(3): 372–398, July 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
117016.html.

Mueller:1987:RMS

[MV87] Robert A. Mueller and Joseph Varghese. Retargetable microcode synthesis. ACM Transactions on Programming Languages and Systems, 9(2):257– 276, April 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
23717.html.

Maassen:2001:EJR

[MVV^{+}01] Jason Maassen, Rob Van Nieuwpoort, Ronald Veldema, Henri Bal, Thilo Kielmann, Ceriel Jacobs, and Rutger Hofman. Efficient Java RMI for parallel programming. ACM Transactions on Programming Languages and Systems, 23(6):747–775, November 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Manna:1980:DAP

[MW80] Zohar Manna and Richard Waldinger. A deductive approach to program synthesis. ACM Transactions on Programming Languages and Systems, 2 (1):90–121, January 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Manna:1984:SCP

[MW84] Zohar Manna and Pierre Wolper. Synthesis of communicating processes from temporal logic specifications. ACM Transactions on Programming Languages and Systems, 6(1):68–93, January 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Mulkers:1994:LSD

[MWB94] Anne Mulkers, William Winsborough, and Maurice Bruynooghe. Live-structure dataflow analysis for Prolog. ACM Transactions on Programming Languages and Systems, 16(2):205– 258, March 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174664.html.

Morrisett:1999:SFT

[MWCG99] Greg Morrisett, David Walker, Karl Crary, and Neal Glew. From System F to typed assembly language. ACM Transactions on Programming Languages and Systems, 21(3):527– 568, May 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/articles/journals/toplas/
                                        1999-21-3/p527-morrisett/
                                        p527-morrisett.pdf;   http:

//www.acm.org/pubs/citations/
journals/toplas/1999-21-3/
p527-morrisett/.

Melicher:2022:BAE

[MXZ^{+}22] Darya Melicher, Anlun Xu, Valerie Zhao, Alex Potanin, and Jonathan Aldrich. Bounded abstract effects. ACM Transactions on Programming Languages and Systems, 44(1): 5:1–5:48, March 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3492427.

McKenzie:1995:ERS

[MYD95] Bruce J. McKenzie, Corey Yeatman, and Lorraine De Vere. Error repair in shift-reduce parsers. ACM Transactions on Programming Languages and Systems, 17(4):672–689, July 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

210193.html. See failure report [BN99].

Myers:1990:CUI

[Mye90] Brad A. Myers. Creating user interfaces using programming by example, visual programming, and constraints. ACM Transactions on Programming Languages and Systems, 12(2): 143–177, April 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

                                        pubs/toc/Abstracts/0164-0925/
                                        78943.html.

Myers:2017:F

[Mye17] Andrew Myers. Foreword. ACM Transactions on Programming Languages and Systems, 39(1): 1:1–1:??, March 2017. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Myers:2018:EFS

[Mye18] Andrew Myers. Editor’s foreword to “Static Backward Slicing of Non-Deterministic Programs and Systems”. ACM Transactions on Programming Languages and Systems, 40(3): 11:1–11:??, August 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [DL18].

Myers:2019:E

[Mye19] Andrew Myers. Editorial. ACM Transactions on Programming Languages and Systems, 41(2): 7:1–7:??, June 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3324782.

Narlikar:1999:SES

[NB99] Girija J. Narlikar and Guy E. Blelloch. Space-efficient scheduling of nested parallelism. ACM Transactions on Programming Languages and Systems, 21(1): 138–173, January 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/

pubs/citations/journals/toplas/
1999-21-1/p138-narlikar/.

Nanevski:2013:DTT

[NBG13] Aleksandar Nanevski, Anindya Banerjee, and Deepak Garg. Dependent type theory for verification of information flow and access control policies. ACM Transactions on Programming Languages and Systems, 35(2): 6:1–6:??, July 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Necula:2005:CTS

[NCH^{+}05] George C. Necula, Jeremy Condit, Matthew Harren, Scott McPeak, and Westley Weimer. CCured: type-safe retrofitting of legacy software. ACM Transactions on Programming Languages and Systems, 27(3):477– 526, May 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Narayanan:2020:SDV

[NcS20] Praveen Narayanan and Chung chieh Shan. Symbolic disintegration with a variety of base measures. ACM Transactions on Programming Languages and Systems, 42(2):9:1– 9:60, May 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3374208.

Norris:2016:PAM

[ND16] Brian Norris and Brian Demsky. A practical approach for model checking C/C++11 code. ACM Transactions on Programming Languages and Systems, 38 (3):10:1–10:??, May 2016. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nelson:1989:GDC

[Nel89] Greg Nelson. A generalization of Dijkstra’s calculus. ACM Transactions on Programming Languages and Systems, 11(4): 517–561, October 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
69559.html.

Nicholson:1989:DSP

[NF89] Tim Nicholson and Norman Foo. A denotational semantics for Prolog. ACM Transactions on Programming Languages and Systems, 11(4):650– 665, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

69564.html. See also remarks in [FA93].

Nguyen:2005:EEA

[NI05] Thi Viet Nga Nguyen and François Irigoin. Efficient and effective array bound checking. ACM Transactions on Programming Languages and Systems, 27(3):527–570, May 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nielson:1985:PTD

[Nie85] Flemming Nielson. Program transformations in a denotational setting. ACM Transactions on Programming Languages and Systems, 7(3):359– 379, July 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3917.html.

Nix:1985:EE

[Nix85] Robert P. Nix. Editing by example. ACM Transactions on Programming Languages and Systems, 7(4):600– 621, October 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4476.html.

Nielson:1986:TCC

[NN86] Flemming Nielson and Hanne Riis Nielson. Technical correspondence: Comments on Georgeff’s “Transformations and Reduction Strategies for Typed Lambda Expressions”. ACM Transactions on Programming Languages and Systems, 8(3): 406–407, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

215007.html. See [Geo84].

Nougrahiya:2024:HDI

[NN24] Aman Nougrahiya and V. Krishna Nandivada. Homeostasis: Design and implementation of a self-stabilizing compiler. ACM Transactions on Programming Languages and Systems, 46 (2):6:1–6:??, June 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3649308.

Nelson:1979:SCD

[NO79] Greg Nelson and Derek C. Oppen. Simplification by cooperating decision procedures. ACM Transactions on Programming Languages and Systems, 1(2): 245–257, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Naik:2008:TSE

[NP08] Mayur Naik and Jens Palsberg. A type system equivalent to a model checker. ACM Transactions on Programming Languages and Systems, 30(5): 29:1–29:24, August 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nanda:2006:ISM

[NR06] Mangala Gowri Nanda and S. Ramesh. Interprocedural slicing of multithreaded programs with applications to Java. ACM Transactions on Programming Languages and Systems, 28(6): 1088–1144, November 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nikolic:2013:RAP

[NS13] Durica Nikolić and Fausto Spoto. Reachability analysis of program variables. ACM Transactions on Programming Languages and Systems, 35(4):14:1– 14:??, December 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nowatzki:2015:SFS

[NSTD^{+}15] Tony Nowatzki, Michael Sartin-Tarm, Lorenzo De Carli, Karthikeyan Sankaralingam, Cristian Estan, and Behnam Robatmili. A scheduling framework for spatial architectures across multiple constraint-solving theories. ACM Transactions on Programming Languages and Systems, 37 (1):2:1–2:??, January 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Nandivada:2013:TFO

[NSZS13] V. Krishna Nandivada, Jun Shirako, Jisheng Zhao, and Vivek Sarkar. A transformation framework for optimizing task-parallel programs. ACM Transactions on Programming Languages and Systems, 35(1): 3:1–3:??, April 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Olderog:1988:FPP

[OA88] Ernst-Rüdiger Olderog and Krzysztof R. Apt. Fairness in parallel programs: The transformational approach. ACM Transactions on Programming Languages and Systems, 10(3): 420–455, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
44504.html.

Odersky:1993:DCD

[Ode93]  Martin Odersky.  Defining

context-dependent syntax without using contexts. ACM Transactions on Programming Languages and Systems, 15(3): 535–562, July 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
174159.html.

Olmedo:2018:CPP

[OGJ^{+}18] Federico Olmedo, Friedrich Gretz, Nils Jansen, Benjamin Lucien Kaminski, Joost-Pieter Katoen, and Annabelle Mciver. Conditioning in probabilistic programming. ACM Transactions on Programming Languages and Systems, 40(1): 4:1–4:??, January 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Oh:2014:GSA

[OHL^{+}14] Hakjoo Oh, Kihong Heo, Wonchan Lee, Woosuk Lee, Daejun Park, Jeehoon Kang, and Kwangkeun Yi. Global sparse analysis framework. ACM Transactions on Programming Languages and Systems, 36(3): 8:1–8:??, September 2014. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ohori:1995:PRC

[Oho95] Atsushi Ohori. A polymorphic record calculus and its compilation. ACM Transactions on Programming Languages and Systems, 17(6):844– 895, November 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
218572.html.

Ohori:2007:PTM

[Oho07] Atsushi Ohori. A proof theory for machine code. ACM Transactions on Programming Languages and Systems, 29(6): 36:1–36:35, October 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ogasawara:2006:EED

[OKN06] Takeshi Ogasawara, Hideaki Komatsu, and Toshio Nakatani. EDO: Exception-Directed Optimization in Java. ACM Transactions on Programming Languages and Systems, 28(1):70– 105, January 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Owicki:1982:PLP

[OL82] Susan Owicki and Leslie Lamport. Proving liveness properties of concurrent programs. ACM Transactions on Programming Languages and Systems, 4 (3):455–495, July 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Oh:2016:SXS

[OLH^{+}16] Hakjoo Oh, Wonchan Lee, Kihong Heo, Hongseok Yang, and Kwangkeun Yi. Selective Xsensitive analysis guided by impact pre-analysis. ACM Transactions on Programming Languages and Systems, 38(2):6:1– 6:??, January 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Odersky:2004:GE

[OP04] Martin Odersky and Benjamin C. Pierce. Guest editorial. ACM Transactions on Programming Languages and Systems, 26(5):767–768, September 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Oppen:1980:P

[Opp80] Dereck C. Oppen. Prettyprinting. ACM Transactions on Programming Languages and Systems, 2(4):465–483, October 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ossefort:1983:CPC

[Oss83] Marty Ossefort. Correctness proofs of communicating processes: Three illustrative examples from the literature. ACM Transactions on Programming Languages and Systems, 5(4):

620–640, October 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

OHearn:2009:SIH

[OYR09] Peter W. O’Hearn, Hongseok Yang, and John C. Reynolds. Separation and information hiding. ACM Transactions on Programming Languages and Systems, 31(3):11:1–11:50, April 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pingali:1985:EDD

[PA85] Keshav Pingali and Arvind. Efficient demand-driven evaluation. part 1. ACM Transactions on Programming Languages and Systems, 7(2):311– 333, April 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3480.html.

Pingali:1986:CFI

[PA86a] Keshav Pingali and Arvind.

Clarification of “Feeding inputs on demand” in Efficient demanddriven evaluation. part 1. ACM Transactions on Programming Languages and Systems, 8(1): 140–141, January 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5003.html.

Pingali:1986:EDD

[PA86b] Keshav Pingali and Arvind. Efficient demand-driven evaluation. part 2. ACM Transactions on Programming Languages and Systems, 8(1):109– 139, January 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Padovani:2019:CFS

[Pad19] Luca Padovani. Context-free session type inference. ACM Transactions on Programming Languages and Systems, 41(2): 9:1–9:??, June 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3229062.

Palsberg:1995:CAC

[Pal95] Jens Palsberg. Closure analysis in constraint form. ACM Transactions on Programming Languages and Systems, 17(1): 47–62, January 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        201001.html.

Palsberg:1998:EBF

[Pal98]  Jens Palsberg.  Equality-

based flow analysis versus recursive types. ACM Transactions on Programming Languages and Systems, 20(6):1251– 1264, November 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-6/
p1251-palsberg/.

Palsberg:2011:E

[Pal11a] Jens Palsberg. Editorial. ACM Transactions on Programming Languages and Systems, 33(1): 1:1, January 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Palsberg:2011:EN

[Pal11b] Jens Palsberg. Editorial note.

ACM Transactions on Programming Languages and Systems, 33 (6):18:1, December 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Palsberg:2012:E

[Pal12] Jens Palsberg. Editorial. ACM Transactions on Programming Languages and Systems, 34(1): 1:1, April 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Palsberg:2013:E

[Pal13] Jens Palsberg. Editorial. ACM Transactions on Programming Languages and Systems, 35(1): 1:1–1:??, April 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Palsberg:2015:E

[Pal15] Jens Palsberg. Editorial. ACM Transactions on Programming Languages and Systems, 37(1): 1:1–1:??, January 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Parnas:1990:TCI

[Par90]  David L. Parnas.  Techni-

cal correspondence: On iterative constructs. ACM Transactions on Programming Languages and Systems, 12(1):139– 141, January 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214517.html;   http://www.
acm.org/pubs/toc/Abstracts/
0164-0925/77613.html.

Patrignani:2015:SCP

[PAS^{+}15] Marco Patrignani, Pieter Agten, Raoul Strackx, Bart Jacobs, Dave Clarke, and Frank Piessens. Secure compilation to protected module architectures. ACM Transactions on Programming Languages and Systems, 37(2): 6:1–6:??, April 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Paulson:2001:MTP

[Pau01] Lawrence C. Paulson. Mechanizing a theory of program composition for UNITY. ACM Transactions on Programming Languages and Systems, 23(5):626– 656, September 2001. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Papadimitriou:1980:PBH

[PB80] Christos H. Papadimitriou and Philip A. Bernstein. On the performance of balanced hashing functions when the keys are not equiprobable. ACM Transactions on Programming Languages and Systems, 2(1): 77–89, January 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pingali:1997:OCD

[PB97] Keshav Pingali and Gianfranco Bilardi. Optimal control dependence computation and the Roman chariots problem. ACM Transactions on Programming Languages and Systems, 19(3): 462–491, May 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-3/p462-pingali/.

Paz:2007:EFC

[PBK^{+}07] Harel Paz, David F. Bacon, Elliot K. Kolodner, Erez Petrank, and V. T. Rajan. An efficient on-the-fly cycle collection. ACM Transactions on Programming Languages and Systems, 29(4):20:1–20:43, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Porter:2015:PFG

[PBR^{+}15] Donald E. Porter, Michael D.

Bond, Indrajit Roy, Kathryn S. Mckinley, and Emmett Witchel. Practical fine-grained information flow control using Laminar. ACM Transactions on Programming Languages and Systems, 37 (1):4:1–4:??, January 2015. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Park:1985:NAL

[PCC85] Joseph C. H. Park, K. M. Choe, and C. H. Chang. A new analysis of LALR formalisms. ACM Transactions on Programming Languages and Systems, 7(1): 159–175, January 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        2527.html.

Preda:2008:SBA

[PCJD08] Mila Dalla Preda, Mihai Christodorescu, Somesh Jha, and Saumya Debray. A semantics-based approach to malware detection. ACM Transactions on Programming Languages and Systems, 30(5):25:1–25:54, August 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pan:2008:PFE

[PE08] Zhelong Pan and Rudolf Eigenmann. PEAK — a fast and effective performance tuning system via compiler optimization orchestration. ACM Transactions on Programming Languages and Systems, 30(3):17:1– 17:43, May 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pearce:2021:LFR

[Pea21] David J. Pearce. A lightweight formalism for reference lifetimes and borrowing in Rust. ACM Transactions on Programming Languages and Systems, 43(1): 3:1–3:73, April 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3443420.

Pemberton:1983:TCT

[Pem83]  Steven Pemberton.  Techni-

cal correspondence: On Tanenbaum, van Staveren, and Stevenson’s “Using Peephole Optimization on Intermediate Code”. ACM Transactions on Programming Languages and Systems, 5 (3):499, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [TvS82, Tan83].

Perrott:1979:LAV

[Per79] R. H. Perrott. A language for array and vector processors. ACM Transactions on Programming Languages and Systems, 1 (2):177–195, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Perry:1990:GEI

[Per90] Dewayne E. Perry. Guest Editor’s introduction. ACM Transactions on Programming Languages and Systems, 12(4):641– 642, October 1990. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Peterson:1982:UAC

[Pet82] Gary L. Peterson. An O(n log n)

unidirectional algorithm for the circular extrema problem. ACM Transactions on Programming Languages and Systems, 4(4): 758–762, October 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). Peterson presents a deterministic distributed algorithm for finding the largest of a set of n uniquely numbered processes in a ring. The algorithm requires O(n log n) messages in the worst case, and is unidirectional. The number of processes is not initially known.

Peterson:1983:CRW

[Pet83a] Gary L. Peterson. Concurrent reading while writing. ACM Transactions on Programming Languages and Systems, 5(1): 46–55, January 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Peterson:1983:NSL

[Pet83b] Gary L. Peterson. A new solution to Lamport’s concurrent programming problem using small shared variables. ACM Transactions on Programming Languages and Systems, 5(1): 56–65, January 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Proebsting:1996:DDR

[PF96] Todd A. Proebsting and Charles N. Fischer. Demand-driven register allocation. ACM Transactions on Programming Languages and Systems, 18(6):683– 710, November 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
235829.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/236117.html.

Pratikakis:2011:LPS

[PFH11] Polyvios Pratikakis, Jeffrey S.

Foster, and Michael Hicks. LOCKSMITH: Practical static race detection for C. ACM Transactions on Programming Languages and Systems, 33(1): 3:1–3:55, January 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Patrignani:2021:RSC

[PG21] Marco Patrignani and Deepak Garg. Robustly safe compilation, an efficient form of secure compilation. ACM Transactions on Programming Languages and Systems, 43(1): 1:1–1:41, April 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3436809.

Poletto:1999:CTL

[PHEK99] Massimiliano Poletto, Wilson C.

Hsieh, Dawson R. Engler, and M. Frans Kaashoek. ’C and tcc: a language and compiler for dynamic code generation. ACM Transactions on Programming Languages and Systems, 21 (2):324–369, March 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-2/p324-poletto/.

Paek:2002:EPA

[PHP02] Yunheung Paek, Jay Hoeflinger, and David Padua. Efficient and precise array access analysis. ACM Transactions on Programming Languages and Systems, 24 (1):65–109, January 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pippenger:1997:PVI

[Pip97] Nicholas Pippenger. Pure versus impure Lisp. ACM Transactions on Programming Languages and Systems, 19(2):223– 238, March 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1997-19-2/p223-pippenger/.

Piquer:1996:IDG

[Piq96]  José M. Piquer.  Indirect

distributed garbage collection: Handling object migration. ACM Transactions on Programming Languages and Systems, 18(5):615–647, September 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL http://www.acm.
                                        org/pubs/toc/Abstracts/0164-
                                        0925/232711.html;   http:/
                                        /www.acm.org/pubs/toc/Abstracts/
                                        toplas/232711.html;   http:
                                        //www.acm.org/pubs/toc/Abstracts/
                                        toplas/235457.html.

Pai:1980:GCR

[PK80] Ajit B. Pai and Richard B.

Kieburtz. Global context recovery: a new strategy for syntactic error recovery by tabledrive parsers. ACM Transactions on Programming Languages and Systems, 2(1):18– 41, January 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Paige:1982:FDC

[PK82] Robert Paige and Shaye Koenig.

Finite differencing of computable expressions. ACM Transactions on Programming Languages and Systems, 4(3): 402–454, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pearce:2007:EFS

[PKH07] David J. Pearce, Paul H. J.

Kelly, and Chris Hankin. Efficient field-sensitive pointer analysis of C. ACM Transactions on Programming Languages and Systems, 30(1):4:1–4:42, November 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Patrignani:2024:UCR

[PKWC24] Marco Patrignani, Robert Künnemann, Riad S. Wahby, and Ethan Cecchetti. Universal composability is robust compilation. ACM Transactions on Programming Languages and Systems, 46(4): 13:1–13:??, December 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3698234.

Park:2004:ORC

[PM04] Jinpyo Park and Soo-Mook Moon. Optimistic register coalescing. ACM Transactions on Programming Languages and Systems, 26(4):735–765, July 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Payet:2006:NIL

[PM06] Etienne Payet and Fred Mesnard. Nontermination inference of logic programs. ACM Transactions on Programming Languages and Systems, 28(2): 256–289, March 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pingali:2009:RTP

[PM09] Keshav Pingali and Kathryn McKinley. Remembrances of things past. ACM Transactions on Programming Languages and Systems, 31(4):17:1– 17:2, May 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Palsberg:1995:TSE

[PO95] Jens Palsberg and Patrick O’Keefe. A type system equivalent to flow analysis. ACM Transactions on Programming Languages and Systems, 17(4): 576–599, July 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
210187.html.

Peng:1991:DFA

[PP91] Wuxu Peng and S. Purushothaman. Data flow analysis of communicating finite state machines. ACM Transactions on Programming Languages and Systems, 13(3):399–442, July 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

117015.html. See [Pur91].

Pinter:1994:POP

[PP94] Shlomit S. Pinter and Ron Y.

Pinter. Program optimization and parallelization using idioms. ACM Transactions on Programming Languages and Systems, 16(3):305–327, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177494.html.

Prywes:1979:UNS

[PPS79] N. S. Prywes, Amir Pnueli, and S. Shastry. Use of a nonprocedural specification language and associated program generator in software development. ACM Transactions on Programming Languages and Systems, 1(2): 196–217, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Park:2008:PLB

[PPT08] Sungwoo Park, Frank Pfenning, and Sebastian Thrun. A probabilistic language based on sampling functions. ACM Transactions on Programming Languages and Systems, 31(1):4:1– 4:46, December 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Podelski:2007:TPA

[PR07] Andreas Podelski and Andrey Rybalchenko. Transition predicate abstraction and fair termination. ACM Transactions on Programming Languages and Systems, 29(3):15:1–15:31, May 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Proebsting:1995:BAG

[Pro95] Todd A. Proebsting. BURS automata generation. ACM Transactions on Programming Languages and Systems, 17(3): 461–486, May 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        203098.html.

Pollock:1992:IGR

[PS92] Lori L. Pollock and Mary Lou Soffa. Incremental global reoptimization of programs. ACM Transactions on Programming Languages and Systems, 14(2): 173–200, April 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
128865.html.

Palem:1993:STC

[PS93] Krishna V. Palem and Barbara B. Simons. Scheduling time-critical instructions on RISC machines. ACM Transactions on Programming Languages and Systems, 15(4):632– 658, September 1993. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
155190.html.

Palsberg:1996:CTT

[PS96] Jens Palsberg and Scott Smith.

Constrained types and their expressiveness. ACM Transactions on Programming Languages and Systems, 18(5):519– 527, September 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
232715.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/232715.html;   http:
//www.acm.org/pubs/toc/Abstracts/
toplas/235454.html.

Poletto:1999:LSR

[PS99] Massimiliano Poletto and Vivek Sarkar. Linear scan register allocation. ACM Transactions on Programming Languages and Systems, 21(5):895– 913, September 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/articles/journals/toplas/
                                        1999-21-5/p895-poletto/p895-
                                        poletto.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        toplas/1999-21-5/p895-poletto/

.

Pottier:2003:IFI

[PS03] François Pottier and Vincent Simonet. Information flow inference for ML. ACM Transactions on Programming Languages and Systems, 25(1):117–158, January 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pearlmutter:2008:RMA

[PS08] Barak A. Pearlmutter and Jeffrey Mark Siskind. Reversemode AD in a functional framework: Lambda the ultimate backpropagator. ACM Transactions on Programming Languages and Systems, 30(2):7:1– 7:36, March 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pottier:2005:SAS

[PSS05] François Pottier, Christian Skalka, and Scott Smith. A systematic approach to static access control. ACM Transactions on Programming Languages and Systems, 27(2):344–382, March 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Pierce:2000:LTI

[PT00] Benjamin C. Pierce and David N. Turner. Local type inference.

ACM Transactions on Programming Languages and Systems, 22 (1):1–44, January 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-1/p1-pierce/.

Purushothaman:1991:CDF

[Pur91] S. Purushothaman. Corrigendum: “Data Flow Analysis of Communicating Finite State Machines”. ACM Transactions on Programming Languages and Systems, 13(4):633, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [PP91].

Purtilo:1994:PSB

[Pur94]  James M. Purtilo.  The

POLYLITH software bus. ACM Transactions on Programming Languages and Systems, 16(1): 151–174, January 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
174629.html.

Pugh:1994:SAU

[PW94] William Pugh and David Wonnacott. Static analysis of upper and lower bounds on dependences and parallelism. ACM Transactions on Programming Languages and Systems, 16(4): 1248–1278, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

                                        pubs/toc/Abstracts/0164-0925/
                                        183525.html.

Pugh:1998:CBA

[PW98] William Pugh and David Wonnacott. Constraint-based array dependence analysis. ACM Transactions on Programming Languages and Systems, 20(3): 635–678, May 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-3/
                                        p635-pugh/.

Palsberg:1995:EIA

[PXL95] Jens Palsberg, Cun Xiao, and Karl Lieberherr. Efficient implementation of adaptive software. ACM Transactions on Programming Languages and Systems, 17 (2):264–292, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        201066.html.

Pagel:2022:SSL

[PZ22] Jens Pagel and Florian Zuleger. Strong-separation logic. ACM Transactions on Programming Languages and Systems, 44(3): 16:1–16:40, September 2022. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3498847.

Palsberg:2005:ADC

[PZJ05] Jens Palsberg, Tian Zhao, and Trevor Jim. Automatic discovery of covariant read-only fields. ACM Transactions on Programming Languages and Systems, 27 (1):126–162, January 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Qian:1995:CRO

[QG95] Xiaolei Qian and Allen Goldberg. Corrigendum: “Referential Opacity in Nondeterministic Data Refinement”. ACM Transactions on Programming Languages and Systems, 17(1): 180, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See ACM Lett. Program. Lang. Syst. 2, 1–4 (Mar.–Dec. 1993), 233–241.

Qian:2000:SFI

[Qia00] Zhenyu Qian. Standard fixpoint iteration for Java bytecode verification. ACM Transactions on Programming Languages and Systems, 22(4): 638–672, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-4/p638-qian/.

Quong:1991:LPI

[QL91] Russell W. Quong and Mark A.

Linton. Linking programs incrementally. ACM Transactions on Programming Languages and Systems, 13(1):1– 20, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        102804.html.

Quillere:2000:OMU

[QR00] Fabien Quilleré and Sanjay Rajopadhye. Optimizing memory usage in the polyhedral model. ACM Transactions on Programming Languages and Systems, 22(5):773–815, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/2000-22-5/
                                        p773-quillere/p773-quillere.
                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-5/p773-quillere/.

Ranganath:2007:NFC

[RAB^{+}07] Venkatesh Prasad Ranganath, Torben Amtoft, Anindya Banerjee, John Hatcliff, and Matthew B. Dwyer. A new foundation for control dependence and slicing for modern program structures. ACM Transactions on Programming Languages and Systems, 29(5):27:1–27:43, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ramalingam:1994:UA

[Ram94] G. Ramalingam. The undecidability of aliasing. ACM Transactions on Programming Languages and Systems, 16(5):1467– 1471, September 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
186041.html.

Ramalingam:1999:ILA

[Ram99] G. Ramalingam. Identifying loops in almost linear time. ACM Transactions on Programming Languages and Systems, 21 (2):175–188, March 1999. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1999-21-2/p175-ramalingam/.

Ramalingam:2000:CSS

[Ram00]  G. Ramalingam.  Context-

sensitive synchronization-sensitive analysis is undecidable. ACM Transactions on Programming Languages and Systems, 22(2): 416–430, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-2/p416-ramalingam/.

Rao:1994:RAP

[Rao94]  Josyula R. Rao.  Reason-

ing about probabilistic parallel programs. ACM Transactions on Programming Languages and Systems, 16(3):798– 842, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177724.html.

Rasch:2024:RCD

[Ras24] Ari Rasch. (De/ re)-composition of data-parallel computations via multi-dimensional homomorphisms. ACM Transactions on Programming Languages and Systems, 46(3):10:1– 10:??, September 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3665643.

Reiter:1994:HSR

[RB94] Michael K. Reiter and Kenneth P. Birman. How to securely replicate services. ACM Transactions on Programming Languages and Systems, 16(3): 986–1009, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        177745.html.

Ramsey:2003:TAB

[RC03] Norman Ramsey and Cristina Cifuentes. A transformational approach to binary translation of delayed branches. ACM Transactions on Programming Languages and Systems, 25(2): 210–224, March 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rogers:1995:SDD

[RCRH95] Anne Rogers, Martin C. Carlisle, John H. Reppy, and L. J. Hendren. Supporting dynamic data structures on distributedmemory machines. ACM Transactions on Programming Languages and Systems, 17(2):233– 263, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201065.html.

Richardson:1993:DPL

[RCS93] Joel E. Richardson, Michael J.

Carey, and Daniel T. Schuh. The design of the E programming language. ACM Transactions on Programming Languages and Systems, 15(3): 494–534, July 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
174157.html.

Reps:1987:SSE

[RD87] Thomas Reps and Alan Demers.

Sublinear-space evaluation algorithms for attribute grammars. ACM Transactions on Programming Languages and Systems, 9 (3):408–440, July 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214529.html.

Rinard:1997:CAN

[RD97] Martin C. Rinard and Pedro C.

Diniz. Commutativity analysis: a new analysis technique for parallelizing compilers. ACM Transactions on Programming Languages and Systems, 19(6): 942–991, November 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1997-19-6/
                                        p942-rinard/.

Rinard:2003:ESB

[RD03] Martin C. Rinard and Pedro C. Diniz. Eliminating synchronization bottlenecks using adaptive replication. ACM Transactions on Programming Languages and Systems, 25(3):316– 359, May 2003. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rossberg:2013:MMM

[RD13] Andreas Rossberg and Derek Dreyer. Mixin’ up the ML module system. ACM Transactions on Programming Languages and Systems, 35(1):2:1– 2:??, April 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rong:2008:RAS

[RDG08] Hongbo Rong, Alban Douillet, and Guang R. Gao. Register allocation for software pipelined multidimensional loops. ACM Transactions on Programming Languages and Systems, 30(4): 23:1–23:68, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Reiss:1983:GCS

[Rei83] Steven P. Reiss. Generation of compiler symbol processing mechanisms from specifications.

ACM Transactions on Programming Languages and Systems, 5 (2):127–163, April 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rem:1981:APN

[Rem81] Martin Rem. Associons: a program notation with tuples instead of variables. ACM Transactions on Programming Languages and Systems, 3(3): 251–262, July 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Reps:1986:GEI

[Rep86] Thomas Reps. Guest Editor’s introduction. ACM Transactions on Programming Languages and Systems, 8(4):417–418, October 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Reps:1998:MMT

[Rep98]  Thomas Reps.  “Maximal-

munch” tokenization in linear time. ACM Transactions on Programming Languages and Systems, 20(2):259–273, March 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-2/
p259-reps/.

Reps:2000:UCS

[Rep00] Thomas Reps. Undecidability of context-sensitive dataindependence analysis. ACM Transactions on Programming Languages and Systems, 22(1): 162–186, January 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-1/p162-reps/.

Ramsey:1997:SRM

[RF97] Norman Ramsey and Mary F. Fernández. Specifying representations of machine instructions. ACM Transactions on Programming Languages and Systems, 19(3):492–524, May 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-3/p492-ramsey/. This

article was written using the noweb literate programming system.

Rosenkrantz:1987:EAA

[RH87] D. J. Rosenkrantz and H. B. Hunt. Efficient algorithms for automatic construction and compactification of parsing grammars. ACM Transactions on Programming Languages and Systems, 9(4):543– 566, October 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
29876.html.

Rhiger:2003:FEL

[Rhi03] Morten Rhiger. A foundation for embedded languages. ACM Transactions on Programming Languages and Systems, 25 (3):291–315, May 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Richter:1985:NSE

[Ric85] Helmut Richter. Noncorrecting syntax error recovery. ACM Transactions on Programming Languages and Systems, 7(3): 478–489, July 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
4019.html.

Roychoudhury:2004:UFT

[RKRR04] Abhik Roychoudhury, K. Narayan Kumar, C. R. Ramakrishnan, and I. V. Ramakrishnan. An unfold/fold transformation framework for definite logic programs. ACM Transactions on Programming Languages and Systems, 26(3):464–509, May 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Renganarayanan:2012:PLT

[RKSR12] Lakshminarayanan Renganarayanan, Daegon Kim, Michelle Mills Strout, and Sanjay Rajopadhye. Parameterized loop tiling. ACM Transactions on Programming Languages and Systems, 34 (1):3:1–3:41, April 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Richter:2023:PTL

[RKW^{+}23] David Richter, David Kretzler, Pascal Weisenburger, Guido Salvaneschi, Sebastian Faust, and Mira Mezini. Prisma : a tierless language for enforcing contractclient protocols in decentralized applications. ACM Transactions on Programming Languages and Systems, 45(3):17:1– 17:??, September 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3604629.

Rinard:1998:DIE

[RL98] Martin C. Rinard and Monica S. Lam. The design, implementation, and evaluation of Jade. ACM Transactions on Programming Languages and Systems, 20 (3):483–545, May 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-3/
p483-rinard/.

Ryder:2001:SIM

[RLS^{+}01] Barbara G. Ryder, William A. Landi, Philip A. Stocks, Sean Zhang, and Rita Altucher. A schema for interprocedural modification side-effect analysis with pointer aliasing. ACM Transactions on Programming Languages and Systems, 23(2):105– 186, March 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2001-23-1/p105-ryder/.

Rival:2007:TPA

[RM07] Xavier Rival and Laurent Mauborgne. The trace partitioning abstract domain. ACM Transactions on Programming Languages and Systems, 29(5): 26:1–26:51, August 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ruggieri:2010:TLC

[RM10] Salvatore Ruggieri and Fred Mesnard. Typing linear constraints. ACM Transactions on Programming Languages and Systems, 32(6):21:1–21:42, August 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rosenkrantz:2006:MMA

[RMH06] Daniel J. Rosenkrantz, Lenore R.

Mullin, and Harry B. Hunt III. On minimizing materializations of array-valued temporaries. ACM Transactions on Programming Languages and Systems, 28 (6):1145–1177, November 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Robertson:1979:CGS

[Rob79] Edward L. Robertson. Code generation and storage allocation for machines with spandependent instructions. ACM Transactions on Programming Languages and Systems, 1(1): 71–83, July 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ryder:1988:IDF

[RP88] Barbara G. Ryder and Marvin C. Paull. Incremental dataflow analysis algorithms. ACM Transactions on Programming Languages and Systems, 10(1): 1–50, January 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
42193.html.

Rugina:2003:PAS

[RR03] Radu Rugina and Martin C. Rinard. Pointer analysis for structured parallel programs. ACM Transactions on Programming Languages and Systems, 25(1): 70–116, January 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rugina:2005:SBA

[RR05] Radu Rugina and Martin C. Rinard. Symbolic bounds analysis of pointers, array indices, and accessed memory regions. ACM Transactions on Programming Languages and Systems, 27 (2):185–235, March 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Rosa:2019:AOT

[RRB19] Andrea Rosà, Eduardo Rosales, and Walter Binder. Analysis and optimization of task granularity on the Java Virtual Machine. ACM Transactions on Programming Languages and Systems, 41 (3):19:1–19:??, July 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3338497.

Rinetzky:2008:CPF

[RRSY08] N. Rinetzky, G. Ramalingam, M. Sagiv, and E. Yahav. On the complexity of partially-flow-sensitive alias analysis. ACM Transactions on Programming Languages and Systems, 30(3): 13:1–13:28, May 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ramanath:1984:JML

[RS84a] M. V. S. Ramanath and Marvin Solomon. Jump minimization in linear time. ACM Transactions on Programming Languages and Systems, 6(4):527–545, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Reif:1984:RTS

[RS84b] John H. Reif and Paul G. Spirakis. Real-time synchronization of interprocess communications. ACM Transactions on Programming Languages and Systems, 6 (2):215–238, April 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). They present probabilistic distributed algorithms for the guard-scheduling problem (Section ??) that guarantee real-time response. A preliminary version of this paper appeared as “Distributed Algorithms for Synchronizing Interprocess Communication in Real Time,” in Proc. 13th Ann. ACM Symp. on Theory of Computing, 1981.

Raja:1997:CFC

[RS97] N. Raja and R. K. Shyamasundar. Combinatory formulations of concurrent languages. ACM Transactions on Programming Languages and Systems, 19(6): 899–915, November 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1997-19-6/
                                        p899-raja/.

Reps:2010:FDL

[RSL10] Thomas Reps, Mooly Sagiv, and Alexey Loginov. Finite differencing of logical formulas for static analysis. ACM Transactions on Programming Languages and Systems, 32(6):24:1–24:55, August 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Runge:2023:IES

[RSPS23] Tobias Runge, Marco Servetto, Alex Potanin, and Ina Schaefer. Immutability and encapsulation for sound OO information flow control. ACM Transactions on Programming Languages and Systems, 45(1): 3:1–3:??, March 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3573270.

Reps:1983:ICD

[RTD83] Thomas Reps, Tim Teitelbaum, and Alan Demers. Incremental context-dependent analysis for language-based editors. ACM Transactions on Programming Languages and Systems, 5(3): 449–477, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Reps:2017:NPA

[RTP17] Thomas Reps, Emma Turetsky, and Prathmesh Prabhu. Newtonian program analysis via tensor product. ACM Transactions on Programming Languages and Systems, 39(2):9:1– 9:??, May 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Reed:1988:SVL

[RY88] Joylyn Reed and Raymond T.

Yeh. Specification and verification of liveness properties of cyclic, concurrent processes. ACM Transactions on Programming Languages and Systems, 10 (1):156–177, January 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
42195.html.

Ryu:2016:TOO

[Ryu16] Sukyoung Ryu. ThisType for object-oriented languages: From theory to practice. ACM Transactions on Programming Languages and Systems, 38(3):8:1– 8:??, May 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Stata:1999:TSJ

[SA99] Raymie Stata and Martin Abadi. A type system for Java bytecode subroutines. ACM Transactions on Programming Languages and Systems, 21(1):90– 137, January 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        1999-21-1/p90-stata/.

Shao:2000:ESS

[SA00] Zhong Shao and Andrew W. Appel. Efficient and safe-for-space closure conversion. ACM Transactions on Programming Languages and Systems, 22(1): 129–161, January 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-1/p129-shao/.

Sager:1986:SPC

[Sag86] Thomas J. Sager. A short proof of a conjecture of DeRemer and Pennello. ACM Transactions on Programming Languages and Systems, 8(2):264– 271, April 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        30850.html.

Sagiv:2007:ISE

[Sag07] Mooly Sagiv. Introduction to special ESOP’05 issue. ACM Transactions on Programming Languages and Systems, 29(5): 23:1–23:2, August 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Samet:1980:CAP

[Sam80] Hanan Samet. A coroutine approach to parsing. ACM Transactions on Programming Languages and Systems, 2(3): 290–306, July 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sands:1996:TCL

[San96] David Sands. Total correctness by local improvement in the transformation of functional programs. ACM Transactions on Programming Languages and Systems, 18(2):175– 234, March 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
227716.html.

Sangiorgi:2009:OBC

[San09] Davide Sangiorgi. On the origins of bisimulation and coinduction. ACM Transactions on Programming Languages and Systems, 31 (4):15:1–15:41, May 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Solovyev:2019:REF

[SBB^{+}19] Alexey Solovyev, Marek S. Baranowski, Ian Briggs, Charles Jacobsen, Zvonimir Rakamarić, and Ganesh Gopalakrishnan. Rigorous estimation of floatingpoint round-off errors with symbolic Taylor expansions. ACM Transactions on Programming Languages and Systems, 41(1): 2:1–2:??, March 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3230733.

Spoto:2019:SII

[SBE^{+}19] Fausto Spoto, Elisa Burato, Michael D. Ernst, Pietro Ferrara, Alberto Lovato, Damiano Macedonio, and Ciprian Spiridon. Static identification of injection attacks in Java. ACM Transactions on Programming Languages and Systems, 41(3): 18:1–18:??, July 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3332371.

Shi:2022:TCP

[SBS22] Kensen Shi, David Bieber, and Rishabh Singh. TF-Coder: Program synthesis for tensor manipulations. ACM Transactions on Programming Languages and Systems, 44(2): 10:1–10:36, June 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3517034.

Schwartz:1980:U

[Sch80]  Jacob T. Schwartz.  Ultra-

computers. ACM Transactions on Programming Languages and Systems, 2(4):484–521, October 1980. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Schneider:1982:SDP

[Sch82] Fred B. Schneider. Synchronization in distributed programs. ACM Transactions on Programming Languages and Systems, 4 (2):125–148, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Schmidt:1985:DGV

[Sch85] David A. Schmidt. Detecting global variables in denotational specifications. ACM Transactions on Programming Languages and Systems, 7(2): 299–310, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3323.html.

Soares:2023:SCE

[SCP23] Luigi Soares, Michael Canesche, and Fernando Magno Quintão Pereira. Side-channel elimination via partial control-flow linearization. ACM Transactions on Programming Languages and Systems, 45(2): 13:1–13:??, June 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3594736.

Sistla:2024:CCF

[SCR24] Meghana Aparna Sistla, Swarat Chaudhuri, and Thomas Reps. CFLOBDDs: Context-free-language ordered binary decision diagrams. ACM Transactions on Programming Languages and Systems, 46(2): 7:1–7:??, June 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3651157.

Stein:2024:IAI

[SCS24] Benno Stein, Bor-Yuh Evan Chang, and Manu Sridharan. Interactive abstract interpretation with demanded summarization. ACM Transactions on Programming Languages and Systems, 46 (1):4:1–4:??, March 2024. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3648441.

Skorstengaard:2020:RAM

[SDB20] Lau Skorstengaard, Dominique Devriese, and Lars Birkedal. Reasoning about a machine with local capabilities: Provably safe stack and return pointer management. ACM Transactions on Programming Languages and Systems, 42(1):1– 53, January 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3363519.

Swalens:2021:CCC

[SDD21] Janwillem Swalens, Joeri De Koster, and Wolfgang De Meuter. Chocola: Composable concurrency language. ACM Transactions on Programming Languages and Systems, 42(4): 17:1–17:56, February 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3427201.

Sampaio:2013:DA

[SdSCP13] Diogo Sampaio, Rafael Martins de Souza, Sylvain Collange, and Fernando Magno Quintão Pereira. Divergence analysis. ACM Transactions on Programming Languages and Systems, 35 (4):13:1–13:??, December 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Strickland:2013:CFC

[SDTF13] T. Stephen Strickland, Christos Dimoulas, Asumu Takikawa, and Matthias Felleisen. Contracts for first-class classes. ACM Transactions on Programming Languages and Systems, 35 (3):11:1–11:58, November 2013. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sethi:1983:CFA

[Set83] Ravi Sethi. Control flow aspects of semantics-directed compiling.

ACM Transactions on Programming Languages and Systems, 5 (4):554–595, October 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/abs/10.1145/69575.
357227.

Stamos:1990:RE

[SG90] James W. Stamos and David K. Gifford. Remote evaluation. ACM Transactions on Programming Languages and Systems, 12 (4):537–565, October 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
88631.html.

Sistla:2004:SRS

[SG04] A. Prasad Sistla and Patrice Godefroid. Symmetry and reduced symmetry in model checking. ACM Transactions on Programming Languages and Systems, 26(4):702–734, July 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sreedhar:1996:ILU

[SGL96] Vugranam C. Sreedhar, Guang R. Gao, and Yong-Fong Lee. Identifying loops using DJ graphs. ACM Transactions on Programming Languages and Systems, 18(6):649–658, November 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/

235827.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/236115.html.

Sreedhar:1997:ICD

[SGL97] Vugranam C. Sreedhar, Guang R.

Gao, and Yong-Fong Lee. Incremental computation of dominator trees. ACM Transactions on Programming Languages and Systems, 19(2):239– 252, March 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-2/p239-sreedhar/.

Sreedhar:1998:NFE

[SGL98] Vugranam C. Sreedhar, Guang R.

Gao, and Yong-Fong Lee. A new framework for eliminationbased data flow analysis using DJ graphs. ACM Transactions on Programming Languages and Systems, 20(2):388–435, March 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-2/
p388-sreedhar/.

Steenkiste:1989:SIR

[SH89] Peter A. Steenkiste and John L.

Hennessy. A simple interprocedural register allocation algorithm and its effectiveness for Lisp. ACM Transactions on Programming Languages and Systems, 11(1):1– 32, January 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        59289.html.

Sharir:1982:SOC

[Sha82] Micha Sharir. Some observations concerning formal differentiation of set theoretic expressions. ACM Transactions on Programming Languages and Systems, 4 (2):196–225, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Stoyle:2007:MMS

[SHB^{+}07] Gareth Stoyle, Michael Hicks, Gavin Bierman, Peter Sewell, and Iulian Neamtiu. Mutatis Mutandis: Safe and predictable dynamic software updating. ACM Transactions on Programming Languages and Systems, 29(4):22:1–22:70, August 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sheard:1991:AGU

[She91] Tim Sheard. Automatic generation and use of abstract structure operators. ACM Transactions on Programming Languages and Systems, 13(4):531– 557, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        115369.html.

Sekiyama:2017:PMC

[SIG17] Taro Sekiyama, Atsushi Igarashi, and Michael Greenberg. Polymorphic manifest contracts, revised and resolved. ACM Transactions on Programming Languages and Systems, 39(1):3:1– 3:??, March 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sijtsma:1989:PRL

[Sij89]  Ben A. Sijtsma.  On the

Productivity of Recursive List Definitions. ACM Transactions on Programming Languages and Systems, 11(4):633– 649, October 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
69563.html.

Sipala:1982:CSB

[Sip82] Paolo Sipala. Compact storage of binary trees. ACM Transactions on Programming Languages and Systems, 4(3): 345–361, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sites:1979:CLI

[Sit79] Richard L. Sites. The compilation of loop induction expressions. ACM Transactions on Programming Languages and Systems, 1(1):50–57, July 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Spoto:2003:CAA

[SJ03] Fausto Spoto and Thomas Jensen. Class analyses as abstract interpretations of trace semantics. ACM Transactions on Programming Languages and Systems, 25(5):578– 630, September 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Scott:2006:RNG

[SJ06] Elizabeth Scott and Adrian Johnstone. Right nulled GLR parsers. ACM Transactions on Programming Languages and Systems, 28(4):577–618, July 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Smans:2012:IDF

[SJP12] Jan Smans, Bart Jacobs, and Frank Piessens. Implicit dynamic frames. ACM Transactions on Programming Languages and Systems, 34(1):2:1– 2:58, April 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Scott:2023:MIP

[SJW23] Elizabeth Scott, Adrian Johnstone, and Robert Walsh. Multiple input parsing and lexical analysis. ACM Transactions on Programming Languages and Systems, 45(3):14:1– 14:??, September 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3594734.

Schwanke:1988:SR

[SK88] Robert W. Schwanke and Gail E. Kaiser. Smarter recompilation.

ACM Transactions on Programming Languages and Systems, 10 (4):627–632, October 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214505.html.  See remarks

[Tic88].

Sangiorgi:2011:EBH

[SKS11] Davide Sangiorgi, Naoki Kobayashi, and Eijiro Sumii. Environmental bisimulations for higher-order languages. ACM Transactions on Programming Languages and Systems, 33(1):5:1–5:69, January 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Skudlarek:1995:NMI

[Sku95] Joseph P. Skudlarek. Notes on “A Methodology for Implementing Highly Concurrent Data Objects”. ACM Transactions on Programming Languages and Systems, 17(1):45– 46, January 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
200999.html.

Shankar:1992:SRH

[SL92] A. Udaya Shankar and Simon S. Lam. A stepwise refinement heuristic for protocol construction. ACM Transactions on Programming Languages and Systems, 14(3): 417–461, July 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        129394.html.

Schultz:2003:APS

[SLC03] Ulrik P. Schultz, Julia L. Lawall, and Charles Consel. Automatic program specialization for Java. ACM Transactions on Programming Languages and Systems, 25 (4):452–499, July 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sloane:1995:EAG

[Slo95] Anthony M. Sloane. An evaluation of an automatically generated compiler. ACM Transactions on Programming Languages and Systems, 17(5):691– 703, September 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        213980.html.

Steensgaard-Madsen:1981:SOA

[SM81]  J. Steensgaard-Madsen.  A

statement-oriented approach to data abstraction. ACM Transactions on Programming Languages and Systems, 3(1):1– 10, January 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remarks [Ell82, SM82].

Steensgaard-Madsen:1982:TCS

[SM82] J. Steensgaard-Madsen. Technical correspondence: Steensgaard-Madsen’s reply. ACM Transactions on Programming Languages and Systems, 4(1):122– 123, January 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [SM81, Ell82].

Steensgaard-Madsen:1989:TRO

[SM89] J. Steensgaard-Madsen. Type representation of objects by functions. ACM Transactions on Programming Languages and Systems, 11(1):67– 89, January 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
77345.html.

Spoto:2010:TAJ

[SMP10] Fausto Spoto, Fred Mesnard, and Étienne Payet. A termination analyzer for Java bytecode based on path-length. ACM Transactions on Programming Languages and Systems, 32(3): 8:1–8:70, March 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Stork:2014:APB

[SNS^{+}14] Sven Stork, Karl Naden, Joshua Sunshine, Manuel Mohr, Alcides Fonseca, Paulo Marques, and Jonathan Aldrich. Æminium: a permission-based concurrent-by-default programming language approach. ACM Transactions on Programming Languages and Systems, 36(1):2:1–2:??, March 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sokolowski:1987:SHL

[Sok87] Stefan Sokołowski. Soundness of Hoare’s logic: An automated proof using LCF. ACM Transactions on Programming Languages and Systems, 9(1): 100–120, January 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        11326.html.

Solworth:1992:E

[Sol92] Jon A. Solworth. Epochs. ACM Transactions on Programming Languages and Systems, 14(1): 28–53, January 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        116785.html.

Sonnenschein:1987:GTS

[Son87] Michael Sonnenschein. Graph translation schemes to generate compiler parts. ACM Transactions on Programming Languages and Systems, 9(4):473– 490, October 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        29874.html.

Sorkin:1989:TCS

[Sor89] Arthur Sorkin. Technical correspondence: Some comments on “A Solution to a Problem with Morel and Renvoise’s “Global Optimization by Suppression of Partial Redundancies””. ACM Transactions on Programming Languages and Systems, 11(4): 666–668, October 1989. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214513.html. See [DS88].

Soundararajan:1984:ASC

[Sou84] N. Soundararajan. Axiomatic semantics of communicating sequential processes. ACM Transactions on Programming Languages and Systems, 6(4):647– 662, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sansom:1997:FBP

[SP97] Patrick M. Sansom and Simon L.

Peyton Jones. Formally based profiling for higher-order functional languages. ACM Transactions on Programming Languages and Systems, 19(2):334– 385, March 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-2/p334-sansom/.

Simonet:2007:CBA

[SP07] Vincent Simonet and François Pottier. A constraint-based approach to guarded algebraic data types. ACM Transactions on Programming Languages and Systems, 29(1):1:1–1:56, January 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Spooner:1986:MAR

[Spo86] C. R. Spooner. The ML approach to the readable allpurpose language. ACM Transactions on Programming Languages and Systems, 8(2):215– 243, April 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        5918.html.

Sekar:1995:FSA

[SR95] R. Sekar and I. V. Ramakrishnan. Fast strictness analysis based on demand propagation. ACM Transactions on Programming Languages and Systems, 17(6):896–937, November 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        218573.html.

Shen:2021:ALI

[SR21] Jiasi Shen and Martin C. Rinard. Active learning for inference and regeneration of applications that access databases. ACM Transactions on Programming Languages and Systems, 42 (4):18:1–18:119, February 2021. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3430952.

Suhendra:2010:SAC

[SRM10] Vivy Suhendra, Abhik Roychoudhury, and Tulika Mitra. Scratchpad allocation for concurrent embedded software. ACM Transactions on Programming Languages and Systems, 32 (4):13:1–13:47, April 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sagiv:1998:SSA

[SRW98] Mooly Sagiv, Thomas Reps, and Reinhard Wilhelm. Solving shape-analysis problems in languages with destructive updating. ACM Transactions on Programming Languages and Systems, 20(1):1–50, January 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-1/
p1-sagiv/.

Sagiv:2002:PSA

[SRW02] Mooly Sagiv, Thomas Reps, and Reinhard Wilhelm. Parametric shape analysis via 3-valued logic. ACM Transactions on Programming Languages and Systems, 24 (3):217–298, May 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Soisalon-Soininen:1982:IEE

[SS82] Eljas Soisalon-Soininen. Inessential error entries and their use in LR parser optimization. ACM Transactions on Programming Languages and Systems, 4(2):

179–195, April 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [LaL84].

Schlichting:1984:UMP

[SS84] Richard D. Schlichting and Fred B. Schneider. Using message passing for distributed programming: Proof rules and disciplines. ACM Transactions on Programming Languages and Systems, 6(3):402–431, July 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Shasha:1988:ECE

[SS88] Dennis Shasha and Marc Snir. Efficient and correct execution of parallel programs that share memory. ACM Transactions on Programming Languages and Systems, 10(2):282– 312, April 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
42277.html.

Skeppstedt:1996:UDA

[SS96] Jonas Skeppstedt and Per Stenström. Using dataflow analysis techniques to reduce ownership overhead in cache coherence protocols. ACM Transactions on Programming Languages and Systems, 18(6):659– 682, November 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/

pubs/toc/Abstracts/toplas/
235828.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/236116.html.

Sagonas:1998:AMT

[SS98] Konstantinos Sagonas and Terrance Swift. An abstract machine for tabled execution of fixed-order stratified logic programs. ACM Transactions on Programming Languages and Systems, 20(3):586–634, May 1998. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

acm.org:80/pubs/citations/
journals/toplas/1998-20-3/
p586-sagonas/.

Schulte:2005:WDB

[SS05a] Christian Schulte and Peter J.

Stuckey. When do bounds and domain propagation lead to the same search space? ACM Transactions on Programming Languages and Systems, 27(3): 388–425, May 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Stuckey:2005:TO

[SS05b] Peter J. Stuckey and Martin Sulzmann. A theory of overloading. ACM Transactions on Programming Languages and Systems, 27(6):1216–1269, November 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Schulte:2008:ECP

[SS08] Christian Schulte and Peter J. Stuckey. Efficient constraint propagation engines. ACM Transactions on Programming Languages and Systems, 31(1): 2:1–2:43, December 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See erratum [SS09].

Schulte:2009:EEC

[SS09] Christian Schulte and Peter J. Stuckey. Erratum to: Efficient constraint propagation engines. ACM Transactions on Programming Languages and Systems, 31(2):1, February 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [SS08].

Staiger-Stohr:2013:PIA

[SS13] Stefan Staiger-Stöhr. Practical integrated analysis of pointers, dataflow and control flow. ACM Transactions on Programming Languages and Systems, 35 (1):5:1–5:??, April 2013. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sneyers:2009:CPC

[SSD09] Jon Sneyers, Tom Schrijvers, and Bart Demoen. The computational power and complexity of constraint handling rules. ACM Transactions on Programming Languages and Systems, 31(2):8:1–8:42, February 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sanchez-Stern:2023:PIA

[SSFZ^{+}23] Alex Sanchez-Stern, Emily First, Timothy Zhou, Zhanna Kaufman, Yuriy Brun, and Talia Ringer. Passport: Improving automated formal verification using identifiers. ACM Transactions on Programming Languages and Systems, 45(2): 12:1–12:??, June 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3593374.

Schonberg:1981:ATS

[SSS81] Edmond Schonberg, Jacob T.

Schwartz, and Micha Sharir. An automatic technique for selection of data structures in SETL programs. ACM Transactions on Programming Languages and Systems, 3(2):126– 143, April 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Sippu:1983:SEH

[SSS83] Seppo Sippu and Eljas Soisalon-Soininen. A syntax-error-handling technique and its experimental analysis. ACM Transactions on Programming Languages and Systems, 5(4): 656–679, October 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Snelting:2000:UCH

[ST00a] Gregor Snelting and Frank Tip. Understanding class hierarchies using concept analysis. ACM Transactions on Programming Languages and Systems, 22(3):540–582, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-3/
p540-snelting/p540-snelting.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-3/p540-snelting/.

Sperber:2000:GLP

[ST00b] Michael Sperber and Peter Thiemann. Generation of LR parsers by partial evaluation. ACM Transactions on Programming Languages and Systems, 22(2): 224–264, March 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-2/p224-sperber/.

Steimann:2018:CBR

[Ste18] Friedrich Steimann. Constraintbased refactoring. ACM Transactions on Programming Languages and Systems, 40(1):2:1– 2:??, January 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Steimann:2022:CPS

[Ste22] Friedrich Steimann. Containerless plurals: Separating number from type in object-oriented programming. ACM Transactions on Programming Languages and Systems, 44(4):21:1– 21:??, December 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527635.

Stone:2004:EOL

[Sto04] Christopher A. Stone. Extensible objects without labels. ACM Transactions on Programming Languages and Systems, 26(5): 805–835, September 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Saha:2003:IAQ

[STS03] Bratin Saha, Valery Trifonov, and Zhong Shao. Intensional analysis of quantified types. ACM Transactions on Programming Languages and Systems, 25 (2):159–209, March 2003. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Shao:2005:TSC

[STSP05] Zhong Shao, Valery Trifonov, Bratin Saha, and Nikolaos Papaspyrou. A type system for certified binaries. ACM Transactions on Programming Languages and Systems, 27(1):1– 45, January 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Smith:1996:PTV

[SV96] Geoffrey Smith and Dennis Volpano. Polymorphic typing of variables and references. ACM Transactions on Programming Languages and Systems, 18(3): 254–267, May 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
229544.html.

Sangiorgi:2019:EBP

[SV19] Davide Sangiorgi and Valeria Vignudelli. Environmental bisimulations for probabilistic higherorder languages. ACM Transactions on Programming Languages and Systems, 41(4):22:1– 22:??, November 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3350618.

Simpson:2020:BEM

[SV20] Alex Simpson and Niels Voorneveld. Behavioural equivalence via modalities for algebraic effects. ACM Transactions on Programming Languages and Systems, 42(1):4:1– 4:??, January 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3363518.

Sabry:1997:RCV

[SW97a] Amr Sabry and Philip Wadler. A reflection on a call-by-value. ACM Transactions on Programming Languages and Systems, 19(6):916–941, November 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-6/
p916-sabry/.

Steckler:1997:LCC

[SW97b] Paul A. Steckler and Mitchell Wand. Lightweight closure conversion. ACM Transactions on Programming Languages and Systems, 19(1):48– 86, January 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-1/p48-steckler/.

Sewell:2010:NPP

[SWU10] Peter Sewell, Paweł T. Wojciechowski, and Asis Unyapoth. Nomadic pict: Programming languages, communication infrastructure overlays, and semantics for mobile computation. ACM Transactions on Programming Languages and Systems, 32 (4):12:1–12:63, April 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Suganuma:2005:DED

[SYK^{+}05] Toshio Suganuma, Toshiaki Yasue, Motohiro Kawahito, Hideaki Komatsu, and Toshio Nakatani. Design and evaluation of dynamic optimizations for a Java just-in-time compiler. ACM Transactions on Programming Languages and Systems, 27(4):732–785, July 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Suganuma:2006:RBC

[SYN06] Toshio Suganuma, Toshiaki Yasue, and Toshio Nakatani. A region-based compilation technique for dynamic compilers. ACM Transactions on Programming Languages and Systems, 28 (1):134–174, January 2006. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Seo:2007:GDW

[SYYH07] Sunae Seo, Hongseok Yang, Kwangkeun Yi, and Taisook Han. Goal-directed weakening of abstract interpretation results. ACM Transactions on Programming Languages and Systems, 29(6):39:1–39:39, October 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Swinehart:1986:SVC

[SZBH86] Daniel Swinehart, Polle Zellweger, Richard Beach, and Robert Hagemann. A structural view of the Cedar programming environment. ACM Transactions on Programming Languages and Systems, 8(4): 419–490, October 1986. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
6466.html.

Sanan:2021:CCT

[SZLY21] David Sanan, Yongwang Zhao, Shang-Wei Lin, and Liu Yang. CSim 2: Compositional topdown verification of concurrent systems using rely-guarantee. ACM Transactions on Programming Languages and Systems, 43 (1):2:1–2:46, April 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3436808.

Terauchi:2008:CCC

[TA08a] Tachio Terauchi and Alex Aiken.

A capability calculus for concurrency and determinism. ACM Transactions on Programming Languages and Systems, 30(5): 27:1–27:30, August 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Terauchi:2008:WSE

[TA08b] Tachio Terauchi and Alex Aiken.

Witnessing side effects. ACM Transactions on Programming Languages and Systems, 30(3): 15:1–15:42, May 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tai:1979:NSG

[Tai79] Kou-Chung Tai. Noncanonical SLR(1) grammars. ACM Transactions on Programming Languages and Systems, 1(2): 295–320, October 1979. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tanenbaum:1983:TCT

[Tan83] Andrew S. Tanenbaum. Technical correspondence: Tanenbaum’s reply. ACM Transactions on Programming Languages and Systems, 5(3):499– 500, July 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [TvS82, Pem83].

Tardieu:2007:DLS

[Tar07]  Olivier Tardieu.  A deter-

ministic logical semantics for pure Esterel. ACM Transactions on Programming Languages and Systems, 29(2):8:1– 8:26, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tsay:1995:DFP

[TB95] Yih-Kuen Tsay and Rajive L. Bagrodia. Deducing fairness properties in UNITY logic — a new completeness result. ACM Transactions on Programming Languages and Systems, 17(1): 16–27, January 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
200997.html.

Tofte:1998:RIA

[TB98] Mads Tofte and Lars Birkedal. A region inference algorithm. ACM Transactions on Programming Languages and Systems, 20(4):724–767, July 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-4/
p724-tofte/.

Trinder:2017:SRI

[TCP^{+}17] Phil Trinder, Natalia Chechina, Nikolaos Papaspyrou, Konstantinos Sagonas, Simon Thompson, Stephen Adams, Stavros Aronis, Robert Baker, Eva Bihari, Olivier Boudeville, Francesco Cesarini, Maurizio Di Stefano, Sverker Eriksson, Viktória Fördős, Amir Ghaffari, Aggelos Giantsios, Rickard Green, Csaba Hoch, David Klaftenegger, Huiqing Li, Kenneth Lundin, Kenneth Mackenzie, Katerina Roukounaki, Yiannis Tsiouris, and Kjell Winblad. Scaling reliably: Improving the scalability of the Erlang distributed actor platform. ACM Transactions on Programming Languages and Systems, 39 (4):17:1–17:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tzannes:2014:LSR

[TCVB14] Alexandros Tzannes, George C.

Caragea, Uzi Vishkin, and Rajeev Barua. Lazy scheduling: a runtime adaptive scheduler for declarative parallelism. ACM Transactions on Programming Languages and Systems, 36 (3):10:1–10:??, September 2014. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Toro:2023:CLT

[TDA^{+}23] Matías Toro, David Darais, Chike Abuah, Joseph P. Near, Damián Árquez, Federico Olmedo, and Éric Tanter. Contextual linear types for differential privacy. ACM Transactions on Programming Languages and Systems, 45(2):8:1–8:??, June 2023. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3589207.

Tip:2011:RUT

[TFK^{+}11] Frank Tip, Robert M. Fuhrer, Adam Kieżun, Michael D. Ernst, Ittai Balaban, and Bjorn De Sutter. Refactoring using type constraints. ACM Transactions on Programming Languages and Systems, 33(3):9:1–9:47, April 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Toro:2018:TDG

[TGT18] Matías Toro, Ronald Garcia, and Éric Tanter. Type-driven gradual security with references. ACM Transactions on Programming Languages and Systems, 40(4):16:1–16:??, December 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See corrigendum [TGT20].

Toro:2020:CTD

[TGT20] Matías Toro, Ronald Garcia, and Éric Tanter. Corrigendum to “Type-driven Gradual Security with References,” by Toro et al., ACM Transactions on Programming Languages and Systems (TOPLAS) Volume 40, Issue 4, Article No. 16. ACM Transactions on Programming Languages and Systems, 42(2): 11:1–11:2, May 2020. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3387725. See

[TGT18].

Thorup:1994:CGA

[Tho94]  Mikkel Thorup.  Controlled

grammatic ambiguity. ACM Transactions on Programming Languages and Systems, 16(3): 1024–1050, May 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177759.html.

Tichy:1986:SR

[Tic86] Walter F. Tichy. Smart recompilation. ACM Transactions on Programming Languages and Systems, 8(3):273– 291, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5959.html.

Tichy:1988:TCT

[Tic88] Walter F. Tichy. Technical correspondence: Tichy’s response to R. W. Schwanke and G. E. Kaiser’s “Smarter Recompilation”. ACM Transactions on Programming Languages and Systems, 10(4):633– 634, October 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214507.html. See [SK88].

Tick:1994:DTN

[TK94] E. Tick and M. Korsloot. Determinacy testing for nondeterminate logic programming languages. ACM Transactions on Programming Languages and Systems, 16(1):3– 34, January 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        174626.html.

Tripakis:2011:TSR

[TLHL11] Stavros Tripakis, Ben Lickly, Thomas A. Henzinger, and Edward A. Lee. A theory of synchronous relational interfaces. ACM Transactions on Programming Languages and Systems, 33 (4):14:1–14:41, July 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tel:1993:DDT

[TM93] Gerard Tel and Friedmann Mattern. The derivation of distributed termination detection algorithms from garbage collection schemes. ACM Transactions on Programming Languages and Systems, 15(1):1– 35, January 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        151647.html.

Thakur:2019:PFP

[TN19] Manas Thakur and V. Krishna Nandivada. PYE: a framework for precise-yet-efficient just-in-time analyses for Java programs. ACM Transactions on Programming Languages and Systems, 41 (3):16:1–16:??, July 2019. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL https://dl.acm.org/ft_
gateway.cfm?id=3337794.

Takisaka:2021:RRS

[TOUH21] Toru Takisaka, Yuichiro Oyabu, Natsuki Urabe, and Ichiro Hasuo. Ranking and repulsing supermartingales for reachability in randomized programs. ACM Transactions on Programming Languages and Systems, 43 (2):5:1–5:46, July 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3450967.

Thammanur:2004:FME

[TP04] Sathyanarayanan Thammanur and Santosh Pande. A fast, memory-efficient register allocation framework for embedded systems. ACM Transactions on Programming Languages and Systems, 26(6):938–974, November 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tratt:2008:DSL

[Tra08]  Laurence Tratt.  Domain

specific language implementation via compile-time metaprogramming. ACM Transactions on Programming Languages and Systems, 30(6):31:1– 31:40, October 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Torp-Smith:2008:LRA

[TSBR08] Noah Torp-Smith, Lars Birkedal, and John C. Reynolds. Local reasoning about a copying garbage collector. ACM Transactions on Programming Languages and Systems, 30(4): 24:1–24:58, July 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tip:2002:PET

[TSL^{+}02] Frank Tip, Peter F. Sweeney, Chris Laffra, Aldo Eisma, and David Streeter. Practical extraction techniques for Java. ACM Transactions on Programming Languages and Systems, 24(6): 625–666, November 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tang:2000:PTR

[TSY00] Hong Tang, Kai Shen, and Tao Yang. Program transformation and runtime support for threaded MPI execution on shared-memory machines. ACM Transactions on Programming Languages and Systems, 22(4):673–700, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/citations/journals/toplas/
                                        2000-22-4/p673-tang/.

Turini:1984:MLO

[Tur84] Franco Turini. Magma2: a language oriented toward experiments in control. ACM Transactions on Programming Languages and Systems, 6(4): 468–486, October 1984. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Turchin:1986:CS

[Tur86] Valentin F. Turchin. The concept of a supercompiler. ACM Transactions on Programming Languages and Systems, 8(3): 292–325, July 1986. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
5957.html.

Thies:2007:STU

[TVA07] William Thies, Frédéric Vivien, and Saman Amarasinghe. A step towards unifying schedule and storage optimization. ACM Transactions on Programming Languages and Systems, 29(6): 34:1–34:45, October 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Tanenbaum:1982:UPO

[TvS82] Andrew S. Tanenbaum, Hans van Staveren, and Johan W. Stevenson. Using peephole optimization on intermediate code. ACM Transactions on Programming Languages and Systems, 4 (1):21–36, January 1982. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See remarks [Pem83].

Thatcher:1982:DTS

[TWW82] J. W. Thatcher, E. G. Wagner, and J. B. Wright. Data type specification: Parameterization and the power of specification techniques. ACM Transactions on Programming Languages and Systems, 4(4):711–732, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Toninho:2018:ISB

[TY18] Bernardo Toninho and Nobuko Yoshida. Interconnectability of session-based logical processes. ACM Transactions on Programming Languages and Systems, 40 (4):17:1–17:??, December 2018. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Toninho:2021:PSF

[TY21] Bernardo Toninho and Nobuko Yoshida. On polymorphic sessions and functions: a tale of two (fully abstract) encodings. ACM Transactions on Programming Languages and Systems, 43 (2):7:1–7:55, July 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3457884.

Tse:2007:RTP

[TZ07] Stephen Tse and Steve Zdancewic. Run-time principals in informationflow type systems. ACM Transactions on Programming Languages and Systems, 30(1):6:1– 6:42, November 2007. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ungar:1992:ATP

[UJ92] David Ungar and Frank Jackson. An adaptive tenuring policy for generation scavengers. ACM Transactions on Programming Languages and Systems, 14 (1):1–27, January 1992. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
116734.html.

Unger:2002:HIL

[UM02] Sebastian Unger and Frank Mueller. Handling irreducible loops: optimized node splitting versus DJ-graphs. ACM Transactions on Programming Languages and Systems, 24(4): 299–333, July 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Ugawa:2018:TSL

[URJ18] Tomoharu Ugawa, Carl G. Ritson, and Richard E. Jones. Transactional Sapphire: Lessons in high-performance, on-the-fly garbage collection. ACM Transactions on Programming Languages and Systems, 40(4):15:1– 15:??, December 2018. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Vera:2005:ACM

[VALG05] Xavier Vera, Jaume Abella, Josep Llosa, and Antonio González. An accurate cost model for guiding data locality transformations. ACM Transactions on Programming Languages and Systems, 27(5):946– 987, September 2005. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

vandenBos:1988:AIT

[van88] Jan van den Bos. Abstract interaction tools: a language for user interface management systems. ACM Transactions on Programming Languages and Systems, 10 (2):215–247, April 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        42191.html.

VanderZanden:1996:CIA

[Van96a] Brad Vander Zanden. Corrigendum: “An incremental algorithm for satisfying hierarchies of multiway dataflow constraints”. ACM Transactions on Programming Languages and Systems, 18(3):354, May 1996. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Van96b].

VanderZanden:1996:IAS

[Van96b] Brad Vander Zanden. An incremental algorithm for satisfying hierarchies of multiway dataflow constraints. ACM Transactions on Programming Languages and Systems, 18(1): 30–72, January 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

225543.html. See corrigendum [Van96a].

Vansummeren:2006:TIU

[Van06] Stijn Vansummeren. Type inference for unique pattern matching. ACM Transactions on Programming Languages and Systems, 28(3):389–428, May 2006. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Vera:2004:FAF

[VBLG04] Xavier Vera, Nerina Bermudo, Josep Llosa, and Antonio González. A fast and accurate framework to analyze and optimize cache memory behavior. ACM Transactions on Programming Languages and Systems, 26 (2):263–300, March 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Venkatesh:1995:ERD

[Ven95] G. A. Venkatesh. Experimental results from dynamic slicing of C programs. ACM Transactions on Programming Languages and Systems, 17(2): 197–216, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201062.html.

VanRoy:1997:MOD

[VHB^{+}97] Peter Van Roy, Seif Haridi, Per Brand, Gert Smolka, Michael Mehl, and Ralf Scheidhauer. Mobile objects in Distributed Oz. ACM Transactions on Programming Languages and Systems, 19(5):804–851, September 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://www.

                                        acm.org:80/pubs/citations/
                                        journals/toplas/1997-19-5/
                                        p804-van_roy/.

vonHanxleden:2000:BCP

[vHK00] Reinhard von Hanxleden and Ken Kennedy. A balanced code placement framework. ACM Transactions on Programming Languages and Systems, 22(5):816–860, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

                                        www.acm.org/pubs/articles/
                                        journals/toplas/2000-22-5/
                                        p816-von_hanxleden/p816-von_
                                        hanxleden.pdf;   http://www.
                                        acm.org/pubs/citations/journals/
                                        toplas/2000-22-5/p816-von_
                                        hanxleden/.

VanDenBrand:2002:CLD

[VHKO02] M. G. J. Van Den Brand, J. Heering, P. Klint, and P. A. Olivier. Compiling language definitions: the ASF+SDF compiler. ACM Transactions on Programming Languages and Systems, 24(4):334–368, July 2002. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

VanderZanden:2001:LLA

[VHM^{+}01] Bradley T. Vander Zanden, Richard Halterman, Brad A. Myers, Rich McDaniel, Rob Miller, Pedro Szekely, Dario A. Giuse, and David Kosbie. Lessons learned about one-way, dataflow constraints in the Garnet and Amulet graphical toolkits. ACM Transactions on Programming Languages and Systems, 23(6):776–796, November 2001. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Verdoolaege:2012:ECS

[VJB12] Sven Verdoolaege, Gerda Janssens, and Maurice Bruynooghe. Equivalence checking of static affine programs using widening to handle recurrences. ACM Transactions on Programming Languages and Systems, 34(3):11:1– 11:35, October 2012. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Vasconcelos:2022:TDM

[VMLY22] Vasco T. Vasconcelos, Francisco Martins, Hugo-Andrés López, and Nobuko Yoshida. A type discipline for message passing parallel programs. ACM Transactions on Programming Languages and Systems, 44(4):26:1– 26:??, December 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3552519.

Volpano:1991:TCS

[Vol91] Dennis M. Volpano. Technical correspondence: Subtypes and quantification. ACM Transactions on Programming Languages and Systems, 13(4):631– 632, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/

214523.html. See [Bur90b, Bur91].

Vukasovic:2023:EPC

[VP23] Maja Vukasovic and Aleksandar Prokopec. Exploiting partially context-sensitive profiles to improve performance of hot code. ACM Transactions on Programming Languages and Systems, 45 (4):20:1–20:??, December 2023. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3612937.

vandenBos:1981:PCB

[vPS81] Jan van den Bos, R. Plasmeijer, and Jan W. M. Stroet. Process communication based on input specifications. ACM Transactions on Programming Languages and Systems, 3(3): 224–250, July 1981. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

VanHentenryck:1995:BTC

[VR95] Pascal Van Hentenryck and Viswanath Ramachandran. Backtracking without trailing in CLP(RLin). ACM Transactions on Programming Languages and Systems, 17(4): 635–671, July 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
210192.html.

Vakar:2022:CCH

[VS22] Matthijs Vákár and Tom Smeding. CHAD: Combinatory Homomorphic Automatic Differentiation. ACM Transactions on Programming Languages and Systems, 44(3):20:1– 20:49, September 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3527634.

VonBank:1994:UMP

[VSS94] D. G. Von Bank, C. M.

Shub, and R. W. Sebesta. A unified model of pointwise equivalence of procedural computations. ACM Transactions on Programming Languages and Systems, 16(6):1842– 1874, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
197402.html.

VanNieuwpoort:2010:SHL

[VWJB10] Rob V. Van Nieuwpoort, Gosia Wrzesińska, Ceriel J. H. Jacobs, and Henri E. Bal. Satin: a high-level and efficient grid programming model. ACM Transactions on Programming Languages and Systems, 32(3):9:1– 9:39, March 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Waddle:1990:PTC

[Wad90] Vance E. Waddle. Production trees: a compact representation of parsed programs. ACM Transactions on Programming Languages and Systems, 12 (1):61–83, January 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        77609.html.

Wallis:1980:ERO

[Wal80] Peter J. L. Wallis. External representations of objects of user-defined type. ACM Transactions on Programming Languages and Systems, 2(2): 137–152, April 1980. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See also corrigendum [Wal81].

Wallis:1981:CER

[Wal81] Peter J. L. Wallis. Corrigendum: “External Representations of Objects of User-Defined Type”. ACM Transactions on Programming Languages and Systems, 3(1):111, January 1981.

CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [Wal80].

Wall:1992:ESD

[Wal92] David W. Wall. Experience with a software-defined machine architecture. ACM Transactions on Programming Languages and Systems, 14(3): 299–338, July 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
129395.html.

Wand:1982:DTC

[Wan82] Mitchell Wand. Deriving target code as a representation of continuation semantics. ACM Transactions on Programming Languages and Systems, 4(3): 496–517, July 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Waters:1983:UFC

[Wat83] Richard C. Waters. User format control in a Lisp prettyprinter. ACM Transactions on Programming Languages and Systems, 5 (4):513–531, October 1983. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Waters:1991:ATS

[Wat91] Richard C. Waters. Automatic transformation of series expressions into loops. ACM Transactions on Programming Languages and Systems, 13(1):52– 98, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
102806.html.

Waters:1994:CBP

[Wat94] Richard C. Waters. Clichébased program editors. ACM Transactions on Programming Languages and Systems, 16(1): 102–150, January 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
174628.html.

Wright:1997:PST

[WC97] Andrew K. Wright and Robert Cartwright. A practical soft type system for Scheme. ACM Transactions on Programming Languages and Systems, 19(1): 87–152, January 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-1/p87-wright/.

Walker:2000:TMM

[WCM00] David Walker, Karl Crary, and Greg Morrisett. Typed memory management via static capabilities. ACM Transactions on Programming Languages and Systems, 22(4):701–771, 2000. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-4/p701-walker/.

Wileden:1990:CEO

[WCW90] Jack C. Wileden, Lori A. Clarke, and Alexander L. Wolf. A comparative evaluation of object definition techniques for large prototype systems. ACM Transactions on Programming Languages and Systems, 12(4): 670–699, October 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

88639.html. See corrigenda [WCW91].

Wileden:1991:CCE

[WCW91] Jack C. Wileden, Lori A. Clarke, and Alexander L. Wolf. Corrigenda: “A Comparative Evaluation of Object Definition Techniques for Large Prototype Systems”. ACM Transactions on Programming Languages and Systems, 13(1):179, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [WCW90].

Webber:1995:OFP

[Web95] Adam Webber. Optimization of functional programs by grammar thinning. ACM Transactions on Programming Languages and Systems, 17(2):293– 330, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201067.html.

Weihl:1989:LAP

[Wei89] William E. Weihl. Local atomicity properties: Modular concurrency control for abstract data types. ACM Transactions on Programming Languages and Systems, 11(2):249– 283, April 1989. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        63518.html.

Weihl:1990:LSA

[Wei90] William E. Weihl. Linguistic support for atomic data types. ACM Transactions on Programming Languages and Systems, 12 (2):178–202, April 1990. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        78944.html.

Wetherell:1982:EDV

[Wet82] C. S. Wetherell. Error data values in the data-flow language VAL. ACM Transactions on Programming Languages and Systems, 4(2):226– 238, April 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Weyuker:1983:ATD

[Wey83] Elaine J. Weyuker. Assessing test data adequacy through program inference. ACM Transactions on Programming Languages and Systems, 5(4):641– 655, October 1983. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wagner:1998:EFI

[WG98] Tim A. Wagner and Susan L.

Graham. Efficient and flexible incremental parsing. ACM Transactions on Programming Languages and Systems, 20(5): 980–1013, September 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-5/
p980-wagner/.

Widom:1992:TBN

[WGS92] Jennifer Widom, David Gries, and F. B. Schneider. Tracebased network proof systems: Expressiveness and completeness. ACM Transactions on Programming Languages and Systems, 14(3):396–416, July 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL http://www.acm.
org/pubs/toc/Abstracts/0164-
0925/129396.html. See corri-

gendum [WGS93].

Widom:1993:CTB

[WGS93] Jennifer Widom, David Gries, and F. B. Schneider. Corrigendum: “Trace-Based Network Proof Systems: Expressiveness and Completeness”. ACM Transactions on Programming Languages and Systems, 15(1): 208, January 1993. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [WGS92].

Whalley:1994:AIC

[Wha94]  David B. Whalley.  Auto-

matic isolation of compiler errors. ACM Transactions on Programming Languages and Systems, 16(5):1648–1659, September 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        186103.html.

Williams:1982:DAF

[Wil82a] John H. Williams. On the development of the algebra of functional programs. ACM Transactions on Programming Languages and Systems, 4(4):733– 757, October 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Williams:1982:FNS

[Wil82b] M. Howard Williams. A flexible notation for syntactic definitions. ACM Transactions on Programming Languages and Systems, 4(1):113–119, January 1982. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Winner:1984:UO

[Win84] Robert I. Winner. Unassigned objects. ACM Transactions on Programming Languages and Systems, 6(4):449–467, October 1984. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wing:1987:WLI

[Win87]  Jeannette M. Wing.  Writ-

ing Larch interface language specifications. ACM Transactions on Programming Languages and Systems, 9(1):1– 24, January 1987. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
10500.html.

Wirth:1988:TE

[Wir88] Niklaus Wirth. Type extensions.

ACM Transactions on Programming Languages and Systems, 10 (2):204–214, April 1988. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

46167.html. See remarks [Coh91, Wir91].

Wirth:1991:TCR

[Wir91] Niklaus Wirth. Technical correspondence: Reply to “Typeextension tests can be performed in constant time”. ACM Transactions on Programming Languages and Systems, 13(4): 630, October 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
214521.html.  See [Wir88,

Coh91].

Wise:1979:MGC

[Wis79] David S. Wise. Morris’s garbage compaction algorithm restores reference counts. ACM Transactions on Programming Languages and Systems, 1(1):115– 120, July 1979. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wright:1998:PSE

[WJ98] Andrew K. Wright and Suresh Jagannathan. Polymorphic splitting: an effective polyvariant flow analysis. ACM Transactions on Programming Languages and Systems, 20(1): 166–207, January 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1998-20-1/
p166-wright/.

Wellings:2000:IOO

[WJS^{+}00] A. J. Wellings, B. Johnson, B. Sanden, J. Kienzle, T. Wolf, and S. Michell. Integrating object-oriented programming and protected objects in Ada 95. ACM Transactions on Programming Languages and Systems, 22(3):506– 539, 2000. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/2000-22-3/
p506-wellings/p506-wellings.
pdf;   http://www.acm.org/
pubs/citations/journals/toplas/
2000-22-3/p506-wellings/.

Wand:2004:SAD

[WKD04] Mitchell Wand, Gregor Kiczales, and Christopher Dutchyn. A semantics for advice and dynamic join points in aspectoriented programming. ACM Transactions on Programming Languages and Systems, 26(5): 890–910, September 2004. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Weihl:1985:IRA

[WL85] William Weihl and Barbara Liskov. Implementation of resilient, atomic data types. ACM Transactions on Programming Languages and Systems, 7(2): 244–269, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
3319.html.

Wagner:2016:TIB

[WLBF16] Gregor Wagner, Per Larsen, Stefan Brunthaler, and Michael Franz. Thinking inside the box: Compartmentalized garbage collection. ACM Transactions on Programming Languages and Systems, 38(3):9:1–9:??, May 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Walicki:1995:CCM

[WM95] Michal Walicki and Sigurd Meidal. A complete calculus for the multialgebraic and functional semantics of nondeterminism. ACM Transactions on Programming Languages and Systems, 17 (2):366–393, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        201070.html.

Wu:2012:STB

[WM12] Xing Wu and Frank Mueller. ScalaExtrap: Trace-based communication extrapolation for SPMD programs. ACM Transactions on Programming Languages and Systems, 34(1):5:1– 5:29, April 2012. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Weimer:2008:ESP

[WN08] Westley Weimer and George C. Necula. Exceptional situations and program reliability. ACM Transactions on Programming Languages and Systems, 30(2): 8:1–8:51, March 2008. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wolf:1992:GEI

[Wol92] Alexander L. Wolf. Guest Editor’s introduction to the special section on the Third International Conference on Computer Languages. ACM Transactions on Programming Languages and Systems, 14(2):145– 146, April 1992. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        128863.html.

Wolfe:1994:DDD

[Wol94] Michael Wolfe. The definition of dependence distance. ACM Transactions on Programming Languages and Systems, 16(4): 1114–1116, July 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
183440.html.

Welch:2010:SCF

[WP10] Peter H. Welch and Jan B. Pedersen. Santa Claus: Formal analysis of a process-oriented solution. ACM Transactions on Programming Languages and Systems, 32(4):14:1–14:37, April 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wang:2008:DSJ

[WR08] Tao Wang and Abhik Roychoudhury. Dynamic slicing on Java bytecode traces. ACM Transactions on Programming Languages and Systems, 30(2):10:1– 10:49, March 2008. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Whitfield:1997:AEC

[WS97] Deborah L. Whitfield and Mary Lou Soffa. An approach for exploring code improving transformations. ACM Transactions on Programming Languages and Systems, 19(6):1053– 1084, November 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-6/
p1053-whitfield/.

Wang:2015:EAS

[WSH15] Farn Wang, Sven Schewe, and Chung-Hao Huang. An extension of ATL with strategy interaction. ACM Transactions on Programming Languages and Systems, 37(3):9:1– 9:??, June 2015. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wall:1985:TCN

[WST85] D. Wall, A. Srivastava, and F. Templin. Technical correspondence: a note on Hennessy’s “Symbolic Debugging of Optimized Code”. ACM Transactions on Programming Languages and Systems, 7(1): 176–181, January 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        215005.html.

Wehr:2011:JIT

[WT11] Stefan Wehr and Peter Thiemann. JavaGI: The interaction of type classes with interfaces and inheritance. ACM Transactions on Programming Languages and Systems, 33(4): 12:1–12:83, July 2011. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wu:2004:ETC

[Wu04] Pei-Chi Wu. On exponentialtime completeness of the circularity problem for attribute grammars. ACM Transactions on Programming Languages and Systems, 26(1):186–190, January 2004. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wu:1995:WCC

[WW95] Pei-Chi Wu and Feng-Jian Wang. A worst case of circularity test algorithms for attribute grammars. ACM Transactions on Programming Languages and Systems, 17(2):228– 232, March 1995. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
201064.html.

Wegman:1991:CPC

[WZ91] Mark N. Wegman and F. Kenneth Zadeck. Constant propagation with conditional branches. ACM Transactions on Programming Languages and Systems, 13 (2):181–210, April 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
103136.html.

Ward:2007:SPT

[WZ07] Martin Ward and Hussein Zedan. Slicing as a program transformation. ACM Transactions on Programming Languages and Systems, 29(2):7:1– 7:53, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Xie:2007:SSF

[XA07] Yichen Xie and Alex Aiken. Saturn: a scalable framework for error detection using Boolean satisfiability. ACM Transactions on Programming Languages and Systems, 29(3):16:1–16:43, May 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Xie:2020:CSA

[XBOS20] Ningning Xie, Xuan Bi, Bruno C. D. S. Oliveira, and Tom Schrijvers. Consistent subtyping for all. ACM Transactions on Programming Languages and Systems, 42(1):2:1– 2:??, January 2020. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL https://dl.acm.org/ft_
                                        gateway.cfm?id=3310339.

Yemini:1985:MVE

[YB85] Shaula Yemini and Daniel M. Berry. A modular verifiable exception-handling mechanism. ACM Transactions on Programming Languages and Systems, 7 (2):214–243, April 1985. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        3320.html.

Yemini:1987:ATE

[YB87] Shaula Yemini and Daniel M. Berry. An axiomatic treatment of exception handling in an expression-oriented language.

ACM Transactions on Programming Languages and Systems, 9 (3):390–407, July 1987. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/

24052.html. See remarks [YB88].

Yemini:1988:TCA

[YB88] Shaula Yemini and Daniel M.

Berry. Technical correspondence: “An Axiomatic Treatment of Exception Handling in an Expression-Oriented Language”. ACM Transactions on Programming Languages and Systems, 10(3):503–504, July 1988. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). See [YB87].

Yiapanis:2016:CDS

[YBL16] Paraskevas Yiapanis, Gavin Brown, and Mikel Luján. Compiler-driven software speculation for thread-level parallelism. ACM Transactions on Programming Languages and Systems, 38(2):5:1–5:??, January 2016. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yang:1998:STE

[YF98] Tao Yang and Cong Fu. Space/ time-efficient scheduling and execution of parallel irregular computations. ACM Transactions on Programming Languages and Systems, 20(6):1195– 1222, November 1998. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic).  URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/toplas/1998-20-6/
                                        p1195-yang/.

Yardimci:2009:MSP

[YF09] Efe Yardimci and Michael Franz. Mostly static program partitioning of binary executables. ACM Transactions on Programming Languages and Systems, 31(5): 17:1–17:46, June 2009. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yip:2023:SDP

[YGRBA23] Eugene Yip, Alain Girault, Partha S. Roop, and Morteza Biglari-Abhari. Synchronous deterministic parallel programming for multi-cores with ForeC. ACM Transactions on Programming Languages and Systems, 45 (2):11:1–11:??, June 2023. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3591594.

Ying:2011:FHL

[Yin11] Mingsheng Ying. Floyd–Hoare logic for quantum programs. ACM Transactions on Programming Languages and Systems, 33 (6):19:1–19:49, December 2011. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yu:1997:NCI

[YK97] Ting Yu and Owen Kaser. A note on “On the conversion of indirect to direct recursion”. ACM Transactions on Programming Languages and Systems, 19(6): 1085–1087, November 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic).  URL http://www.
acm.org:80/pubs/citations/
journals/toplas/1997-19-6/
p1085-yu/.

Yang:1997:SMC

[YMW97] Jin Yang, Aloysius K. Mok, and Farn Wang. Symbolic model checking for event-driven realtime systems. ACM Transactions on Programming Languages and Systems, 19(2):386– 412, March 1997. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-2/p386-yang/.

Yoshida:2022:ISI

[Yos22]  Nobuko Yoshida.  Introduc-

tion to the special issue on ESOP 2021. ACM Transactions on Programming Languages and Systems, 44(3): 13:1, September 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3524730.

Yu:1994:LTS

[YR94] Lin Yu and Daniel J. Rosenkrantz. A linear-time scheme for version reconstruction. ACM Transactions on Programming Languages and Systems, 16(3): 775–797, May 1994. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
177705.html.

Yellin:1991:ILI

[YS91] Daniel M. Yellin and Robert E. Strom. INC: a language for incremental computations. ACM Transactions on Programming Languages and Systems, 13(2): 211–236, April 1991. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
103137.html.

Yellin:1997:PSC

[YS97] Daniel M. Yellin and Robert E. Strom. Protocol specifications and component adaptors. ACM Transactions on Programming Languages and Systems, 19(2): 292–333, March 1997. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/citations/journals/toplas/
1997-19-2/p292-yellin/.

Young:1999:SCB

[YS99] Cliff Young and Michael D. Smith. Static correlated branch prediction. ACM Transactions on Programming Languages and Systems, 21(5): 1028–1075, September 1999. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic). URL http://

www.acm.org/pubs/articles/
journals/toplas/1999-21-5/
p1028-young/p1028-young.pdf;
 http://www.acm.org/pubs/
citations/journals/toplas/
1999-21-5/p1028-young/.

Yahav:2010:VSP

[YS10] Eran Yahav and Mooly Sagiv.

Verifying safety properties of concurrent heap-manipulating programs. ACM Transactions on Programming Languages and Systems, 32(5):18:1–18:50, May 2010. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yang:2002:EEB

[YUW02] Minghui Yang, Gang-Ryung Uh, and David B. Whalley. Efficient and effective branch reordering using profile data. ACM Transactions on Programming Languages and Systems, 24(6): 667–697, November 2002. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Yang:2022:DDZ

[YW22] Albert Mingkun Yang and Tobias Wrigstad. Deep dive into ZGC: a modern garbage collector in OpenJDK. ACM Transactions on Programming Languages and Systems, 44(4):22:1– 22:??, December 2022. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3538532.

Zave:1985:DAF

[Zav85] Pamela Zave. A distributed alternative to finite-state-machine specifications. ACM Transactions on Programming Languages and Systems, 7(1):10– 36, January 1985. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

                                        URL  http://www.acm.org/
                                        pubs/toc/Abstracts/0164-0925/
                                        2365.html.

Zhao:2007:FFS

[ZCG^{+}07] Peng Zhao, Shimin Cui, Yaoqing Gao, Raúl Silvera, and José Nelson Amaral. Forma: a framework for safe automatic array reshaping. ACM Transactions on Programming Languages and Systems, 30(1):2:1–2:30, November 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Zhang:2005:CPT

[ZGZ05] Xiangyu Zhang, Rajiv Gupta, and Youtao Zhang. Cost and precision tradeoffs of dynamic data slicing algorithms. ACM Transactions on Programming Languages and Systems, 27(4): 631–661, July 2005. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Zhou:1996:PPC

[Zho96] Neng-Fa Zhou. Parameter passing and control stack management in Prolog implementation revisited. ACM Transactions on Programming Languages and Systems, 18(6):752– 779, November 1996. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/toplas/
235832.html;   http://www.
acm.org/pubs/toc/Abstracts/
toplas/236120.html.

Zic:1994:TCB

[Žic94] J. J. Žic. Time-constrained buffer specifications in CSP+T and timed CSP. ACM Transactions on Programming Languages and Systems, 16(6):1661– 1674, November 1994. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  http://www.acm.org/
pubs/toc/Abstracts/0164-0925/
197322.html.

Zhang:2017:SSH

[ZMVPJ17] Danfeng Zhang, Andrew C. Myers, Dimitrios Vytiniotis, and Simon Peyton-Jones. SHErrLoc: a static holistic error locator. ACM Transactions on Programming Languages and Systems, 39 (4):18:1–18:??, September 2017. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Zhuang:2007:AAR

[ZP07] Xiaotong Zhuang and Santosh Pande. Allocating architected registers through differential encoding. ACM Transactions on Programming Languages and Systems, 29(2):9:1–9:46, April 2007. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Zhuang:2010:OFE

[ZP10] Xiaotong Zhuang and Santosh Pande. An optimization framework for embedded processors with auto-addressing mode. ACM Transactions on Programming Languages and Systems, 32 (4):11:1–11:41, April 2010. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Zhong:2009:PLA

[ZSD09] Yutao Zhong, Xipeng Shen, and Chen Ding. Program locality analysis using reuse distance. ACM Transactions on Programming Languages and Systems, 31(6):20:1–20:39, August 2009. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Zhang:2021:CP

[ZSO21] Weixin Zhang, Yaozhu Sun, and Bruno C. D. S. Oliveira. Compositional programming. ACM Transactions on Programming Languages and Systems, 43(3): 9:1–9:61, September 2021. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

                                        tronic). URL https://dl.acm.
                                        org/doi/10.1145/3460228.

Zhao:2020:DLS

[ZSS20] David Zhao, Pavle Subotić, and Bernhard Scholz. Debugging large-scale Datalog: a scalable provenance evaluation strategy.

ACM Transactions on Programming Languages and Systems, 42 (2):7:1–7:35, May 2020. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL  https://dl.acm.org/
doi/abs/10.1145/3379446.

Zhou:2022:RIR

[ZZO22] Yaoda Zhou, Jinxu Zhao, and Bruno C. D. S. Oliveira. Revisiting iso-recursive subtyping. ACM Transactions on Programming Languages and Systems, 44 (4):24:1–24:??, December 2022. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (elec-

tronic). URL https://dl.acm.
org/doi/10.1145/3549537.